A Complete Bibliography of Publications in Journal of the History of Biology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

08 October 2025 Version 1.22

Title word cross-reference

$1100 [Bel05]. $112.50 [Ban22]. $14.99 [Loz19]. $15.95 [Tob09]. $17.00 [Gia18]. $17.95 [Irm05]. $179.00 [Mét14]. $20.00 [Zal05]. $22.95 [Ano04a, Irm05]. $23.40 [Goo19]. $24.95 [Bru18, McG18]. $25 [Can13, Sel05]. $25.95 [Yan18]. $27.00 [Ano06b, Ste05]. $27.50 [Kin20, Ile05]. $27.95 [You09]. $28.00 [Des05]. $29.95 [Som21]. $30 [dC14]. $30.00 [Ank20, Mat21, Kei11b]. $32.99 [Coo20]. $34.95 [Hub20, Zal05]. $35.00 [Gor19, Gru18, Pow18, Sch18, Shi18, Ben09, Kev05, Kle11, Mil11]. $37.50 [Ano04a]. $37.62 [SD19a]. $39.00 [Mar09]. $40.00 [Lan20, McG21, Mon18, Sat20]. $45.00 [Cau19, Kim18, Nyh18, Tan20b, Wu19, Wei11]. $47.50 [Lig18]. $49.00 [Rab11]. $49.95 [Ber19, Ano05a, Cro05]. $50.00 [HS18]. $55.00 [Bra20, Mes19]. $59.95 [Far05]. $65.00 [Bri09]. $67.72 [Ank18]. $85.00 [Ben18]. $90 [dC14]. $99 [Wal20]. $99.95 [Lin18]. $99.99 [Cre19]. _{2} [Sin97]. X [Des01].

* [Ile05, Joh07].

-Fixation [Sin97].

0 [Ank18, Kin20, Ray19, SD19a].

0-226-82301-6 [Bar23a].

1 [Bru18, Loz19, Mat21, Sat20].

1-5343-0154-2 [Loz19]. 16.99 [Bar05]. 1793 [Por25]. 1800 [Bra20]. 1819 [Ste05]. 1820 [May23]. 1830 [Ste19]. 1830s [FE20].

1834 [Tan20a]. 1840s [FE20]. 1857 [Bel05]. 1858 [Nov23]. 1860s [Moo91]. 18th [Eli14, KFP09, dC23]. 19 [Ras24a]. 1914 [Shm23a, Shm23b]. 1920s [Erl09, Erl16, Hal76]. 1930s [Arm12, Car08, Har12b, Lam15, Ras99, San10, Thi01]. 1933 [WY23]. 1936 [Luk22]. 1940s [Kle99, San16]. 1941 [Kle13]. 1950 [Har22b, Ona23]. 1950s [Dan21, De 96, Dua20, Lam15, Mil08, Ras95, dC06]. 1960 [Tyr23]. 1960s [Car08, Dan21, Den16, Gro13, Pal91, San16, Sto15]. 1963 [Joh09]. 1970s [Pal91]. 1980 [Wu19]. 1980s [Gro13]. 1984 [Ric25]. 19halftones [dC14]. 19th [dMP23, Wie06, Bre02, LB04, Sch04, Tam15].

2 [Cau19, Goo19, HS18, Loz19, Moh18, Shi18, Som21, Wal20]. 2002/Princeton [Ano04a]. 2010 [Mét14]. 2010s [Sto15]. 2017 [Ano18, Ton18]. 2019 [Ano19]. 2020 [Mar24, RR20a]. 2021 [RR21a]. 2022 [RR22a]. 2023 [RR23]. 2024 [SR24]. 2025 [Smo25a]. 20th [Kev05, Eli14, Gli06, LMB03, Sep19, Sla03, Wie06, ZS15]. 2nd [Bod23].

3 [Ban22, Bra20, Lan20, McG18, Tan20b]. 336 [Ras24b].

4 [Ber19, Hub20, Yan18]. 40pp [Cau19, Gor19, SD19a].

5 [Gor19, Mar24, Mes19, Muk19].

6 [Gia18, Kim18, Lig18, McG21, Mon18, Sch18]. 66 [Bar23a].

7 [Ank20, Ben18, Lin18, Nyh18, Pow18, Web20, Wu19].

8 [Coo20, Cre19, Gru18].

978 [Ank20, Ank18, Ban22, Ben18, Ber19, Bra20, Bru18, Cau19, Coo20, Cre19, Gia18, Goo19, Gor19, Gru18, HS18, Hub20, Kim18, Kin20, Lan20, Lig18, Lin18, Mar24, Mat21, McG21, McG18, Mes19, Moh18, Mon18, Muk19, Nyh18, Pow18, Ray19, Sat20, Sch18, Shi18, Som21, SD19a, Tan20b, Wal20, Web20, Wu19, Yan18]. 978-0-14-218082-2 [Moh18]. 978-0-14-312584-6 [Gia18]. 978-0-226-11029-5 [Mes19]. 978-0-226-15193-9 [Fre25]. 978-0-226-19275-8 [Gru18]. 978-0-226-39008-6 [Kim18]. 978-0-226-39543-2 [Cau19]. 978-0-226-41647-2 [Shi18]. 978-0-226-43690-6 [Lig18]. 978-0-226-44998-2 [HS18]. 978-0-226-52079-7 [Nyh18]. 978-0-226-53832-7 [Pow18]. 978-0-226-54608-7 [Ank20]. 978-0-226-56312-1 [Mat21]. 978-0-226-56990-1 [Sat20]. 978-0-226-61798-5 [Gor19]. 978-0-226-63504-0 [SD19a]. 978-0-226-63809-6 [McG21]. 978-0-226-65088-3 [Lan20]. 978-0-226-68508-3 [Ban22]. 978-0-226-70629-0 [Kin20]. 978-0-226-71361-8 [End23]. 978-0-226-73031-8 [Bar24]. 978-0-226-74485-8 [Gue23]. 978-0-226-81923-5 [Liu23]. 978-0-226-82272-3 [Gli24]. 978-0-226-82413-0 [Jia24]. 978-0-226-82532-8 [Mor23]. 978-0-226-82877-0 [Nat24]. 978-0-226-83448-1 [You25]. 978-0-231-18070-2 [Hei19]. 978-0-231-18071-9 [Hei19]. 978-0-262-03586-6 [Sch18]. 978-0-262-04646-6 [Lin23]. 978-0-300-20961-7 [Ben18]. 978-0-300-24791-6 [Cao24]. 978-0-300-26722-8 [Hag24]. 978-0-520-29778-4 [Hub20].

978-0-520-39040-9 [Bis23]. 978-0-674-08905-1 [Bru18]. 978-0-674-28136-3 [Tan20b]. 978-0-674-97942-0 [Kei23]. 978-0-674-98347-2 [Stu24]. 978-0-691-18188-2 [Som21]. 978-0-691-23379-6 [Fic24]. 978-0-691-24736-6 [Gre25]. 978-0-7735-5312-5 [Muk19]. 978-0-8070-5551-9 [Pau23]. 978-0-8139-5219-2 [Por25]. 978-0-8173-2089-8 [Kim23b]. 978-0-8223-5894-7 [Lin18]. 978-0-8223-6345-3 [McG18]. 978-0-8223-6872-4 [Yan18]. 978-0-8229-4472-0 [Ray19]. 978-0-8229-4474-4 [Ber19]. 978-0-8229-4527-7 [Wu19]. 978-0-8229-4553-3 [Bra20]. 978-0-8229-4627-4 [Rad23]. 978-0-8229-4630-4 [Boo23]. 978-0-8229-4654-0 [Str23]. 978-0-8229-4661-8 [Tyr23]. 978-0-8229-4708-0 [Bro24]. 978-0-8229-6634-0 [Nov23]. 978-1-032-03548-2 [Ric25]. 978-1-032-16223-2 [Gue24]. 978-1-032-36532-9 [Kle24]. 978-1-032-75714-8 [Jin25]. 978-1-107-19144-0 [Ank18]. 978-1-108-45737-8 [Coo20]. 978-1-108-48415-2 [Wal20]. 978-1-108-49009-2 [Shm23a, Shm23b]. 978-1-108-84216-7 [May23]. 978-1-4214-2658-7 [Web20]. 978-1-4214-4405-5 [Bod23]. 978-1-4214-4655-4 [Gib24]. 978-1-4214-4678-3 [Ras24a]. 978-1-4314-2425-2 [Goo19]. 978-1-4696-7590-9 [Koh24]. 978-1-4780-2753-9 [Som24]. 978-1-80511-265-5 [Mar24]. 978-1-942130-81-9 [Ras24b]. 978-3-030-74861-6 [Fla24].

978-3-031-06129-5 [Ste24]. 978-3-031-40164-0 [Hal25]. 978-3-319-64336-6 [Mon18]. 978-3-319-69717-8 [Cre19].

A. [Ank00, Ano78e, Mit88, Ste84]. A.-P. [Ste84]. Abandoned [Mat06]. Abbott [Kin78]. ABCC [Hat21]. Abena [Coo20]. Abigail [Mon18]. abiogenesis [Far72c]. Abnormal [Mos98]. Absentia [Sto15]. academic [Pau84]. Academy [Hea20, Por79]. acceptance [Ano76b, Ano91a]. Access [Ano20]. Accessory [Bra11b]. Accord [Hol17]. Achievements [Ege83b, Ege85]. Across [CdC25, LK25, Ste24, Cre03, May23]. Action [Swa70a, Ano75a, Swa68b]. Activist [Gor09]. activities [Bro71]. Acts [Hic21]. Adams [Ile99]. Adaptation [Hag17, Kim23b, Loi25, Rai03, Eds72]. Adaptationist [Tam20a]. adaptedness [Mic86]. Adaptive [Gro17a]. adaptiveness [Kim86]. addendum [Coo90]. Additions [Mor97b, Bro73]. adenosine [Ano91d]. Adjustment [vW14]. Admixture [Bur22]. Adolf [RWE13]. Adrian [Mun05]. Advancement [ZS15]. Advantage [Kim86]. Advent [Yi08]. Advice [Sum04]. advocacy [Hal07, McG21]. advocate [Hal76]. Advocating [Mai01]. Aesthetic [Ste02]. Affair [Cas12, Köh11]. Affinity [vP21]. Africa [Ben18, Coo20, Ank04, Ank18, Bro11, HR11]. African [Car08, Dan21]. Africanizing [Lew18]. After [Bel05, Hag15, Hag17, Mit21, Veu19, Cog02, Coo20, Lin18, Osp79, Tay88]. Aftermath [Far16, Sul82b]. Against [Hod05, Jin25, Mor97b, Rue84]. Agassiz [Irm05, Irm05, JK99, Mor97b, Wei68]. Age [CS06, Kin09b, McG21, Por04, Thu02, Bar00, Par21, Str23]. ageing [Wal20]. ages [Por85]. Aging [Jia14]. Agnes [Hen03, Str24]. Agricultural [Kim06, Par15, Wie06]. Agriculture [Ack06, Bon06, Har06b, Hen97, All91, Fit19]. Agrostology [Hen03]. aid [Ste04]. Ain’t [Smo05b]. Air [Por09]. Alabama [Kim23b]. Alberch [Art24]. Albert [Sin00, SS17, Thi01]. Albino [Log01]. Albrecht [Ano75j]. alcohol [Pau96]. Aldo [You25]. Aleksandr [Kre07]. Alexander [FG78, FE20, Gre25, Mir22, Rad06, Ste97]. Alfonso [LMB03]. Alfred [Fic24, Gre04, Hic21, Rhe00, Shm20, Smi15, Sta24]. Alfredo [Ras24a]. Algae [Hut22, Kol22, Zal93a]. Algology [Hut22]. Alisha [Gue23]. Allee [Mit88]. allelopathy [Wil85]. Allen [Ric23, Bar97, Die16b, Kle23, Mai16, Mai23b, Men16, SR23b]. allergy [Smi92]. Alliance [SD19c, SD19b]. Alliances [Oak15]. Allied [Ile99]. Alligator [Bar09]. Allostatic [Arm16]. Aloi [Hei19]. Alpheus [Boc88]. Alphonse [Pir24, ÁTBHBJB24]. Alternative [Hal98]. Altruism [Har11a]. am [Raa76]. Amateur [Gal22, Rad04]. Amateurs [Alb01, Des01]. Amazon [vW14]. ambiguous [Mar92]. Amboseli [Lew18]. America [Bar24, Ona23, Rab11, Ano81f, Ano00c, App88, Bar00, Ben09, Bru02, Hen97, Kim18, KFP09, Lin23, Pau84, Rum12, Sel12, Som21, Stu24, Thu02]. American [Bar19, Dav97a, Fit19, Joh09, Led04, Mar04, Por09, Ano81b, Ano81d, Ano81h, Bar09, Ben85, Ben88, CC18, Coo98, Coo02, Har03, Hen03, Ile99, Joh15, Kei17, Lad08, Lar87, Lud69, Man09, McC69, Mit17, Par21, Pat09, Pau16, Por04, Sep05, Whi20]. Americans [Rai01]. Ammal [Ric25]. Amoebae [Rey08]. Among [PH22]. Amoy [Luk22, Luk21]. Analogies [CC18, KK13]. Analogy [Bec05, Cor84, The12, dC23, Eva84]. Analysis [ÁTBHBJB24, Elw07, Gru22, Par15, SD14, Ano75d, Bow71, Fal91, Par82]. analytical [Kay85]. Anatomical [Gam21, Ray22]. Anatomy [Cos97, Ric89a, App88, Raa76]. Anaximander [KK13]. Ancestor [GG97, dV09]. ancestral [Ano78b]. Ancestry [Veu19]. Ancient [Hug11, Nat24, Som05, Ano75m]. Anctil [Muk19]. and/or [Gru22]. Anderson [Kle99, Kle13, Kle16]. André [Por25]. Andrea [For20]. Andreas [Gre25]. Andrei [Bar24]. Andrew [Gue24, Mat21]. Angela [Mon18, Joh19b]. Animal [Ank20, Bla07, BJ97, Esp13, FG17, Gor19, Gro19, Hei19, KFP09, Lev18, dMP23, Ray19, Tho05, Vic18, Vic23, App80, Ege68a, Ege68b, Fra91, Hal71, Kev93, Rhe96, But18, Rad04]. animalcules [San73]. Animals [Ack04, Bod23, De 19, Gam21, Gib24, GM13, Hag17, Har13, Led04, Log02, Lyn19, Mel13, dMP23, Pre70, Tho05, Vor69, Mon18]. Anita [Bod23]. Ann [Ste05]. Anne [Bar24, Cau19, Kim18]. Annie [Ste97]. Annotated [NL11, Bed88b]. Announcement [Ano82a, RR20b, Ano88a, Ano88b, Ano89b]. announcements [Ano88c]. Announcing [Ano85a]. Anomaly [Sch99a]. Anonymous [Cor21]. Anschauung [Sch04]. Ant [Buh00]. Antebellum [Par21]. Anthrax [Gor97]. Anthropocene [Hei19, Lyn19]. anthropological [Atr87]. Anthropologists [Hum08]. Anthropology [Hag10, Hag17, LK25, Mar24, Som05, Som08, Wei06, dR23]. Anti [Bel05, Erl02, Ona23, Par00, Wal20, Wei96]. Anti-ageing [Wal20]. Anti-Asian [Ona23]. Anti-Creationism [Par00]. Anti-Evolution [Bel05]. Anti-Haeckelian [Erl02]. anti-progressive [Wei96]. Antibiotic [San16]. antibodies [KO91]. antibody [Söd94]. Antievolution [Sha08]. Antiquity [Byn84]. antiwar [Cro89]. Anton [GG97, Mat10]. Anxieties [Bur22]. appearance [Pau84]. apple [OV68]. Application [Alt07, Ano76d]. Applications [Swi14]. applied [Pal91]. appointment [Raa76]. Appraisal [Ste02]. appraisals [Wer71]. Appreciation [Kle23, Men16]. Approach [ÁTBHBJB24, Com11, Dre23, ML10, Pir23, Ras24a, KC94]. Approaches [Fin08, Jia14, Pre13]. Appropriating [Cai02]. Appropriations [Cec24]. April [Ton18]. Aquarium [LM14, Pow18, Ham86]. aquatic [Boc90]. Arber [Str24]. Arbor [Ste05]. archival [Ano82m]. Archive [Ano88d, Wol11, Ano78e, Fra79, Ano82i, BA81, FG78]. Archives [Bur16, Ano75c, Ano82i]. Arctic [Hag17, Lue21]. Area [Sch25]. argument [Bow77, Lyo93]. Arguments [Mor97b]. Arid [HR11]. Aristotelian [Rie90]. Aristotle [Bod23, Des05, Gue24, Ano82b, Bod83, Boy84, De 19, Fra91, Gaz12, Got99, Gre92, Hen06, Hor76, Len80, Mor79, Mun93, Pre70, Pre77, SK22, Sha72, Zwi18]. arithmetic [Bro80]. Ark [Pow18]. Arkansas [Ste05]. Armstrong [Ano05a]. Aron [Pas21]. Art [Bae19a, Bae19b, Har11b, Hei19, Mes19]. Arteries [Boy07]. Arthrobalanus [Buc17]. arthropoda [Bow94]. Arthur [Ano75a, Eng97, Rus15]. Articles [Ano20]. Artificial [The12, Eva84]. Arturo [Kle04]. Aryan [Rie12]. Asajiro [Sul11]. Asare [Coo20]. Ascent [Ste05]. Ashkenazi [Bur22]. Asia [Joh19b]. Asiacentrism [pY14]. Asian [Ona23]. Asilomar [CC25, CdC25, Cob25, Sch25, Yi25]. aspect [Bur93a]. Aspects [Bur93b, Dex70, GT69]. Assay [Sch14]. Assembling [Hun19c]. Assimilation [Ser16]. Assistants [AdC24]. Assumptions [Nov23, Ano81h]. Astrobiology [Cro05]. Atavism [Hob15]. Athens [Por09]. Atlantic [May23, CdC25, Cre03]. Atlas [Str10]. Atmungsferment [Koh73]. atom [Boc95]. Atomic [Cre06, CS06, Ham07, Hoo97, Kin09b, Kli98, SP25, Bea93, Coo20, Hat21]. ATP [All02, Pre13]. Attempt [Mat06]. Attempts [Dre20]. Auckland [Goo19]. August [Chu68, Chu10, Sta03, Wei10, Win01b].

Augustin [Ste04]. Augustin-Pyramus [Ste04]. Aurum [Bar05]. Australia [Som24]. Australian [Ben11a]. Authority [Gue23, Mir11, Sap83, Cre19]. Authorship [Cor21]. Autistic [Pau23]. Autobiography [Ham99, Col85]. Autochthonous [Ras22]. Axel [Die04]. Axiomatic [vdBD20]. Axis [Des01]. Ayo [Hub20]. Azara [Ano75g, Ano75l, Bed83b].

B [Ada00, Ano76b, Ano76d, Ano82i, BA81, FG78, Kle24, Pau83, RN11, Ben18, Cre19, Som21]. b&w [Ank20, Ber19, Cau19, Coo20, Hub20, Kin20, Lan20, McG18, Yan18]. B. [AKB89, ASR^{+}93, Ano69a, Ano78e, Ano78f, Ano80a, Ano80b, Ano82j, Ano82k, Ano84b, Ano85b, Ano85c, Ano91g, Ano92, Ano96, Ano98b, BBF^{+}91, CRBD90, CHCC97, FHP^{+}91, FRS^{+}95, FMR^{+}98, FH77, FH78, Fra79, FH79, GLB^{+}94, HS72, HF73, HF74, Hal74, HLM^{+}95, KNM^{+}93, MAC^{+}94, MFM^{+}96, NSMB92, PKS^{+}95, RHW^{+}97, RG76, RFL^{+}98, RGH^{+}81, Roe82, RS83, Roe83, Roe84, RBFC84, RRC^{+}85, Roe86, RCK^{+}86, RBK^{+}86, RBHK87, RRE^{+}87, RPCH87, RRCB88, RSBM88, RMR^{+}88, RKMB89, Roe89, Roe90, RBWB90, SJB^{+}93, SE71, SHB^{+}92, Swa68a, Swa70b, Swa70c, Swa71, TS^{+}96]. b/w [Ben18, Cre19, Som21]. Babies [Hal13]. Baby [Shm20]. Back [Com11, Fol21]. Backdrop [Kon23]. Background [AdC24, PSBS^{+}22, WO05, Hal71, Koh71, Koh73]. Bacterial [Biv00, Boo23, Mét16, O’M18, San16, Ams91, Sum91, Van94b]. Bacteriology [Sin00]. Bacteriophage [Bil16, Sum93, Van94b]. Bacteriophages [San10]. bad [Cor87]. Baer [Ano76d]. Bailey [Bar19, Zir68]. Bajema [Bow89, Coo90]. Baker [Eng97]. Balance [Hon16, Kim23b, Swi14, Bea87]. Balancing [Kim23a]. Balanoglossus [Pet08]. Baldwin [Cec19]. Baltimore [Bod23, Gib24, Man13, Ras24a, Sug23, Web20]. Baltzer [Bra22a].

Baluchitherium [Man15]. Bangham [Soo21]. Banking [Hub20]. Baptiste [Tes23]. Barbara [Kle24, Com99, Kei99]. Barbosa [GM18]. Barcroft [Hol69]. Barnacle [Lov02]. Bary [Mat10]. Based [SG18]. bases [Lew69, Wil85]. Basic [Zal05, Mac93]. Basis [SD19b, SD19c]. Bastard [Kle04]. Bastian [Str99]. Bates [vW14]. Bateson [Coc73, Dar77, Pet08]. Baton [Ano05e]. Battle [Gli24, Gue23]. battleground [Ano91c]. Battles [Mun05]. Bautista [Piñ88]. Bay [Rum12, Ber70]. Bayscape [Eri17]. BBC [McG19]. Be [Col98, Har12a, Bec16, Hoq16, Jia14, Sum93]. Beacon [Pau23]. Beadle [Kev05, Kay89]. Beagle [Bri10b, Bri10a, Sul82b, Sul83]. Beagles [Bol22a]. Beanbag [RN11]. Beatley [SP25]. Beaumont [Num79]. Beautiful [Gro17b, LaB17]. Became [Gor12, Kra24, Mac01b, vHS19, Bru02]. become [Pau96]. Becoming [Ste19]. Bedrock [Buh00]. Bee [Mun05, Ano91b]. Beecher [Par21]. beekeeping [Ano91b]. Been [Esp20]. Beer [Wai88]. bees [Pre90]. Beetles [Ric24]. Before [Bel05, Coo98, Cor05, Eng97, Gal17, Lar87, Log02, PSBS^{+}22, Tan20a, Ste76, Ano06b]. Beginnings [Ada23, CK92, The96]. Behavior [Ank20, Gro19, Lev18, Vic23, Bar95, Ric81, Van94a, Wei87]. Behind [Gru18]. Beholder [Buc17]. Beijerinck [The96]. being [App80]. Belgium [De 08]. belief [Bur72a]. Belknap [Sac09]. Ben [Pow18]. Beneficial [Wei06]. Benjamin [LD20]. Benson [Kin20]. Bentley [Wol11]. Berg [Kev05, Yi08]. Bergmann [Hag17]. Berkeley [Cre96, Raa76, Sun13]. Bermuda [JACD18]. Bernard [Hal10, Ano76a, Grm68]. Bertoloni [Mes19]. Beside [Moh18]. Between [AN08, Arm12, BB22, Bra11a, CK92, Cec19, CBM86, Dru23, Gam21, Gru22, KK13, Nic22, Pir23, Pir24, San16, Som08, Ste05, Sur22, The12, BA07, Eva84, LMB03, Ric89a, Wei10].

Beurlen [Rie12]. Beyond [DT07, Esp20, Fis10, Hoq24, Pow18, Rad06, Str23, Vic23]. Biased [Dre23]. bibl [Ank18, Bru18, Kim18]. Bibliometric [Gin10]. Big [SD19a, Mat87]. Bihari [Bar23a]. Biliothèque [Ste04]. Bill [Tan20a]. Bind [Var09]. Bio [dC14]. Bio-informatics [dC14]. Bioactivists [Bar09]. Biochemical [Arm12, Ric07, Eds72, Kay89]. Biochemistry [BA07, Sin00, Ano75h, Ano84a, Cre96]. Biochemists [Arm12, Bur96, De 96, DG96, Gau96, Kay96]. Biocracy [Arm16]. Biodiversity [SG18]. Bioenergetics [All02]. Bioethics [Zal05]. biogenetic [Ano78b, Rin81]. Biogeochemical [Pir23]. Biogeographical [Fat17, Pir24]. Biogeography [Ano81a, Mir08, Vet06, Nel78]. Biographical [Smo99b]. Biography [Ala04, Com11, Die11, Gre07, Har11a, Har11b, Smo99b, Gre25]. Bioheritage [Pal22]. Bioi [Söd11]. Biological [AM14, Art17, Cec19, Dru23, Gor97, Gra20, Gre92, Ha11, Ham07, Luk21, Luk22, Mac01a, Mit21, Pre13, Rad23, Riz20, Sch17, Sch14, SK22, Vet12, ZS15, All90, All84, Bur93a, Ege69, Fry96, Gas91, Kin86, Lew69, Mac93, May68, Ric89a, Scr69, Van94a, Wig69, Ras24b, Str19]. biologii [Ano69b]. Biologist [Lin23, Ghi05]. Biologists [Bor87, Hag99, PH22, Ras99, Ano78g, Bur96, De 96, DG96, Eds80, Far74, Gau96, Kay96]. Biology [Ada00, AMW00, Ala24, Alb01, All16, Ano75b, Bae19a, BB22, BA07, Bec16, Cre03, Des05, Des01, Die98, Die17c, Esp20, Gay98, Gib19, Gro19, Hag03, Hag15, Hag17, Ham99, Har06b, JRR18, Koh12, Kra04, Lad08, LMB03, Lew04, Lew05, Lin18, LGB17, Loi25, Lon09, Lue21, Mai01, McL02, Men17, MD16, Mit03, NG14, Nic22, Nyh95, NL11, Nyh18, O’M18, Oak15, PBDL17, Rad06, RR22b, RC11, Rud00, Sch12, Sep19, Sha69, Sha08, Shi18, Smo91, Smo24, Som24, Str20, Tan20b, Uch93, Yi08, vdBD20, Ano69b, Ano76a, Ano81e, Ano81h, Ano81c, Ano82b, Ano91c, Bea88, Ben85, Ben88, Bow94, BGZ88, Bur90, Cam83, Car71, Cas73, CBM86, Cro90, Fut86, Gai73, Gau93, Gla69, Gre90, Hul69, Jun96, Kay85, Kel90, LT93, McC69, Mor79]. biology [Mun93, Pal91, Pau84, Ras95, RH84, Sap92, Sch69, Sch74a, Slo85, Smo92, Zal93b, Gli24, Liu23, Mat21, Pie12, SD19a, Bae19b, Die17a, MAR21, Ric23, SR23a]. Biomedical [Ano75c, Gau06, Led04, San06, Sin89]. Biomedicine [Cre06]. biometric [Nor73]. biometrician [De 74]. biometrician-Mendelian [De 74]. Biometry [Hag03, Ano75b, Coc73]. Biomolecular [Chi09]. Bionomics [Lar99]. Biophilosophical [Rus15]. Biophysics [Abr04, Luk15]. Biotechnology [Ras99, Sch25]. Bioterrorism [Bar04a]. Biotic [Ona23]. Bird [Bar05, Hic21, Lau19, Rad04, dB11, Bar95]. Birders [Ben18]. Birds [Bar19, Dav97b, Kre07, Mac01a, Por09, Bed83b]. Birdsong [Whi19]. Birdwatching [Bar05]. Birkhead [Rad04]. Birth [AD18, Alt07, Arm16, Bri10a, Bri10b, D’H12, Hon16, Rad23, Bor87, Gau96]. Black [Ras24a, Tan20b, Get91]. Blackburn [Tob09]. Blainville [App80]. Blakeslee [Kim06]. Bleek [dG02]. Blending [Hoq24]. Blight [Mat10]. Blind [Ham97a]. Blobel [Liu23, LaB17]. Blood [Ben09, Boy07, Eds72, Fag10, Web20, Ano75f, Gue24, Soo21]. Blowing [Cha22]. Blyth [Bed72, Bed73, BJ97, Sch74b]. Boca [Kle24]. Bocage [GM18]. Boddice [Shm23a, Shm23b]. Bodies [Kei11b]. Body [Gur02, Mos98, Nat24, Pal22, Par21]. Bomb [Hat21, Bea93]. Bombs [Hoo97]. Bone [Mat87]. Bonelli [For20]. Bonnet [Bow73, Rie88]. Bonnevie [SM07]. Bono [Bez97]. Bont [Tyr23]. Book [Ack04, Ala13, All09, All12, All13a, All22, Ank20, Ank18, Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano01a, Ano01b, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano04a, Ano05a, Ano06a, Ano06b, Ban22, Bar23a, Bar04b, Bar04a, Bar05, Bar10a, Bar11b, Bar19, Bar24, Bar10b, Bel05, Bel11, Ben10a, Ben18, Ben09, Ben10b, Ben13, Ber19, Bis23, Bod23, Boo23, Bor11, Bow18, Bra20, Bre15, Bri09, Bri17, Bro17, Bru18, Bur14, Bur17, Cab17, Can13, Cao24, Car09, Cau19, Che14, Chi18, Com12, Coo20, Cre19, Cro05, Cro12, Cro15, Dav97a, Dav12, Dav15, Des05, Die04, Die10, Dre17, Ell09, Ell11, End23, Far05, Far11b, Fit19, Fla24, Fra10, Fre25, Gia18, Gib18, Gib19, Gli04, Gli24, Goo17, Goo19, Gor19, Gre04, Gre25, Gro09, Gru18, Gue23, Hag24]. Book [Hal09, Hal12a, Hal12b, Hal17, Hal25, Ham97b, Hay17, Hei19, Hen10, HS18, Her10, Hub20, Ile05, Irm05, Ive10, Jac09, Jac10, Jin25, Joh19b, Kei11a, Kei11b, Kei23, Kev05, Kim18, Kin20, Kin09c, Kir13, Kle04, Kle10, Kle11, Koh04, Koh24, Kro97, Lan20, Lar08a, Lar06, Lar07c, Lar07a, Lar07b, Lar08b, Lar08c, Lar09, Lau19, Led04, Lig18, Lin23, Lin18, Lin15, Liu23, LD12, Loz19, Luk18, Lyn09, Lyn10, Mac12, Mad17, Mai17, Mai23b, Man13, Mar04, Mar24, Mar09, Mat21, May23, McG21, McG18, Mes19, Mét14, Mil11, Mil12, Mil14, Moh18, Mon11, Mon18, MS18, Mor12, Mor23, Muk17, Muk19, Nat24, Nic11, Nov23, Nyh18, Ona10a, Ona23, Pau23, Por04, Por09, Por25, Pow18, Rab11, Rad04, Rad12, Rad23, Ras24b, Ray19, Ren12, Ric09, Ric10]. Book [Ric11, Ric25, Roo17, Rud09, Rum10, Rum13, Rus12, RRM^{+}05, Sac09, San13, Sat20, Sch18, Sel05, Sep11, Sep15, Sha11, Shi18, Shm23a, Som21, Soo21, SW12, Ste19, Ste05, Ste04, Ste14, Ste15, Ste24, Str19, Str23, SD19a, Sug23, Sum04, Tam17, Tan20a, Tan20b, Tob09, Tyr23, Ule13, Ule14, Van10, Wad11, Wal20, Web20, Wei09, Wei11, Whi19, Wil04, Woi10, Wu19, Yan13, Yan18, You09, You10a, You10b, You25, Yud10, Zal05, dC14, Bro24, Jia24, Ras24a, Shm23b, Som24, Stu24]. Books [Gia18, Moh18, Ras24b, Str00]. Books/Putnam [Moh18]. Bookshelf [Ano98b, CRBD90, CHCC97, FRS^{+}95, FMR^{+}98, FH78, GLB^{+}94, HF73, Hal74, HLM^{+}95, KNM^{+}93, MAC^{+}94, MFM^{+}96, RHW^{+}97, RG76, RFL^{+}98, RBFC84, RRE^{+}87, Roe90, RBWB90, SJB^{+}93, AKB89, ASR^{+}93, Ano69a, Ano78f, Ano80a, Ano80b, Ano82j, Ano82k, Ano84b, Ano85b, Ano85c, Ano91g, Ano92, Ano96, BBF^{+}91, FHP^{+}91, FH77, FH79, HS72, HF74, NSMB92, PKS^{+}95, RGH^{+}81, Roe82, RS83, Roe83, Roe84, RRC^{+}85, Roe86, RCK^{+}86, RBK^{+}86, RBHK87, RPCH87, RRCB88, RSBM88, RMR^{+}88, RKMB89, Roe89, SE71, SHB^{+}92, Swa68a, Swa70b, Swa70c, Swa71, TS^{+}96]. Border [Kin09b, Mit03, Han93]. Borders [DM97, Ona23]. Bordet [CT97]. Bordogna [Mar09]. Born [KK13]. Boss [DT07]. Boston [Pau23, Sch25]. Botanical [Fla24, Hoq16, Kle04, SP25, Bro80, Neg95, Ste84]. Botanist [Kle13]. Botanists [Lus01]. Botany [Bel06, Gos23, Hag24, Köh11, Rob11, SP25, dC23, Ano81g, DCS01, Hag84]. Bottleneck [BB22]. bottles [Lew87]. Bouillabaisse [Kei17]. Boundaries [Bra20, Hea20, Liu23, Lue21, Mar09, Zal93b]. Boutan [Tho05]. Bowler [Coo90]. Box [End23, Tan20b]. Boyd [Eng97]. Boyer [All02, Pre13]. Boys [DT07]. Brad [Ros18]. Bradley [Ege69]. Brady [Tob09]. Brain [Mar04, War23, Mil11]. Brand [Rad04]. Brazil [Art17, Löw97, dMP23]. Brazilian [WdLdS21a, WdLdS21b]. break [Chu68]. Breed [Web20]. Breeding [Bon06, But18, Kim18, The08, Thu02, Vor69, Wie06, WO05, Kle11]. Bresalier [Mon18]. Bridge [Som05]. bridges [Fic77]. Brief [Eli14, Bow71, Cor89, HG76]. Bringing [Luk21, Luk22, Kle16]. Brinkman [Rab11]. Bristol [Bel05]. Britain [All20, Bar97, BJ97, Bru02, Cro89, Erl09, HA19, Hol17, Jac84, Jac03, LD19, Mac01b, Pau83, Wal20, Web20, dC06]. British [Koh24, Nov23, Ano91b, Bor76, But88, Far72c, Fis00, Hal76, Har03, Har12b, Joh04, Kin80, Mur23, Reh79, Ric81, Str99, Wal01]. Broads [CM11]. Bronn [Gli07]. Brooks [Kei98, McC69]. Brothers [dC23]. Brought [Den16]. Brown [Fla24, Bar23a, Wal21]. Brown-Séquard [Wal21]. Browne [Ano04a]. Bru [Piñ88]. Brücke’s [Liu24]. Bruno [SD19a]. Brunswick [Cro05]. Brutish [Mad20]. Bruyninckx [Whi19]. Bruynoghe [Bil16]. Brzezinski [Hal25]. Buchenau [Ray19]. Buchner [Koh71, Koh72]. Buffon [Ano76c, Bow73, Cau19, Edd23, Far72a, dMP23]. bug [Whe81]. Bugs [WJ22]. Building [Buh00, Cre03, FG17, Gau06, Luk15, Man15, Uch93]. Builds [Kim06]. Bungener [Ste04]. Burdens [Har08]. Bureau [Hut22]. Bureaucracy [Whi20]. Burgess [Bar11a]. Burnet [AM14, San10, TP94]. Bush [Bar05]. Business [WJ22]. Bussey [Rad98]. Butler [Tur18]. Bütschli [Chu89a]. Buytendijk [Gru22].

C [Ank20, Gor09, Har22b, Ile99, Mit88, Por04, Rhe96, Sel05, Str99]. C$49.95 [Muk19]. C. [Har03, Har06a, Ile99]. ca [Mar24]. Cacogenic [MM17]. Cain [Som21]. Cajal [Sum04]. Calcutta [Bed73]. Calhoun [Rad25]. California [Bis23, Hub20, Ala04, Ala12, Cre96]. Caltech [Ras99]. Cambrian [Str00]. Cambridge [Ank18, Bru18, Coo20, Kei23, Lin23, Mar24, May23, Mil11, Sac09, Sch18, Shm23a, Shm23b, Stu24, Tan20b, Tob09, Wal20, You09, Zal05, Ano78e, Bla07, Rhe96, Ric06]. Came [Hun19a, Sum93]. Camels [Hag15]. Campaign [Bar09, Rum12, Rue84]. Campaigns [Dan21]. Can [Ano91b]. Canal [Kei17]. Cancer [Cro14, Gau98, Mor97a, Yi08, Löw94].

Candaux [Ste04]. Candolle [Ste04, ÁTBHBJB24, Nel78, Ste84]. Canguilhem [Gay98]. cannibal [PW92]. Cannon [Arm16]. Canon [FG17]. Can’t [Zal05]. Capacity [You98]. Capital [Bol22a]. Capitalism [Kle04]. Capitalist [Hod09]. captain [Ste90]. Captivity [Koh24]. cardiac [Sha72]. cardiocentric [Fra91]. cardiology [Fre70]. Care [Dru23]. Career [GM18, Muk14, SM07, SP25]. Careers [Hen97, Kim06, Koh04]. Caribou [Lue21]. Carl [AN08, Bow89, Di 95]. Carleton [Jac01]. Carlson [Die04]. Carnegie [Wil12]. Carol [Ano05a]. Carolina [Koh24]. Carpenter [Ano76d, LD20, Mon05]. Carruthers [Ank18]. Carrying [You98]. Carson [SP25]. Cartographies [MM17]. Case [BS08, End23, Gli06, Har11a, Lep15, Str20, Wei06, Bea93, Bor89, Sil94, Wel73, Zal93b]. Caspar [Roe79]. Cassidy [Mon18]. cast [Raa76]. Casualty [Hat21, Bea93]. Cataloging [Hun19c]. Catalytic [Bra11b]. Categories [FC21]. Catharine [Par21]. Catherine [Ano05a, Tob09]. cats [Dan85]. Cattell [Wei09]. Cattle [The08]. Causa [Hic21]. Causation [Gue13, Mat10]. Cause [Whe81]. Causes [Jia14, Rie90]. cave [Boc88]. Cell [All22, Ano78b, BA07, Gur02, Jia14, Liu17, Liu23, Liu24, Mat21, Men03, Ric00, San16, Sur22, Tru97, Cor89, Jac89, Jac84, KC94, Koh71, Koh72, Ras95, Ric89b, Sum91]. cell-free [Koh71, Koh72]. Cells [All22, Biv00, Fag10, Rey08]. cellular [Jac83]. Centennial [Bil16]. Center [Ano05a, NL11, Ano88d]. Centered [Bae19a, Bae19b]. Centers [But88]. Central [Hoq16, Ano78c, Hut22, Kei17, WO05]. Centric [Shi18]. centuries [Ano91b]. Century [Ank18, Ben09, Bro11, Bro24, Gue13, Jen16, Joh04, Kim18, KFP09, Luk15, Mos98, dMP23, Sho13, Stu24, Wol13, dC23, Ano76f, Ano81e, Ano81f, Ano82d, Ano82l, Ano91c, Ano91f, App80, Ben85, Bor89, Bow71, Bro74, Cam83, CB82, Cor71, Cor89, Cor83, Cro90, Deb70, Del83, DCS01, Ege69, Gai73, Ger70, Gro79, Hag84, Hag86, Kin80, Lar86, Pau84, Pió79, Ric89b, RH84, Ste90, Whe81, Wig69, Bol22a, BJ97, Bre02, CC18, Die16b, Dre20, Eli14, Eri17, FC21, For20, Gli06, Ha11, HA19, Joh09, Kev05, Koh12, LMB03, LD19, LB04, Man15, Mir08, Qui17, Rab11, Sch04, Sch99b, Sep19, Sla03, Tam15, Tam20b, Vet12, VWP20a, VWP20b, Wie06, Wit16, WO05, ZS15, vP21]. cerebral [Swa70a]. Chadarevian [Ban22]. chain [App80]. Challenge [Bor03, Cha22]. Challenges [Bae19a, Bae19b, Hag99, Boy84]. Challenging [Tam20a]. Chalmers [Cro89]. Cham [Fla24, Hal25]. Chamber [Hod72]. Chambers [Sch99b]. Champion [Sel23]. Chance [Hal13, Har14, Veu19]. Chances [SM07]. Change [Mac21a, Mac21b, Get91, Lin92, Rho87]. Changed [End23]. Changes [HR11]. Changing [CS06, Eli14, Jia14, Kin09b, San10]. Channel [Tru97]. Chaos [Kei01]. Chapel [Koh24]. Chapin [Joh22]. character [Sch80b]. Characterized [Mad20]. Characters [Hum08, KC97, Sta95]. Charismatic [Lag03]. Charles [Mét14, AN08, Ano75d, Ano78c, Ano04a, Baj88, Bed73, BJ97, Bri10a, Bri10b, Byn84, Col85, Hal13, Hol14, JK99, Joh19a, Mir22, Nov19, Qui17, Rie88, Sch74b, Sch95, Sch89, Smi78, Smi90, Smi92]. Charlottesville [Por25]. Chase [Hen03]. Che [Joh19b]. Chemiosmotic [Pre01]. Chemosynthesis [Ack06]. Chen [Luk22, Luk21]. Chester [Sin00]. Chetverikov [Ada68]. Cheyne [Gue85]. Chicago [Ank20, Ban22, Bar23a, Bar24, Bri09, Cau19, Des05, End23, Fre25, Gli24, Gor19, Gru18, Gue23, HS18, Ile05, Jia24, Kim18, Kin20, Kle11, Lan20, Lig18, Liu23, Mar09, Mat21, McG21, Mes19, Mor23, Nat24, Nyh18, Ona23, Pow18, Rab11, Ray13, Sat20, Shi18, SD19a, Wei09, Wei11, You25, dC14, Zal92]. Chicken [BK10]. chief [Sch95]. Child [Tho05]. Children [Fol21]. Chilean [Wal15]. China [Luk22, Hub20, JF03, Luk15, Luk21, Sch12, Yi21]. Chip [RF18]. Choice [Hop24, VWP20a, VWP20b, Bur93a, Cla93]. Christa [Goo19]. Christian [Han93, Fre25, Mor14]. Chromosomal [Lam15]. Chromosome [Bru02, Wil16, All74b, Ano76b, Ric25, dC18]. Chromosomes [Ban22, Hat21, Kle24, Led89]. Chrulew [Lau19]. Churchill [Ano78e]. Cinchona [May23]. Cinematic [Gib19]. Circadian [Bec16]. Circular [Cog02]. Circulation [Cre19, Gro17b, Gue24]. circulatory [Ano82b]. Circumstances [Pea10]. Cirripedia [AdC24, Lov02]. Citations [ddAM20, dd22]. civet [Dan85]. Civic [Sha08, dB11]. Civil [Pin02]. Civilize [Lin24]. civilizing [Fin91]. Cladograms [Kra24]. Claim [Bre02]. Clara [Pau23]. Clarence [Mon05]. Clashing [Esp13]. Class [Muk14, Thu02]. Classical [Die04, Bea87]. classical/balance [Bea87]. Classification [Mir11, Mor97b, Nov16, Bor89, Irm05]. Claude [Ack04, Ano76a, Grm68]. Cleavage [Gur02]. Clements [vdV14]. Climate [Ber19, Cha22, HR11, LD19, Mac21a, Mac21b, vH19, Osp77]. Climax [vdV14]. Clinical [Gut97, Löw94]. Clock [Mor98]. Clones [Rad23]. Cloning [Mai01, Rad23, Yi25]. Closer [Den16]. Cloth [Ban22, Ben18, Bra20, Cau19, Gor19, Gru18, HS18, Lan20, Lig18, Lin18, McG21, Mes19, Muk19, Nyh18, Som21, Wu19, Kim18, dC14]. clothing [Di 95]. Co [Cai02, Tho05, Ham84, Sin97]. co-discoverer [Ham84]. Co-invention [Tho05]. Co-opting [Cai02]. Coats [Bar23a]. Cobb [Tan20b]. Code [Pie12]. Cognition [Ray19, Tau97]. cognitive [Kel90]. Cold [Ber19, Hec21, Kev05, Lin23, Sel05, Som21, SD19b, Cas15, GS15, SP25, SD19c, Wol12, dLK12]. Colder [Lan20]. Coleopterists [Eli14]. Coleridge [Smi99]. Colin [Nat24]. collaborate [Mai93]. Collaboration [AN08, GG93, Koh22b, Mat06, Bea93, Ste18]. Collaborative [AM17]. Collaborators [AdC24, ddAM20, dd22]. Collapse [LGB17, Loi25]. Collapse-of [Loi25]. Colleague [Bro23, Sch23]. Colleagues [Cai02]. Collecting [Cao24, cH17, Str10, SD19a]. Collection [RR22c, Sta25, Fra79, dR23]. Collections [Koh24, Kol22, RR20b]. College [Ano78e]. Collins [Pow18, Oga01]. Collision [Sep06]. Colonial [Dan21, Vet06, Lew18]. Colony [Wil12]. color [Ben18, Cre19, Gru18, Mes19]. coloration [Kin78]. Columbia [Hei19, Cai02]. Combining [vP21]. Comes [Stu23]. Comics [Loz19]. Coming [Bea82]. Commemoration [Ras25]. Comment [Koh12, Bur96]. Commentary [MKC69, MN69]. Comments [Kay96, Nel78, Bea88]. Commerce [Bis23]. Commercial [WJ22, Yi25]. Commission [Hat21, All90, Bea93, Cre06]. Committee [Fis00, Mit17, Oak15]. Committees [Par00]. Common [GG97]. commune [Swa70a]. Communication [Kei99, Mon05]. Communist [Cas12]. Communities [Ste24]. Community [You25]. Company [GZ98]. Comparative [Arn19, CS06, Bor89, Ste84]. Comparing [Rhe96, Str10]. Competition [Tam95]. complete [Ano76c]. Completely [Moh18]. Complex [Ste07]. Complexity [Ank20, Bar11a, Dre23]. Complicating [Pie19]. Complication [Pea10]. Compromises [Sch17]. Computers [Nov18]. Computing [Str10]. comte [Cau19]. Concept [Alt07, Bri05, D’H12, Gay98, Kli98, Wit16, Ano76f, Chu74, Far72a, Jac89, May68, Swa68b, Wil85]. Conception [Har03, LaB17, Rus04, Ano75n, Boy84, Boy86, Koh73, Pre77]. Concepts [Bil71, Cor05, Jax98, KK13, Li04, PH22, Pre01, RWE13, Ada70, Atr87, Cam83, IB71, Jun96, Par71b, Sht92]. Conceptual [Kay85]. conceptualization [Oud90]. Concerning [Tau97]. Concise [Gre25]. concussion [Por79]. Conditioned [HA19]. Condor [Ala04]. Conflict [Cro14, IB71, LMB03, Som05, Ano91c, Lew69]. Conflicting [Hag17]. conformation [De 96]. Confusion [VHL20]. Congress [Joh09]. Conklin [AM68, Atk85, Coo02]. Connecting [HR21]. Connection [Wal21, Ano82c, Bed88a, Gor89]. consciousness [Smi78]. consensus [Ano91c]. Consequences [Jax98, Pir23]. Conservation [Ala04, All13b, Car08, Dru23, Lew05, Sch17, SG18, Whi20, Str92, Pow18]. consider [Ano78g]. considered [Biz92]. Consolidation [Str04]. Conspiracy [Por25]. constitution [KC94]. Constraints [Tam20a]. Constructing [Boc12, Rad25]. Construction [BP19, GM18, AA82, Van94a, Van94b]. Constructions [Tam20a]. Consuming [Rad25]. Container [Liu17]. Contaminated [Lue21]. Contamination [VHL20]. Contemporary [Söd11]. Content [SK22]. Contest [Bra11a]. Context [Hal25, Lep15, Thu02, Wol12, Hay88, Hul05, Out86, Wel73]. Contexts [Hod09, ML10]. Contingency [Hal13]. continuation [Ano81b]. continued [Bar95]. continuity [Get91]. continuous [Sta95]. Contrasting [Hal76, Pre13, Sla03]. Contribution [Bil16, Sch10, Ser16, Cow72, Sta95]. Contributions [Ada68, KFP09, LGB17, Loi25, Vic23, ddAM20, dd22, Car71, Ric81, Wai88, Dav97a]. Contributions-and [Loi25]. Control [Ber19, Com99, Gut97, Sch18, Bor87]. Controlling [Com99, Gau06, Lin24, Mai09]. Controversies [Bar11a]. Controversy [Ala04, Boc12, Fat17, Gor97, Gro17a, Kei17, Mit17, Mun05, Mur23, Ser21, Stu23, Wei09, Ams91, Ano91b, Bea87, Far72b, Far72c, Hei72, Kre96, dLK12]. conundrum [Pre90]. Converged [Lou09]. Converging [Tru97]. Conversation [Har22a, Mac01b]. Conversion [Ank00, Led89, Sul82b]. Conwy [Arn19]. Coon [Jac01]. cooperative [Mit88]. Cope [Cec19]. copy [Bed88b]. Coral [Mir22, vP20]. Cormorants [Hag15]. Corn [Kle99, Mat06, Kle04]. Correction [Ano20, Bae19b, BH21a, Har22b, Kim23b, Luk22, Mac21a, Meu21a, Shm23b, VWP20a, WdLdS21a, dd22]. Correia [DCS01]. Correspondence [Lig22, Par00, Sch88, WP06, Sch95]. Correspondents [Sch99b]. corrupted [Ano78g]. Cosmic [Str04]. Cosmology [Mit21]. Cosmopolitan [And16]. Cossar [But18]. Costa [Fic24]. Counting [Sto15]. County [Alb01]. Courtney [Fit19]. Courtship [Bar95]. cover [Hei19]. COVID [Ras24a]. COVID-19 [Ras24a]. Cowdry [Par08]. Crab [Bar04a]. Crabs [Bar04a]. Cradle [LB04]. Craft [Har11a, Jia24]. Craniological [dR23]. CRC [Kle24]. Created [Rad04, vP21]. Creating [Rai03, Str04]. Creation [LM14, Mac01b, Mat21, Sch90, Sep06, Cor83]. Creationism [Par00, Bel05, Jin25]. Creationist [FE20]. Creativity [Bea16, Bea19]. Creatures [Muk19, Som21]. Creighton [Buh00]. Creole [Wal15]. Crick [Cob25]. Crisis [BB22, Gli76]. CRISPR [Bod23]. criterion [Gla74]. critic [Hal76]. Critical [Ano76a, Arm16]. Criticism [Ric00, Ano76e, Pre77]. critics [Byn84, Ghi73, Ric89b]. Critique [Coo02, Cos97, Gam21, Ano82l, Rai85]. Croizat [Nel78]. Crossing [Liu23]. crossings [Han93]. Crow [Pat09]. Crowe [Rad23]. crusade [Lar87]. Crystallization [Str04]. Crystallography [Ste84]. Culling [All13b]. Cult [Wal20]. Cultivation [Eri17]. Cultural [Jum02, Lew05, Thu02]. Culture [JRR18, Jin25, Kol22, LK25, Rai03, Sch80a, Sur22, Ile05, McE92, Sum91]. Cultures [Bon06, Sta25]. Culturing [Pas21]. Cunningham [Gue24]. Curators [Gru18]. Curious [Cob25, Joh07]. Current [Bro71]. Curry [Kim18]. Cushny [Ano75a]. Cuvier [Bor89, Eig97, Out86]. Cycle [Ack07]. Cycles [Jon17]. Cyclical [Ser21]. cyclogenic [Ams91]. Cytogenetics [Hat21]. Cytology [Har06a, Har22b]. cytometry [KC94].

D [Ank00, Ber19, Can13, Fre25, Har03, Har06a, Har22b, Rab11, Ste04]. Dairy [The08]. Dalcq [Thi01]. Dalrymple [Mir22]. Dalton [Bel01, End11]. Dance [Mun05]. Daniel [Ste04, McG21, Vic23]. Darbishire [Ank00]. Dark [Smo24]. Darkness [Mad20]. Darlington [Har22b, Har03, Har06a]. Darwin [Ano04a, Ano06b, Bel05, Bri10b, Far05, Hal25, Tan20a, dd22, Alt07, AdC24, AN08, And24, Ano75d, Ano75e, Ano75i, Ano81a, App88, Baj88, Bar92, Bea16, Bed68, Bed72, Bed73, Bed88a, Bed88b, Biz92, Bow89, Bra11a, BJ97, Bri10a, Bro86, Bro80, Bro82, Bro22, Buc17, Cog02, Col85, Coo90, Cor84, Cor87, Cor05, Ege68b, Ege70, End03, Eng97, Eva84, Fag07, Far09a, Far74, Fre70, Gal17, Gam21, Ghi73, Goo19, Gor89, Got99, Gre77, Gre92, Gri74, Gri85, Hai91, Har22a, Her71, Her74, Her77, Hol14, Hoq24, Hul05, Joh07, Joh19a, Kly14, Lap90, Len05, Len87, Lig18, Lig22, Lov02, Lyo95, May92, Mir22, NSRG09, Nov19, Osp79, Par82, Par18, Pea10, PKS18, Por93, Pre90, RR21b, RM84, Rho87, Ric81, RR22c, Sch74b, Sch84]. Darwin [Sch90, Sch95, Sch79, Sch80b, Sch88, Sch89, Ser16, SP81, Slo86, Slo07, Smi78, Smi90, Smi92, Smo24, Sta96, Ste90, Sul82b, Sul82a, Sul83, Sul09, Tam15, Tam95, The12, Vor69, Vor77, Wal21, Whi22, Win00, ddAM20, vP20, vP21]. Darwinian [Bro24, Kim23b, Ala24, Arc09, Bea82, Des84, FE20, Ghi05, Her05, Hod09, cH17, KK13, LK25, Meu21a, Meu21b, Osp80, Rho87, Rus05, Smo05b, Söd94, Sta24, Ste90, Sul11, Wel71]. Darwinism [Bea16, Bow77, Bow05, Har90, Lar99, Moo91, Mor97b, Nor73, Sch95, Ser16, Str99]. Darwinists [Ray13]. Darwins [Ano76e]. Darwinsche [Gli04]. Data [JACD18, Sep13, SD14, SD19a, dC14, SP81, Shi18]. Data-Centric [Shi18]. Data-Driven [dC14]. Database [Ste18]. dating [Sul83]. Daubenton [Lov06]. Daughter [Pau23]. Daum [Gre25]. David [Bar97, Ber19, Bor03, Bri09, Ile05, Str23]. Dawson [Cor83, Pri24]. Dayhoff [Str10]. De-Stalinize [Dua20]. Dead [Har12a, PH22]. Dealer [Bar00]. Dean [Ogi07]. Death [Bri10a, Bri10b, Che13, D’H12, Lau19, Lus00, Ras24a, Get91]. Debate [Bar11a, Far16, Jum02, Pir24, Vic23, Wei10, Boc88, Coo90, De 74, Kin78, Mil08]. Debates [Bel05, FE20, GS15, LD19, Mir11, Sch25, Sch10, Str99, Ste90]. Debating [CT97]. Deborah [Lau19]. debt [Sch74b]. Debut [Gin10]. Decades [PSBS^{+}22, Uch93]. Decapitation [Che13]. Decay [LD19]. Deciphering [Tob09]. decline [Ano91c]. Decolonizing [Gos23]. Deconstructing [Hul05, Moo91]. Decree [Mun23]. Defence [Shm23a, Shm23b]. Defending [Sel12]. Defense [Lar99, Nor73, Sch95]. Deficiency [Fre25]. Defined [Sel03]. Defining [Pal91, You98]. Definition [vHS19, Bur93b]. Deforestation [Hug11]. Degeneration [dMP23]. degradation [Mar92]. Delbrück [Kay85]. Delft [The96]. Delicate [vW14]. delineator [Rud90]. Demarcating [Kel88]. Democracy [Sel12]. Demographics [Eli14]. demography [Ege68a]. demonstrated [Wai88]. Demythologizing [Wei92]. Dentinger [Mon18]. Department [Wil12, Ben85, Cre96, But18]. Deployed [Gis23]. Descartes [Ano75n, Hal70]. Descent [Gam21, Gli98, Har22a, Sch84]. description [Piñ88]. Descriptions [CC18]. Descriptive [Ano81f]. Design [Kei23, Par15, Rhe00, Ric04, Bow77, Cor83, Lyo93, Osp80]. Desire [Smo99a, Tam20b]. Destiny [Buc17]. determinism [All84]. Developing [Jan24]. Development [All04, Ano69b, Bla07, Bon06, Bra22a, Bul99, Chi09, Cro05, Cro14, GS10, Hop24, Lev18, Log01, Mai16, Mon05, Pin02, Sch09, Sun10, Swi14, You98, Ano75j, Ano81f, Bed72, Bil71, But88, Her74, Her77, Kin80, Lus74, Rho87, Rin81, Rin88, Sch74a]. Developmental [Gur02, Jan24, Ray22, Rhe00, All74b, Chu69]. Developments [RR21b, Bro82]. Diagnosis [Mit17]. Diagrammatic [Cec24]. Diagrammatics [Mar24]. Dialectical [Yi21]. Diauxic [BH21b, BH21a]. Dick [Cro05]. did [Ano82n, Pau96]. Dietrich [Sat20]. Difference [Ha11, RF18]. Differences [CC18, Mil11]. Differentiation [Buc17]. Difficult [Ben11b]. Diffusion [AMW00, GZ98, Gli76]. digestive [Ano82b]. dimension [Kre96]. Dimensional [vP21]. Dimensions [vP21]. Dinosaur [Rab11, Ano82c]. Dioscorides [Ano81g]. diplomacy [Bea93]. Diplomats [Tyr23]. dir [Ros18]. Disadapted [Vic18]. Disappearance [Sch15]. Disappointed [Som05]. disciple [Sch95]. Disciplinary [Ano91c, Har06a, Har22b, Gau93]. Discipline [Dav97b, Dhe21, Iid10, Str04, Bur96, DG96]. disciplines [Kay96]. Discolored [Eri17]. Discontinuity [Pet08]. Discourse [Ano76c, Kel88]. Discourses [Art17]. Discover [Cog22, Ano82n]. Discovered [Gue24]. discoverer [Ham84]. discoveries [Gla74]. Discovering [Dav97b, Tom22]. discovers [Dan85]. Discovery [Ack06, BK10, Fun12, Ras22, VHL20, Ano91d, Bed73, Bor76, Grm68, Koh71, Koh72]. Discrete [Chi09]. Discrimination [Gor09]. Discussion [Ano82d, Gru22, Sta03]. Disease [Hon16, Jon17, Mat10, Mes19, SD19b, SD19c]. Diseases [Ack04]. disguises [Bla82]. Disinterested [Har11c]. Disorders [Ank20]. Dispassionate [Ham07]. Disproportionate [Fol21]. Dispute [RN11]. Disputed [Gli24]. Disruptors [Kei11b]. Dissemination [Sch99b]. Distance [Sch15, Smo99b, Sto15]. Distinguishing [Mil08]. Distress [Bar09]. Distribution [ÁTBHBJB24, Mir22, Kin80]. Diverged [Lou09]. Divergence [Sch80b, Bed88a, Bro80, May92, Tam95]. Divergent [Hag86]. Diversity [Log01, dC23]. Divide [Sep06]. Division [D’H12, DT07, Tam95]. DNA [GS10, Gia93, Sch25, Sel03, Tob09, Tom22, VHL20, Yi08]. Do [Zal05]. Dobzhansky [Cai02, Cro87, Far16, Gor12, Joh22, Kon23]. Does [Hoq16, PKS18]. Dog [Dan21, Web20]. Dog-Owners [Dan21]. Dogs [Bol22a]. Dohrn [GG97]. Doing [Stu23]. Dollo [Gou70]. Dolman [Ack04]. Domenico [Mes19]. Domesticated [dMP23]. Domestication [ÁTBHBJB24, Lyn19, Ric06, Bar92]. Dominance [Can15, Cam83, Fal91, Kle04]. Don [Ano75l]. Donaldson [Kli98]. Donna [Ham97b]. Dooren [Lau19]. Dordrecht [Mét14]. Dorée [LM14]. dormancy [Cam83]. Dose [GS15]. Double [Gin10, Log07, De 74]. double-edged [De 74]. Doubt [Fis10]. Dove [Coo20]. Dover [Irm05]. draft [Ano75e]. Dragons [Bar09]. Drawing [Ano05a]. Drawings [Yan18, Pow18, dC14]. Dream [Coo02, Kim06, Cre96]. Dreamers [Sat20]. dredgers [Reh79]. Drift [Mil08]. Driven [dC14]. Drosophila [Ano78a, Car74, DT07, Koh93]. Drouin [Ste04]. drug [Ano75a]. Drugs [Bol22a, Gue23]. Dubos [Ras22]. Dubrova [GS15]. Duke [Lin18, McG18, Som24, Yan18]. Dulbecco [Kev93]. Dull [End11]. Dunn [Gor09]. Durham [Lin18, McG18, Som24, Yan18]. During [Mos98, Wei06, Ams91, Ano76f, Cro89, WO05]. Durkheim [Gis23]. Dutch [KFP09]. Duty [Coo02, Str92]. Dynamics [AdC24, Bra22b, vH19].

E. [Atk85, RH84]. Earliest [Vin22]. Early [Bri05, Elw07, Eri17, FC21, For20, Gli06, GM13, Ha11, Har12b, Ile99, LM14, LMB03, LB04, Man15, Mel13, Mos98, Ray19, Rhe93, Ric07, SM07, Thi01, Vet12, Wie06, ZS15, Ano75e, Ano81g, Ano91c, Bor89, DCS01, Ger70, Gro79, Hal76, IB71, Kin80, Por79, Reh79, Sch80a, Sta95, Hec21, Jen16, Mir08, WO05]. Early-Nineteenth-Century [For20]. Early-Twentieth [Man15]. Earth [Fit19, McG19, Hen03]. East [Joh19b, Ano75m]. Eastern [Vet06]. ebb [Ano75f]. eclipse [Bow05, Meu21a, Meu21b]. Eclipsing [Meu21a, Meu21b]. Ecological [Hea20, Hut22, Jon17, Kei23, Pal91, Col86a, Col86b, Ege69, Tay88, Rum12]. Ecologies [And16, CM11]. Ecologists [Hag86, Boc95]. Ecology [Ack07, Ank04, Bea88, DM97, Fis00, Fut86, Hen97, Hon16, Hun19a, Ile99, MD16, SP25, Boc90, Col86a, Col86b, CBM86, Ege83a, Ege83b, Ege85, Kim86, Kin86, Kin94, Lus74, Pal91, Hag24]. Economic [Kra04]. economists [Sch80b]. Economization [McG18]. Economy [Dan21, Pea10, Tay98, Gor89, Ham86]. Ecosystem [Jax98, Lew18, vdV14]. Ecosystems [Boc95]. ed [Bel05, Bod23, Bro24, Hal25, You09]. Edgar [Kle99, Kle13, Kle16]. edged [De 74]. Edinburgh [Tan20a, Ano91e, But18, Cor21]. edition [Baj88]. Editor [Ano82e, Ano82f, Ano82g, Ano03c, Hun19b, MAR21, Ric23, SR23a]. Editorial [AM99, Die12, Far06, Far07, Far09a, Far11a, JRR18, Men68, RR18, RS23, RS24, Ras25, Smo24]. editors [Ano89a, AM01a, AM01b, Ano04b, Ano05b, Far09b, GD18, MSA69]. Edmund [Ano76b, Par08]. eds [Ano05a, Bri09, Dav97a, Joh19b, Kro97, Lau19, Lin18, Mét14, Ray19, Ste04, Str19, Wu19]. Eduard [Koh71, Koh72]. education [Lar87]. Edward [Irm05, Ric23, Ano82i, Arc09, Bed72, Bed73, BJ97, Gib13, Joh22, Sch74b]. Edwards [Bor03]. Edwin [AM68, Coo02]. Eels [KFP09]. effect [De 74]. Effects [GS15, Ham07, Pin02, Pau96]. Effort [AM17, Ham07]. Ehrenberg [Chu89a]. Ehrlich [CT97]. Eighteenth [Mos98, Wol13, Ano82l, Ano91b, Bor89, Bro74, Cor71, Del83, Ege69, Lar86]. eighteenth- [Bor89]. eighteenth-century [Ano82l, Bro74, Ege69]. Electricity [Hom70, KFP09]. Electron [vHS19]. Electrophoresis [Chi09]. Elementarorganismen [Liu24]. Elementary [BL24, Rey08]. Elements [Com99, Kle24]. Elephant [PKS18]. Elias [GG97]. Elice [Hal25]. Elisabeth [SM07]. Elites [Lin24]. Elizabeth [Irm05, Gia18, Tob09]. Ella [Lad08]. Ellen [Mun23]. Ellis [Ala24]. Elliston [Bar97]. Elof [Die04]. Eloge [Gue21]. ELSI [Wol01]. Elusive [Gre04]. Elwick [Nov23]. Emancipation [Lou09]. EMBL [CdC25]. EMBO [Cas15, CdC25]. Embrace [Nov18]. Embraces [Chu10]. Embracing [Hec21]. Embryo [All04, ML10]. Embryogenesis [Thi01]. Embryological [Gal21, Gam21, Ano81f, Gil78]. embryologist [Atk85]. Embryologists [Arm12]. Embryology [Ano81h, Kot81, Ray22, Ste19, Wil12, Ano76d, Chu70, Rin88, Roe79, Ano75j]. Embryonic [Arm12]. Emergence [All02, Bro11, Dav97b, Dre23, Kev05, Par15, Pin02, Wit16, vdBD20, Ben88, Boc90, Bor85, Kel90]. Emerging [Hal25, LD20, Mit17, SM07]. Emerson [Mit88]. Emile [Mor98]. Emma [Som24]. Empire [Hod09, LM14]. Empirical [Log02]. Empirically [Mil08]. Empiricist [Cos97]. Encyclopédie [Lla00]. End [Tam15, Swe95]. Endangered [Bar24, Hic21]. Endocrine [Kei01]. Endocrinologists [Oud90]. Endocrinology [HG76, Jan24, Log07, San06, Bor85, Gli76, Hal76]. endogenous [Van94b]. Ends [Tob09]. endurance [Hei72]. Enduring [dMP23]. Energy [Cre06]. Enfant [Erl16]. Engagement [Nov19]. Engelbert [CB82]. Engineered [Rad04]. Engineering [Gra20, KC94, Ber19]. English [Ano76c, Bro74, GT69, Kra04]. Enigmas [Cob25]. Enlightenment [Mét14, Han93]. Enrico [Bec05]. entelechy [Chu69]. Entire [Wal15]. Entomological [Pas21]. Entomology [Hen97, Whe81]. Entrepreneurial [Bar00]. Environment [Ber19, Ell07, Mun23, Tab08, Fry96, Hol69, Pla69]. Environmental [Boc95, Boc12, DM97, Kei01, Mar23, Mar92, RC11, You25]. environmentalism [Wei92, Kin20]. Environments [Kin20]. envisioning [All22]. Enzymatic [Loi25]. Ephestia [Rhe00]. Epidemics [Mét16]. Epidemiological [Mét16]. Epidemiologies [Bro11]. Epidemiology [Löw97]. epigenesis [Roe79]. Epigenetic [Ric00, Yan18]. Epigenetics [Bru18]. Epilepsy [Wal21]. Epilogue [GS92, Ano81c]. episode [Bow94, Whe81]. Epistemic [Yi21]. Epistemological [Bur93a]. Epistemologies [Sep13]. Epistemology [Chi09, CT97, Rus15, Ste19]. Epochs [Cau19]. Equation [JK99]. Era [GM13, Mun23, Pat09, San16]. Erika [Som21]. Erlanger [Fra79]. Ernest [Rhe96]. Ernst [Ano76d, Di 95, Gre92, Gre94, JW13, Jun96, Kle13, Kol22, Liu24, RN11, Rin81, Slo85, Smo05a, dG02]. Err [All02]. Erratum [Ano05c, Bri10b, SD19b]. Errol [Sel05]. Erwin [Wil16]. Escaping [End03]. Escherich [WY23]. Essay [AA82, All84, All91, Ano68, Ano84a, Ano91f, Bed83a, Bon84, Bro82, Bro88, Bur90, BR20, Cai99, Cre03, Cre10a, DM97, Eds80, Ege83a, Ell07, FS90, Fin91, Gau98, Har13, Har89, Har90, Kin94, Koh95, Len87, Lus01, Mac01b, Mor97b, Nyh95, Pau93, Por85, Por95, Pri24, Rai01, Rud00, Rus07, Sap92, Sch80a, Sch88, Sel03, Slo85, Slo07, Smo91, Smo99a, Smo05a, SS99, Str00, Wol01, YD00, Irm05, Rad07, Wei05]. Essays [Bri09]. Essentialism [Var09]. Establishing [Cla93]. establishment [Ano91d]. estimation [Cor87]. Ethic [You25]. ethical [Str92]. Ethics [Far05, Num79]. Ethno [pY14]. Ethno-Nationalism [pY14]. ethological [Kal83]. Ethologist [Rad07]. Ethology [Bur99, Dhe21, Gil90]. Etienne [Kin20, Gal21]. Eugen [Vin22]. Eugenic [Har12b, MM17, Wei87]. Eugenicist [BS08, BA81]. Eugenicists [Har12b]. Eugenics [All13b, Cec24, Coo98, Coo02, Far05, Lou09, Pau16, Pau93, Ser16, Sta24, Wal15, Ano91h, Lud69, Ogi07, Wei06]. Eukaryotic [Yi08]. Euler [Kly14]. EUR33.00 [Ste04]. EUR40.50 [Tan20b]. EURATOM [Cas15]. Euripus [Ano75f]. Europe [Cas15, Dan85, WO05]. European [CC18, Cor05, Lep15]. Euthenics [Mun23]. Evangelism [Par21]. Evans [Raa76]. Evelleen [Lig18]. Evelyn [Pir23, Sla03]. Event [GH19]. Everett [Ano18, Ano19, Die17b, RR20a, RR21a, RR22a, RR23, SR24, Smo25a, Ada23, Ank23, Bar23b, Bro23, Bur23, Die16a, Har23, Mai23a, Sch23, Sel23, SR23a, Sta23]. Everyone [Win01a]. Evidence [Kei99, Mit21, Mos98, Ray22]. evlution [Cor83]. Evolution [Ano82h, Bel05, Bro24, Cam01, Cec19, Col86a, Den16, Die98, For20, Gal17, Gli98, Gli06, Gre04, Hag99, Hal10, Hal13, Hal25, Har06a, Har15, Har22b, Hod05, KK13, Lar87, Len05, LK25, Lus00, Mac01b, Mor14, O’M18, Pie19, Pow18, Ric24, Sep06, Smi82, SD19b, SD19c, Sul11, Tam15, WdLdS21a, WdLdS21b, Ano75d, Atk85, Bed72, Boc88, Bro86, Bur70, Bur72a, Die94, Eds72, Hay88, May85, Moo91, Ric81, Sht92, Smo94, Sul82a, Swe95, Jin25, Kim18, Man13, Tan20a]. Evolutionary [Ala24, All20, ÁTBHBJB24, Art24, Buc17, Cai09, Col86b, De 08, Die98, Esp20, Far05, Gal21, Gur02, Hag17, Hal25, Hul05, Jen16, Kle99, KK13, Mar24, May97, Mor98, O’M18, Sep19, SC14, Tab04, Tan17, pY14, Ada70, Ano82c, Bea88, Bow94, CBM86, Die95, Fut86, Gou70, Gre90, Har94, Kel88, Kim86, Mic86, Moo95, Rho87, Smo92, Wai88]. evolutionism [Cro89]. Evolutionist [FE20, Gre77]. Evolving [Sep13]. Ewart [But18]. Ex [Ano81g]. Examination [Gam21, San73]. Example [Str20, ZS15]. examples [Hal70]. Exclusion [Hoo97]. Excursion [CM11]. Excursions [BA07]. Exemplar [Sta23]. Exemplary [Rey08]. Exercise [Joh15, Sch15]. existence [Bow71]. Exobiology [Str04]. exogenous [Van94b]. Expedition [cH17]. Experience [Irm05]. Experiences [Tay21]. Experiment [Gue23, Har12b, Mét16, Rad05, Rhe93, dC06]. Experimental [Art24, Bon06, Erl09, GZ98, Hag84, Koh04, Löw94, O’M18, Par15, Rhe00, Ric24, Shm23a, Shm23b, Ano76a, Bur93a, Hol93, RM84, Rhe96, Rin88, Sil94, Sti94]. Experimentalists [Hag84]. Experimentation [FG17, Gli06, Joh15, Mel13, San06, Wol13, Num79, Mor23]. Experimenters [HS18]. Experimenting [GM13, Bod23, Bra20]. Experiments [Ank20, Che13, Gue13, Hag10, Log01, Mat10, Ona10b, PSBS^{+}22, SD19a, vH19, Di 91]. Expert [Buc17]. Expertise [CdC25, Whi20]. Explaining [Har13]. Explanation [Gla69, Kal83, Mat16, Scr69, Gro69, Lew69, Mic86]. Explanations [Hod72, Kin80, Sch69]. Exploration [Lan20]. Explorations [Fla24]. Exploring [DM97, Lov02, Mit03]. Explosion [Str00]. Exposition [Gib24]. Exposure [Fol21, HR21]. Extinction [Win24, Lau19]. Extraordinary [Fun12, Ste97]. Extreme [Lan20]. Eye [Buc17].

F [Rhe96, Wal20]. F.

[AM14, Dex70, Gru22, Ste90]. Faced

[Har11b]. Facing [BB22]. Fact

[SG18, CK92]. Fact-Based [SG18]. factor [Deb70]. Factors [Bra11b, Jun96, Ano78d, Biz92]. Facts [Jac03, Kin86]. failure [RH84]. Faith [Har23, Mor14, Wei11]. Fall [Gro17b, Gur02, Kra04, Gri74, Gri85]. Fallacy [Har12a]. Familial [Ric06]. Family [Cec24, MM17]. Far [Str23]. Farber [Dav97b, Gue21, MAR21]. Farewell [Far11a]. Farmer [Kev05]. Fascism [vH19]. Fashioning [AM14]. Fast [PKS18]. Fate [Mun23, Rud18a, Yi25, BGZ88]. Fates [JK99]. Father [Wol12]. Fatigue [Heg15, Joh15, Sch15]. Fauna [Pir24, Boc88]. Feathered [Ala04]. Feed [WJ22]. Feeling [KG21]. feelings [Gro79]. Félix [Ano75l, Ano75g, Bed83b]. Female [Des05, Neg95, Sta95, Van94a]. females [Ano91b]. Feminine [Hut22]. femininis [Ano81g]. femininity [Ano91a]. Feminism [Ham97b, Van94a]. ferment [Bon84]. fermentation [Koh71, Koh72]. Fern [Kli98]. Fernel [Bon90]. Ferri [Bec05]. Fertilization [Cog22, San73]. Festetics [PSBS^{+}22]. Fever [Löw97]. Few [Sta23]. fiber [Bil71]. Fibre [Jac03]. Fichman [Gre04]. Fictional [KC97]. Field [Adl14, All16, Gor19, Gut97, Irm05, Kin09b, Koh12, Lew18, Lin24, Mit03, Mon05, Par08, Sto15, Vet06, Vet12, Whi19, Sap83, DC96, Kro97]. Fields [Thi01]. Fiers [Pie12]. Fifty [Die17c, PBDL17, SK22]. figments [Kin86]. Figures [vP21]. Film [Ano00c, Ros18]. Fin [Coe06]. Fin-de-Siècle [Coe06]. Finance [Hod09]. finches [Sul82a]. Finland [Lep15]. First [Dre20, Eig97, Gal21, Grm68, Gue13, Ham07, Hun19c, Rad04, WdLdS21a, WdLdS21b, Ano76c, Ano76f, Baj88, CB82, Gri74, Wer71, Die17b]. fish [All90, Bor89, CB82]. Fisher [Hal07, Moo95, Tab08]. Fisheries [FC21, Kli98]. Fishery [Tay98]. fishes [Dex70]. fitness [Fry96, Mic86]. Fittest [Pau88]. Five [Rus05, YD00]. Fixation [Sin97]. Fixed [Bel06]. fixity [Hol69]. Flannery [Cao24]. Flaws [Mil11]. Flesh [Ben09]. Flora [Ras22, SP25]. Flow [KC94, Ano75f]. Flowers [Ano05a]. Fludd [Deb70]. fluid [Hom70]. Fly [Koh22a]. Follow [Gue24]. Fontenelle [Cor71]. Food [Bra11b, LD19]. Foods [Bis23]. Foot [Har15]. Forbes [Boc90, Fat17, FG78]. Force [Cog22]. Forces [Edd23]. Forebears [SS17]. Foreign [Lin24]. Forel [Pir24]. foreword [Men68, MSA69]. Forgotten [Bil16, cH17, Ras99, Riz20, SC14, Rad23]. Form [Dre20, Gra20, Har11b, Mat16, Ste19]. Formation [Iid10, vdV14, Söd94, vM71]. formative [Uch93]. forming [Dre20]. Forms [BP19, Tru97]. Forward [MM24]. Fossil [Bri10a, Bri10b, LK25, Rai89]. Fossils [Lyo93, Mat87]. Found [vP20]. Foundation [Cas15, GM18, Yi21, vM71, Löw97, Zal92]. Foundational [RWE13]. Foundations [LK25, Ano75b, Ano82m, Har94]. founder [Ege68a]. Founding [Smo94, SR23a, Ada68, All90, Car71]. Four [PSBS^{+}22, Rie90]. Fowler [Moh18]. Fractures [Pau16]. Frame [Hag03]. Framework [Rob11]. France [Bon06, Gau06, GZ98, Sch10]. Francesca [Mar09]. Francis [De 74, Ric25, Wal01, Bul99, Cow72, Gök94, Sch14]. Franco [For20]. Francois [Pir24]. Frank [San10, TP94]. Franklin [Sel03]. Franz [Gli98]. Fred [Mét16]. Frederic [Koh04]. free [Cre96, Koh71, Koh72]. free-standing [Cre96]. Freedom [Sel12]. Freidberg [Sac09]. French [BGZ88, Far74, Gau93, Gro79, Loi11, LB04, Mor97b, Tho05, Wal21]. Frequency [SK22]. Fresh [Sac09]. Freud [Bra24, Sla08, Ste76]. Friedberg [Sel05]. Friedrich [Hon16, Roe79, VHL20, dG02]. Frisch [Dhe21, Mun05]. Frits [Kin09b]. Fritz [Bra22a]. frog [Hol93]. front [Kre96]. fronts [Pau83]. Fullilove [Fit19]. Function [Bec69, Bra11b, Har06a, Har22b, Grm68, Sha72, Swa70a]. Functional [Rud18b, Eds72, Rai89]. functionalism [Rie90]. Fundamental [Ile99]. Fungi [Kol22]. Further [Bed72, Biz92, Sul83]. Future [Gli24, Tom22]. Futures [Rad25].

G [Atk85, Coo02, Gli07, Hal10]. G.

[DC96, Kay89, Sla03, Smo99b]. GA [Por09]. Galápagos [Lin24, Sul09]. Gale [Rhe96]. Galen [Boy86, Boy07, Cos97, Pre77, Sha85]. Galenic [Boy84]. Galileo [You09]. Galton [Bul99, Cow72, De 74, Gök94, Wal01]. Game [Shm20]. Gänger [May23]. Gap [Har13]. García [Ste24]. García-Sancho [Ste24]. Garden [Fla24]. Garland [Die16b, Kle23, Mai16, Mai23b, Men16, Ric23, SR23b]. Garstang [Esp20]. Gegenbaur [Di 95]. Gel [Chi09]. Gender [Biv00, Fin91, Koh22b, Li04, Muk14, Ano91b, Joh19b]. Gendered [Hut22]. Gene [Pie12, Som08, Yi25, Car71, Car74, Gil78]. genera [Len80]. General [Rhe96]. Generality [Log01, Log02]. Generation [De 19, Mat10, Pre70, Str99, Zwi18, Bow73, Cor71, Far72b, Far72c, Rue84]. Generational [Ano78d]. Generations [Lau19]. Generous [Bar23b]. Genes [Led89, Zal05]. Genesis [Civ13, Ano81a]. Genespeak [San10]. Genetic [Bur22, GS15, KG21, Kle24, Mar24, Pie12, PSBS^{+}22, Ste07, Fal91, Gla74]. Genetically [Rad04]. Geneticist [BS08, Joh22]. Geneticists [Rad04, Ric06, Lud69]. Genetics [Ank20, BPA05, Biv00, Bon06, DT07, Fal03, Gor09, Gro17b, Har06a, Har22b, Iid10, Kev05, Kim06, Kle16, McG21, Mit17, Pin02, Rad98, RF18, RN11, Rhe00, Ric07, San06, Sel12, Ser21, SM07, VHL20, Veu19, Wei11, Wei05, Wei06, Wol12, dC06, Ada68, All74b, Ano78a, BGZ88, Bur93b, Cow72, Fal95, Gai80, Kay89, Kre96, May73, Olb74, Sap83, Sta95, Sum91, The96, Die04, Soo21, Kon23]. Genève [Ste04]. Genitalia [Fun12]. genius [Ano75g, Bed83b, Gök94]. Genome [GD18, JACD18, Pie12, dC18, Ban22, Lin18]. Genomics [Nov18, Sch18, Ste18, Ste24, GD18]. Genotype [Lam15, RH09, Tab08, Chu74]. Gentleman [AN08]. Gentlemanly [Wal01]. genus [Atr87, Piñ88]. Geochronometric [Mar23]. Geoffroy [Gal21, Lag03]. Geographical [ÁTBHBJB24, Fin08, Kin80]. Geographies [Ile05]. Geography [Hoq16, cH17, Lar86, Mar09, Col86a, Hag86, RPM23]. Geology [Bel05, Ano78c]. Geopolitics [Sur22]. Georg [Kol22, Ste04]. George [Hal03, Bel81, Cog22, Gue85, Hal10, Har11a, Kev05, Pir23, Sch95, Tab04]. Georges [Cau19, Eig97, Out86]. Georges-Louis [Cau19]. Georgia [Por09]. Germ [Win01b]. German [Nyh18, BS08, Bur72b, Far72c, Rei86, Ric89b, Ric00]. Germany [Bra20, Far05, Har12b, Lew04, Mac93, Wei87, Wie06]. Germs [Pat09]. Gerontology [Par08, Par13]. Gestation [Nyh18, Hod72]. Ghent [Pie12]. Ghost [Bru18]. Gilbert [Gia18, Mul69]. Gilded [Bar00, Par21, Thu02]. Giovanni [Hei19]. Giraffe [Lag03]. Girard [JK99, Qui17]. Glacier [Rum12]. Gladfelter [Irm05]. Glance [Cec24]. Gland [Log07]. Glass [Jia24, Wol11]. Global [Ben11a, Fit19, Gos23, Kle04, Mac21a, Mac21b, Mar23, Wu19]. Globalizing [Ste18]. glycogenic [Grm68]. goals [Bor89]. God [Ano06b, Bez97, Cor87, Osp80]. Goes [Sch25, You09]. Goethe [Rob11, Ste02]. going [Raa76]. Golden [Wol12].

Goldschmidt

[All74b, Die95, Die11, May97]. Good [Hub20]. Goodman [Hag10]. Goodrich [Wai88]. Goodsir [Jac83]. Gordon [Fun12]. Got [Wil19]. Gould [Sep05]. Government [Sto15, Vor92]. Gowan [Pri24]. Grace [Sla03]. Gradients [Thi01]. Gradual [Bri10a, Bri10b]. gradualism [Rho87]. Graham [Bru18]. Grand [And24]. Grande [Gru18]. Grant [AM68, Ano91e, Des84]. Granville [Sla03]. Gray [Por93]. Great [Mil08, Pea10, Boc90, Cro89]. Greater [Sch25]. Greek [Nat24]. Green [Zal93a]. Greene [Por04]. Greenhouse [Kin09b]. Gregg [Ano00c]. Gregory [Gli24, Rai89, Sug23]. Grew [Kle04]. Griffith [Mét16]. Grinnell [SG18]. Groningen [Bei08]. Group [Cre10a, Sum93, Bow94, Car74, Wai88]. Grow [PKS18]. Growth [Bri09, Gre92, Par13, Pau93, Sin89]. Guerrini [Bod23]. Guide [Var09]. Gulf [CB82]. Günter [Liu23]. guts [Chu89a]. Györgyi [Bei08].

H [AKB89, ASR^{+}93, Ano69a, Ano78e, Ano78f, Ano80a, Ano80b, Ano82c, Ano82i, Ano82j, Ano82k, Ano84b, Ano85b, Ano85c, Ano91g, Ano92, Ano96, Ano98b, BBF^{+}91, BA81, CRBD90, CHCC97, FHP^{+}91, FRS^{+}95, FMR^{+}98, FH77, FG78, FH78, Fra79, FH79, Gil90, GLB^{+}94, HS72, HF73, HF74, Hal74, HLM^{+}95, KNM^{+}93, Lyo95, MAC^{+}94, MFM^{+}96, NSMB92, Nyh18, PKS^{+}95, RHW^{+}97, RG76, RFL^{+}98, Ric00, RGH^{+}81, Roe82, RS83, Roe83, Roe84, RBFC84, RRC^{+}85, Roe86, RCK^{+}86, RBK^{+}86, RBHK87, RRE^{+}87, RPCH87, RRCB88, RSBM88, RMR^{+}88, RKMB89, Roe89, Roe90, RBWB90, RH84, SJB^{+}93, SE71, SHB^{+}92, Swa68a, Swa70b, Swa70c, Swa71, TS^{+}96, Wei09]. H. [Car71, Gli07, Hal10, Moo95, Rus15, Str99, Tay88]. habits [Sul83]. Haeckel [Di 95, Esp20, Rin81, WdLdS21a, WdLdS21b, dG02]. Haeckelian [Erl02]. Hagen [Kim23b]. Hagiography [Lag03]. Haldane [Ada00, Pau83, RN11, Tei07, Wer71]. Half [Dre20, Gue13, Ano76f]. halftones [Ban22, Lig18, McG21, Mes19, Pow18]. Hallam [Lin18, dC14]. Haller [Ano75j]. Hamburger [All04]. Hamilton [BA81, Can13]. Handmaid [Ton18]. Hans [Bra22a, Bra22b, Ham99, Mor23, Rin88]. Hans-Jörg [Mor23]. Harbor [Kev05, Sel05, Cas15]. hard [Hei19]. Hardback [Ank18, Bru18, Tan20b, Wal20]. Hardcover [Ber19, Cre19, Moh18, Mon18, Sch18, Web20]. Hardening [Hum08]. Harman [Sat20]. Harnessing [Thu02]. Harrison [Sla03]. Harry [BA81]. Harvard [Bru18, Kei23, Mil11, Sac09, Stu24, Tan20b, You09, Ank23, Bro23, Heg15, Joh15, Oak15, Par71a, Sch15]. Harvey [Gue24, Bon90, Deb70, Fra72, Sha85]. Harway [Ham97b]. Hastings [Ala12]. Haunting [Som24]. Haushofer [Bis23]. Haüy [Ste84]. Havelock [Ala24]. Haven [Ano06b, Ben18, Cao24, Hag24, Kei11b]. Hazard [Sch25]. head [Raa76]. Health [Cre19, Dru23, Joh19b, Pat09, Ras24a, Smi90, Smi92, Bar04a]. heat [Ano78c, Hal71]. Heath [Mon18]. Heggie [Lan20]. Heidelberg [Mét14]. Heinrich [BS08]. Helen [Kim18, Ogi07]. Helical [Har11a]. Helix [Gin10]. Helmont [Bon84]. Helpers [KC97]. hemoglobin [Eds72]. hemolysis [Sil94]. Henderson [Fry96, Par71a, Par71b]. Henri [App80]. Henrika [Kro97]. Henry [Bel81, Hal03, Sch99b, Eng97, Hay88]. Hensen [Lus74]. Her [Pau23, Win01a]. herbae [CC18]. herbal [Sca78]. herbals [Sca78]. Herbarium [Cao24]. Herbert [Ano82h, Ano91h, Bec05, Hal03, Raa76]. herbis [Ano81g]. Herd [All13b]. Hereditary [Gök94]. Heredity [Bra24, Bra22a, Bru02, Bul99, Cec19, Coo98, GZ98, Hum08, LD19, LD20, LGB17, LB04, Ric06, Thu02, Tur18, Wei68, Wei06, Win00, All91, Gai73, Sap83, Ste76, Wel71, Ban22]. heresies [Die95]. heritable [Lin92]. Heritage [Rob11]. heroes [AA82]. Herrera [LMB03]. Herschel [Sch89]. Herz [Han93]. Hesketh [Bro24]. heterosis [Lew87]. Heterotrophic [Sin97]. Heuristic [Mat16, Sil94]. Hey [Pri24]. Hidden [Cao24]. Higgins [Irm05]. High [Lad08]. Higher [Lan20]. Hilaire [Gal21]. Hilde [Ham84]. Hilgartner [Sch18]. Hill [Koh24, Ano78e]. Hincks [Cog02]. Hippocratic [Boy84]. Hiroshima [Lin92]. Histoire [Ano76c, Ste04, Lov06]. Historian [SR23b, Win01a]. Historians [Lus01, FS90]. Historical [Bro24, Chi09, Dex70, Har11a, HR11, JK99, Jax98, Joh07, PSBS^{+}22, Str19, Sum91, Tan17, Tau97, CKT94, Wel73, Wil85]. histories [Get91, Mon18]. Historiographic [KBCT97]. Historiographical [vH19, Har94]. Historiography [Bro22, Meu21a, Meu21b, Pau16, SS99, Tan17, VHL20, May73]. History [Ala12, Alb01, All91, All16, Ano76e, Bar97, Bar00, Bar05, Bar24, Ben11a, Ben18, Bur99, Bur22, Cor05, Cre10b, DM97, Dav97a, Die16b, Die17a, Die17c, Eli14, Fag07, Fin08, Fla24, Gli07, Gru18, Gur02, Hal25, HR21, Har08, Hei19, HS18, Hol14, Joh04, Joh07, Joh19b, JACD18, Kin20, Kle11, Koh24, Lad08, Lan20, Lla00, Lyo95, MM24, Mai01, ML10, MAR21, Man13, Mar24, Mar23, Men17, Mes19, Mét14, Mir08, Muk19, PBDL17, RR22b, Ras24b, Ric23, RC11, Sch10, Sch90, Sep13, SR23a, Smo24, Som08, Ste24, Str23, Sug23, Sun13, Tan20b, Tom22, VWP20a, VWP20b, Whi20, Wit16, dC14, AJR94, Ano75h, Ano78c, Ano82d, Ber70, Bow94, Bro71, Bro73, BGZ88, Chu70, Chu89b, Ege83a, Ege83b, Ege85, Gil87, Jun96]. history [Kel90, Kin94, Koh95, Lar86, McC69, Nel78, Por95, RM84, SP81, Sin89, Slo85, Sac09]. Hitchcock [Arc09]. Hitler [Far05, JW13]. Hitoshi [Iid10]. hive [Ano91b]. Hogben [Erl16, Tab08]. Hogue [ZS15]. Holds [Hen03]. Hole [Luk22, ZS15, Luk21]. Holism [Bra22b]. Holistic [All04, Par71b]. Holists [Bec16]. Hollaender [Rad06]. Holliday [Jia14]. Holmes [Koh04, Sch14]. Holocoen [Jax98]. Home [cH17, Raa76]. Homeostasis [Kim23b]. Homogeneity [Wal15]. Homology [Gur02]. Homonym [Wit16]. Honey [Mun05, Ano91b, Pre90]. Hooker [Bel01, Bel06, End11, Por93]. Hopkin [Ras24a]. Hopkins [Ben85, Bod23, Gib24, Man13, Sug23, Web20, Lon09]. Horace [Eds80]. horizon [All91]. Hormonal [Kei01]. Hormone [Kei11b]. Hormones [Ras99]. Horseshoe [Bar04a]. Hospital [Rhe96]. Host [And16, MD16]. Hours [Sta23]. House [Kim06]. Howard [Fis10]. Howling [Man09]. HSTM [Far09b]. Hughlings [Smi82]. Hugo [All69, Sta03, Zir68]. Hulu [Ton18]. Human [Art17, Ban22, Bar04a, Bra24, Gam21, Goo19, Gor09, Gor19, GD18, Hal03, Hop24, KG21, Mar24, McG21, Mit17, Oak15, RF18, Som21, Soo21, Stu24, Sul11, Vet06, Vic18, Wei11, Wei06, dC18, KC94, Maz75, Num79, JACD18, Ray19]. Humane [Shm23a, Shm23b]. Humanities [BR20]. Humans [Bod23, Har13]. Humboldt [Gre25, Ege70, RPM23]. Humoral [CT97]. Hunch [Goo19]. Hundred [Gau98]. Hunt [All68, Led89, Mai16, Som21, Sun10]. Hunterian [Pad95]. Hunters [Som05, Sug23]. Hussey [Eng97]. Hutchinson [Pir23, Sla03]. Hutton [Ste90]. Huxley [Ano82c, Bar95, Gas91, Gre90, Hal03, Lyo93, Lyo95, Ric00, Sch90, Sch99b, Swe95, Wai88, Wer71]. Hybrid [Kle11]. Hybridization [Kle16, Gia93]. hybridoma [CK92]. hybrization [OV68]. Hyde [Zir68]. Hydrobiology [Riz20]. Hyena [Fun12].

Hygiene [JW13, Har89]. Hypnozoite [Mar11]. Hypotheses [Hal98]. Hypothesis [Fis10, For12, Hol14, Kei01, LaB17, Smi92]. Hyung [Jin25].

Ian [Bro24]. iatromechanical [Hal71]. Ibex [vH19]. Ibis [Joh04]. Icelandic [Erl02]. Ida [Muk14]. Idea [LaB17, Swi14, Tur18, WdLdS21a, WdLdS21b, CT90, Cor83, Osp80]. ideal [Cla93]. Ideals [Bel01]. Ideas [Bul99, Gli98, Hon16, KK13, Ano81a, Bon84, Ege69, Rue83, Wel71]. Identification [Sch09]. Identifying [Lin92]. Ideology [JF03, Gil87, Kal83, Wei09]. Idiotypic [Civ13]. If [Har12a]. Ignored [BH21a, BH21b]. II [McG19, Bea19, Hag15, Hag17, Her77, MKC69, Smi82, Tay88]. III [Por09, Ric23]. ill [Smi90]. Illiger [May68]. Illinois [Wei09]. illus [Ank20, Ank18, Ano04a, Ano05a, Ano06b, Bar05, Bel05, Ben18, Ber19, Bri09, Bru18, Can13, Cau19, Coo20, Cre19, Cro05, Goo19, Hub20, Ile05, Irm05, Kei11b, Kev05, Kim18, Kin20, Lan20, Mar09, McG18, Mét14, Mil11, Rab11, Sat20, Sel05, Som21, Ste04, Tob09, Yan18, Zal05]. illus. [Por09]. illustration [Por95, Mes19]. Image [Loz19, Mét14]. Images [Rue83, Tru97, Ano91a]. Imagination [Bra11a]. Imagining [Bro24]. Immanuel [Ano76a]. Immediate [Sch04]. Immigrant [Lin23]. immigration [Ano91h]. Immune [Swi14, Tau97, Sil94, TP94]. Immunity [CT97, CT90, Maz75]. Immunochemists [CT97]. Immunodiffusion [Hag10]. Immunological [AM14]. Immunology [CKT94, For12, KC97, Löw97, SS99, AJR94, Löw94, Sti94]. immunotherapy [Löw94]. Impact [Gal17, Gin10, Pat09, Ben88, DCS01]. Impacts [Fol21]. impediment [Pre90]. Imperial [Sta25, Man15]. imperialism [Nyh95]. Implications [O’M18]. Importance [Rai03, Wai88]. Impresions [Eig97]. Impressions [Ano05a]. Improve [Mat06]. Improving [Rai01]. Inaugural [RR18, RS23, WY23]. Inaugurating [Die16a]. Inbreeding [Mat06, Ogi07, PSBS^{+}22]. Incident [Cob25, Gor97]. incipient [Par82]. Incomplete [Dre23]. Indeed [Gro17b]. independence [Fal95, Coo20]. Index [Ano98a, Ano99d, Ano00d, Ano01c, Ano02d, Ano03d, Ano05d, Kim18]. India [Bar23a, BJ97]. Indian [Mar04, Lew05]. Indians [Por09]. Indigeneity [Som24]. individual [Ano81f, Car74, Rin81]. Individualism [Coo02]. Individuality [AM14, Gay98, Str19]. Individuals [NL11]. Individuen [NL11]. Indonesia [Gos23]. Indricotherium [Man15]. Induction [Arm12, Log07]. Industrial [Oak15, Thu02]. Industry [Hod09, Sch25, Smo25b, WJ22, Len87, Pau93, Smo25b]. Inevitable [LD19]. Infamous [Cec24]. Infancy [SS99]. Infection [Den16, MD16]. Infections [Hon16]. Inference [Gro17a]. infestans [Mat10]. Influence [NSRG09, All90, Ano76d, Cor87, Hai91]. Influential [San06]. informatics [dC14]. information [De 96]. Infusoria [Chu89a]. ing [CC25]. Inheritance [Cec19, Gli24, Hen06, Hoq24, NSRG09, Ona10b, Ric24, Bar92]. Inhibition [BH21a, BH21b]. Initial [Ano76c, Far74]. Innovation [Kim18, Nov18]. Insect [BR20, Ona23]. Insight [GS10]. Insights [FE20]. inspiration [Bur72a]. instability [Ams91]. Instinct [Bri05, Bur72b, Ric81, Stu24, Vic23]. Instinctual [Sul11]. Institute [Lon09, Wei06]. Institution [Rai03, Wil12, Rad98]. Institutional [GH19]. Institutionalization [BPA05, Erl09, LMB03, Wei87]. Institutions [Kim06]. Instrumentalisation [Som05]. Integrated [ML10]. Integrating [Str19]. Integration [Hag15, Lam15]. integrative [Swa68b]. intellectual [Chu70, Out86]. Intelligence [Har08, Ser21, Pla69, Ric81]. Intelligent [Pau23]. Interaction [Fag10, Lou09, Tab08, Gre90]. Interactions [MD16]. Interactive [Fal03]. Interdisciplinary [Rad25]. Interest [Mai01]. Interesting [End11]. Interface [MD16, Gil90]. Interfield [BA07]. Interlinked [Alt07]. internal [Bor76, Hol69]. International [Gib24, Joh09, Kol22, Man15, Kre96, CM11, Sch17, Ste18]. Internationalism [Tyr23, Bea93]. interplay [Ric89a]. Interpretation [Bec05, Gru22, Ric00, vP20, Gre92, Maz75, San73]. Interpretations [Esp13]. Interpreting [Bez97, HA19]. Intertwined [Sel03]. Intervention [Tru97]. Interwar [BB22, Bra22a, Bra22b, Mét16, Ams91]. Intracellular [Sta03]. Introducing [JRR18]. Introduction [All74a, AM01a, AM01b, Ano81d, BPA05, Bol22b, CKT94, CC25, Chu89b, CBM86, Die12, Die16b, Gra92, GD18, GM13, Har11b, Har06b, Heg15, Hun19b, KBCT97, Kin09a, LB93, Liu24, Ras25, RC11, Smo24, HG76, Irm05]. Introductory [Ano93, Rud00]. Introgressive [Kle16]. Invalid [Col98]. Invasion [And24]. Invasive [And24]. Invention [Web20, Tho05]. inventory [Ege83a]. Invertebrate [All20]. Investigation [Ack06]. Investigative [Koh04, Pas21]. Investigator [Sum04]. Iowa [Sin00]. IQ [Ser21]. Irene [Wil16]. irrational [Deb70]. Irreversibility [Gou70]. Isaac [LMB03, Rud90]. ISBN [Ank20, Ank18, Ban22, Bar23a, Bar24, Ben18, Ber19, Bis23, Bod23, Boo23, Bra20, Bro24, Bru18, Cao24, Cau19, Coo20, Cre19, End23, Fic24, Fla24, Fre25, Gia18, Gib24, Gli24, Goo19, Gor19, Gre25, Gru18, Gue23, Gue24, Hag24, Hal25, Hei19, HS18, Hub20, Jia24, Jin25, Kei23, Kim23b, Kim18, Kin20, Kle24, Koh24, Lan20, Lig18, Lin23, Lin18, Liu23, Loz19, Mar24, Mat21, May23, McG21, McG18, Mes19, Moh18, Mon18, Mor23, Muk19, Nat24, Nov23, Nyh18, Pau23, Por25, Pow18, Rad23, Ras24a, Ras24b, Ray19, Ric25, Sat20, Sch18, Shi18, Shm23a, Shm23b, Som21, Som24, Ste24, Str23, Stu24, SD19a, Sug23, Tan20b, Tyr23, Wal20, Web20, Wu19, Yan18, You25]. ISBN-10 [Loz19]. Isfahānī [Gam21]. Ishi [Mar04]. Island [DC96, Mit21]. Islands [Lin24]. isoantibodies [Maz75]. isolated [Ano75g, Bed83b]. isomerism [Ano75a]. Isotopes [Gau06]. Issue [GD18, Har11b, Har06b, Hun19b, Pir24, WP06]. Issues [KBCT97, Mit17, dLK12]. Italian [Cas12]. Italy [For20, Pir24]. iv [Mét14]. Ivanovic [Pir23]. ix [Ben09, Irm05, Kev05]. Izumi [Joh19b].

J [Ack04, Ben18, Bez97, Car71, Cau19, Cro05, Fry96, Fun12, Gru22, Ham97b, Kei23, Moo95, Par71a, Par71b, Rus04, SD19a, WP06]. J. [Ada00, AKB89, ASR^{+}93, Ano69a, Ano78e, Ano78f, Ano80a, Ano80b, Ano82i, Ano82j, Ano82k, Ano84b, Ano85b, Ano85c, Ano91g, Ano92, Ano96, Ano98b, BBF^{+}91, BA81, Bur93b, CRBD90, CHCC97, FHP^{+}91, FRS^{+}95, FMR^{+}98, FH77, FG78, FH78, Fra79, FH79, GLB^{+}94, HS72, HF73, HF74, Hal74, HLM^{+}95, KNM^{+}93, MAC^{+}94, MFM^{+}96, NSMB92, PKS^{+}95, Pau83, RHW^{+}97, RN11, RG76, RFL^{+}98, RGH^{+}81, Roe82, RS83, Roe83, Roe84, RBFC84, RRC^{+}85, Roe86, RCK^{+}86, RBK^{+}86, RBHK87, RRE^{+}87, RPCH87, RRCB88, RSBM88, RMR^{+}88, RKMB89, Roe89, Roe90, RBWB90, RH84, SJB^{+}93, SE71, SHB^{+}92, Swa68a, Swa70b, Swa70c, Swa71, TS^{+}96]. Jacana [Goo19]. Jack [Sep05]. Jackson [Dav97a, Smi82]. Jacob [Boc90]. Jacobs [Ben18]. Jacques [BH21a, BH21b, Mai09]. Jail [You09]. James [Cro05, Mar09, Pow18, Sel05, Ste24, Bez97, But18, Fic24, GS15, Gue85, Mor14, Nov23, Sel03, Wal20]. Jameson [Ano91e, Jen16]. Jan [Sin00, SS17]. Janaki [Ric25]. Jane [Pow18, ZS15, Ank18]. Janet [Ano04a]. Janice [SP25]. Janina [Ras24b, Ste19]. January [McG19]. Japan [Ona10b, Uch93]. Japanese [Eri17, Iid10]. Jay [Sep05]. Jean [Ste04, Tes23]. Jean-Baptiste [Tes23]. Jean-Daniel [Ste04]. Jean-Marc [Ste04]. Jeannie [Ona23]. Jefferson [Por25, Por04]. Jeffries [App88]. Jenkin [Hoq24]. Jenkins [Tan20a]. Jennings [Ano91h, Rus15]. Jenny [Soo21]. Jerne [Söd94]. Jerome [Dav97a]. Jesup [Kle13]. Jesús [Cre19]. Jewish [BS08, Bur22]. JHB [AM17, RR20b, RR21b, RR22c, RR22b]. Jim [Pat09]. Joan [Bra20]. job [Zal93a]. Jobs [Bol22b]. Joel [Fla24, Kim23b]. Joeri [Whi19]. Johann [Han93]. Johanna [Kol22]. Johannes [dB11]. Johannsen [Chu74, RH09]. John [Loz19, Sch95, Bel05, Gil87, Jac83, Nyh18, Pie19, Por04, Rad25, Sch89, Slo72, Smi82]. Johns [Bod23, Gib24, Man13, Ras24a, Sug23, Web20, Ben85, Lon09]. Joint [Far09b]. Jordan [Mil11]. Jordan-Young [Mil11]. Jörg [Mor23]. Joseph [Bel01, Bel06, End11, Fra79, Hol69]. Journal [Die17c, Joh04, Men17, PBDL17, Ano85a, Die17a, MAR21, Ric23, SR23a]. Journals [Far09b, SP81]. Joy [Moh18]. Jr [Dav97a]. Jr. [Smo99b]. Juan [Piñ88]. Jubilee [Wol12]. Judgment [Ada00]. Judith [DC96]. Judson [Eds80]. Juke [Cec24]. Jules [CT97]. Julian [Gre90, Bar95, Gas91, Swe95]. Julie [Web20]. Julie-Marie [Web20]. Junction [Coo20]. June [Ton18]. Junker [Gli04]. Jurassic [Rab11]. Jussieu [dC23]. Just [Tho05]. Justus [WH02]. Jutta [HS18].

K. [Gru22]. Kaempfer [CB82]. Kaiser [Wei05, Wei06]. Kametaro [Ona10b].

Kammerer [Gli06, Log07]. Kangaroo [Hag15]. Kant [Ano76a]. Karen [Led04, Moh18]. Kärin [Str23]. Karl [Ano76d, Dhe21, Hon16, Liu23, Mun05, Rie12, WY23]. Kass [Kle24]. Kathleen [Koh24]. Keep [Har23]. Keill [Gue85]. Keith [Ano06b]. Kellogg [Lar99, Ona10b]. Kenney [Loz19]. Kenya [Lew18]. Keogh [End23]. Keys [Sch09]. Ki [Joh19b]. Kihara [Iid10]. Killer [Hal98, Stu24]. Kind [Bar23b]. King [Rai89, Ogi07]. Kingsbury [Kle11]. Kingsland [Hag24]. Kingsley [Hal13]. Kirby [Var09]. Kluyver [Sin00, SS17]. Knapp [Gib24]. Knopf [Ano04a]. Know [Mor14]. Knowing [vHS19]. Knowledge [Bru02, Jon17, Nic22, Sch18, Eds72, Sin89, Ile05, Mar09]. Known [dC23, Ano75m]. Knox [Ric89a]. Kohler [Gor19, Kro97, Mit03]. Konrad [Bri05, Kal83]. Korean [Jin25]. Kowal [Som24]. Krech [Por09]. Krimsky [Kei01]. Kristine [SM07]. Kuêma [De 19]. Kühn [Rhe00]. Kuklick [Kro97]. Kuljian [Goo19].

L [You09, You25]. L. [Fry96, Gor09, Par71a, Par71b]. Lab [Hag24, Kin09b, Mit03, Pie12]. Labels [dLK12]. Labor [DT07, Hal03, Tam95]. Laboratories [Kay96, Mac21a, Mac21b, Bur93b]. Laboratory [Adl14, Chi09, Hag24, Kev05, Luk21, Luk22, Sel05, ZS15, All90, DC96, Koh93, Boc95, Erl09, Heg15, Joh15, Mar23, Pie12, Sch15]. Labour [D’H12]. Labs [Vet12]. Labscape [Kin09b]. Labscapes [Mit03]. Lacepède [Sch10, Bor89]. Lack [Bor03]. Lady [Str24]. Lagrange [Hal71]. Lake [Kli98]. Lakes [Boc90]. Lamarck [Ano82n, Bur70, Bur72a, Ege68b, For20, Gal17, May72, Tes23]. Lamarckian [Gis23, LGB17]. Lamarckians [Ano91e, Cor21]. Lamarckism [Gai80, HA19, Loi25, Rin88, Sla08]. Lamarckisms [Loi11]. Lance [Gru18].

Lancelot [Erl16, Tab08]. Land [Hod09, HR11, Ste05, Wil04, You25, Fic77]. Landscape [Har06a, Har22b, Wil04]. Landscapes [Gal22, Mit03, Yan18]. Landsteiner [Maz75]. Langston [Kei11b, Wil04]. Language [Biv00, Mun05, Rad05, dG02]. Languages [Bez97, Bon90]. Lashley [Gru22]. Lasker [KG21]. Last [Ada00, Bei08, Mai23b, Mar04, SK22]. Late [Mat10, Wie06, Ben85, Bor89, CB82, Gai73, Lar86, Pau84, Sch04, Alb01, Bla07]. late-19th-Century [Sch04]. Late-Victorian [Bla07]. latency [Cam83]. Latent [Hon16]. Later [Fat17, KK13, Rus05, WO05, YD00]. Laughlin [BA81]. Laura [Kei23]. Laureate [Kle24]. Lauren [Kli98]. law [Ano76d, Ano78b, Cor87, Gou70, Rin81, Esp20]. Lawrence [Dav97b, Gue21, Koh04, Ric24, Wel71]. Lawrie [Ano82i]. Laws [PSBS^{+}22, Gou70, Out86, Sta95]. Lawson [Ste05]. Leadership [Koh22b]. Leclerc [Cau19]. lecture [Pad95, WY23]. Lectures [Cai02, Kle13]. Led [Sla03, Lap90]. Lederer [Ben09]. Ledyard [Smo99b]. Lee [Kle24]. Leeuwenhoek [Cas73, Ege68a, Rue83, Rue84]. Left [Ank04, Lep15]. Left-Wing [Lep15]. Legacy [Die04, Irm05, Kei11b, NG14, Sch25, Sch17]. Legallois [Che13]. Legend [Sul09, Sul82a]. Legislation [Hic21, Ano91h]. legislator [Out86]. Lehrman [Vic23]. Leibniz [Rie88]. Leiden [Mir08]. Lenny [Zal05]. Lens [Mat21, Stu23]. Leo [Can15]. Leonard [Ser16]. Leonelli [Shi18]. Leopold [You25]. less [Len80]. lessened [Gro79]. Lessons [Ser21]. Let [Hat21]. Letter [Ano89a]. Letters [Bed73, Whi22]. Leuckart [NL11]. Leung [Joh19b]. Level [Kei17]. levels [Gro69]. Lewes [Bel81]. L’Hérédité [LB04]. Liberty [Zir68]. Lice [Boc12]. Lick [Mat87].

Lidgard [Str19]. Liebig [WH02]. Life [Ack07, Arc09, BP19, Bol22a, Bra20, Bur99, Can13, Che13, D’H12, Die16b, Elw07, End11, Erl16, Fic24, Gur02, Hal25, Kim23b, Kle24, KK13, Kol22, Liu23, Mai09, Man13, Mét14, Nov23, Ric25, Sat20, Sch18, Sel05, SS17, Smi99, SR23b, Söd11, Str99, Str00, Str23, Sul11, Wal01, Wil19, Wol13, Wu19, dR23, vP21, Ano75n, Jac89, Kin80, Koh93, Olb71, Pod96, McG18, Ras24b, dC14]. Liggers [Gal22]. Light [Fre25, Muk19, Ano76d, Ano82m, Zal93a]. Liguus} [Gal22]. Likely [Jia14]. limitations [Ano76a]. Limits [Che13, Coo98, Rho87]. Limnology [Pir24]. Line [Vet06, Pow18, dC14]. Lineage [Gur02, Ano78b]. Lines [Sur22]. lining [Get91]. Linkage [Fal03]. Linking [Van94b]. Linnaean [Rob11]. Linnaeus [dC23]. Linus [Mor98]. Lisa [Bis23]. Lisbon [GM18]. Listening [Whi19]. Literature [Sep06, Kot81, Sht92]. Little [Pie19]. Live [Mel13]. liver [Grm68]. Lives [Gal22, Men03, Sel03, Whi22, Sch89]. Livestock [Gib24]. Living [Coe06, Cro05, HR21, LK25, Muk19, Oga01, Ric04, Sch04, Smo99b]. Livingstone [Ile05]. Lloyd [Arn19, FG17]. load [Pau87]. Loblolly [Ber70]. Local [Rai03, Gau93]. localization [Swa70a]. Locating [Lue21, PH22]. Locke [Slo72]. locomotion [Fra91]. locus [Gro79]. Loeb [Mai09]. Logic [Art24]. London [Bar05, Bri09, Cre19, Des05, Gue24, Ile05, Jin25, Kim18, Kin20, Kle11, Loz19, Mét14, Pow18, Ric25, Ste24]. Long [Lew18, Sch25, SD14, Fre25, Gla74]. Look [Wol11]. Looking [Die17c, MM24]. Lords [Koh22a]. Loren [Bru18]. Lorenz [Bri05, Kal83]. Lorraine [Som21]. Lost [Lad08]. Louis [San83, Cau19, Fra79, Irm05, JK99, Mor97b, Tho05]. Love [Pau23]. Lovejoy [Rus15]. Low [GS15]. Low-Dose [GS15]. Lowe [Ste24]. Lucas [NSRG09]. Luis [LMB03]. Luke [End23]. Luminous [Muk19]. Luria [Lin23]. Lurie [Irm05]. Lyell [And24, Ano78c, Byn84, Osp77]. Lyman [Lar99]. Lynch [Bel05]. Lynn [Str19]. Lyons [All22]. Lysenko [Bru18, Cas12, Dua20, Har03, Köh11, Kre96, Sel12, Tei07, dLK12]. Lysenkoism [Gor12, Pau83, Riz20, Yi21, dL12]. Lysogeny [LGB17].

M

[Bel05, Mil11, Muk14, Ste97, Ste05, Tan20b]. MA [Bru18, Mil11, Sch18, Tan20b, Tob09, You09]. MAB [Sch17]. Macfarlane [AM14, San10, TP94]. machine [Chu69, PW92]. machine-theory [Chu69]. Machines [Gib24, Tho05]. Macleay [Nov19, VWP20a, VWP20b]. Macmillan [Cre19, Far05, Ste24]. macroevolutionary [Rei86]. Made [Kim18, Smo05a, Wil19, Che09]. magnificent [Cor87]. Maienschein [Pow18]. Making [Adl14, And16, Cra03, Des01, Eri17, Gos23, Hen97, Lig18, Pal22, Par08, RPM23, Sch14, Soo21, Str10, SD19a, Tay98, Tur18, Wal21, Jac83, Gib24, Led04, Rad22]. Malacology [Gal22]. Malaria [Den16, Mar11]. Male [Neg95, Van94a]. Malformations [Hoo97, Ray22]. Malthus [Her71, Kly14, Osp79, Sch74b]. Mammal [Lag03, Cla93]. mammalian [Cla93]. Man [Ack04, Bel01, Bez97, Den16, For20, Hal13, KK13, Ray13, Sch84, Wil12, Wol11, Baj88, Bel81, Biz92, Bow89, Coo90, Her74, Her77, OM71, Har22a, pY14, Byn84, Gam21, Har22a, Pri24]. Management [Car08, Kei98, Kli98, Sta25, Sto15, You98, McE92]. Mangold [Ham84]. Manipulation [Ric24]. Manning [Wu19]. Manson [Li04]. mantis [PW92]. Manton [Wil16]. Manuals [Sch09]. manuscript [Ano75e]. manuscripts [Sul83]. Many [Har11b, Jia14, Whi22]. Maps [vP21, Ste05]. Maralinga [Tay21].

Marañón [Gli76]. Marc [Ste04]. Marcus [Han93]. Marga [Pau23]. Margaret [Bar19, KG21, Str10]. Marginalized [PH22]. Maria [Hal25, Cre19]. Marian [Moh18]. Marianne [Mar24]. Marie [Web20]. Marilyn [Bar19]. Marine [AD18, Luk21, Luk22, All90, Ben88, DC96, ZS15]. Martin [Kei23, Gre04]. Martinus [The96]. martyr [Hol93]. Marvel [Ank00, Gib19]. Marvelling [Ank00]. Marxism [Kle23]. Marxist [Bec05]. Mary [Bar24, ZS15]. Maryland [Kei98]. masculinity [Ano91a]. Mason [Mon18]. Mass [Yi08, Zal05]. Massachusetts [Rhe96]. Mastermind [JW13]. Mat [Wu19]. Materia [dC23]. Material [Chi09]. Materialism [AMW00, Yi21]. Materials [Pet08, AM68, Par71a]. Mateusz [Cau19]. Mathematical [Abr04, Kin86, Pal91]. Matlin [Liu23]. Matter [Gro13, Smo24, Tur18, Chu89a, Fry96, Slo86]. matters [Bur93a]. Matthew [Lau19, Wel73]. Maupertuis [Ano82l, San83]. Maura [Cao24]. Max [Kay85]. Maximizing [Gro17b]. Maxine [Kev05]. May [PH22]. Mayhew [Des05]. Maynard [Pie19]. Mayr [Gre92, Gre94, Jun96, Kle13, RN11, Slo85]. MBL [Mai23a]. McClintock [Com99, Kei99, Kle24]. McCollum [Lon09]. McCosh [Mor14]. McElroy [Lon09]. McGill [Muk19]. McLean [Raa76]. Me [cH17]. Mean [Hoq16, Mur23]. Meaning [Get91, dG02]. Means [HA19]. Measuring [Cha22, OM71]. Mechanical [Par21]. Mechanism [Bul99, Cra03, Hen06, Nic22, Pre13, AA82, Ano75a, Bro74, Hei72]. mechanism-vitalism [Hei72]. Mechanisms [BA07]. Mechanistic [AMW00, All04, Ano75n, RH84]. Mechanists [Bec16]. Media [Goo19]. Medica [dC23]. medical [Ano81g]. Medicine [Bez97, Fra79, Kev93, Mai09, MD16, Mon18, Shm23a, Shm23b, Bon90, Har89, Nat24, Ray19]. Medieval [Gaz12, Ano81g, Get91]. Medusa [Fre70]. Meet [Wil04]. Megatherium [Piñ88]. Meiosis [BF79]. Meli [Mes19]. Mellen [Muk14]. Membranes [Gro13, Tru97]. Memoires [Ste04]. Memoirs [Fla24]. Memories [Har23]. Memory [Cra03, Die11, MAR21, Pal22, Tur18]. Men [Hal10, Hal13, Wal01]. Mendel [Gli24, BF79, Di 91, Die04, OV68, PSBS^{+}22, San83, Sta95, Wei77, Wei68, Wol12, WO05]. Mendelian [All74b, Car74, De 74, SC14]. Mendelian-chromosome [All74b]. Mendelian-Mutationism [SC14]. Mendelism [Bon06, Ank00, Coc73, Dar77, Ser16, The08, Zir68]. Mendelsohn [Ano18, Ano19, Die16a, Die17b, RR20a, RR21a, RR22a, RR23, SR24, Smo25a, Ada23, Ank23, Bar23b, Bro23, Bur23, Mai23a, Sch23, Sel23, SR23a, Sta23]. Mendelson [Har23]. Mentor [Bar23b, Sel23, Ser16, Sta23]. Mentors [Sin97]. Merck [Ras99]. Merogones [Bra22a]. Merrill [Yan18]. Message [Bur16]. messenger [Gau96]. Metabolism [San16]. Metagons [Hal98]. Metamorphosis [Rob11]. Metaphor [LK25, Mat21, Gre94, Tay88, Yan18]. Metaphorical [Bra22b]. metaphors [Mit88]. Metaphysical [Har94]. metaphysics [CT90, Cor87]. Metazoa [Lus00, Ric89b]. Metazoan [GG97]. Metchnikoff [CT90]. Method [Har06a, HS18, Ile99, Rud18b, Rud18a, Sch09, Hal70, Wig69, Har22b]. Methodological [BB22, Gin10, Hoo97, Len05, Rob11, Ano75b]. Methodology [Arn19, Gaz12, Mon05, Zwi18]. Methods [Gro17b, Har08, Bor89, Kev93, Lus74]. Methylation [Tom22]. Metschnikoff [GG97]. Mexican [Mat06]. Mexico [BPA05, LMB03, Mat06]. Mexico/United [Mat06]. Meyer [Hon16]. Mice [Hal10, Led04, dC06, Rad22]. Michael [Bri09, Mon18, Sat20, Ric04, Web20]. Michaux [Por25]. Michel [Muk19, Tan20b]. Michelle [McG18]. Michigan [Ste05]. Michurinist [Sch12]. Microbes [Ack06, Rhe96]. Microbial [Kol22]. Microbiologist [Ack04]. Microbiology [Ack07, AD18, Pat09, Sta25]. Microbiome [Ras22]. Microcosm [Bre02]. Microorganisms [Sta25, The96]. Microphotography [Bre02]. Microscope [Ban22, Gro17a, vHS19]. Microscopic [Jac03]. microsomes [Rhe95]. Mid [Des01, KFP09, Luk15, dMP23, Elw07, Whe81]. Mid-18th-Century [KFP09]. Mid-19th [dMP23]. Mid-Century [Luk15]. mid-nineteenth-century [Whe81]. mid-Victorian [Elw07]. Middle [Thu02]. Midrashic [Kot81]. Miescher [VHL20]. Might [Esp20]. Migration [Ona23, Yi08]. Miguel [Ste24]. Milam [Som21]. Mile [Mit21]. Military [Mac01a]. Millennium [You25]. Millstein [You25]. Milon [Cau19]. Mimush [PSBS^{+}22]. Mind [Cec19, Hag03, Hal13, Ano82h, Smi82]. Minds [FG17]. Mineola [Irm05]. Minnesota [Koh22b]. Minor [Hoo97]. Minteer [Pow18]. Mirror [Bar24]. missing [Pad95]. Mitchell [Cro89, Pre01, Pre13, WP06]. Mitman [Ano00c]. Mitochondrial [BA07, Mat16, Ras95]. Mixed [Har12b]. Mobile [Kle24]. Mobility [Cha22]. Mobilizing [McG21]. Model [Arm16, Cre22, Gal21, Hop24, FS90, Fra91, Sin89, Ank20]. Modeling [Bur22, Wil12, vH19]. Models [Kei99, Sha72, Sla03, Kay85, Pal91]. Modern [Bez97, Bri09, GH19, Gro19, GM13, Hag24, Hon16, Hun19a, Hun19b, Joh19b, Mat21, Mel13, Mon18, O’M18, Par15, Pin02, Ray19, Som05, Wal20, Web20, vHS19, Fre70]. Modernizing [Sun13]. Modest Witness [Ham97b]. Moisture [KK13]. Molecular [Chi09, Die98, Gau93, Gau96, Gro13, Hag99, Hag10, Jia14, Kra24, Lew04, Liu23, LGB17, Loi25, Mor98, Pie12, San06, Som08, SD14, SD19b, SD19c, Tan20b, Uch93, Bur93b, Bur96, Car71, De 96, DG96, Die94, Eds80, Kay96, Kel90, LT93, Olb74, Sap92, Sch74a, Zal93b]. molecules [Kev93]. Monist [Erl02]. Monitoring [Sto15]. Monkey [Wil12, Pri24]. Monkeys [Den16, Hal13, Pri24, Sur22]. Monod [BH21a, BH21b, Bur93b]. Monstering [Che09]. Monsters [Art24, Mos98, Ray22]. Montagu [Far16]. Month [Bei08]. Montreal [Muk19]. Monument [Rum12]. Moon [Str23]. Moore [Nov18]. Moraba [Gro17a]. Morabia [Ras24a]. Moral [Jac03, Ric89a, Ham86, Str92]. Morals [Hal13]. Morange [Tan20b]. Moreau [San83]. Mores [Nov18, Thu02]. Morgan [Sug23, All68, Arn19, Bru02, Esp13, FG17, Led89, Mai16, Sun10]. Moritz [Wei10]. Moritzi [FE20]. Morphogenesis [Tam20a, vM71]. morphological [Ano81b]. morphologists [Ano81d]. Morphology [Ano81e, Bla07, Cam01, De 08, Dre20, Gra20, Mat16, RWE13, Rie20, Rud18b, San16, Ste02, Tam20b, Ano81f, Ben85, Rai89, Ste84, Wai88]. Morris [Hag10, Hal03, Hal10]. Morse [Bar19]. Mosaic [Lew04, Sch14]. Moscona [Pas21]. Moss [Bar05, Zal05]. Most [Loz19, Wal15]. Mother [Pau23]. Motion [Ras24b]. Motivation [Kly14]. motives [De 74]. Mountain [Vet12, RPM23]. Mouse [Rad98]. Moved [End23]. Movement [All13b, Sha08, Lud69, Pió79, Str92]. Movements [War23]. Moving [SL18]. Mr. [Bed73, Buc17, Ghi73, Kim06, Oga01]. MS2 [Pie12]. Muhammad [Gam21]. Muhammad-Rida [Gam21]. Muka [Jia24]. Mukharji [Bar23a]. Muller [Car71, Cro87, Moo95, dG02]. Multidisciplinary [Par08]. Multimodal [Com11]. Multum [Mul69]. Mungo [Fla24]. Munns [Ber19, Str23]. Muriel [Ric07]. Murine [Rad98]. Murphy [Koh24, McG18]. Muscles [War23]. Museum [GM18, Joh22, Sun13, GG93]. Museums [Bar24, Koh95, Rab11, Gru18]. Muslim [Gam21]. Must [Bec16]. Mutant [San10]. Mutation [Hoo97, All69, Lin92]. Mutationism [SC14]. Mutationists [Bea16]. Mutations [McG21, Pau87]. Mutualism [Hut22]. Mutually [Wei06]. My [RPM23]. Myrtle [Hea20]. Mysterious [Oga01]. Mystery [Hec21]. Mysticism [Rie12]. Myth [Bed83b, Har11c, Pau23, Wal15, Wig69, AA82, Ano75g, Hoq24]. Myths [dMP23, You09].

N [Ile05, Ona23]. N.J [Cro05]. N.Y [Irm05, Kev05, Sel05]. Nadine [Stu24]. Naef [RWE13]. Nagasaki [Lin92]. Nair [Ric25]. Nakayama [Joh19b]. Naked [CC18]. name [Rot89]. Naming [McL02, Ber70]. Nancy [Ben18, Kei11b, Wil04]. Naples [Ben88]. Narrative [Fol21, NL11]. Narratives [Bro24]. NASA [Cro05]. Nascent [War23]. Nathan [Rad23]. Nation [Sul11]. Nation-State [Sul11]. National [Boc95, Car08, Rum12, Sta25, Ank18, GM18, Mac93]. Nationalism [Sul11, pY14]. Natural [Ala12, Alb01, Bar00, Bar24, Bea16, Bea19, Bla07, Bla82, Bur99, CM11, Cor05, Eng97, Fag07, FC21, Fla24, For12, Gil87, Gro17a, Gru18, Hea20, Hei19, Joh04, Joh19a, Kly14, Koh24, Lla00, Mir08, Por79, Sch09, Sch10, Sch90, Sch99b, Sep13, Smi15, Sun13, The12, VWP20a, VWP20b, Whi20, Wit16, vP21, vdBD20, All68, Ano82d, Ano82n, Bed68, Bow94, Col86b, Cor87, Eva84, Gil90, KO91, Koh95, Lar86, McE92, Osp80, RM84, Ric81, SP81, Slo72, Wel73]. naturalism [Ric89a]. Naturalist [Bar97, Erl02, dG02, DCS01, Rud90]. naturalista [Ano75l]. Naturalistic [Har12a]. Naturalists [Bar09, Eng97, Hag99, vP21, Hag84]. Naturalization [Hal03]. naturalizing [Reh79]. Naturally [Stu23]. naturalness [Pet08]. Nature [Ano00c, Bez97, Bra24, Cau19, Coo98, Fic24, Gli07, Hal03, Pea10, Rey08, Rie12, San10, Som21, Str92, Stu24, Vic18, Vor92, Bla82, Cor84, Hod72, Out86, Sht92, Ano05a, Ano06b, Bar24, Can13, Tyr23, dC23]. Naturelle [LB04, Ano76c, Lov06]. Naturphilosophie} [Cog22, Han93, Ano78d]. Navon [McG21]. Nazi [Wei11, Har89, Har12b]. Nazis [BS08]. NC [Som24]. Neal [Gib24]. Neanderthal [Mad20]. Near [Ano75m]. Necessary [Sla03]. Necessity [Har14]. needs [OM71]. Neel [GS15]. Neeraja [Boo23]. neglect [Gla74]. Negotiating [Die98, Ham07]. Negotiation [Jac03]. Negotiations [Som08]. Neil [Web20]. Nelson [Ank20]. Neo [Gis23, Loi11, Meu21a, Meu21b, Rin88]. Neo-Darwinian [Meu21a, Meu21b]. Neo-Lamarckian [Gis23]. Neovitalism [Rus15]. Neptunism [Jen16]. Nerve [Tru97, Bil71]. nervous [Hom70]. Netherlands [Mir08, The08]. Network [Civ13, Gau96, Ben18]. Networks [Jon17, Man15, Kay96]. Neuroembryology [All04]. Neurophysiology [Gru22]. neutral [Die94]. Nevada [SP25]. Newport [Cog22]. Newspeak [Str20]. Newton [Hic21]. Newtonian [Gil87, Gue85]. Next [Die17c]. Nice [Bar19]. Nickelsen [Str23]. Nicolas [Abr04]. Nicole [Ank20]. Niels [Söd94]. Night [Cob25]. Night-Time [Cob25]. Niko [Bur16, Vic18]. Nineteenth [Bro24, Jen16, Mir08, VWP20a, WO05, Ano76f, Ano82d, Ano91f, App80, Ben85, Bor89, Cam83, Cor83, DCS01, Gai73, Ger70, Gro79, Kin80, Pau84, Ste90, Whe81, BJ97, For20, LD19, Qui17, Sch99b, VWP20b, Wit16, vP21]. Nineteenth-Century [Jen16, App80, Bor89, Cam83, DCS01, Ger70, Kin80, Pau84, Ste90, BJ97, LD19, vP21]. no [Fla24, Zal05]. Nobel [Kle24, All02]. Noel [Kle11]. Noisy [Tho05]. Nomad [Ric25]. Nomenclatural [Wit16]. Nomenclature [Mir11]. nominalist [Sta96]. Non [Sul11]. Nonhuman [Dru23]. Nonpolitical [dG02]. Norfolk [CM11]. Normal [Gau06]. Normandin [Mét14]. North [Dav97a, Koh24, Hen97, Lue21, Whi20, Yi21]. Norway [Log01]. Norwegian [Sto15]. Note [Ano03c, Ano04b, Ano82e, Ano82f, Ano82g, Ano05b, Gin10, Har94, Led89]. notebooks [Vor77]. Notes [AM68, Col85, Men98, Par71a, Ano93, Bed72, Bed73]. notice [Ano86, Ano87a, Ano87b]. Notions [Str24]. Novel [Gia18, Moh18]. November [McG19]. Nuclear [Coo20, Cre06, Cro14, Lue21]. Nuclein [VHL20]. Nucleotide [Ste18]. nudae [CC18]. Numbers [You09]. Numerical [Vin22, Par82]. Nurse [Li04]. Nurture [Coo98]. nutrition [Ano75k, Bis23]. Nuttall [Ste05]. Nyhart [Str19].

O [Gib13, Rus15, Sch14, Str10, Bae19a, Bae19b]. Oak [Boc95, Rad06]. Oakland [Bis23, Hub20]. Object [Liu17, CKT94]. Objective [Ham07]. Objectivity [Bre02]. Objects [Ray22]. Observatory [dB11]. Observe [AdC24]. Occupations [Eli14]. Ocean [Kei17, Mac01a, Ano05a]. Oceanography [Rai03]. Oceans [Jia24]. Ochoterena [LMB03]. Odds [Mat06]. Odum [Tay88]. Oedipus [Ste76]. Offspring [Hoo97, Lin92]. Ogilvie [Bar19]. Oka [Sul11]. Old [Bae19b, Lew87, Ghi73, Hol93, Por85, Bae19a, CC18]. Oncogene [Mor97a]. One [Alb01, GM18, Har15, Mon18, Pre90, Sch23, Ege83b, Wal15]. Ongoing [NL11]. Onslow [Ric07]. Ontology [San16]. Open [Ano20, Mar24]. opium [Ano75m]. opponents [RH84]. opportunities [Ege83b, Ege85]. Opposition [All74b, Sta24]. optical [Ano75a]. optimism [Tay88]. opting [Cai02]. Oracle [Can13]. Orangutan [Shm20]. Order [Kim18, Gil87]. Ordinary [End11, PH22]. Oren [Sat20]. Organ [Ben09, Gut97]. Organic [Bec05, Bra20, PSBS^{+}22, Rin81]. Organicism [Ste07]. Organicist [Gib13]. Organism [Bae19a, Bae19b, BP19, Nat24, Pla69, Rey08, Bur93a, Sum91, Zal93a]. Organismal [Hag15, Som08]. Organismic [Par71b]. Organisms [All22, BL24, Can15, Cre22, Bol22b, Muk19]. Organization [Lam15, Liu24, Ric06, Wol11]. Organizational [Gro69]. organizer [Ham84, Bra22b]. Organizers [Bra22a]. Organizing [Smo94]. Organotherapy [Bor76, Bor85]. organs [Mun93, vM71]. orientation [Rhe93]. Origin [Atr87, Biz92, Bra11a, Die04, Eng97, For20, Gal21, Hal25, Jac01, Jax98, Kly14, KK13, Kon23, Mar11, Por93, Sha08, Str99, Str00, Tab08, ddAM20, dd22, vW14, Ano78d, Ano82m, Far72b, Gia93, Gou68, Kin80, Pod96, Sch77, Smi78, Söd94, Baj88, Bed88b, Bow89, Cog02, Coo90, Far74, Joh07, Kle13, Oga01, Sch77, Tan17, ddAM20]. origin-of-life [Pod96]. original [Ano75l]. Originality [WP06]. Origins [Ala12, Alt07, Arn19, But18, Hag10, Kei01, Liu23, Lov02, Mad20, Mét16, Mit17, Nov16, Pir24, Pre01, Rud18b, Ste18, Wei10, pY14, Yi25, Die94, Gai80, Gil78, LT93, Lyo95, Mon70, Olb74, Ste76, Sum91, Goo19]. Orin [Mar04]. Ornithologist [Bar19]. Ornithology [Dav97b, dB11, Dav97a]. Osiris [Kro97]. Osseo [Coo20]. Osseo-Asare [Coo20]. osteogenesis [Del83]. Other [Cob25, Gam21, Har13, HA19, LT93, Vet06, You09, vP21, Jen16]. Otto [Koh73]. Ought [Har12a]. Our [SS17, You25, Fol21, Pau87]. Ours [KC94]. Ourselves [Moh18]. outline [Ano75e]. Outsider [Muk14]. OVATOOMB [LT93]. Overcoming [Eig97]. overdominance [Cro87]. Overfishing [FC21]. Overheated [Log07]. Overlooked [Bil16]. overtones [Söd94]. overview [Cor89]. Owen [Ano76d, Cam01, Di 95, Joh19a, Pad95]. Own [GM18, Kle99, Win01a]. Owners [Dan21]. Oxford [Ben09, Can13, Irm05, Wai88]. Oxidative [Mat16, WP06]. Oyster [Kei98].

P [Ber19, Pow18, WP06]. P. [Bur93b, Ste84]. P.D [Str23]. Pacific [Mac01a]. Packard [Boc88]. Pact [All04]. Palaeontological [Ano76d]. Palaetiology [Elw07]. Palais [LM14]. Paleobiological [Bri09]. Paleobiology [Dre23, Rud18b, Rud18a, Sep05]. Paleontology [All20, Man15, Rab11, Rai85, Rie12, Sep13, Sep19, Tam15, Ano81b, Ger70, Rei86, Bri09]. Palgrave [Cre19, Far05, Mon18, Ste24]. Palladino [Hen97]. Pangenesis [Hol14, Sta03]. Panpsychic [Ste07]. panpsychism [Smi78]. Paolo [Hen97]. Paper [Kin20, Mat21, Pow18, Sat20, SD19a, Ano82d, Ben88, Irm05, Tob09, Wei77, dC14, Ano04a, Ste04, Zal05]. Paperback [Ank20, Coo20, Gia18, Hei19, Hub20, Moh18, Shi18, Yan18, Goo19]. Papers [Cor21, Eng97, SK22, Wol11, vP20, AM68, Ano78e, Ano82i, BA81, FG78, Par71a]. Paradigm [Mor97a, Ray22, Rud18b, Rud18a, Tam20a]. Paradox [Cre10a, Die98]. Paraguay [Bed83b]. parallel [Sch89]. Parallels [Boo23, KK13]. Paramecia [Hal98]. Parasite [And16, Den16, MD16]. Parasites [Li04]. parasitic [Far72b]. Parasitism [And16]. Parasitological [Li04]. Pardi [Can15]. Paris [Hoq16, LM14]. Park [Ank18, Fla24, Goo19, Jin25, Pau23]. Parks [Car08]. Part [Mor14, Ege83b, Ano82h, Bea16, Bea19, Ege85, Her77, MKC69, MN69, Smi82]. partial [Tay88]. Particular [Kre07].

Particulate [Fal03]. Parts [NL11]. Party [Cas12]. Parvo} [Mul69]. Passing [Ano05e]. Passports [Bro88]. Past [SL18]. Pasteurian [AD18, LGB17, Loi25]. Pasts [Rad25]. Path [Loz19]. Pathological [Mes19]. pathology [Whe81]. Pathway [Pas21]. Pathways [Gau06, Koh04]. Patient [McG21]. Patrick [Ste04, Hea20, Li04, Por25, Wel73, Wu19]. Patron [Ste97]. Patronage [Kra04]. Pattern [Vin22, Rho87]. Patterns [Den16, Kei99, Koh04, Gau93]. Paul [All02, CT97, Dav97b, Gli06, Gue21, Kev05, Log07, MAR21, Pre13, Rab11, Rhe96, Yi08]. Pauling [Mor98]. Pavesi [Pir24]. Pavlov [HA19]. pbk [McG18]. Peace [San06]. Pearl [Eri17]. Peas [PSBS^{+}22]. Pedagogy [Lig22]. Peking [pY14]. Pelagic [Pir24]. Pemberton [Web20]. Penguin [Pie19]. Penicillin [Cre19]. Pentosuria [KG21]. People [Rad98, Rhi25, Sch12]. Perception [Sch04, Rue83]. Pere [Art24]. Period [Bra22a, Bra22b, Cai09, Mét16, Ams91, Par21]. Periodical [Gök94]. peripherality [Sch74a]. peripheries [But88]. Perishable [Sac09]. Permit [Ben11b]. Persian [CB82]. persistence [Lyo93]. Personal [JK99]. Perspective [CS06, Gaz12, Atr87]. Perspectives [Hag17, Hut22, Lin18, PBDL17, Tau97, Sap83, Bec16, Str19]. Persuasion [Die98]. Pet [Shm20]. Peter [Cro89, Pre13, WP06]. Peyton [BK10, Ros18]. Phage [Cre10a, Pie12, Sum93]. Phenograms [Kra24]. Phenology [Mac21a, Mac21b]. Phenomena [SK22]. Phenomenology [Mor23]. Phil [Kin09a]. Philadelphia [Hea20, Por79]. Philosophical [Pre01, Rob11, Shi18, Str19, Tau97, App88, Ghi05]. Philosophy [All04, Ano84a, Gay98, Hal25, Mar09, Mét14, NG14, Nyh18, Ray19, Rob11, Rud00, SG18, Ste07, Str24, vdBD20, Ano75b, Bur90, Gre94, Hul69, Pre70, Rai85, Rie88]. Phoenix [Joh09]. Phosphorylation [Mat16, WP06]. photos [Muk19]. Photosynthesis [Nic22, Zal93a, Zal93b]. phyllotaxis [Mon70]. Phylogenetic [Kra24, SD14, dV09]. Phylogeny [Rie20, vP21]. Physical [Hag17, Mar24]. Physicians [LB04]. physicist [Kay85]. Physicochemical [Nic22]. Physics [Kel90]. Physiological [D’H12, Hag24, Rhe00, Ric00, All74b, Hal70]. Physiologists [WH02, Fre70]. Physiology [Arm16, Hag17, Joh15, Lan20, Log02, Nyh18, Par21, Sch04, Bor76, Bro74, Bro71, Bro73, But88, GT69, Gro79, Gue85, Hol93, Oak15]. Phytogeographical [CM11, Hun19c]. Phytophthora [Mat10]. Phytotrons [Ber19]. Pickford [Sla03]. Pictorial [Kei99]. Pictures [Rhi25, Sch04]. Pierre [San83]. Pietro [Pir24]. Pietsch [Man13]. Pig [Ala04]. Pigeons [The12]. Pioneer [Bil16]. Pittsburg [Ray19]. Pittsburgh [Ber19, Boo23, Bra20, Bro24, Nov23, Rad23, Ray19, Str23, Tyr23, Wu19]. Place [Boc12, Bur99, Die98, Ile05, Koh12, MM17, Mon05, Ano78c, Cro90, Her74, Her77, Ano04a]. Placing [Hob15]. plague [Get91]. plan [Lar86]. Planet [McG19]. Plans [Gur02]. Plant [Bon06, Che09, Hag24, cH17, Kim18, Kle11, LD19, Ona23, Ras99, RPM23, Sch09, WH02, Wie06, Col86a, Hag86, Whe81, vM71]. Plants [End23, Mac21a, Mac21b, Rhi25, Rob11, Str24, vM71, Cao24, cH17]. Plasm [Win01b]. plates [Gru18, Mes19]. Platypus [Hob15]. Playback [Rad05]. Plying [Kli98]. Plymouth [Erl09]. Poetry [dB11]. Point [Com99, DC96]. Poison [Gue23, Hob15]. Poland [dL12]. Policies [CS06, JACD18, Pat09]. Policy [Man09]. Polio [Sur22]. Polish [Köh11, Str20]. Polistes [Can15]. Political [Dan21, Tay98, Gor89, Kel90, Sch80b]. Politics [Ank04, CdC25, Ell07, Hal03, Iid10, Jum02, Kei98, Kei17, Kra04, Par00, Sla08, Tay21, Wei11, Wei06, WY23, Ano91f, Bur70, McE92, Moo91, Smo91]. Poll [BS08]. Pollen [Cha22]. Polygenist [Hum08]. Polymorphism [Lew87]. Polymorphismus [NL11]. Pond [DC96]. poppy [Ano75m]. Popular [Hec21, LK25, Pie19, Stu24, Atk85]. Popularization [Sch99b]. Population [Ada70, Gro17b, Gro19, Jon17, Kin86, Lam15, PKS18, Ras24a, Veu19, vH19, Ada68, Ege70, Mit88]. populations [Ege68b]. Porte [LM14]. Portion [Ham99]. Portrait [Muk14]. Portraying [Sap92]. Portuguese [DCS01]. Position [Hal70]. Positive [For12]. Post [Ala24, Bel06, Gli98, Lew18, Mét14, Plu19, Rob11]. Post-1848 [Gli98]. Post-colonial [Lew18]. Post-Darwinian [Ala24]. Post-Enlightenment [Mét14]. Postcolonial [And16]. Postgenomics [Lin18]. Postwar [Gau06, Rad06, Bea88]. Potato [Mat10, Whe81]. Power [Ano04a, Coo20, Sch15]. pp [Ank20, Ank18, Ano04a, Ano05a, Ano06b, Ban22, Bar05, Bel05, Ben18, Ben09, Ber19, Bra20, Bri09, Bru18, Can13, Coo20, Cre19, Cro05, Des05, Far05, Gia18, Goo19, Gru18, Hei19, HS18, Hub20, Ile05, Irm05, Kei11b, Kev05, Kim18, Kin20, Kle11, Lan20, Lig18, Lin18, Loz19, Mar09, Mat21, McG21, McG18, Mes19, Mét14, Mil11, Moh18, Mon18, Muk19, Nyh18, Por09, Pow18, Rab11, Ras24b, Ray19, Sat20, Sch18, Sel05, Shi18, Som21, Ste05, Ste04, Tan20b, Tob09, Wal20, Web20, Wei11, Wu19, Yan18, You09, Zal05, dC14, Boo23, End23, Kim23b, Liu23, Ona23, Tyr23]. pp. [Bar23a, Bar24, Bis23, Bod23, Bro24, Cao24, Fic24, Fla24, Fre25, Gib24, Gli24, Gre25, Gue23, Gue24, Hag24, Hal25, Jia24, Jin25, Kei23, Kle24, Koh24, Lin23, May23, Mor23, Nat24, Nov23, Pau23, Por25, Rad23, Ras24a, Ric25, Shm23a, Shm23b, Som24, Ste24, Str23, Stu24, Sug23, Wei09, You25]. Practice [Boc12, Fag07, Iid10, Koh12, Mon05, The08, Vet06, Wie06, You98, Bur93a, Ras95]. Practices [CS06, ML10, Sep13, Löw94].

Practitioner [Win01a]. Pragmatics [JACD18]. Pragmatism [Kra04]. Pragmatist [Rus15]. Pratt [Lon09]. Pray [AdC24]. praying [PW92]. Pre [Arc09, FE20, KK13, Bow71, Des84, Wel71]. Pre-Darwinian [FE20, Des84, Wel71]. pre-existence [Bow71]. Precisely [Coe06]. Precision [dB11]. precursor [San83]. precursors [Ric89b]. Predicament [Vic18, Des84]. Preetha [Ric25]. preexistence [Ano82l]. Preface [Joh07, MCB86]. Preformation [Bow71]. preformationist [Ano76b]. Pregnancy [Fol21]. prematurity [Gla74]. Preparations [Ray22, Sch04]. Prepares [Kle13]. Present [Bel06, Bro24]. presentations [Tay21]. Preservation [Hic21, Rum12, Tyr23, Sht92]. Preserving [Cao24]. Presidential [WY23]. Press [Ank20, Ank18, Ano04a, Ano05a, Ano06b, Ban22, Bar23a, Bar05, Bar24, Bel05, Ben18, Ben09, Ber19, Bis23, Bod23, Boo23, Bra20, Bri09, Bro24, Bru18, Can13, Cao24, Cau19, Coo20, Cro05, Des05, End23, Fic24, Fre25, Gib24, Gli24, Gök94, Gor19, Gre25, Gru18, Gue23, Hag24, Hei19, HS18, Hub20, Ile05, Irm05, Jia24, Kei11b, Kei23, Kev05, Kim23b, Kim18, Kin20, Kle11, Kle24, Koh24, Lan20, Lig18, Lin23, Lin18, Liu23, Man13, Mar09, Mat21, May23, McG21, McG18, Mes19, Mil11, Mor23, Muk19, Nat24, Nov23, Nyh18, Ona23, Pau23, Por09, Por25, Pow18, Rab11, Rad23, Ras24a, Ray19, Sac09, Sat20, Sch18, Sel05, Shi18, Shm23a, Shm23b, Som21, Som24, Ste05, Str23, Stu24, SD19a, Sug23, Tan20b, Tob09, Tyr23]. Press [Wal20, Web20, Wei09, Wei11, Wu19, Yan18, You25, Zal05, dC14]. Pressure [Riz20]. Prestes [Hal25]. Presti [Ray19]. Price [Har11a]. Primary [Sel23]. Primate [FS90, Mon05, Rad05]. Primates [Dru23]. Primatology [Mon05]. Princeton [Ano04a, Ano05a, Fic24, Gre25, Som21, AM68]. principle [Ano82n, Bro80, May92, Tam95].

Principles [JACD18, RWE13, Hal70]. Pringsheim [Kol22]. Prior [Lov02]. Priority [Joh19a, WP06]. Privacy [Smo99b]. private [Kel88]. Prize [All02, Ano18, Ano19, Die16a, Die17b, RR20a, RR21a, RR22a, RR23, SR24, Smo25a]. Problem [Bur99, Che09, Dre20, FC21, Gro17a, Log07, Lue21, Pre13, Sch99a, vW14, All68, Ano82h, Bow73, Cor71, Far72c, Fic77, Han93, KO91, Olb71, Rin81, Slo72, Smi82]. Problems [Ano81f, Cai99, Gai73, Hal98, Ile99, Smo99b, Tan17, Smi92]. Process [ÁTBHBJB24, Ben11b, Vin22, Bed73, Fin91]. Processes [Sch04]. Production [Log02, Muk19, Pal22, Riz20, Sch04]. Productivism [Yi21]. Productivity [Riz20, Ege69]. Professional [Bel01, Gal22]. Professionalization [Gor09, Wal01]. Professionals [Alb01, Des01]. Professions [Shm23a, Shm23b]. Professor [Ank23]. Profit [Fit19]. Program [Cre06, Gur02, Hea20, Ano78a, OV68, Mac01a, Sch17]. programme [Jac84]. programs [Zal92]. Progress [Cai99, Des01, WdLdS21a, WdLdS21b, Ano75i, Bar95, Gas91]. Progressive [Mun23, Pat09, Ano91h, Wei96]. progressivism [Wei96]. Project [Hat21, ML10, Mai23b, Ras99, Vin22, AdC24, GD18, JACD18, Lig22]. Projects [Ste24]. Projit [Bar23a]. Promise [Ras99, Dar77]. Promotion [GZ98, Bor87]. Promptov [Kre07]. Propaganda [San06]. properties [Fry96]. propre [Swa70a]. Prosper [NSRG09]. Protean [Rey08]. protection [Str92, Vor92]. protective [Kin78]. Protein [Rhe96, Str10, Bur93b, Rhe93]. Proteins [GS10, Tru97, Ano84a]. protista [Rot89]. Proto [Lus00]. Proto- [Lus00]. protoctista [Rot89]. Protoplasm [Liu17]. Protozoa [Chu10, Ric89b, Rot89, Sch99a]. Protozoology [Jac89, Chu89b]. protozoon [Cor89]. Provirus [Fis10]. Proximity [DC96]. Proxy [Cha22]. Pseudo [Ano81g]. Pseudo-Dioscorides [Ano81g]. Pseudoscience [Gor12]. Psychiatric [Ank20]. psychoanalysis [Ste76]. Psychobiology [Wei96]. Psychology [Arn19, Tho05, Bro71]. Public [Dru23, Eng97, Mai01, Pat09, Ras24a, Kel88]. Publication [AdC24, Pre90]. Publications [Gin10, Por79]. Publisher [Bae19b, Ano86, Ano87a, Ano87b, Ano88c, Ano88a, Ano88b, Ano89b]. Publishers [Mar24]. Pulse [Boy07]. Pump [Gro13]. Pure [Har08, Kay89]. Puritanism [Har08]. Purple [Gro13]. purpose [Maz75]. Putnam [Moh18, Dex70]. Putting [Ile05]. Puzzle [KG21, Wil16]. Pyramus [Ste04].

Quality [Hub20]. Quantifying [Hum08]. Quantitative [Hag03, PBDL17, Sep05]. Quarrelsome [Mur23]. Queen [Muk19]. Quest [Ber19, Cog22]. Question [Ano78g, Hol17, Joh19a, Kei98, Lew05, Hai91]. Questions [Bea88, Gaz12, Vor69]. Quiet [Gin10]. Quinarian [Nov16]. Quinarianism [Cog02].

R. [Fun12, Hal07, Moo95, Tab08]. Rabid [Bro11]. Rabies [Bro11, Dan21]. Race [Alt07, Bar23a, Far16, Goo19, Har89, Log07, Mar24, Ser21, Ste76, Wei09, dR23]. Races [Art17, Jac01, Wal15]. Rachel [Mon18]. Racial [JW13, Sla08, Som05, Wal15, Wei05]. Racism [Art17, Far05, Ona23]. Rader [Led04]. Radiation [Fol21, GS15, Ham07, HR21, Hec21]. Radical [Ank04, Fic24]. Radick [Gli24]. Radiobiology [CS06]. Radioisotope [Cre06]. Radioisotopes [San06]. Raf [Tyr23]. Rainfall [HR11]. Ramón [Sum04]. Rampage [Ros18]. randomization [Hal07]. Range [You98]. Rankin [Gue23]. Rashevsky [Abr04]. rat [Cla93]. rate [Rho87]. rational [Deb70]. Rationale [Hoo97]. Rationalism [Roe79]. Rationalist [Cos97]. rationalization [Wei87, Des05]. Rationalizing [Thi01]. Raton [Kle24]. Rats [Hag15, Log01, Log07]. Ravenel [Hay88]. Ray [Mon05, Gil87, Slo72]. Razvitie [Ano69b]. Re [All22, Dre20, Tay21, San73, Log01]. Re-envisioning [All22]. re-examination [San73]. Re-forming [Dre20]. Reaction [Har03, Bur72b]. Reactions [Eng97, Sin97, Sta03, Far72c, Far74]. Reactor [dC06]. Reader [Men98]. Readers [Mac01b]. Reading [RR22b, Rus07, SP81, Vor77]. Real [Com99]. reality [Bed83b, Jac83]. really [Sta96]. Reap [Lus01]. Reappraisal [Nov19, Ano76d, Ano78c]. Reason [Bra11a, Des05]. Reasonable [Fis10]. Reasoning [Elw07, Nov23]. reasons [Ano76b]. Rebecca [Mil11]. Recapitulation [Bra24, Gur02, WdLdS21a, WdLdS21b]. recapitulationism [Ano91c]. Receiving [dV09]. Reception [BK10, Com99, Gli98, Hol14, Jac01, JF03, Rob11, All69, App88, Gök94, Jac84, Koh72, Rie88]. reciprocal [Hai91]. Recombinant [Sch25, Yi08]. Reconciling [Ano06b]. Reconsidered [Smi25, Gor89]. Reconsidering [Mad20, Win24]. reconstruction [Ano82m]. record [Ano82m]. Recording [Whi19]. Recremental [Jan24]. Red [Eri17, Man09]. Redefining [Des01, Hea20, Gau93]. rediscovery [Zir68]. Redrawing [Zal93b]. reductionism [Sch74a]. reductionist [Ano78a]. Redux [Col98, Fra72, RR21b]. Reefs [Mir22, Mir22]. Reel [Ano00c]. Reevaluating [Ano91h]. reevaluation [Bow77]. Refinement [LaB17]. Reflections [All16, Ano75i, Bro22, Far17, Fut86, Gre92, Irm05, Men17, Pau16, Rad22, dG02, Bur93a, GS92]. reflex [PW92]. Reflexes [HA19]. Reform [Bon90, Gli98, Par21, Tes23, Por79]. Reformation [Buh00]. Refrigerator [Pau23]. Regeneration [Esp13, MM24, dMP23, Sun10]. Regenerative [Mai09]. Region [HR11, Boc90]. Registers [BP19]. Regulatory [Mor97a, Ano84a]. Rehabilitate [Dre20]. Reich [Wei11, Wei06]. Reichenbach [Rob11]. Reighard [Boc90]. Reil [Han93]. reinterpretation [Ano82c, Di 91]. Reinventing [Die11]. Rejection [JF03]. Relatedness [Mar24]. Relation [Lev18]. Relations [Soo21, WdLdS21a, WdLdS21b]. Relationship [Gur02, Hal03, Pir23]. Relationships [Qui17]. Relevance [Tan17]. Relic [AMW00]. Religion [Jin25, Mat87, You09]. Religious [PW92]. remarks [AA82, Sul83]. remedies [Sca78]. Remedy [May23]. Remembering [Bar23b, Bur23, Kin09a, Ric23, SS17]. Remembrances [Sch23]. Remind [cH17]. reminiscence [Ano78b]. Rena [Lin23]. Renaissance [Gue23, Bon90]. Renato [Kev93]. René [Ste04, Ras22]. Reordering [Sch18]. Repertoire [For12]. reply [Bow89]. Report [Des01, Ano78e, Ano82i, BA81, CB82, FG78, Fra79]. Representation [Bre02, Qui17, Tru97, Rhe95, War23]. Representations [Cec24, LaB17]. Representatives [Som05]. Reproducibility [Hal98]. Reproduction [Edd23, Bod83, Pau96]. reproductive [Bor85]. Republic [Sch12, Tho05]. Reputation [BK10, Die11, RPM23]. Research [Ank18, Boo23, But18, CS06, Elw07, Gau06, Gro13, Gur02, Hop24, Led89, Lew04, Lew18, Li04, Lov02, ML10, Par15, Pas21, Rai03, RF18, Sla03, Sur22, Yi08, dC18, Ano78a, Boc95, Bor87, Mac93, Zal92, Zal93a, Led04]. Researcher [BS08]. Reservation [Ala12]. Reserve [Ala12]. Reshaping [Par15]. Residence [Ano88d]. Resistance [Wol13, Loz19]. resistant [Sin89]. Resolution [Gor97, Ser21]. Resource [Kei98]. Resources [Wei06, McE92].

Respondents [Mac01b]. Response

[Far09b, HR11, Lep15, May97, Sel12, vdV14, Ano81e, Pau83]. Responses [Rus15]. Restoration [Kei23]. Restroom [Rus07]. restructuring [Pau93]. Resurgence [Bro11]. Rethinking [Cai09, NG14, Smo05b]. Retrospective [Ano20]. Return [Joh09, Joh22, Ser21]. Revenge [Wol01]. Review [Ack04, Ala13, All09, All12, All13a, All22, Ank20, Ank18, Ban22, Bar23a, Bar04a, Bar10a, Bar11b, Bar19, Bar24, Bar10b, Bel11, Ben88, Ben10a, Ben18, Ben09, Ben10b, Ben13, Ber19, Bis23, Bod23, Boo23, Bor11, Bow18, Bra20, Bre15, Bri09, Bri17, Bro17, Bro24, Bru18, Bur14, Bur17, BR20, Cab17, Cai99, Can13, Cao24, Car09, Cau19, Che14, Chi18, Com12, Coo20, Cre03, Cre10a, Cre19, Cro12, Cro15, DM97, Dav97a, Dav12, Dav15, Die04, Die10, Dre17, Ell07, Ell09, Ell11, End23, Far11b, Fit19, Fla24, Fra10, Fre25, Gau98, Gia18, Gib18, Gib19, Gli04, Gli24, Goo17, Goo19, Gor19, Gre04, Gre25, Gro09, Gru18, Gue23, Hag24, Hal09, Hal12a, Hal12b, Hal17, Hal25, Ham97b, Har13, Hay17, Hei19, Hen10, HS18, Her10, Hub20, Ive10, Jac09, Jac10]. Review [Jia24, Jin25, Joh19b, Kei11a, Kei11b, Kei23, Kim18, Kin20, Kin09c, Kir13, Kle04, Kle10, Kle11, Koh04, Koh24, Kro97, Lan20, Lau19, Led04, Lig18, Lin23, Lin18, Lin15, Liu23, LD12, Loz19, Luk18, Lyn09, Lyn10, Mac12, Mac01b, Mad17, Mai17, Man13, Mar04, Mar09, Mat21, May23, McG21, McG18, Mes19, Mét14, Mil11, Mil12, Mil14, Moh18, Mon11, Mon18, MS18, Mor12, Mor23, Muk17, Muk19, Nat24, Nic11, Nov23, Nyh18, Ona10a, Ona23, Pau23, Por04, Por09, Por25, Pow18, Pri24, Rab11, Rad04, Rad07, Rad12, Rad23, Rai01, Ras24a, Ras24b, Ray19, Ren12, Ric09, Ric10, Ric11, Ric25, Roo17, Rud00, Rud09, Rum10, Rum13, Rus07, Rus12, Sac09, San13, Sat20, Sch18, Sel03, Sep11, Sep15, Sha11, Shi18, Shm23a, Shm23b, Slo07, Smo99a]. Review [Smo05a, SS99, Som21, Som24, Soo21, SW12, Ste19, Ste04, Ste14, Ste15, Ste24, Str00, Str19, Str23, Stu24, SD19a, Sug23, Sum04, Tam17, Tan20a, Tan20b, Tob09, Tyr23, Ule13, Ule14, Van10, Wad11, Wal20, Web20, Wei09, Wei11, Wei05, Whi19, Wil04, Woi10, Wol01, Wu19, Yan13, Yan18, You09, You10a, You10b, You25, YD00, Yud10, Zal05, dC14, AA82, All84, All91, Ano68, Ano84a, Ano91f, Bed83a, Bon84, Bro82, Bro88, Bur90, Eds80, Ege83a, FS90, Fin91, Har89, Har90, Kin94, Koh95, Len87, Lus01, Nyh95, Pau93, Por85, Por95, Sap92, Sch80a, Sch88, Slo85, Smo91]. Reviews [Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano01a, Ano01b, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano04a, Ano05a, Ano06a, Ano06b, Bar04b, Bar05, Bel05, Cro05, Des05, Far05, Ile05, Irm05, JRR18, Kev05, Lar08a, Lar06, Lar07c, Lar07a, Lar07b, Lar08b, Lar08c, Lar09, Sel05, Ste05, RRM^{+}05]. Revis [CC25]. Revisited [Esp13, Fis10, Har14, Har22a, Her05, Hug11, Koh22a, Meu21a, Meu21b, Tei07, The12, May72, Pau87, Sch77, The96]. Revisiting [Bow05, Die16b, Gin10, Hun19b, Koh95, Lep15, Ser21]. Revival [Sho13]. Reviving [Bil16, Hop24]. Revolt [Ano81d]. Revolution [Bri09, Bro24, Gli98, Gli04, Hag24, Her05, Hod05, cH17, Rus05, Sep05, Smo05b, Bea82, Ghi05, Por79, Rad23, Sch18, Tan20b]. Revolutionaries [Sat20]. Revolutionary [Fic24, Get91]. Rewriting [Mar23]. Reynolds [Mat21]. Rheinberger [Mor23]. Rhesus [Sur22]. Rhode [DC96]. Rhodesia [Dan21]. Rhodopsin [Gro13]. Rhythm [Ste19]. ribosomes [Rhe95]. Richard [Ack04, Ano76d, All74b, Bil16, Cam01, Di 95, Die95, Die11, Ege69, Far05, Joh19a, Pad95]. Richards [Lig18, Mun23, Rus04]. Richardson [Lin18]. Rickets [Fre25, Sin89]. Rida [Gam21]. Riddle [Ha11]. Ridge [Boc95, Rad06]. Ridiculously [Dre23]. right [Cla93, Bol22b]. Riparian [HR11]. Rise [Gro17b, Gur02, Hag24, Kei23, Kra04, Ona23, Gri74, Gri85, Sap83]. Risks [Hec21, Bea87]. Rivalry [CT97]. Riverhead [Gia18]. Rivers [HR11, Ste05]. RNA [Gau96, Gia93, Pie12]. Road [Smi15, SD14, Tay98, Por93]. Rob [Shm23a, Shm23b]. Robert [Gre25, Kro97, Ano91e, Des05, Des84, Fla24, Gor19, Ham86, Jen16, Mit03, Ric89a, Rus04, Sch99b, WP06]. Roberta [You25]. Roberto [Ray19]. Robin [Jia14]. Rockefeller [Ano82i, Ano88d, Löw97, Zal92]. Rocky [Vet12]. Rodent [Rad25]. Rogue [AN08]. Role [Ack06, Cec24, Cre10b, Fis00, GZ98, Log02, Str99, Tam15, Tes23, vHS19, Mar92, McC69, Mic86, Pod96, San73, Zir68]. roles [Ano91b]. Roman [Nat24]. Romana [Del97]. Romance [Ano00c]. Romanes [Sch95]. Romantic [Rob11, Rus04, Jac84]. Romantische [Ano78d]. Ronald [You09]. Roosth [Wil19]. Roots [Loi11, Lyn19, Pet08, Pir23, All84]. Rosalind [Sel03]. Rose [Lau19]. Ross [Sla03]. Rossitten [dB11]. rot [Whe81]. Rothamsted [Par15]. Rous [BK10]. Routinization [Hub20]. Routledge [Gue24, Jin25, Ric25]. Routledge/Taylor [Ric25]. Roux [Ham97a]. Royal [Fla24]. Rüdin [JW13]. Rudolf [NL11]. Rule [Hag17, Ano91b]. Ruminology [WJ22]. Running [Sch15]. Ruse [Bri09, Ric04]. Rush [Rab11]. Russel [Fic24, Gre04, Smi15, Sta24]. Russell [RH84, Ste05]. Russia [Bru18]. Russian [Ada70, Gai73, GS92, Har90, Kon23, Str92]. Rutgers [Cro05]. Ruth [Hea20].

S [Ada00, Gru22, Jac01, Kin20, Koh24, Lin18, Liu23, Mat21, Pau83, RN11, RH84, Rus15, Tab08]. Sabatini [LaB17]. Sabina [Shi18]. Safaris [Man09]. Sailor [Hoq24]. Saint [Gal21]. Saint-Hilaire [Gal21]. Salmon [Tay98, Boc12]. saltationism [Lyo95]. Salvador [Lin23]. Samantha [Jia24]. Sambucana [Pal22]. sampling [Lus74]. Samuel [Tur18]. Sancho [Ste24]. Sanctuary [Ala12]. Sanger [GS10]. Sankaran [Boo23]. Santesmases [Cre19]. Santiago [Sum04]. Sara [Loz19]. Sarah [Lin18]. Sarcoma [BK10]. Sargent [Bar04a]. Savage [Gre25]. Save [Bar09]. Saved [Bar24]. Savelli [Wu19]. Savithri [Ric25]. Scale [And24]. Scenes [Gru18]. Schelling [Nyh18]. Schema [Liu24]. Schickore [HS18]. Schiemann [SM07]. Scholar [Ano88d, Gam21]. School [Fra79, Lad08, Rai03, Sha08, Ada68]. Schools [Sla03]. Schrödinger [Olb71, Wil16]. Schwartz [Fla24]. Science [Adl14, AD18, Ala12, Ank18, Ano78g, Ano91f, Bar23a, Bel01, Bez97, Bis23, Boc12, Chi09, Cre10b, Dua20, FC21, Fin08, FG17, Gau98, GZ98, Goo19, Gor19, Gos23, Gre94, Har03, Har11c, Hec21, Iid10, Ile05, JF03, Jin25, Joh15, Kei98, Kei17, Kle11, Koh24, Lew05, Lew18, LD20, Luk15, Mac01a, Man09, Mar09, Mat87, McE92, Mét14, Mil11, Mit21, Muk19, Par15, Pie19, Por04, Pre70, Rum12, Sch99b, SG18, Sla08, SR23b, Som24, Stu24, Sul11, Tay98, Tyr23, Wal01, Wei09, Wei05, Whi19, Win01a, You09, vdBD20, Ano75g, Ano75k, Bed83a, Bur70, CT90, Deb70, FS90, Ger70, Gil90, Gre90, Hol93, Kay89, Mar92, Sch80a, Sha69, Uch93, Gor19, Gue23, Kro97]. Sciences [Bla07, Bur99, Die16b, Gal22, Hal25, Hea20, Ile99, Jia24, Mét14, Nov23, Por79, Ste04, Wal01, War23, Wu19, Koh95, Scr69, Sat20]. Scientific [Art17, BB22, Bea93, Bec05, BK10, Bre02, CdC25, Com11, Dav97b, Gin10, Gor09, Gre07, Hen97, Hol17, Ile05, Kei01, Mét14, Muk14, Par08, Sel12, Str19, Wie06, WO05, ZS15, Ano91a, App88, Fin91, Gau96, Por95, Ric89a, Har11b]. scientifically [Pau96, HS18]. Scientist [Gor09, Irm05, Ric25]. Scientists [Che09, Lin24, Söd11, AA82, Koh04].

Scotland [Jen16]. Scott [Str19]. Screen [Stu23]. Screening [Mit17]. Scripps [Rai03]. Scriptural [Bel05]. scurra [Gro17a]. Sea [Adl14, AD18, Boc12, Kei17, Jia24]. Sea-Level [Kei17]. seacoast [DC96]. Search [BA07, Fag10, Goo19, Ile99, Mar04, Ano81c, De 74, PW92, Rei86]. seas [Reh79]. Seaside [Koh22b]. Season [Ton18]. Seasons [Hag24]. Seaweed [vP20]. Sebastian [Mét14]. Second [Rab11, Ric23, Gri85, Kre96]. Second Millennium.FemaleMan\begin{matrix}c \\ ◯ \end{matrix} MeetsOncoMouse^{T M} [Ham97b]. secondary [Lar87]. Secret [Cog22]. secretions [Bor76]. Section [Heg15]. Security [LD19]. seductive [PW92]. Seed [De 19]. Seeds [Fit19]. Seeing [Kei99, vHS19]. seen [Rus05]. Segerstrale [Can13]. Selborne [Mul69]. Selection [Ala24, Alt07, Bea16, Bea19, Bor03, Eng97, For12, Gro17a, Joh19a, Kly14, Mil08, Pet08, Smi15, The12, All68, Ano82n, Bed68, Col86b, Cor87, Eva84, Her71, Osp80, Pau88, Söd94, Wel73, Lig18]. Self [AM14, Joh15, TP94]. Self-Experimentation [Joh15]. Selling [Kay89]. Selya [Lin23]. Semen [Jan24]. Semi [HR11]. Semi-Arid [HR11]. Senescence [Par13]. Sentiment [SG18]. Separated [Alt07]. Separation [Chi09]. Sepkoski [Bri09, Sep05]. Séquard [Wal21]. Sequence [Dre23, Ste18, Str10, dC14]. Sequences [De 96, Hag10, Str10]. Sequencing [GS10]. Sergei [Ack06, Ack07]. Series [Bel05, Ile05, Zal05, App80]. Serra [DCS01]. Serves [Yi21]. Service [Cre06, Kle16]. Seven [Söd11]. Seventeenth [Gue13, Bow71, CB82, Deb70, Wig69]. seventeenth-century [Wig69]. Sewall [Cro90, Ste07]. Sex [Biv00, Ha11, Jan24, Lus00, Mil11, Moo95, Oud90]. sexist [Mor79]. Sexology [Ala24]. Sexual [Ala24, Alt07, Buc17, Ha11, Lig18, Wei87].

Sexuality [Log01]. Shadow [End03, Fre25, Lyo95]. Shale [Bar11a]. Shaped [Bar24]. Shaping [Mon18, vHS19, Jun96]. Shared [Nov23, Veu19]. Sharing [JACD18]. sharks [Bod83]. Sharon [Hag24]. Sharp [Nov19]. Shaw [Hal10]. Sheep [PSBS^{+}22, Di 95]. Sheila [Wei11]. Sheldon [Kei01]. Shelf [dR23]. Shepard [Por09]. Sherrie [All22]. Sherrington [Swa68b]. Shifting [Ano81h]. Shinozuka [Ona23]. Ship [Adl14]. Shipwrecked [Hoq24]. Shorebirds [Whi20]. Siam [Ona10b]. Siècle [Coe06]. Sign [Del97]. Signal [For12, LaB17]. Signature [Gia18]. Significance [Art24, Mét16, Mir08, Fry96, Moo95]. Sigrist [Ste04]. Silkworm [Ona10b]. silver [Get91]. Simians [Sur22]. Similarities [Gam21]. Simple [End23]. Simpson [Lap90, Lap94]. Since [Bea19]. Singer [Kev05]. Singular [May23, BGZ88]. Sinocentric [pY14]. Sinophone [BR20]. Sir [Cor83, De 74, Gre90, Wal01, Wel71]. Sisters [Söd11]. Site [SP25]. sites [Koh95]. Situating [Ras25]. Six [Bro73]. Sixteenth [Sho13, Ano91b, Pió79, CC18]. sixteenth-century [Pió79]. sixteenth-eighteenth [Ano91b]. Sketch [Joh07]. Sketches [vP20]. Skins [Bar23a]. skull [Di 95]. Skulls [dR23]. Slave [Koh24]. Slips [AdC24]. Smirnov [Vin22]. Smith [Ack04, Lad08, Pie19]. Smithsonian [Mac01a]. Snail [Gal22, Mil08]. Snake [HS18]. Social [Bar97, Bar05, Bur22, Can15, Cec19, Fag10, FC21, Gal22, Hol17, Kei01, ML10, Mit17, Par21, Tes23, WY23, Des84, Gil87, Gre77, Hay88, Mar92, Ric89a]. Socialism [Bec05]. Socialist [Hal10, Mac93, Yi21]. Societies [Can15]. Society [Mit17, Rum12, Sel12, Mit88, Smo94, Fla24, Wei05]. Sociobiology [JF03, Jum02, Lep15, Lev18, Stu23, YD00, Stu23]. Sociopolitical [Hoo97]. Sodium [Tru97]. Sodium-Channel [Tru97]. Soil [Ack07].

Sollas [Som05]. Solution [For20, Gor97]. Somatic [Log07]. Some [AA82, Ano76b, Eng97, GT69, Ric81]. Sommer [Mar24]. Sonneborn [Sch99a]. Sophia [Wil19]. Sophie [Cau19]. Soraya [Ban22]. Source [Sel23, AM68, Par71a]. Sources [Com11, RH09]. South [Ank18, Bro11, Jin25, Ank04, Car08, HR11, KFP09, Man09, Pat09, Por09]. Southern [Dan21]. Souvenirs [Ste04]. Soviet [Gai80, Har03, Kre96, Sht92]. Sow [Lus01]. Space [Adl14, Eri17, GM18, MM17, Str23]. Spain [Cre19, Gli76, Pin02, Pió79, San06]. Spallanzani [Cas73, San73]. Spanish [Ano75g, Bed83a, Bed83b, Pin02]. Sparrow [Hol17]. Spasmodic [Dre23]. Spatial [Fin08]. Speak [Hat21]. Special [Har11b, Har06b, Smo99b, GD18, Heg15, Hun19b]. specialty [Gli76]. Speciation [Kre07, Plu19]. Species [And24, Bri10a, Bri10b, Edd23, Hic21, Hop24, Sch99a, Ste24, Tan20a, Tes23, Wei10, vW14, Ano82m, Atr87, Biz92, Bow73, Far72a, Gou68, Lap94, Len80, May68, Par82, Slo86, Sta96, Bar24, dd22, Baj88, Bed88b, Bow89, Bra11a, Coo90, Far74, Joh07, Kle13, Kon23, Oga01, Tan17, ddAM20]. Specimen [Bar00, Shm20]. Specimens [Hun19c]. Spectre [PSBS^{+}22]. spectroscopy [Zal92]. Speculative [Hei19]. spelling [Sul83]. Spemann [Bra22a, Bra22b, Ham99, Rin88]. Spencer [Ano91h, Ano82h, Bec05, Hai91]. Sperm [Cog22, Hub20]. Sperm-Force [Cog22]. Sperma [De 19]. spermatic [San73]. spermatozoa [Rue83]. Spermatozoan [Cas73]. Spero [Por25]. spiral [Mon70]. Spirit [Mur23]. Spirits [Por09]. Splendid [Sel23]. Splice [Mor23]. Split [Mor23]. Spontaneous [Mat10, Str99, Zwi18, Far72b, Far72c, Rue84]. Spot [Ham97a]. Spotted [Fun12]. Sprague [Rud90]. Spring [Boc88, Cas15, Kev05, Sel05]. Springer [Fla24, Hal25, Mét14]. Spur [Hob15].

Squier [Yan18]. SSSR [Ano69b]. St [Fra79]. Stabilizing [Ams91]. Stages [Koh04]. Stahl [Nyh18]. Stake [RPM23]. Stalemate [Bea87]. Stalinist [Dua20]. Stalinize [Dua20]. Standardization [Dru23]. standardized [Cla93]. Standardizing [Led04]. Standards [Fag10, Log02, Cla93]. standing [Cre96]. Stanley [Cre96]. Stark [Wal20]. Starn [Mar04]. Starved [Fre25]. State [Bel06, Dan21, Sul11, Hal76, Sin00]. statement [Bod83]. States [All13b, DT07, Gor09, Hun19c, Mat06, Ona10b, Par08, Par13, SP25, WJ22]. Station [Lin24, Par15, Koh22b]. Stations [Vet12]. Statistical [Gro17a, Hag03]. Statistics [Lou09, Par15, Tam15]. status [Gou70]. Stazione [Ben88]. Stebbins [Smo99b]. Steel [Buh00]. Stefanie [May23, Ray19]. Steigerwald [Bra20]. Stem [Fag10]. Stems [Fag10]. Stephen [Bar05, Boc90, Sch18, Sep05]. steps [Grm68]. Steven [Cro05]. Stevens [Lin18, dC14]. Still [SS99]. Stone [Por85]. Stories [Gor19, Lau19]. Storm [Mil11]. Story [Pau23, Pie19, Tob09, Var09]. Strange [Rad25, Web20]. Strasser [SD19a]. Strategies [Rhe95]. strategy [Col86a]. Stratigraphic [Dre23]. Stratigraphy [Dre23]. Strick [Cro05]. Strictly [Mac01a]. Structuralism [Rie90]. Structure [Ric00, Str10, Wil16, Ano91d, Ras95, Sha72, Mir22]. Struggle [Jin25, Fal95, Gai80, Sap83]. Students [Sin97]. Studies [Bar92, Cai09, Ege68b, Gut97, Lau19, Lev18, RR21b, Sto15, Vic23, Wal21, Ano75i, Bar95, Bro82, Chu69, Dex70, IB71]. Study [ÁTBHBJB24, Arm12, Ban22, Ham07, Hoo97, MM17, Nic22, O’M18, Pet08, dC18, All91, Bed73, Bor89, Ege69, Sch89, Smo94, Wel71, Shi18]. Studying [MM24, MD16]. Styles [Elw07, Van94b, Nov23]. Subgenres [Söd11]. Subject [KG21, Smo99b]. Sublime [Bra11a]. Subscribed [Hun19c]. Subscribing [Hun19c]. subsets [KC94].

Substance [Liu17]. Subversive [Tam15]. success [Bro88]. Succession [Bri10a, Bri10b]. Suffering [Mit21]. Suggestions [Bow77]. supplicant [PW92]. Support [Coo02, Str23]. Supposed [Ank00]. Suppressing [Wit16, Ack04]. Suppressors [KC97]. Surfaces [Hei19]. Surgeon [Loz19]. Surroundings [Kin20]. Survey [Fis00, Hun19c, Mac01a, Ano75h]. Survival [Pau88, Di 95, Mac93]. Survivors [Hoo97, Lin92]. Susan [Ben09, Yan18]. Susanne [Sac09]. Süssmilch [Kly14]. Sverdlovsk [Gor97]. Sweat [Joh15]. Swiss [FE20]. Switzerland [Pir24]. Symbiosis [Wei11]. Syncretic [Sch10]. synergy [Kel90]. Synonymy [Wit16]. Synthase [All02, Pre13]. Synthesis [Ada70, Bea19, Bor03, Cai09, Elw07, GH19, Gro19, Hag15, Hun19a, Hun19b, Kim23a, Kle99, Kre07, May97, O’M18, Plu19, Sep19, SC14, Tab04, Veu19, Bur93b, Die95, Har94, Rhe93, Rhe96, Smo92, Wai88]. Synthetic [Gra20, Wil19]. System [Ala12, Cog02, Nov16, vP21, Eds72, Pla69, Slo72, Sti94]. Systematic [Bel06, Hag03, RWE13, Dex70]. Systematics [Buh00, Kle13, Mur23, Qui17, SL18, Ste84]. Systems [Lam15, Ste07, Str23, Ano82b, Löw94, Rhe93, Rhe96, Sil94]. Szent [Bei08].

T [Fic24, Mét14, Tay88, KC94]. T-cell [KC94]. T. [Ano82c, Lyo95, Ric00]. tables [Hub20, Pow18]. Taken [Tay98]. Takes [Bel06]. Taking [Fol21]. Tale [Bar04a, Boo23, dMP23, Rad25, Ton18]. Talk [Com11]. Talmudic [Kot81]. Tamarisk [Che09]. Tammes [Sta95, SM07]. Tank [Jia24]. Tanks [LM14]. Tansley [vdV14]. Tantalizing [Sul09]. tarnished [Whe81]. task [Bur93b]. Tatum [Ano82i]. Taxidermists [Bar24]. Taxidermy [Hei19]. taxonomists [Hag86]. Taxonomy [JK99, Joh22, Kle16, Tan17, Vin22, Hag84]. Taylor [Ric25]. Teacher [Sta23].

Technique [Bur93b, CK92, Rin88]. Techniques [Cro14, Mit17, Tru97]. Technocratic [Tay88]. Technological [Kim18]. Technologies [RF18, Sto15]. Technology [Chi09, Gut97, Luk15, Nat24, Yi08, CK92, Cor84, Mar92]. Technoscience [Ham97b]. Teleology [Str24, Ano76a, Bec69, Chu69]. Television [McG19, Ton18]. Telomeres [Tob09]. Temin [Fis10]. Temminck [Mir08, Mir11]. Temporal [BP19]. Tensions [ÁTBHBJB24, Pau16]. Teratology [Art24]. Term [Lew18, Mar11]. terms [Bea82]. Terrible [Erl16]. Test [Log02, SP25]. Testing [Har08]. Textbooks [Lad08]. Texts [Rud00, Ano91b]. Thayer [Kin78]. Thea [Lad08]. Their [Gli98, Som05, ddAM20, dd22, AA82, Mic86, Out86]. theism [Bro86]. theistic [Cor87]. Theobald [Ack04]. Theodore [Man13]. Theodosius [Joh22]. theology [Bla82, Gil87, Gil90, Ric81]. Theophrastus [Neg95, Sca78]. Theoretical [BB22, GH19, Bon90]. Theories [Bow73, Bru02, Del83, Ha11, Sch69, Sul11, Tan20a, WP06, Ams91, Ano91a, Pal91, Par18]. Theorists [Jen16]. theorizing [Pod96]. Theory [AMW00, Bru02, Civ13, For12, Gal17, Gro17b, Hal25, Hod09, Kle99, Kly14, KFP09, Mét14, dMP23, Pie19, Ric00, Riz20, RH09, Smi99, Tam15, Tes23, The08, WdLdS21a, WdLdS21b, Wie06, Zwi18, All69, All74b, Ano75d, Ano76b, Ano78c, Ano91c, Bed68, Bed72, Boy84, Boy86, Bur72b, Car71, Chu69, Col86b, Di 95, Die94, Fry96, Gil78, Gri74, Gri85, Hal71, Her74, Her77, Hul05, Jac89, Jac84, Kal83, Mon70, Osp77, Pre90, Pre77, Rei86, Ric81, Ric89b, Roe79, Söd94, Hal25, All22]. There [Log02]. Thermoregulation [Hag17]. Thesis [WdLdS21a, WdLdS21b]. Thiamin [Ras99]. Thienemann [dB11]. Things [Gia18, Log01]. Think [Wol12]. Thinking [All20, Lam15, Ras24a]. Third [MAR21, Mat21, Tho05, Wei11, Wei06]. thirties [Kin86]. Thoemmes [Bel05]. Thom [Lau19]. Thomas [Por25, Ste05, All68, Eng97, Gli04, Hal03, Led89, Lyo93, Mai16, Sch99b, Sun10]. Thomson [Ano06b]. Thoroddsen [Erl02]. Thorpe [Gil90]. Thorstein [Ray13]. Thorvaldur [Erl02]. Thou [Bae19a, Bae19b]. Thought [Gre92, Tan17, Ada70, Par71b, Ric89a, Van94b, Wei87, Wig69]. Threads [Var09]. Threat [Far09b]. Three [Cor21, vP21, Mit21]. Three-Dimensional [vP21]. Threshold [Lue21]. Thymectomy [Sti94]. Tide [Eri17]. Ties [Var09]. Tigers [Lew05]. til [Smo05b]. Tim [Rad04]. Time [AdC24, Ben11b, BP19, Cob25, HR11, Lau19, Ray13, Bel81]. Times [End11]. Tinbergen [Bur16, Vic18]. Tine [Sta95, SM07]. Tipping [Hon16]. Tissue [Sur22, Rhe96]. Tissues [Pas21]. Tobacco [Lew04, Sch14]. Together [Den16, Hen03]. Too [Bec16]. Tool [vH19]. Tools [Gis23, Bur96, DG96, Kay85, Nat24]. Topical [RR20b]. Topographies [Gro17a]. Tormenting [Smo99a]. torpedo [CB82]. Tortoises [Sul09]. Tournefort [dC23]. Tower [Ric24]. Toxic [Kei11b]. Toyama [Ona10b]. Tracy [Sch99a]. Trade [BJ97, VWP20a, VWP20b, Koh24]. Tradition [Gib13, Ano81b, Chu68, Gau93, Wei96, The96]. traditions [Gau93, Hag86]. traits [Fal91]. Trajectories [Boo23]. Trans [Gre25, Tan20b]. Transacting [VWP20a, VWP20b]. Transatlantic [Gru22, Koh24, Pet08, Ger70]. Transfer [Nic22]. Transformation [Erl02, Lon09, Mét16, Tes23, Ano75k, Ano82d, Gro79, Tay88]. Transformations [Joh04, Wu19]. Transformed [Wil04]. Transforming [Mét16]. Transformism [Jen16, Slo86]. Transformist [Cor05, Tes23]. Transfusion [Ben09, Soo21]. Transit [San16]. Transition [Sun13, Bur93b, Car74].

Translated [Cau19]. Translation [Ham99, Mor97b, NL11, Ano76c]. Translocal [CM11]. Transmutation [Bel06, Tan20a, Ano81a, Gri74, Her74, Her77]. Transmutationist [FE20, Des84]. Transnational [Bur22]. Transpacific [Ona23]. Transplantation [Ben09, Cro14, Gut97, Mun93]. Travels [Fla24, LK25]. Tree [Arc09, Som05, dV09]. Trees [Ben11a, Kra24, vP20, vP21, OV68, Man13]. trend [Str92]. Trends [Pau16]. Trials [Gue23]. Triangle [JACD18]. Trigonia [Gou68]. triphosphate [Ano91d]. Tristram [Eng97]. Tropical [LM14]. Tropics [Vet06]. Trunkless [Som05]. Trust [Smo99b]. Tucker [Wei09]. Tumor [Boo23, Löw94, Sug23]. Turn [Fin08, Rab11, dC18, Ano81f, Ric89b]. Tuscaloosa [Kim23b]. Twentieth [Ben09, Kim18, Rab11, Stu24, Ano81e, Ano91c, Cor89, Cro90, Hag84, Hag86, RH84, Bol22a, Bro11, Die16b, Dre20, Eri17, FC21, Ha11, HA19, Joh04, Joh09, Koh12, Man15, Tam20b, Vet12]. Twentieth-Century [Ben09, Kim18, Stu24, Ano81e, Ano91c, Cor89, Cro90, Hag84, Hag86, RH84, FC21, HA19, Koh12, Tam20b]. Twenty [Rus05, YD00]. Twenty-Five [YD00]. Twin [BS08]. Two [Ano76e, Boo23, Chu69, Dre20, For12, Jax98, Kei17, Nic22, Rad04, Slo07, Str24, vP21, Ege85, Par18, Pau83, RH84]. Two-Ocean [Kei17]. Type [Sta25, Wit16, Ano76f]. type-concept [Ano76f]. Types [Eig97].

U.S. [Cre06, McE92]. Udny [Tab04]. UK [Coo20, Mar24, Shm23a, Shm23b]. Ullica [Can13]. Ultimate [Kly14]. Ultra [Sul11]. Ultra-Nationalism [Sul11]. Unacademical [Jac01]. unacknowledged [Car71]. Unbound [Cre22]. Uncertain [Out86]. Uncommon [Kev05]. Unconscious [Alt07, Tur18].

Underground [Sch25]. Understanding [Bra11b, Can15, Ste07, Ege69, Sin89, Hal25]. UNESCO [Cas15, Gos23]. Unfinished [Sep19]. Unfold [Har12b]. Unger [Gli98]. Unification [Sch99a]. Uniform [Wal15]. Unifying [Smo92]. uninteresting [Pau96]. Unique [Smo05a]. unit [Car74, Jac89]. United [SP25, All13b, DT07, Gor09, Hun19c, Mat06, Ona10b, Par08, Par13, WJ22]. Unity [SS17, Smo99a, Jac89, Sha69]. Universal [Hod72]. Universe [Cro05, Gib19]. University [Ank20, Ank18, Ano82i, Ano04a, Ano05a, Ano06b, Ban22, Bar23a, Bar24, Ben18, Ben09, Ber19, Bis23, Bod23, Boo23, Bra20, Bri09, Bro24, Bru18, Can13, Cao24, Cau19, Coo20, Cro05, Des05, End23, Fic24, Fra79, Fre25, Gib24, Gli24, Gor19, Gre25, Gru18, Gue23, Hag24, Hei19, HS18, Hub20, Ile05, Irm05, Jia24, Kei11b, Kei23, Kim23b, Kim18, Kin20, Kle11, Koh24, Lan20, Lig18, Lin18, Liu23, Man13, Mar09, Mat21, May23, McG21, McG18, Mes19, Mil11, Mor23, Muk19, Nat24, Nov23, Nyh18, Ona23, Por09, Por25, Pow18, Rab11, Rad23, Ras24a, Ray19, Sac09, Sat20, Shi18, Shm23a, Shm23b, Som21, Som24, SM07, Ste05, Str23, Stu24, SD19a, Sug23, Tan20b, Tyr23, Wal20, Web20, Wei09, Wei11, Wu19, Yan18, You09, You25]. University [dC14, Ala12, AM68, Ben85, Cre96, Luk15, Ray13, Ric06]. Unknown [dC23, Wei77]. unnatural [AJR94]. Unsatisfactory [Bel06]. Untangled [Fat17]. Urbana [Wei09]. ureters [Sha85]. Urphänomene [Ste02]. USA [Kei23, Lin23]. Use [HR11, Hop24, Log01, Col86b, Eva84, Rai89]. used [Sum93]. Uses [Gue13, Mos98, San06]. Using [vH19]. USSR [Ano69b, Bro71, Bro73, Riz20, Vor92]. Utility [Mir22]. Utopia [Hal10]. Utrecht [Zal92]. Utter [Win24].

V [Ano78e, GS15, Ano69b]. Vaccination [Dan21]. Value [Irm05, Kol22, Sil94]. Values [Lew05, Kel88]. Valuing [Whi20]. Vanessa [Lan20]. Variance [Par15]. Variation [Pet08, Veu19, Win00, Win01b, Ams91, Gai73, Wel71]. Variations [RF18]. Varieties [Par82]. Veblen [Ray13]. Vegetation [Fis00, HR11, SP25]. veins [Sha85]. Velikovsky [Gor12]. Venom [HS18]. Venture [Yi08]. Vera [Hic21]. Vernadsky [Pir23]. Vernon [Lar99, Ona10b]. versa [Lew87]. Versus [Esp13, Lam15, Log01, Ano76e, Ano78c, Bon90]. vertebrae [Pad95]. vertebral [Di 95]. Vertebrate [Bri10a, Bri10b, Ger70, Sun13, GG93]. Vertebrates [Gal21, Rai89]. Vesalian [Pió79]. Vestiges [Hod72, Mac01b, Sch90]. vice [Lew87]. Vicedo [Pau23]. Victim [BS08]. Victor [Lus74]. Victorian [Alb01, Gre04, Web20, Ano78g, Bla07, Des01, Elw07, Gök94, Jac03, Mac01b, Ric89a, Sch80a]. Vienna [Coe06, Gli98]. Vietnam [Yi21]. View [Bra24, AA82, Ano82c, Gre90]. viewed [Ghi05]. Viewpoint [Kle13]. Views [Die16b, Har12b, Hay88, Ano75j, Hal76, Rho87]. viii [Can13]. Viktor [All04]. Vilmorin [GZ98]. Vincent [Par08]. Vinogradskii [Ack06, Ack07]. Virginia [Por25]. virology [Kev93, Sug23]. Virus [BK10, Lew04, Sch14, Pod96, Van94b, Sug23]. Viruses [Boo23, Yi08, vHS19, LT93, Boo23]. Visible [Sch14]. Vision [Gra20, Mor97a, Rad06, Cor84]. Visionaries [Sat20]. Visionary [Ada00, Ham97a]. visions [FS90]. Visual [Kei99, LaB17, Man13, Sch04, Tay21]. Visualizing [Mar24, Mes19]. vital [Slo86]. Vitalism [Bra22b, Ham99, Wol13, Bro74, Hei72, Mét14]. Vitalist [Erl02]. Vitality [Bra20]. Vitamin [Fre25, IB71]. vitro [Rhe93]. Vivisection [Gue13, Sho13]. Vladimir [Pir23, Riz20]. Vogt [AN08]. Vol [Loz19]. vols [Bel05]. Volume [Mét14, Hal25]. Volumes [Slo07].

Voluntarist [Yi21]. Voyage

[Bri10a, Bri10b, Sul82b, Sul83]. Vries [All69, Sta03, Zir68]. Vrolik [dR23]. Vulnerabilities [Tay21].

W [Can13, DC96, Dex70, Kay89, Man13, Ste90, Ben18, Cre19, Som21]. W. [Gil90, Kei98, McC69, Mit88]. Wagner [Wei10]. Wahlberg [Hub20]. Waiting [Hag10]. Wall [San16]. Wallace [Bed68, Bed72, Bed88a, Bed88b, Eng97, Fag07, Fic77, Fic24, Sch84, Sch90, Shm20, Smi15, Smi25, Smo25b, Sta24, Vet06, vW14, Gre04]. Walter [Esp20, Pie12]. War [Gau98, Hag15, Hag17, Lin23, Som21, SP25, SD19b, Tay88, Bea88, Pau83, Ber19, Cro89, GS15, Hec21, Pin02, SD19c, Wol12, dLK12]. Warburg [Koh73]. Wardian [End23]. Warington [Ham86]. Warman [Kle04]. Warming [Mac21a, Mac21b, Col86a]. Warren [Fre25]. Wars [BB22, Bar04a, Nic22, SL18]. wartime [Kay89]. Was [Mor79, Sta96, Gro17b, Wei77]. Washington [Fra79, Wil12]. Wasps [Can15]. Watching [Har12b]. Water [Eri17, Hal13, Wil04]. Waters [Kli98, LM14]. Watkiss [Loz19]. Watson [Fat17, Sel03, Sel05]. Way [MM24]. Ways [Jac01]. Wealth [Cre19]. Weapons [Gor97]. Web [Bar11a]. Webster [Nat24]. Week [Sta23]. Weidman [Stu24]. Weighing [Bea87]. Weikart [Far05]. Weismann [Chu68, Chu10, Esp13, May85, Sta03, Wei10, Win01b]. Weismannism [Bur72b]. Weiss [Wei11]. Wellman [Ras24b]. Wellmann [Ste19]. Wells [Hal03, Hal10]. Wendell [Cre96]. Wenner [Mun05]. Went [AD18, Kin09b]. We’re [Pri24, Ano75m, Ano81d, Log02]. Werkman [Sin97, Sin00]. Westerdijk [Kol22]. Western [Wil04]. Wheldale [Ric07]. Where [Bae19a, Bae19b, Wil04]. Which [Bol22b, Lap90]. White [Bar23a, Hoq24, Mul69]. Wholes [NL11]. Whose [Mir11, dC18]. Wild [Mac21b, Mar04, Hic21, Kei23, Mac21a]. Wilderness [Man09]. Wildlife [Ano00c, Car08]. Wilhelm [BS08, Wei05, Wei06, Ham97a, RH09, dG02]. Will [Tes23]. William [Ano76d, Cor83, Gue24, Wel71, Bar04a, Buh00, Chu74, Coc73, Cog02, Dar77, Dav97a, Hal03, Hal10, LD20, Lon09, Mar09, Nov19, Num79, Pet08, Rai89, Ric24, Som05, Var09, Wei09]. Williams [WP06]. Wilson [Ano76b, Gib13]. Win [All02]. Wind [Cha22]. Winding [SD14]. Window [Sun10]. wine [Lew87]. Wing [Lep15]. Winter [HR11]. Wistar [Cla93]. within [Die98]. Without [The08, Vin22, Lar86]. Witness [Mit21]. Wolf [Sto15, Di 95]. Wolfe [Ack04, Mét14]. Wolff [Roe79]. Wolves [Man09]. Woman [Ano78g, Ric25, Hor76]. Women [DT07, Fol21, SM07, ZS15]. Wonder [Gue23, Bis23]. Wood [Moh18, Sin97]. Woodger [NG14, RH84]. Woods [Luk22, Mon18, ZS15, Luk21]. Wook [Jin25]. Worboys [Web20]. Word [Bez97, Ano75n, Bea82]. Work [BH21a, BH21b, Can13, Can15, Erl16, Hal03, Kei99, Kle24, Mir08, Nov19, Rad98, Sch15, Slo07, WO05, dC23, Bur93b]. works [Neg95]. World [Bed83a, Cao24, Kim23b, May23, Nic22, Rad07, Wal15, Gre90, Lap90, CC18, End23, Hag15, Hag17, Tay88]. Worldviews [Sep06]. worldwide [Ege83a]. worms [Far72b]. Wright [Cro90, Ste07]. Writing [Gre07, Hec21, HS18, RR22b, Sel05, Smo91, Tom22]. Writings [SK22, Atk85, Gre94, Rie88]. Wyman [App88]. Wynne [Bor03]. Wynne-Edwards [Bor03].

X [Loz19, Kim18, Mar09, Tyr23, You09]. XI [Cre19, Des05, Far05, Wei09]. xii [Gue24, Ile05]. xiii [Cro05]. xiv [Ano06b, Mil11, Rab11, Sug23]. xv [Kle11, Ste04]. xvi [Por09]. xvii [Kei11b, Sel05]. xx [Zal05]. xxxiii [Irm05].

Yale [Ano06b, Ben18, Cao24, Hag24, Kei11b, Mar23, Sla03]. Year [Far09a]. Years [Die17c, Fat17, Gau98, Har22a, LM14, Men17, Oga01, PBDL17, Rus05, SK22, YD00, Bea88]. Yellow [Löw97]. York [Ano04a, Ano05a, Ben09, Coo20, Far05, Fla24, Gue24, Hei19, Irm05, Mét14, Ras24b, Ric25]. Yorkshire [Alb01]. Young [Gis23, Mil11, Sum04, Sch79, Sch88]. Youth [Wal20]. Yugoslavia [Dua20]. Yule [Tab04]. Yuri [GS15].

Zalasiewicz [Cau19]. Zamecnik [Bur93b, Rhe96]. Zammito [Nyh18]. Zegher [Ano05a]. Zhadin [Riz20]. Ziying [Luk22, Luk21]. Zone [Ras24b]. Zoo [Pow18]. Zoogeography [Fic77]. Zoologica [Ben88]. Zoological [Mir11]. Zoologists [Joh09]. Zoology [Erl09, GM18, Gaz12, Hag03, Joh09, Str20, Sun13, Ano76f, GG93]. Zuckerkandl [Mor98]. zweite [Gli04].

References

Abir-Am:1982:ERH

[AA82]  Pnina G. Abir-Am.  Es-

say review: How scientists view their heroes: Some remarks on the mechanism of myth construction. Journal of the History of Biology, 15 (2):281–315, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                                     tronic). URL http://link.
                                                     springer.com/article/10.
                                                     1007/BF00233019.

Abraham:2004:NRM

[Abr04] Tara H. Abraham. Nicolas Rashevsky’s mathematical biophysics. Journal of the History of Biology, 37(2): 333–385, June 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000038267.
09413.0d;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000038267.
09413.0d.pdf.

Ackert:2004:BRC

[Ack04]  Lloyd Ackert.  Book re-

view: Claude E. Dolman and Richard J. Wolfe, Suppressing the Diseases of Animals and Man: Theobald Smith, Microbiologist. Journal of the History of Biology, 37(3): 597–598, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2092-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2092-5.pdf.

Ackert:2006:RMA

[Ack06] Lloyd T. Ackert, Jr. The role of microbes in agriculture: Sergei Vinogradskii’s discovery and investigation of chemosynthesis, 1880–1910. Journal of the History of Biology, 39(2):373–406, July 2006. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-006-0008-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-0008-2.pdf.

Ackert:2007:CLE

[Ack07] Lloyd T. Ackert,, Jr. The “cycle of life” in ecology: Sergei Vinogradskii’s soil microbiology, 1885–1940. Journal of the History of Biology, 40(1): 109–145, March 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9104-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9104-6.pdf.

Adler:2018:WPS

[AD18] Antony Adler and Erik Dücker. When Pasteurian science went to sea: The birth of marine microbiology. Journal of the History of Biology, 51 (1):107–133, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9477-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9477-8.pdf.

Adams:1968:FPG

[Ada68]  Mark B. Adams.  The

founding of population genetics: Contributions of the Chetverikov school 1924– 1934. Journal of the History of Biology, 1(1):23– 39, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00149774.

Adams:1970:TSP

[Ada70] Mark B. Adams. Towards a synthesis: Population concepts in Russian evolutionary thought, 1925–1935. Journal of the History of Biology, 3(1): 107–129, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00569308.

Adams:2000:LJV

[Ada00] Mark B. Adams. Last judgment: The visionary biology of J. B. S. Haldane. Journal of the History of Biology, 33 (3):457–491, December 2000. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004891323595;
http://link.springer.com/
content/pdf/10.1023/A:1004891323595.
pdf.

Adams:2023:BEM

[Ada23] Mark B. Adams. Beginnings: Everett Mendelsohn, 1963–1973. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09745-5.

AlvesValverde:2024:POH

[AdC24] Bruno Alves Valverde and Cristina de Campos. “Pray observe how time slips by:” collaborators, assistants, and the background dynamics in the publication of Darwin’s Cirripedia Project. Journal of the History of Biology, 57 (3):349–377, September 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09781-9.

Adler:2014:SLM

[Adl14] Antony Adler. The ship as laboratory: Making space for field science at sea. Journal of the History of Biology, 47(3): 333–362, August 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9367-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9367-7.pdf.

Anderson:1994:TUH

[AJR94] Warwick Anderson, Myles Jackson, and Barbara Gutmann Rosenkrantz. Toward an unnatural history of immunology. Journal of the History of Biology, 27(3): 575–594, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058997.

Alborn:1989:JHB

[AKB89] Timothy L. Alborn, Elizabeth B. Keeney, and Keith R. Benson. The J. H. B. bookshelf. Journal of the History of Biology, 22(2):361– 371, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139519.

Alagona:2004:BFP

[Ala04] Peter S. Alagona. Biography of a ‘feathered pig’: The California condor conservation controversy. Journal of the History of Biology, 37(3): 557–583, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2083-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2083-6.pdf.

Alagona:2012:SSH

[Ala12] Peter S. Alagona. A sanctuary for science: The Hastings Natural History Reservation and the origins of the University of California’s natural reserve system. Journal of the History of Biology, 45 (4):651–680, November 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9298-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9298-0.pdf.

Alaniz:2013:BR

[Ala13] Rodolfo John Alaniz. Book review. Journal of the History of Biology, 46(2): 323–325, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9357-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9357-9.pdf.

Alaniz:2024:HES

[Ala24] Rodolfo John Alaniz. Havelock Ellis, sexology, and sexual selection in post-Darwinian evolutionary biology. Journal of the History of Biology, 57 (1):89–112, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09760-0.

Alberti:2001:APO

[Alb01] Samuel J. M. M. Alberti. Amateurs and professionals in one county: Biology and natural history in Late Victorian Yorkshire. Journal of the History of Biology, 34(1): 115–147, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010373912743;
http://link.springer.com/
content/pdf/10.1023/A:1010373912743.
pdf.

Allen:1968:THM

[All68] Garland E. Allen. Thomas Hunt Morgan and the problem of natural selection. Journal of the History of Biology, 1(1):113–139, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00149778.

Allen:1969:HVR

[All69] Garland E. Allen. Hugo de Vries and the reception of the “mutation theory”. Journal of the History of Biology, 2 (1):55–87, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137268.

Allen:1974:I

[All74a] Garland E. Allen. Introduction. Journal of the History of Biology, 7(1): 1–3, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00179290.

Allen:1974:OMC

[All74b] Garland E. Allen. Opposition to the Mendelian-chromosome theory: The physiological and developmental genetics of Richard Goldschmidt. Journal of the History of Biology, 7(1):49–92, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00179293.

Allen:1984:ERR

[All84] Garland E. Allen. Essay review: The roots of biological determinism. Journal of the History of Biology, 17(1): 141–145, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00397505.

Allard:1990:FCL

[All90] Dean C. Allard. The fish commission laboratory and its influence on the founding of the marine biological laboratory. Journal of the History of Biology, 23(2):251–270, ???? 1990. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00141471.

Allen:1991:ERH

[All91] Garland E. Allen. Essay review: History of agriculture and the study of heredity — a new horizon. Journal of the History of Biology, 24(3): 529–536, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00156324.

Allchin:2002:EWN

[All02] Douglas Allchin. To err and win a Nobel Prize: Paul Boyer, ATP synthase and the emergence of bioenergetics. Journal of the History of Biology, 35(1):149– 172, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014583721788;
http://link.springer.com/
content/pdf/10.1023/A:1014583721788.
pdf.

Allen:2004:PEV

[All04] Garland E. Allen. A pact with the embryo: Viktor Hamburger, holistic and mechanistic philosophy in the development of neuroembryology, 1927–1955. Journal of the History of Biology, 37(3): 421–475, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-004-4448-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-4448-2.pdf.

Allen:2009:BR

[All09] Garland E. Allen. Book review. Journal of the History of Biology, 42(3):593– 598, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9195-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9195-y.pdf.

Allen:2012:BR

[All12]  Garland E. Allen.  Book

review. Journal of the History of Biology, 45(2): 353–356, May 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9322-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9322-z.pdf.

Allen:2013:BR

[All13a]  Garland E. Allen.  Book

review. Journal of the History of Biology, 46(2): 319–322, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9356-x;

http://link.springer.com/
content/pdf/10.1007/s10739-
013-9356-x.pdf.

Allen:2013:CHE

[All13b] Garland E. Allen. “Culling the herd”: Eugenics and the conservation movement in the United States, 1900– 1940. Journal of the History of Biology, 46(1):31– 72, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9317-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9317-1.pdf.

Allen:2016:RHB

[All16] Garland E. Allen. Reflections on the history of biology as a field: 1966–2014. Journal of the History of Biology, 49 (4):733–742, December 2016. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9458-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
016-9458-3.pdf.

Allmon:2020:IPE

[All20] Warren D. Allmon. Invertebrate paleontology and evolutionary thinking in the US and Britain, 1860–1940. Journal of the History of Biology, 53 (3):423–450, September 2020.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09599-1.

Allen:2022:BRS

[All22] Garland E. Allen. Book review: Sherrie Lyons, From Cells to Organisms: Reenvisioning Cell Theory. Journal of the History of Biology, 55(1):181–184, March 2022. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09667-8.

Alter:2007:SBI

[Alt07] Stephen G. Alter. Separated at birth: The interlinked origins of Darwin’s unconscious selection concept and the application of sexual selection to race. Journal of the History of Biology, 40(2): 231–258, June 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9113-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-9113-5.pdf.

Allen:1968:NSM

[AM68] Garland E. Allen and Dennis M. McCullough. Notes on source materials: The Edwin Grant Conklin papers at Princeton University. Journal of the History of Biology, 1 (2):325–331, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351924.

Allen:1999:E

[AM99] Garland Allen and Jane Maienschein. Editorial. Journal of the History of Biology, 32(1):1, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004443922228;
http://link.springer.com/
content/pdf/10.1023/A:1004443922228.
pdf.

Allen:2001:EIa

[AM01a] Garland Allen and Jane Maienschein. Editors’ introduction. Journal of the History of Biology, 34(1): 1–2, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010305711432;
http://link.springer.com/
content/pdf/10.1023/A:1010305711432.
pdf.

Allen:2001:EIb

[AM01b] Garland Allen and Jane Maienschein. Editors’ introduction. Journal of the History of Biology, 34(2): 227–228, June 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1010509702059;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1010509702059.
                                        pdf.

Anderson:2014:FIS

[AM14] Warwick Anderson and Ian R. Mackay. Fashioning the immunological self: The biological individuality of F. Macfarlane Burnet. Journal of the History of Biology, 47(1): 147–175, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9352-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9352-1.pdf.

Allen:2017:JCE

[AM17] Garland E. Allen and Jane Maienschein. JHB as a collaborative effort. Journal of the History of Biology, 50(3): 469–471, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9480-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9480-0.pdf.

Amsterdamska:1991:SIC

[Ams91] Olga Amsterdamska. Stabilizing instability: The controversy over cyclogenic theories of bacterial variation during the interwar period. Journal of the History of Biology, 24 (2):191–222, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209429.

Agutter:2000:DTB

[AMW00] Paul S. Agutter, P. Colm Malone, and Denys N. Wheatley. Diffusion theory in biology: a relic of mechanistic materialism. Journal of the History of Biology, 33(1): 71–111, March 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004745516972;
http://link.springer.com/
content/pdf/10.1023/A:1004745516972.
pdf.

Amrein:2008:GRC

[AN08] Martin Amrein and Kärin Nickelsen. The gentleman and the rogue: The collaboration between Charles Darwin and Carl Vogt. Journal of the History of Biology, 41 (2):237–266, June 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9144-6;
 http://link.springer.com/

                                        content/pdf/10.1007/s10739-
                                        007-9144-6.pdf.

Anderson:2016:PEP

[And16] Warwick Anderson. Postcolonial ecologies of parasite and host: Making parasitism cosmopolitan. Journal of the History of Biology, 49(2): 241–259, April 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9407-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9407-6.pdf.

Anderson:2024:IGS

[And24] Eric Burns Anderson. Invasion on so grand a scale: Darwin, Lyell, and invasive species. Journal of the History of Biology, 57(2): 207–229, June 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09772-w.

Ankeny:2000:MMS

[Ank00] Rachel A. Ankeny. Marvelling at the marvel: The supposed conversion of A. D. Darbishire to Mendelism. Journal of the History of Biology, 33 (2):315–347, September 2000. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1023/A:1004750216919;
http://link.springer.com/
content/pdf/10.1023/A:1004750216919.
pdf.

Anker:2004:PES

[Ank04] Peder Anker. The politics of ecology in South Africa on the radical left. Journal of the History of Biology, 37(2): 303–331, June 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000038258.
63985.aa;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000038258.
63985.aa.pdf.

Anker:2018:BRJ

[Ank18] Peder Anker. Book review:

Jane Carruthers, National Park Science: a Century of Research in South Africa (Cambridge: Cambridge University Press, 2017), 554 pp., illus., bibl., $67.72 Hardback, ISBN 978-1-107-19144-

0. Journal of the History of Biology, 51(3):617– 619, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9531-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9531-1.pdf.

Ankeny:2020:BRN

[Ank20] Rachel A. Ankeny. Book review: Nicole C. Nelson, Model Behavior: Animal Experiments, Complexity, and the Genetics of Psychiatric Disorders (Chicago: University of Chicago Press, 2018), 272 pp., 6 b&w illus., $30.00 Paperback, ISBN: 978-0-226-54608-7. Journal of the History of Biology, 53(4):657– 658, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09625-2.

Anker:2023:EMH

[Ank23] Peder Anker. Everett Mendelsohn: The Harvard Professor. Journal of the History of Biology, 56(4):??, December 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09748-2.

Anonymous:1968:ER

[Ano68] Anonymous. Essay review. Journal of the History of Biology, 1(1):155–161, ???? 1968. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00149780.

Anonymous:1969:JHB

[Ano69a] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 2(2):445– 455, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125028.

Anonymous:1969:RBV

[Ano69b] Anonymous. Razvitie biologii v SSSR (development of biology in the USSR). Journal of the History of Biology, 2 (2):439–444, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125027.

Anonymous:1975:ACO

[Ano75a] Anonymous. Arthur Cushny, optical isomerism, and the mechanism of drug action. Journal of the History of Biology, 8(2):145–165, ???? 1975. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130436.

Anonymous:1975:BPM

[Ano75b] Anonymous. Biology and philosophy: The methodological foundations of biometry. Journal of the History of Biology, 8(1):85–93, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129596.

Anonymous:1975:BA

[Ano75c]  Anonymous.  Biomedical

archives. Journal of the History of Biology, 8(2):319– 320, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00130442.

Anonymous:1975:CDT

[Ano75d] Anonymous. Charles Darwin’s theory of evolution: An analysis. Journal of the History of Biology, 8(2): 219–241, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00130439.

Anonymous:1975:EDM

[Ano75e] Anonymous. An early Darwin manuscript: The “outline and draft of 1839”. Journal of the History of Biology, 8 (2):191–217, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00130438.

Anonymous:1975:EEF

[Ano75f] Anonymous. Euripus; or, the ebb and flow of the blood. Journal of the History of Biology, 8(2):321–350, ???? 1975. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF00130443.

Anonymous:1975:FAM

[Ano75g] Anonymous. Félix de Azara:

The myth of the isolated genius in Spanish science. Journal of the History of Biology, 8(1):67–83, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129595.

Anonymous:1975:HBS

[Ano75h] Anonymous. The history of biochemistry: a survey. Journal of the History of Biology, 8(2):275–318, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130441.

Anonymous:1975:RPD

[Ano75i] Anonymous. Reflections on the progress of Darwin studies. Journal of the History of Biology, 8(2):243– 273, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130440.

Anonymous:1975:DAH

[Ano75j] Anonymous. The development of Albrecht von Haller’s views on embryology. Journal of the History of Biology, 8(2):167–190, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130437.

Anonymous:1975:TSN

[Ano75k] Anonymous. The transformation of the science of nutrition. Journal of the History of Biology, 8(1):135–144, ???? 1975. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129599.

Anonymous:1975:NOD

[Ano75l] Anonymous. “Un naturalista original”: Don Félix de Azara, 1746–1821. Journal of the History of Biology, 8 (1):15–66, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129594.

Anonymous:1975:WOP

[Ano75m] Anonymous. Were the opium poppy and opium known in the ancient Near East? Journal of the History of Biology, 8(1):95–114, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129597.

Anonymous:1975:WAD

[Ano75n] Anonymous. A word about Descartes’ mechanistic conception of life. Journal of the History of Biology, 8(1): 1–13, ???? 1975. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129593.

Anonymous:1976:CTI

[Ano76a] Anonymous. Critical teleology: Immanuel Kant and Claude Bernard on the limitations of experimental biology. Journal of the History of Biology, 9(1):59–91, ???? 1976. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129173.

Anonymous:1976:EBW

[Ano76b]  Anonymous.  Edmund B.

Wilson as a preformationist: Some reasons for his acceptance of the chromosome theory. Journal of the History of Biology, 9(1): 29–57, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129172.

Anonymous:1976:IDB

[Ano76c] Anonymous. The “initial discourse” to Buffon’s histoire naturelle: The first complete English translation. Journal of the History of Biology, 9 (1):133–181, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129176.

Anonymous:1976:IKE

[Ano76d] Anonymous. The influence of Karl Ernst von Baer’s embryology, 1828–1859: a reappraisal in light of Richard Owen’s and William B. Carpenter’s “Palaeontological application of ‘von Baer’s law”’. Journal of the History of Biology, 9(1):1–28, ???? 1976. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129171.

Anonymous:1976:TDH

[Ano76e] Anonymous. Two Darwins: History versus criticism. Journal of the History of Biology, 9(1):121–132, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129175.

Anonymous:1976:TCZ

[Ano76f]  Anonymous.  The type-

concept in zoology during the first half of the nineteenth century. Journal of the History of Biology, 9(1): 93–119, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129174.

Anonymous:1978:DGR

[Ano78a] Anonymous. Drosophila genetics: a reductionist research program. Journal of the History of Biology, 11(1): 159–210, ???? 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127774.

Anonymous:1978:CLA

[Ano78b] Anonymous. Cell lineage, ancestral reminiscence, and the biogenetic law. Journal of the History of Biology, 11(1): 129–158, ???? 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127773.

Anonymous:1978:CLV

[Ano78c] Anonymous. Charles Lyell versus the theory of central heat: a reappraisal of Lyell’s place in the history of geology. Journal of the History of Biology, 11(1):101–128, ???? 1978. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127772.

Anonymous:1978:GFO

[Ano78d]  Anonymous.  Generational

factors in the origin of Romantische Naturphilosophie. Journal of the History of Biology, 11(1):57–100, ???? 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127771.

Anonymous:1978:JHBa

[Ano78e] Anonymous. The J. H. B. archive report the A. V. Hill papers at Churchill College, Cambridge. Journal of the History of Biology, 11(1): 211–214, ???? 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127775.

Anonymous:1978:JHBb

[Ano78f] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 11(1): 215–221, ???? 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127776.

Anonymous:1978:SCV

[Ano78g] Anonymous. Science corrupted: Victorian biologists consider “the woman question”. Journal of the History of Biology, 11(1):1–

55, ????  1978.  CO-

DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127770.

Anonymous:1980:JHBa

[Ano80a] Anonymous. The J. H. B.

bookshelf. Journal of the History of Biology, 13(1): 159–168, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125358.

Anonymous:1980:JHBb

[Ano80b] Anonymous. The J. H. B.

bookshelf. Journal of the History of Biology, 13(2): 347–356, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125747.

Anonymous:1981:BGD

[Ano81a] Anonymous. Biogeography and the genesis of Darwin’s ideas on transmutation. Journal of the History of Biology, 14(1):1–41, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127513.

Anonymous:1981:CMT

[Ano81b] Anonymous. The continuation of the morphological tradition: American paleontology, 1880–1910. Journal of the History of Biology, 14(1): 129–158, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127518.

Anonymous:1981:SNB

[Ano81c] Anonymous. In search of the new biology: An epilogue. Journal of the History of Biology, 14(1):177–191, ???? 1981. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127520.

Anonymous:1981:IWA

[Ano81d] Anonymous. Introduction: Were American morphologists in revolt? Journal of the History of Biology, 14 (1):83–87, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127515.

Anonymous:1981:MTC

[Ano81e]  Anonymous.  Morphology

and twentieth-century biology: a response. Journal of the History of Biology, 14(1): 159–176, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127519.

Anonymous:1981:PID

[Ano81f] Anonymous. Problems of individual development: Descriptive embryological morphology in America at the turn of the century. Journal of the History of Biology, 14(1): 115–128, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127517.

Anonymous:1981:PDE

[Ano81g] Anonymous. Pseudo-Dioscorides’ Ex herbis femininis and early medieval medical botany. Journal of the History of Biology, 14(1):43–81, ???? 1981. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127514.

Anonymous:1981:SAA

[Ano81h] Anonymous. Shifting assumptions in American biology: Embryology, 1890–1910. Journal of the History of Biology, 14(1):89–113, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127516.

Anonymous:1982:A

[Ano82a] Anonymous. Announcement.

Journal of the History of Biology, 15(2):i, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00233021.

Anonymous:1982:DCS

[Ano82b] Anonymous. The digestive and “circulatory” systems in Aristotle’s biology. Journal of the History of Biology, 15 (1):89–118, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132006.

Anonymous:1982:DCR

[Ano82c] Anonymous. The dinosaur connection: a reinterpretation of T. H. Huxley’s evolutionary view. Journal of the History of Biology, 15(3): 397–418, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00133144.

Anonymous:1982:DPT

[Ano82d] Anonymous. Discussion paper the transformation of natural history in the nineteenth century. Journal of the History of Biology, 15(1): 145–152, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132008.

Anonymous:1982:ENa

[Ano82e] Anonymous. Editor’s note.

Journal of the History of Biology, 15(1):iii–iv, ???? 1982. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132003.

Anonymous:1982:ENb

[Ano82f] Anonymous. Editor’s note.

Journal of the History of Biology, 15(2):i–ii, ???? 1982. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00233012.

Anonymous:1982:ENc

[Ano82g] Anonymous. Editor’s note.

Journal of the History of Biology, 15(3):iii–iv, ???? 1982. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00133142.

Anonymous:1982:EPM

[Ano82h] Anonymous. Evolution and the problem of mind: Part I. Herbert Spencer. Journal of the History of Biology, 15 (1):55–88, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132005.

Anonymous:1982:JHBa

[Ano82i]  Anonymous.  The J. H.

B. Archive report: the Edward Lawrie Tatum papers at the Rockefeller University Archives. Journal of the History of Biology, 15(1): 153–154, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132009.

Anonymous:1982:JHBb

[Ano82j] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 15(1): 155–161, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132010.

Anonymous:1982:JHBc

[Ano82k] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 15(3): 467–473, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00133147.

Anonymous:1982:MEC

[Ano82l] Anonymous. Maupertuis and the eighteenth-century critique of preexistence. Journal of the History of Biology, 15 (1):119–144, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132007.

Anonymous:1982:NLF

[Ano82m] Anonymous. New light on the foundations of the origin of species: a reconstruction of the archival record. Journal of the History of Biology, 15 (3):419–442, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00133145.

Anonymous:1982:WLD

[Ano82n] Anonymous. Why Lamarck did not discover the principle of natural selection. Journal of the History of Biology, 15 (3):443–465, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00133146.

Anonymous:1984:ERP

[Ano84a] Anonymous. Essay review:

Philosophy and biochemistry of regulatory proteins. Journal of the History of Biology, 17(3):429–431, ???? 1984. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126371.

Anonymous:1984:JHB

[Ano84b] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 17(3): 433–439, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00126372.

Anonymous:1985:ANJ

[Ano85a] Anonymous. Announcing a new journal. Journal of the History of Biology, 18 (1):161, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127962.

Anonymous:1985:JHBa

[Ano85b] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 18(1): 155–160, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127961.

Anonymous:1985:JHBb

[Ano85c] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 18(2): 289–294, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00120113.

Anonymous:1986:PN

[Ano86] Anonymous. Publisher’s notice. Journal of the History of Biology, 19(3):505, ???? 1986. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138289.

Anonymous:1987:PNa

[Ano87a] Anonymous. Publisher’s notice. Journal of the History of Biology, 20(1):137, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125262.

Anonymous:1987:PNb

[Ano87b] Anonymous. Publisher’s notice. Journal of the History of Biology, 20(3):iii, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139456.

Anonymous:1988:PAb

[Ano88a] Anonymous. Publisher’s announcement. Journal of the History of Biology, 21 (2):i, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146984.

Anonymous:1988:PAc

[Ano88b] Anonymous. Publisher’s announcement. Journal of the History of Biology, 21 (3):iii, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144086.

Anonymous:1988:PAa

[Ano88c] Anonymous. Publisher’s announcements. Journal of the History of Biology, 21 (1):iii, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125792.

Anonymous:1988:SRR

[Ano88d] Anonymous. Scholar in residence at the Rockefeller Archive Center 1989–1990. Journal of the History of Biology, 21(3):527, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144095.

Anonymous:1989:LE

[Ano89a] Anonymous. Letter to the editors. Journal of the History of Biology, 22(2):357– 359, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00139518.

Anonymous:1989:PA

[Ano89b] Anonymous. Publisher’s announcement. Journal of the History of Biology, 22 (1):iii, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209602.

Anonymous:1991:AST

[Ano91a] Anonymous. The acceptance of scientific theories and images of masculinity and femininity: 1959–1985. Journal of the History of Biology, 24 (1):19–49, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130473.

Anonymous:1991:CFR

[Ano91b] Anonymous. Can females rule the hive? The controversy over honey bee gender roles in British beekeeping texts of the sixteentheighteenth centuries. Journal of the History of Biology, 24 (1):113–144, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130476.

Anonymous:1991:DRE

[Ano91c]  Anonymous.  The decline

of recapitulationism in early twentieth-century biology: Disciplinary conflict and consensus on the battleground of theory. Journal of the History of Biology, 24(1): 51–89, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00130474.

Anonymous:1991:DAT

[Ano91d]  Anonymous.  The discov-

ery of adenosine triphosphate and the establishment of its structure. Journal of the History of Biology, 24(1): 145–154, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00130477.

Anonymous:1991:ELR

[Ano91e]  Anonymous.  Edinburgh

Lamarckians: Robert Jameson and Robert E. Grant. Journal of the History of Biology, 24(1):1–18, ???? 1991. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00130472.

Anonymous:1991:ERS

[Ano91f] Anonymous. Essay review: Science and politics in the nineteenth century. Journal of the History of Biology, 24 (1):155–161, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130478.

Anonymous:1991:JHB

[Ano91g] Anonymous. The J. H. B.

bookshelf. Journal of the History of Biology, 24(1): 163–170, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130479.

Anonymous:1991:RPE

[Ano91h]  Anonymous.  Reevaluating

progressive eugenics: Herbert Spencer Jennings and the 1924 immigration legislation. Journal of the History of Biology, 24(1):91–112, ???? 1991. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00130475.

Anonymous:1992:JHB

[Ano92] Anonymous. The J. H. B.

bookshelf. Journal of the History of Biology, 25(1): 157–173, Spring 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01947509.

Anonymous:1993:IN

[Ano93]  Anonymous.  Introductory

notes. Journal of the History of Biology, 26(3):385, Fall 1993. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01062054.

Anonymous:1996:JHB

[Ano96] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 29(3): 463–479, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127385.

Anonymous:1998:I

[Ano98a] Anonymous. Index. Journal of the History of Biology, 31 (3):463–475, September 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1017131922677;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1017131922677.
                                        pdf.

Anonymous:1998:JHB

[Ano98b] Anonymous. The J. H. B. bookshelf. Journal of the History of Biology, 31(2): 289–302, June 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1023/A:1017188804880;
http://link.springer.com/
content/pdf/10.1023/A:1017188804880.
pdf.

Anonymous:1999:BRa

[Ano99a]  Anonymous.  Book re-

views. Journal of the History of Biology, 32(1):217– 230, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017150727680;
http://link.springer.com/
content/pdf/10.1023/A:1017150727680.
pdf.

Anonymous:1999:BRb

[Ano99b]  Anonymous.  Book re-

views. Journal of the History of Biology, 32(2):395– 420, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017265501742;
http://link.springer.com/
content/pdf/10.1023/A:1017265501742.
pdf.

Anonymous:1999:BRc

[Ano99c]  Anonymous.  Book re-

views. Journal of the History of Biology, 32(3):557– 582, December 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1023/A:1017291029970;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1017291029970.
                                        pdf.

Anonymous:1999:I

[Ano99d] Anonymous. Index. Journal of the History of Biology, 32 (3):583–590, December 1999. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1017132302260;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1017132302260.
                                        pdf.

Anonymous:2000:BRa

[Ano00a]  Anonymous.  Book re-

views. Journal of the History of Biology, 33(1):189– 220, March 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1017277312797;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1017277312797.
                                        pdf.

Anonymous:2000:BRb

[Ano00b]  Anonymous.  Book re-

views. Journal of the History of Biology, 33(3):585– 614, December 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1017317607168;
                                        http://link.springer.com/

content/pdf/10.1023/A:1017317607168.
pdf.

Anonymous:2000:GMR

[Ano00c]  Anonymous.  Gregg Mit-

man, reel nature: America’s romance with wildlife on film. Journal of the History of Biology, 33(2):385– 424, September 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017257331362;
http://link.springer.com/
content/pdf/10.1023/A:1017257331362.
pdf.

Anonymous:2000:I

[Ano00d] Anonymous. Index. Journal of the History of Biology, 33 (3):615–624, December 2000. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017218628950;
http://link.springer.com/
content/pdf/10.1023/A:1017218628950.
pdf.

Anonymous:2001:BRa

[Ano01a] Anonymous. Book reviews.

Journal of the History of Biology, 34(2):395–421, June 2001. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017351907599;
http://link.springer.com/

content/pdf/10.1023/A:1017351907599.
pdf.

Anonymous:2001:BRb

[Ano01b]  Anonymous.  Book re-

views. Journal of the History of Biology, 34(3):593– 618, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017445823873;
http://link.springer.com/
content/pdf/10.1023/A:1017445823873.
pdf.

Anonymous:2001:I

[Ano01c] Anonymous. Index. Journal of the History of Biology, 34 (3):619–628, December 2001. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017333408565;
http://link.springer.com/
content/pdf/10.1023/A:1017333408565.
pdf.

Anonymous:2002:BRa

[Ano02a]  Anonymous.  Book re-

views. Journal of the History of Biology, 35(1):173– 206, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1017422704238;
http://link.springer.com/
content/pdf/10.1023/A:1017422704238.
pdf.

Anonymous:2002:BRb

[Ano02b] Anonymous. Book reviews.

Journal of the History of Biology, 35(2):385–420, June 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016029911083;
http://link.springer.com/
content/pdf/10.1023/A:1016029911083.
pdf.

Anonymous:2002:BRc

[Ano02c]  Anonymous.  Book re-

views. Journal of the History of Biology, 35(3):595– 623, September 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021110505387;
http://link.springer.com/
content/pdf/10.1023/A:1021110505387.
pdf.

Anonymous:2002:I

[Ano02d] Anonymous. Index. Journal of the History of Biology, 35 (3):625–634, September 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021157809908;
http://link.springer.com/
content/pdf/10.1023/A:1021157809908.
pdf.

Anonymous:2003:BRa

[Ano03a]  Anonymous.  Book re-

views. Journal of the History of Biology, 36(1):197– 224, March 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022585403854;
http://link.springer.com/
content/pdf/10.1023/A:1022585403854.
pdf.

Anonymous:2003:BRb

[Ano03b] Anonymous. Book reviews. Journal of the History of Biology, 36(2):405–431, June 2003. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1024418128316;
http://link.springer.com/
content/pdf/10.1023/A:1024418128316.
pdf.

Anonymous:2003:EN

[Ano03c] Anonymous. Editor’s note. Journal of the History of Biology, 36(3):435, September 2003. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004717.
44209.fd;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004717.
44209.fd.pdf.

Anonymous:2003:I

[Ano03d] Anonymous. Index. Journal of the History of Biology, 36 (3):631–640, September 2003.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004670.
32386.d8;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004670.
32386.d8.pdf.

Anonymous:2004:BRJ

[Ano04a] Anonymous. Book reviews:

Janet Browne, Charles Darwin: The Power of Place (New York: Knopf, 2002/ Princeton: Princeton University Press 2003) 591 pp., illus., $37.50, $22.95 paper. Journal of the History of Biology, 37 (2):387–418, June 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000038377.
32375.1c;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000038377.
32375.1c.pdf.

Anonymous:2004:EN

[Ano04b] Anonymous. Editors’ note.

Journal of the History of Biology, 37(1):1, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020449.
16057.20;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020449.
16057.20.pdf.

Anonymous:2005:BRC

[Ano05a] Anonymous. Book reviews: Carol Armstrong and Catherine de Zegher, eds., Ocean Flowers: Impressions from Nature (New York: The Drawing Center; Princeton: Princeton University Press, 2004), 336 pp., illus., $49.95. Journal of the History of Biology, 38(3):615– 648, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-1851-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-1851-2.pdf.

Anonymous:2005:EN

[Ano05b] Anonymous. Editors’ note. Journal of the History of Biology, 38(1):1, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-6505-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-6505-2.pdf.

Anonymous:2005:E

[Ano05c] Anonymous. Erratum. Journal of the History of Biology, 38(2):403–421, Summer 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-8043-y;

http://link.springer.com/
content/pdf/10.1007/s10739-
005-8043-y.pdf.

Anonymous:2005:I

[Ano05d] Anonymous. Index. Journal of the History of Biology, 38 (3):649–657, November 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4048-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
005-4048-9.pdf.

Anonymous:2005:PB

[Ano05e] Anonymous. Passing the baton. Journal of the History of Biology, 38(3):423– 424, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4047-x;
http://link.springer.com/
content/pdf/10.1007/s10739-
005-4047-x.pdf.

Anonymous:2006:BR

[Ano06a]  Anonymous.  Book re-

views. Journal of the History of Biology, 39(1):199– 234, Spring 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0001-9;
http://link.springer.com/

content/pdf/10.1007/s10739-
006-0001-9.pdf.

Anonymous:2006:BRK

[Ano06b] Anonymous. Book reviews: Keith Thomson, Before Darwin: Reconciling God and Nature (New Haven: Yale University Press, 2005), xiv + 314 pp., illus., $27.00. Journal of the History of Biology, 39 (2):407–424, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0009-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-0009-1.pdf.

Anonymous:2018:EMP

[Ano18] Anonymous. 2017 Everett Mendelsohn Prize. Journal of the History of Biology, 51 (1):5–6, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9504-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9504-9.pdf.

Anonymous:2019:EMP

[Ano19] Anonymous. 2019 Everett Mendelsohn Prize. Journal of the History of Biology, 52 (1):1–2, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-019-9554-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9554-2.pdf.

Anonymous:2020:CRO

[Ano20] Anonymous. Correction to:

Retrospective Open Access Articles. Journal of the History of Biology, 53(1):209– 210, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09582-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-09582-5.pdf.

Appel:1980:HBA

[App80] Toby A. Appel. Henri de Blainville and the animal series: a nineteenth-century chain of being. Journal of the History of Biology, 13(2): 291–319, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125745.

Appel:1988:JWP

[App88]  Toby A. Appel.  Jef-

fries Wyman, philosophical anatomy, and the scientific reception of Darwin in America. Journal of the History of Biology, 21(1):69–94, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125794.

Archibald:2009:EHP

[Arc09] J. David Archibald. Edward Hitchcock’s pre–Darwinian (1840) “tree of life”. Journal of the History of Biology, 42(3):561–592, August 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9163-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9163-y.pdf.

Armon:2012:BBE

[Arm12] Rony Armon. Between biochemists and embryologists — the biochemical study of embryonic induction in the 1930s. Journal of the History of Biology, 45(1):65– 108, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9266-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9266-0.pdf.

Arminjon:2016:BAM

[Arm16] Mathieu Arminjon. Birth of the allostatic model: From Cannon’s biocracy to critical physiology. Journal of the History of Biology, 49(2): 397–423, April 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9420-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9420-9.pdf.

Arnet:2019:CLM

[Arn19] Evan Arnet. Conwy Lloyd Morgan, methodology, and the origins of comparative psychology. Journal of the History of Biology, 52(3):433– 461, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09577-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-09577-2.pdf.

Arteaga:2017:BDH

[Art17] Juanma Sánchez Arteaga. Biological discourses on human races and scientific racism in Brazil (1832–1911). Journal of the History of Biology, 50 (2):267–314, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9445-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
016-9445-8.pdf.

Arteaga:2024:LMP

[Art24] Juanma Sánchez Arteaga.

“The logic of monsters:” Pere Alberch and the evolutionary significance of experimental teratology. Journal of the History of Biology, 57(3):379– 401, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09783-7.

Allen:1993:JHB

[ASR^{+}93] Garland E. Allen, V. B. Smocovitis, Ronald Rainger, Lynn K. Nyhart, Keith R. Benson, Peter G. Sobol, and Angela Creager. The J. H. B. bookshelf. Journal of the History of Biology, 26(1): 147–163, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01060684.

Alvarez-Tostado:2024:SGD

[ÁTBHBJB24] Miriam Álvarez-Tostado, Alfredo Bueno-Hernández, Ana Barahona, and Fabiola Juárez-Barrera. The study of geographical distribution in the analysis of domestication as an evolutionary process: Tensions in Alphonse de Candolle’s approach. Journal of the History of Biology, 57 (3):445–475, September 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10739-024-09785-5.

Atkinson:1985:GCE

[Atk85] J. W. Atkinson. E. G. Conklin on evolution: The popular writings of an embryologist. Journal of the History of Biology, 18(1):31–50, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127956.

Atran:1987:OSG

[Atr87] Scott Atran. Origin of the species and genus concepts: An anthropological perspective. Journal of the History of Biology, 20(2):195– 279, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138437.

Bird:1981:JHB

[BA81] Randall D. Bird and Garland Allen. The J. H. B. Archive report: the papers of Harry Hamilton Laughlin, eugenicist. Journal of the History of Biology, 14(2): 339–353, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141098.

Bechtel:2007:SMM

[BA07] William Bechtel and Adele Abrahamsen. In search of mitochondrial mechanisms: Interfield excursions between cell biology and biochemistry. Journal of the History of Biology, 40(1):1–33, March 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9103-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-9103-7.pdf.

Baedke:2019:OWA

[Bae19a] Jan Baedke. O organism, where art thou? Old and new challenges for organism– centered biology. Journal of the History of Biology, 52(2): 293–324, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9549-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9549-4.pdf. See pub-

lisher correction [Bae19b].

Baedke:2019:PCO

[Bae19b] Jan Baedke. Publisher correction to: O Organism, Where Art Thou? Old and New Challenges for Organism–Centered Biology. Journal of the History of Biology, 52(4): 747, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09581-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-09581-6.pdf.  See

[Bae19a].

Bajema:1988:CDM

[Baj88] Carl J. Bajema. Charles Darwin on man in the first edition of the Origin of Species. Journal of the History of Biology, 21(3):403–410, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144088. See reply

[Bow89].

Bangham:2022:SCE

[Ban22] Jenny Bangham. Book review: Soraya de Chadarevian, Heredity Under the Microscope: Chromosomes and the Study of the Human Genome (Chicago: University of Chicago Press, 2020), 272 pp, 36 halftones, $112.50 Cloth, ISBN 978-0-226-68508-3. Journal of the History of Biology, 55(3):609– 611, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09690-9.

Bartley:1992:DDS

[Bar92] Mary M. Bartley. Darwin and domestication: Studies on inheritance. Journal of the History of Biology, 25(2): 307–333, ???? 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00162844.

Bartley:1995:CCP

[Bar95] Mary M. Bartley. Courtship and continued progress: Julian Huxley’s studies on bird behavior. Journal of the History of Biology, 28(1): 91–108, Spring 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061247.

Barrow:1997:DEA

[Bar97] Mark V. Barrow, Jr. David Elliston Allen, the naturalist in Britain: a social history. Journal of the History of Biology, 30(3):493– 494, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004252007958;
http://link.springer.com/
content/pdf/10.1023/A:1004252007958.
pdf.

Barrow:2000:SDE

[Bar00] Mark V. Barrow, Jr. The specimen dealer: Entrepreneurial natural history in America’s Gilded Age. Journal of the History of Biology, 33(3):493– 534, December 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004856813466;
http://link.springer.com/
content/pdf/10.1023/A:1004856813466.
pdf.

Barrow:2004:BRW

[Bar04a] Mark V. Barrow, Jr. Book review: William Sargent, Crab Wars: a Tale of Horseshoe Crabs, Bioterrorism, and Human Health. Journal of the History of Biology, 37(3):608– 609, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2098-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2098-z.pdf.

Barrow:2004:BR

[Bar04b] Mark V. Barrow, Jr. Book reviews. Journal of the History of Biology, 37(1):181– 212, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020457.
97208.c0;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020457.
97208.c0.pdf.

Barrow:2005:BRS

[Bar05] Mark V. Barrow, Jr. Book reviews: Stephen Moss, A Bird in the Bush: a Social History of Birdwatching (London: Aurum Press, 2004), 375 pp., illus, £16.99. Journal of the History of Biology, 38 (2):392, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-4231-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-4231-z.pdf.

Barrow:2009:DDN

[Bar09] Mark V. Barrow, Jr. Dragons in distress: Naturalists as bioactivists in the campaign to save the American alligator. Journal of the History of Biology, 42(2):267–288, May 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9178-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9178-z.pdf.

Barrow:2010:BR

[Bar10a] Mark V. Barrow, Jr. Book review. Journal of the History of Biology, 43(1):195– 197, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1007/s10739-010-9223-y;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9223-y.pdf.

Barton:2010:BR

[Bar10b] Michael D. Barton. Book review. Journal of the History of Biology, 43(2): 403–405, May 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9232-x;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9232-x.pdf.

Baron:2011:WCC

[Bar11a] Christian Baron. A Web of controversies: Complexity in the Burgess shale debate. Journal of the History of Biology, 44(4):745– 780, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9248-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9248-2.pdf.

Barrow:2011:BR

[Bar11b] Mark V. Barrow, Jr. Book review. Journal of the History of Biology, 44(3): ??, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-011-9303-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9303-7.pdf.

Barrow:2019:BRM

[Bar19] Mark V. Barrow, Jr. Book review: Marilyn Bailey Ogilvie, For the Birds: American Ornithologist Margaret Morse Nice. Journal of the History of Biology, 52(3):489– 490, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09572-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-09572-7.pdf.

Barbosa:2023:BRP

[Bar23a] Thiago Pinto Barbosa. Book review: Projit Bihari Mukharji, Brown Skins, White Coats: Race Science in India, 1920– 66, Chicago: University of Chicago Press, 2022, ISBN: 0-226-82301-6, 348 pp. Journal of the History of Biology, 56(3):571–573, October 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09735-7.

Barrow:2023:REM

[Bar23b] Mark V. Barrow, Jr. Remembering Everett Mendelsohn, a kind and generous mentor.

Journal of the History of Biology, 56(4):??, December 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09747-3.

Barrow:2024:BRM

[Bar24] Mark V. Barrow, Jr. Book review: Mary Anne Andrei, Nature’s Mirror: How Taxidermists Shaped America’s Natural History Museums and Saved Endangered Species, Chicago: University of Chicago Press, 2020, ISBN: 978-0-226-73031-8, 250 pp. Journal of the History of Biology, 57(1):157– 159, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09762-y.

Baedke:2022:BWF

[BB22] Jan Baedke and Christina Brandt. Between the wars, facing a scientific crisis: The theoretical and methodological bottleneck of interwar biology. Journal of the History of Biology, 55(2):209– 217, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09689-2.

Barrow:1991:JHB

[BBF^{+}91] Mark V. Barrow, Jr., Keith R.

Benson, Paula Findlen, Michael Fortun, Shirley A. Roe, and Joel B. Hagen. The J. H. B. bookshelf. Journal of the History of Biology, 24(2): 339–351, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00209435.

Beatty:1982:WWC

[Bea82] John Beatty. What’s in a word? Coming to terms in the Darwinian revolution. Journal of the History of Biology, 15 (2):215–239, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00233015.

Beatty:1987:WRS

[Bea87] John Beatty. Weighing the risks: Stalemate in the classical/balance controversy. Journal of the History of Biology, 20(3):289–319, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00139457.

Beatty:1988:EEB

[Bea88] John Beatty. Ecology and evolutionary biology in the war and postwar years: Questions and comments. Journal of the History of Biology, 21(2): 245–263, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146988.

Beatty:1993:SCI

[Bea93] John Beatty. Scientific collaboration, internationalism, and diplomacy: The case of the atomic bomb casualty commission. Journal of the History of Biology, 26(2):205– 231, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061966.

Beatty:2016:CNS

[Bea16] John Beatty. The creativity of natural selection? Part I: Darwin, Darwinism, and the mutationists. Journal of the History of Biology, 49(4):659– 684, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9456-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9456-5.pdf.

Beatty:2019:CNS

[Bea19]  John Beatty.  The cre-

ativity of natural selection? Part II: The synthesis and since. Journal of the History of Biology, 52(4):705– 731, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09583-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09583-4.pdf.

Beckner:1969:FT

[Bec69] Morton Beckner. Function and teleology. Journal of the History of Biology, 2(1): 151–164, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137271.

Beck:2005:EFS

[Bec05] Naomi Beck. Enrico Ferri’s scientific socialism: a Marxist interpretation of Herbert Spencer’s organic analogy. Journal of the History of Biology, 38(2):301– 325, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-3796-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-3796-2.pdf.

Bechtel:2016:MMH

[Bec16] William Bechtel. Mechanists must be holists too! Perspectives from circadian biology. Journal of the History of Biology, 49(4):705– 731, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9439-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9439-6.pdf.

Beddall:1968:WDT

[Bed68] Barbara G. Beddall. Wallace, Darwin, and the theory of natural selection. Journal of the History of Biology, 1 (2):261–323, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351923.

Beddall:1972:WDE

[Bed72] Barbara G. Beddall. Wallace, Darwin, and Edward Blyth: Further notes on the development of evolution theory. Journal of the History of Biology, 5(1):153– 158, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113489.

Beddall:1973:NMD

[Bed73] Barbara G. Beddall. “Notes for Mr. Darwin”: Letters to Charles Darwin from Edward Blyth at Calcutta: a study in the process of discovery. Journal of the History of Biology, 6(1):69–95, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137299.

Beddall:1983:ERS

[Bed83a] Barbara G. Beddall. Essay review: Spanish science and the new world. Journal of the History of Biology, 16 (3):433–440, Fall 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00582410.

Beddall:1983:ISG

[Bed83b] Barbara G. Beddall. The isolated Spanish genius — myth or reality? Félix de Azara and the birds of Paraguay. Journal of the History of Biology, 16 (2):225–258, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00124699.

Beddall:1988:DDW

[Bed88a] Barbara G. Beddall. Darwin and divergence: The Wallace connection. Journal of the History of Biology, 21 (1):1–68, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125793.

Beddall:1988:WAC

[Bed88b] Barbara G. Beddall. Wallace’s annotated copy of Darwin’s Origin of Species. Journal of the History of Biology, 21(2): 265–289, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146989.

Beintema:2008:LMS

[Bei08] Jaap J. Beintema. The last month of Szent–Györgyi in Groningen. Journal of the History of Biology, 41(1): 159–165, March 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9135-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9135-7.pdf.

Bell:1981:GHL

[Bel81] Srilekha Bell. George Henry Lewes: a man of his time. Journal of the History of Biology, 14(2):277–298, ???? 1981. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141095.

Bellon:2001:JDH

[Bel01] Richard Bellon. Joseph Dalton Hooker’s ideals for a professional man of science. Journal of the History of Biology, 34(1):51–82, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010368323221;
http://link.springer.com/
content/pdf/10.1023/A:1010368323221.
pdf.

Bellon:2005:BRJ

[Bel05] Richard Bellon. Book reviews: John M. Lynch, ed., Creationism and Scriptural Geology, 1817–1857, Series on Evolution and Anti-Evolution: The Debates Before and After Darwin (Bristol: Thoemmes Press, 2002), 7 vols., 3171 pp., illus., $1100. Journal of the History of Biology, 38(2): 398–399, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4234-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4234-9.pdf.

Bellon:2006:JHT

[Bel06]  Richard Bellon.  Joseph

Hooker takes a “fixed post”: Transmutation and the “present unsatisfactory state of systematic botany”, 1844–1860. Journal of the History of Biology, 39(1):1–39, Spring 2006. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

1007/s10739-004-3800-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-3800-x.pdf.

Bellon:2011:BR

[Bel11] Richard Bellon. Book review.

Journal of the History of Biology, 44(3):??, August 2011. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9306-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9306-4.pdf.

Benson:1985:AML

[Ben85] Keith R. Benson. American morphology in the late nineteenth century: The biology department at Johns Hopkins University. Journal of the History of Biology, 18(2): 163–205, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00120109.

Benson:1988:RPN

[Ben88] Keith R. Benson. Review paper: The Naples Stazione Zoologica and its impact on the emergence of American marine biology. Journal of the History of Biology, 21(2): 331–341, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF00146991.

Benya:2009:BRS

[Ben09] Frazier Benya. Book review: Susan E. Lederer, Flesh and Blood: Organ Transplantation and Blood Transfusion in Twentieth-Century America (New York: Oxford University Press, 2008), ix + 224 pp., $35.00. Journal of the History of Biology, 42(4):829– 832, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9214-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9214-z.pdf.

Benson:2010:BR

[Ben10a] Keith R. Benson. Book review. Journal of the History of Biology, 43(4):805– 807, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9261-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9261-5.pdf.

Benya:2010:BR

[Ben10b] Frazier Benya. Book review. Journal of the History of Biology, 43(4):817– 819, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9264-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9264-2.pdf.

Bennett:2011:GHA

[Ben11a]  Brett M. Bennett.  A

global history of Australian trees. Journal of the History of Biology, 44(1):125– 145, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9243-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9243-7.pdf.

Benson:2011:DTP

[Ben11b] Etienne Benson. A difficult time with the permit process. Journal of the History of Biology, 44(1):103– 123, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9244-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9244-6.pdf.

Benya:2013:BR

[Ben13] Frazier Benya. Book review.

Journal of the History of Biology, 46(2):327–329, May 2013. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9358-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9358-8.pdf.

Benson:2018:BRN

[Ben18] Etienne S. Benson. Book review: Nancy J. Jacobs, Birders of Africa: History of a Network (New Haven: Yale University Press, 2016), 352 pp., 16 color + 49 b/w illus., $85.00 Cloth, ISBN: 978-0-300-20961-7. Journal of the History of Biology, 51(2): 391–394, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9519-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9519-x.pdf.

Berkeley:1970:HNL

[Ber70] Edmund Berkeley. The history of the naming of the Loblolly Bay. Journal of the History of Biology, 3(1): 149–154, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00569311.

Berry:2019:BRD

[Ber19] Dominic J. Berry. Book review: David P. D. Munns, Engineering the Environment: Phytotrons and the Quest for Climate Control in the Cold War (Pittsburgh: University of Pittsburgh Press, 2017), 360 pp., 38 b&w illus., $49.95 Hardcover, ISBN 978-0-8229-4474-4. Journal of the History of Biology, 52(1): 203–205, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9547-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9547-6.pdf.

Bezzola:1997:JJB

[Bez97] Ladina Bezzola. James J.

Bono, the word of God and the languages of man: Interpreting nature in modern science and medicine. Journal of the History of Biology, 30 (3):489–492, September 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004362830685;
http://link.springer.com/
content/pdf/10.1023/A:1004362830685.
pdf.

Baxter:1979:MM

[BF79] Alice Baxter and John Farley.

Mendel and meiosis. Journal of the History of Biology, 12 (1):137–173, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00128138.

Burian:1988:SFG

[BGZ88] Richard M. Burian, Jean Gayon, and Doris Zallen. The singular fate of genetics in the history of French biology, 1900–1940. Journal of the History of Biology, 21(3): 357–402, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00144087.

Blaiseau:2021:CDI

[BH21a] Pierre Louis Blaiseau and Allyson M. Holmes. Correction to: Diauxic inhibition: Jacques Monod’s ignored work. Journal of the History of Biology, 54 (2):197, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09644-7.

See [BH21b].

Blaiseau:2021:DIJ

[BH21b] Pierre Louis Blaiseau and Allyson M. Holmes. Diauxic inhibition: Jacques Monod’s ignored work. Journal of the History of Biology, 54(2): 175–196, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10739-021-09639-4.

See correction [BH21a].

Billings:1971:CNF

[Bil71] Susan M. Billings. Concepts of nerve fiber development, 1839–1930. Journal of the History of Biology, 4(2): 275–305, ???? 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138314.

Billiau:2016:CBR

[Bil16] Alfons Billiau. At the centennial of the bacteriophage: Reviving the overlooked contribution of a forgotten pioneer, Richard Bruynoghe (1881– 1957). Journal of the History of Biology, 49(3):559– 580, August 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9429-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9429-0.pdf.

Bishop:2023:BRL

[Bis23] Joseph Bishop. Book review: Lisa Haushofer, Wonder Foods: The Science and Commerce of Nutrition, Oakland: University of California Press, 2023, ISBN: 978-0-520-39040-9, 270 pp. Journal of the History of Biology, 56 (4):??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09756-2.

Bivins:2000:SCG

[Biv00] Roberta Bivins. Sex cells: Gender and the language of bacterial genetics. Journal of the History of Biology, 33(1): 113–139, March 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004779902860;
http://link.springer.com/
content/pdf/10.1023/A:1004779902860.
pdf.

Bizzo:1992:DMO

[Biz92] Nelio Marco Vincenzo Bizzo. Darwin on man in the origin of species: Further factors considered. Journal of the History of Biology, 25(1): 137–147, Spring 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01947507.

Brandon-Jones:1997:EBC

[BJ97] Christine Brandon-Jones. Edward Blyth, Charles Darwin, and the animal trade in Nineteenth-Century India and Britain. Journal of the History of Biology, 30(2): 145–178, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004209901090;
http://link.springer.com/
content/pdf/10.1023/A:1004209901090.
pdf.

Becsei-Kilborn:2010:SDS

[BK10] Eva Becsei-Kilborn. Scientific discovery and scientific reputation: The reception of Peyton Rous’ discovery of the chicken Sarcoma virus. Journal of the History of Biology, 43(1):111–157, February 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9171-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9171-y.pdf.

Brucke:2024:EO

[BL24] Ernst Brücke and Daniel Liu.

The elementary organisms. Journal of the History of Biology, 57(2):305–330, June 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09773-9.

Blaisdell:1982:NTN

[Bla82] Muriel Blaisdell. Natural theology and nature’s disguises. Journal of the History of Biology, 15(2):163–189, ???? 1982. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00233013.

Blackman:2007:NSD

[Bla07] Helen J. Blackman. The natural sciences and the development of animal morphology in Late-Victorian Cambridge. Journal of the History of Biology, 40(1):71– 108, March 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9102-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9102-8.pdf.

Bocking:1988:ASP

[Boc88] Stephen Bocking. Alpheus Spring Packard and cave fauna in the evolution debate. Journal of the History of Biology, 21(3):425–456, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00144090.

Bocking:1990:SFJ

[Boc90] Stephen Bocking. Stephen Forbes, Jacob Reighard, and the emergence of aquatic ecology in the Great Lakes region. Journal of the History of Biology, 23(3):461–498, ???? 1990. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136379.

Bocking:1995:EEA

[Boc95] Stephen Bocking. Ecosystems, ecologists, and the atom: Environmental research at Oak Ridge National Laboratory. Journal of the History of Biology, 28(1):1–47, Spring 1995. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061245.

Bocking:2012:SSS

[Boc12] Stephen Bocking. Science, Salmon, and sea lice: Constructing practice and place in an environmental controversy. Journal of the History of Biology, 45(4):681– 716, November 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9299-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9299-z.pdf.

Bodson:1983:ASR

[Bod83] Liliane Bodson. Aristotle’s statement on the reproduction of sharks. Journal of the History of Biology, 16 (3):391–407, Fall 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00582408.

Boddice:2023:BRA

[Bod23]  Rob Boddice.  Book re-

view: Anita Guerrini, Experimenting with Humans and Animals: From Aristotle to CRISPR, 2nd ed., Baltimore: Johns Hopkins University Press, 2022, ISBN: 978-1-4214-4405-5, 208 pp. Journal of the History of Biology, 56(2):403–405, August 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09731-x.

Bolman:2022:DLB

[Bol22a] Brad Bolman. Dogs for life: Beagles, drugs, and capital in the Twentieth Century. Journal of the History of Biology, 55(1):147–179, March 2022. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09649-2.

Bolman:2022:IWE

[Bol22b] Brad Bolman. Introduction: What Right? which Organisms? why Jobs? Journal of the History of Biology, 55 (1):3–13, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09676-7.

Bono:1984:ERF

[Bon84] James J. Bono. Essay review: The ferment of Van Helmont’s ideas. Journal of the History of Biology, 17(2): 291–294, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00143736.

Bono:1990:RLR

[Bon90] James J. Bono. Reform and the languages of Renaissance theoretical medicine: Harvey versus Fernel. Journal of the History of Biology, 23(3): 341–387, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136375.

Bonneuil:2006:MPB

[Bon06] Christophe Bonneuil. Mendelism, plant breeding and experimental cultures: Agriculture and the development of genetics in France. Journal of the History of Biology, 39 (2):281–308, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0005-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-0005-5.pdf.

Bootcov:2023:BRN

[Boo23] Michelle Bootcov. Book review: Neeraja Sankaran, A Tale of Two Viruses: Parallels in the Research Trajectories of Tumor and Bacterial Viruses, Pittsburgh: University of Pittsburgh Press, 2021, ISBN: 978-0-8229-4630-4, 312 pp. Journal of the History of Biology, 56(2): 415–417, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09727-7.

Borell:1976:OBP

[Bor76]  Merriley Borell.  Organ-

otherapy, British physiology, and discovery of the internal secretions. Journal of the History of Biology, 9(2): 235–268, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209884.

Borell:1985:OER

[Bor85] Merriley Borell. Organotherapy and the emergence of reproductive endocrinology. Journal of the History of Biology, 18(1):1–30, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127955.

Borell:1987:BPB

[Bor87] Merriley Borell. Biologists and the promotion of birth control research, 1918–1938. Journal of the History of Biology, 20(1):51–87, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125258.

Bornbusch:1989:LCC

[Bor89] Alan H. Bornbusch. Lacépède and Cuvier: a comparative case study of goals and methods in late eighteenth- and early nineteenth-century fish classification. Journal of the History of Biology, 22(1): 141–161, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209606.

Borrello:2003:SSW

[Bor03] Mark E. Borrello. Synthesis and selection: Wynne-Edwards’ challenge to David Lack. Journal of the History of Biology, 36(3):531– 566, September 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004569.
71224.65;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004569.
71224.65.pdf.

Borrello:2011:BR

[Bor11] Mark E. Borrello. Book review. Journal of the History of Biology, 44(2):??, May 2011. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9296-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9296-2.pdf.

Bowler:1971:PPE

[Bow71] Peter J. Bowler. Preformation and pre-existence in the seventeenth century: a brief analysis. Journal of the History of Biology, 4(2): 221–244, ???? 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138311.

Bowler:1973:BBT

[Bow73] Peter J. Bowler. Bonnet and Buffon: Theories of generation and the problem of species. Journal of the History of Biology, 6(2):259– 281, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127610.

Bowler:1977:DAD

[Bow77] Peter J. Bowler. Darwinism and the argument from design: Suggestions for a reevaluation.

Journal of the History of Biology, 10(1):29–43, ???? 1977. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126093.

Bowler:1989:DMO

[Bow89] Peter J. Bowler. Darwin on man in the Origin of Species: a reply to Carl Bajema. Journal of the History of Biology, 22(3):497–500, Fall 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00401579.  See

[Baj88].

Bowler:1994:ANG

[Bow94] Peter J. Bowler. Are the arthropoda a natural group? An episode in the history of evolutionary biology. Journal of the History of Biology, 27(2):177–213, Summer 1994. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062562.

Bowler:2005:RED

[Bow05] Peter J. Bowler. Revisiting the eclipse of Darwinism. Journal of the History of Biology, 38(1):19–32, ???? 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-6507-0;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-6507-0.pdf.

Bowler:2018:BR

[Bow18] Peter J. Bowler. Book review. Journal of the History of Biology, 51(2):399–401, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9511-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9511-5.pdf.

Boylan:1984:GHC

[Boy84] Michael Boylan. The Galenic and Hippocratic challenges to Aristotle’s conception theory. Journal of the History of Biology, 17(1):83– 112, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00397503.

Boylan:1986:GCT

[Boy86] Michael Boylan. Galen’s conception theory. Journal of the History of Biology, 19(1): 47–77, Spring 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00346617.

Boylan:2007:GBP

[Boy07] Michael Boylan. Galen: On blood, the pulse, and the arteries. Journal of the History of Biology, 40(2): 207–230, June 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9116-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-9116-2.pdf.

Berry:2019:LTO

[BP19] Dominic J. Berry and Paolo Palladino. Life, time, and the organism: Temporal registers in the construction of life forms. Journal of the History of Biology, 52(2): 223–243, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9513-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9513-3.pdf.

Barahona:2005:IIG

[BPA05] Ana Barahona, Susana Pinar, and Francisco J. Ayala. Introduction and institutionalization of genetics in Mexico. Journal of the History of Biology, 38(2):273– 299, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-3798-0;
 http://link.springer.com/

                                        content/pdf/10.1007/s10739-
                                        004-3798-0.pdf.

Burton-Rose:2020:TSI

[BR20] Daniel Burton-Rose. Towards a sinophone insect humanities: a review essay. Journal of the History of Biology, 53 (4):667–678, December 2020. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09624-3.

Bradley:2011:DSC

[Bra11a] Benjamin Sylvester Bradley. Darwin’s sublime: The contest between reason and imagination in On the Origin of Species. Journal of the History of Biology, 44(2): 205–232, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9210-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9210-3.pdf.

Braun:2011:AFF

[Bra11b] Robyn Braun. Accessory food factors: Understanding the catalytic function. Journal of the History of Biology, 44(3): 483–504, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9255-3;

http://link.springer.com/
content/pdf/10.1007/s10739-
010-9255-3.pdf.

Bradley:2020:BRJ

[Bra20]  Rory Bradley.  Book re-

view: Joan Steigerwald, Experimenting at the Boundaries of Life: Organic Vitality in Germany Around 1800 (University of Pittsburgh Press, 2019), 460 pp., $55.00 Cloth, ISBN: 978-0-8229-4553-3. Journal of the History of Biology, 53(1): 195–197, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-020-09595-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
020-09595-5.pdf.

Brandt:2022:DHI

[Bra22a] Christina Brandt. Development and heredity in the interwar period: Hans Spemann and Fritz Baltzer on organizers and merogones. Journal of the History of Biology, 55(2): 253–283, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09681-w.

Brandt:2022:VHM

[Bra22b]  Christina Brandt.  Vital-

ism, holism, and metaphorical dynamics of Hans Spemann’s “Organizer” in the interwar period. Journal of the History of Biology, 55(2): 285–320, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09682-9.

Branding:2024:RHF

[Bra24]  Jonah Branding.  Reca-

pitulation, heredity, and Freud’s view of human nature. Journal of the History of Biology, 57(3):403– 422, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09784-6.

Breidbach:2002:RMC

[Bre02] Olaf Breidbach. Representation of the microcosm — the claim for objectivity in 19th Century scientific microphotography. Journal of the History of Biology, 35(2): 221–250, June 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016044427910;
http://link.springer.com/
content/pdf/10.1023/A:1016044427910.
pdf.

Bretscher:2015:BR

[Bre15] Peter A. Bretscher. Book review. Journal of the History of Biology, 48(4):679– 680, November 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9422-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9422-7.pdf.

Brigandt:2005:ICE

[Bri05] Ingo Brigandt. “The instinct concept of the early Konrad Lorenz”. Journal of the History of Biology, 38(3):571– 608, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-6544-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-6544-3.pdf.

Brinkman:2009:BRD

[Bri09] Paul D. Brinkman. Book review: David Sepkoski and Michael Ruse, eds., The Paleobiological Revolution: Essays on the Growth of Modern Paleontology (Chicago and London: The University of Chicago Press, 2009), 584 pp., illus., $65.00. Journal of the History of Biology, 42(4):819– 821, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9211-2;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9211-2.pdf.

Brinkman:2010:CDB

[Bri10a] Paul D. Brinkman. Charles Darwin’s Beagle voyage, fossil vertebrate succession, and “the gradual birth & death of species”. Journal of the History of Biology, 43(2): 363–399, May 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9189-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9189-9.pdf. See erratum

[Bri10b].

Brinkman:2010:ECD

[Bri10b] Paul D. Brinkman. Erratum to: Charles Darwin’s Beagle Voyage, Fossil Vertebrate Succession, and “The Gradual Birth & Death of Species”. Journal of the History of Biology, 43(2):401, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9228-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9228-6.pdf. See [Bri10a].

Brinkman:2017:BR

[Bri17] Paul D. Brinkman. Book review. Journal of the History of Biology, 50(3):685– 687, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9484-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9484-9.pdf.

Brozek:1971:UCA

[Bro71] Josef Brozek. USSR: Current activities in the history of physiology and psychology. Journal of the History of Biology, 4(1):185– 208, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356982.

Brozek:1973:SRA

[Bro73] Josef Brozek. Six recent additions to the history of physiology in the USSR. Journal of the History of Biology, 6 (2):317–334, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127612.

Brown:1974:MVE

[Bro74] Theodore M. Brown. From mechanism to vitalism in eighteenth-century English physiology. Journal of the History of Biology, 7(2): 179–216, Fall 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00351203.

Browne:1980:DBA

[Bro80]  Janet Browne.  Darwin’s

botanical arithmetic and the “principle of divergence,” 1854–1858. Journal of the History of Biology, 13(1): 53–89, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125354.

Browne:1982:ERN

[Bro82] Janet Browne. Essay review: New developments in Darwin studies? Journal of the History of Biology, 15(2): 275–280, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00233018.

Brown:1986:EDT

[Bro86] Frank Burch Brown. The evolution of Darwin’s theism. Journal of the History of Biology, 19(1):1–45, Spring 1986. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00346616.

Browne:1988:ERP

[Bro88] Janet Browne. Essay review: Passports to success. Journal of the History of Biology, 21 (2):343–349, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146992.

Brown:2011:REE

[Bro11] Karen Brown. Rabid epidemiologies: The emergence and resurgence of rabies in Twentieth Century South Africa. Journal of the History of Biology, 44(1):81– 101, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9241-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9241-9.pdf.

Browne:2017:BR

[Bro17] Janet Browne. Book review. Journal of the History of Biology, 50(4):993– 995, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9497-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
017-9497-4.pdf.

Browne:2022:RDH

[Bro22] Janet Browne. Reflections on Darwin historiography. Journal of the History of Biology, 55(2):381–393, August 2022. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09686-5.

Browne:2023:EMH

[Bro23]  Janet Browne.  Everett

Mendelsohn, the Harvard colleague. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09753-5.

Browne:2024:BRI

[Bro24] Janet Browne. [Book review:] Ian Hesketh, ed., Imagining the Darwinian Revolution: Historical Narratives of Evolution from the Nineteenth Century to the Present, Pittsburgh: University of Pittsburgh Press, 2022, ISBN: 978-0-8229-4708-0, 352 pp. Journal of the History of Biology, 57(2):335–337, June 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09777-5.

Brush:2002:HTB

[Bru02] Stephen G. Brush. How theories became knowledge: Morgan’s chromosome theory of heredity in America and Britain. Journal of the History of Biology, 35(3):471– 535, September 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021175231599;
http://link.springer.com/
content/pdf/10.1023/A:1021175231599.
pdf.

Bruno:2018:BRL

[Bru18] Andy Bruno. Book review:

Loren Graham, Lysenko’s Ghost: Epigenetics and Russia (Cambridge, MA: Harvard University Press, 2016), 224 pp., illus., bibl., $24.95 Hardback, ISBN 978-0-674-08905-1. Journal of the History of Biology, 51(4):883– 885, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9542-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9542-y.pdf.

Braund:2008:CHW

[BS08] James Braund and Douglas G.

Sutton. The case of Heinrich Wilhelm Poll (1877–1939): a German–Jewish geneticist, eugenicist, twin researcher, and victim of the Nazis. Journal of the History of Biology, 41(1):1–35, March 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-007-9122-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9122-z.pdf.

Buchanan:2017:DMA

[Buc17] Roderick D. Buchanan. Darwin’s “Mr. Arthrobalanus”: Sexual differentiation, evolutionary destiny and the expert eye of the beholder. Journal of the History of Biology, 50 (2):315–355, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9444-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9444-9.pdf.

Buhs:2000:BBW

[Buh00]  Joshua Buhs.  Building

on bedrock: William Steel Creighton and the reformation of ant systematics, 1925–1970. Journal of the History of Biology, 33(1):27–70, March 2000. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004712911037;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004712911037.
                                        pdf.

Bulmer:1999:DFG

[Bul99] Michael Bulmer. The development of Francis Galton’s ideas on the mechanism of heredity. Journal of the History of Biology, 32(2):263– 292, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004608217247;
http://link.springer.com/
content/pdf/10.1023/A:1004608217247.
pdf.

Burkhardt:1970:LEP

[Bur70] Richard W. Burkhardt, Jr.

Lamarck, evolution, and the politics of science. Journal of the History of Biology, 3 (2):275–298, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137355.

Burkhardt:1972:ILB

[Bur72a] Richard W. Burkhardt, Jr.

The inspiration of Lamarck’s belief in evolution. Journal of the History of Biology, 5 (2):413–438, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346666.

Burnham:1972:ITG

[Bur72b] John C. Burnham. Instinct theory and the German reaction to Weismannism. Journal of the History of Biology, 5(2):321–326, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00346662.

Burian:1990:ERT

[Bur90] Richard M. Burian. Essay review: Toward a new philosophy of biology. Journal of the History of Biology, 23(2): 321–328, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00141474.

Burian:1993:HCE

[Bur93a] Richard M. Burian. How the choice of experimental organism matters: Epistemological reflections on an aspect of biological practice. Journal of the History of Biology, 26(2): 351–367, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01061974.

Burian:1993:TTD

[Bur93b] Richard M. Burian. Technique, task definition, and the transition from genetics to molecular genetics: Aspects of the work on protein synthesis in the laboratories of J. Monod and P. Zamecnik. Journal of the History of Biology, 26(3):387–407, Fall 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF01062055.

Burian:1996:TDB

[Bur96] Richard M. Burian. “The tools of the discipline: Biochemists and molecular biologists”: a comment. Journal of the History of Biology, 29 (3):451–462, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127384.

Burkhardt:1999:ENH

[Bur99] Richard W. Burkhardt, Jr.

Ethology, natural history, the life sciences, and the problem of place. Journal of the History of Biology, 32(3):489– 508, December 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004780803794;
http://link.springer.com/
content/pdf/10.1023/A:1004780803794.
pdf.

Burian:2014:BR

[Bur14] Richard M. Burian. Book review. Journal of the History of Biology, 47(2): 323–327, May 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9379-y;
 http://link.springer.com/

                                        content/pdf/10.1007/s10739-
                                        014-9379-y.pdf.

Burkhardt:2016:NTM

[Bur16] Richard W. Burkhardt, Jr. Niko Tinbergen: a message in the archives. Journal of the History of Biology, 49(4): 685–703, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9450-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9450-y.pdf.

Burkhardt:2017:BR

[Bur17] Richard W. Burkhardt, Jr. Book review. Journal of the History of Biology, 50(1):227– 229, February 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9467-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9467-x.pdf.

Burton:2022:AAT

[Bur22] Elise K. Burton. Ashkenazi anxieties: A transnational social history of Jewish genetic admixture modeling, 1971–1986. Journal of the History of Biology, 55(3): 411–442, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10739-022-09693-6.

Burkhardt:2023:REM

[Bur23] Richard W. Burkhardt. Remembering Everett Mendelsohn. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09746-4.

Butler:1988:CPD

[But88] Stella V. F. Butler. Centers and peripheries: The development of British physiology, 1870–1914. Journal of the History of Biology, 21(3): 473–500, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144092.

Button:2018:JCE

[But18] Clare Button. James Cossar Ewart and the origins of the Animal Breeding Research Department in Edinburgh, 1895–1920. Journal of the History of Biology, 51 (3):445–477, September 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9500-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9500-0.pdf.

Bynum:1984:CLA

[Byn84] W. F. Bynum. Charles Lyell’s Antiquity of Man and its critics. Journal of the History of Biology, 17(2):153– 187, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00143731.

Cable:2017:BR

[Cab17]  Kele Cable.  Book re-

view. Journal of the History of Biology, 50(4):997– 999, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9495-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9495-6.pdf.

Cain:1999:ERP

[Cai99] Joe Cain. Essay review: Progress and its problems. Journal of the History of Biology, 32(1):197–204, March 1999. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1004687316894;
                                         http://link.springer.com/
                                        content/pdf/10.1023/A:1004687316894.
                                        pdf.

Cain:2002:COC

[Cai02] Joe Cain. Co-opting colleagues: Appropriating Dobzhansky’s 1936 lectures at Columbia. Journal of the History of Biology, 35(2):207–219, June 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016008821530;
http://link.springer.com/
content/pdf/10.1023/A:1016008821530.
pdf.

Cain:2009:RSP

[Cai09] Joe Cain. Rethinking the synthesis period in evolutionary studies. Journal of the History of Biology, 42(4):621– 648, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9206-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9206-z.pdf.

Campbell:1983:CDL

[Cam83] Margaret Campbell. The concepts of dormancy, latency, and dominance in nineteenthcentury biology. Journal of the History of Biology, 16 (3):409–431, Fall 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00582409.

Camardi:2001:ROM

[Cam01] Giovanni Camardi. Richard Owen, morphology and evolution. Journal of the History of Biology, 34(3):481– 515, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012946930695;
http://link.springer.com/
content/pdf/10.1023/A:1012946930695.
pdf.

Caniglia:2013:BRU

[Can13] Guido Caniglia. Book review: Ullica Segerstrale, Nature’s Oracle. The Life and Work of W. D. Hamilton (Oxford: Oxford University Press, 2013), viii + 441 pp., illus., $25. Journal of the History of Biology, 46(4):757– 759, November 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9368-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9368-6.pdf.

Caniglia:2015:USI

[Can15]  Guido Caniglia.  Under-

standing societies from inside the organisms. Leo Pardi’s work on social dominance in Polistes wasps (1937– 1952). Journal of the History of Biology, 48(3):455– 486, August 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-015-9401-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9401-z.pdf.

Caomhanach:2024:BRM

[Cao24] Nuala Caomhánach. Book review: Maura Flannery, In the Herbarium: The Hidden World of Collecting and Preserving Plants, New Haven: Yale University Press, 2023, ISBN: 978-0-300-24791-6, 335 pp. Journal of the History of Biology, 57(3):485– 488, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09787-3.

Carlson:1971:UFM

[Car71] Elof Axel Carlson. An unacknowledged founding of molecular biology: H. J. Muller’s contributions to gene theory, 1910–1936. Journal of the History of Biology, 4(1): 149–170, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356980.

Carlson:1974:DGT

[Car74]  Elof Axel Carlson.  The

Drosophila group: The transition from the Mendelian unit to the individual gene. Journal of the History of Biology, 7(1):31–48, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00179292.

Carruthers:2008:CWM

[Car08] Jane Carruthers. Conservation and wildlife management in South African national parks 1930s–1960s. Journal of the History of Biology, 41 (2):203–236, June 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9147-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9147-3.pdf.

Carson:2009:BR

[Car09]  John Carson.  Book re-

view. Journal of the History of Biology, 42(3):599– 604, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9196-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9196-x.pdf.

Castellani:1973:SBL

[Cas73] Carlo Castellani. Spermatozoan biology from Leeuwenhoek to Spallanzani. Journal of the History of Biology, 6 (1):37–68, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137298.

Cassata:2012:ICP

[Cas12] Francesco Cassata. The Italian Communist Party and the “Lysenko Affair” (1948– 1955). Journal of the History of Biology, 45(3):469– 498, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9286-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9286-4.pdf.

Cassata:2015:CSH

[Cas15]  Francesco Cassata.  “A

Cold Spring Harbor in Europe.” EURATOM, UN-ESCO and the foundation of EMBO. Journal of the History of Biology, 48(4):539– 573, November 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9408-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9408-5.pdf.

Caulkins:2019:BRG

[Cau19] Tamara Caulkins. Book review: Georges-Louis Leclerc, Le comte de Buffon, Translated by J. A. Zalasiewicz, Mateusz Zalasiewicz, and Anne–Sophie Milon, The Epochs of Nature (Chicago: University of Chicago Press, 2018), 288 40pp., 7 b&w illus., $45.00 Cloth, ISBN: 978-0-226-39543-2. Journal of the History of Biology, 52(4):743– 745, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09589-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-09589-y.pdf.

Carrubba:1982:EKF

[CB82] Robert W. Carrubba and John Z. Bowers. Engelbert Kaempfer’s first report of the torpedo fish of the Persian Gulf in the late seventeenth century. Journal of the History of Biology, 15(2): 263–274, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00233017.

Collins:1986:IBE

[CBM86] James P. Collins, John Beatty, and Jane Maienschein. Introduction: Between ecology and evolutionary biology. Journal of the History of Biology, 19 (2):169–180, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00138874.

Cermakova:2018:NON

[CC18] Lucie Cermáková and Jana Cerná. Naked in the Old and the New World: Differences and analogies in descriptions of European and American herbae nudae in the Sixteenth Century. Journal of the History of Biology, 51(1): 69–106, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9468-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9468-9.pdf.

Campos:2025:IRI

[CC25] Luis A. Campos and Francesco Cassata. Introduction: Revis(it)ing Asilomar. Journal of the History of Biology, 58 (1):5–7, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09817-8.

Cassata:2025:AAA

[CdC25] Francesco Cassata and Soraya de Chadarevian. Asilomar across the Atlantic: EMBO, EMBL, and the politics of scientific expertise. Journal of the History of Biology, 58 (1):67–93, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-025-09808-9.

Ceccarelli:2019:BSB

[Cec19] David Ceccarelli. Between social and biological heredity: Cope and Baldwin on evolution, inheritance, and mind. Journal of the History of Biology, 52(1):161– 194, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9522-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9522-2.pdf.

Ceccon:2024:GRD

[Cec24]  Andrea Ceccon.  “At a

glance:” the role of diagrammatic representations in eugenics appropriations of the “Infamous Juke Family”. Journal of the History of Biology, 57(1):51–87, March 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09755-3.

Hung:2017:PRM

[cH17] Kuang chi Hung. “Plants that remind me of home”: Collecting, plant geography, and a forgotten expedition in the Darwinian revolution. Journal of the History of Biology, 50(1):71–132, February 2017. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9431-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9431-6.pdf.

Charenko:2022:BWP

[Cha22] Melissa Charenko. Blowing in the wind: Pollen’s mobility as a challenge to measuring climate by proxy, 1916– 1939. Journal of the History of Biology, 55(3):465– 493, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09677-6.

Cooper:1997:JHB

[CHCC97] Alix Cooper, Elizabeth Hanson, Kathy J. Cooke, and Angela N. H. Creager. The J. H. B. bookshelf. Journal of the History of Biology, 30(1): 135–144, March 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004235032011;
http://link.springer.com/
content/pdf/10.1023/A:1004235032011.
pdf.

Chew:2009:MTH

[Che09] Matthew K. Chew. The monstering of tamarisk: How scientists made a plant into a problem. Journal of the History of Biology, 42(2): 231–266, May 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9181-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9181-4.pdf.

Cheung:2013:LLD

[Che13] Tobias Cheung. Limits of life and death: Legallois’s decapitation experiments. Journal of the History of Biology, 46 (2):283–313, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9335-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9335-7.pdf.

Chettiar:2014:BR

[Che14]  Teri Chettiar.  Book re-

view. Journal of the History of Biology, 47(1):181– 183, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9375-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9375-2.pdf.

Chiang:2009:LTD

[Chi09] Howard Hsueh-Hao Chiang.

The laboratory technology of discrete molecular separation: The historical development of gel electrophoresis and the material epistemology of biomolecular science, 1945–1970. Journal of the History of Biology, 42(3): 495–527, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9169-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9169-5.pdf.

Chiang:2018:BR

[Chi18] Howard Chiang. Book review.

Journal of the History of Biology, 51(2):395–397, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9515-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9515-1.pdf.

Churchill:1968:AWB

[Chu68] Frederick B. Churchill. August Weismann and a break from tradition. Journal of the History of Biology, 1(1): 91–112, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF00149777.

Churchill:1969:MTE

[Chu69] Frederick B. Churchill. From machine-theory to entelechy: Two studies in developmental teleology. Journal of the History of Biology, 2(1): 165–185, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137272.

Churchill:1970:HEI

[Chu70] Frederick B. Churchill. The history of embryology as intellectual history. Journal of the History of Biology, 3(1): 155–181, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00569312.

Churchill:1974:WJG

[Chu74] Frederick B. Churchill. William Johannsen and the genotype concept. Journal of the History of Biology, 7(1): 5–30, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00179291.

Churchill:1989:GMI

[Chu89a] Frederick B. Churchill. The guts of the matter. Infusoria from Ehrenberg to Bütschli: 1838–1876. Journal of the History of Biology, 22(2): 189–213, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139512.

Churchill:1989:ITH

[Chu89b] Frederick B. Churchill. Introduction: Toward the history of protozoology. Journal of the History of Biology, 22 (2):185–187, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139511.

Churchill:2010:AWE

[Chu10] Frederick B. Churchill. August Weismann embraces the protozoa. Journal of the History of Biology, 43(4):767– 800, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9230-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
010-9230-z.pdf.

Civello:2013:GIN

[Civ13] Andrea Civello. On the genesis of the idiotypic network theory. Journal of the History of Biology, 46(1):125– 158, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9346-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9346-4.pdf.

Cambrosio:1992:BFT

[CK92] Alberto Cambrosio and Peter Keating. Between fact and technique: The beginnings of hybridoma technology. Journal of the History of Biology, 25(2):175–230, ???? 1992. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00162840.

Cambrosio:1994:IIH

[CKT94] Alberto Cambrosio, Peter Keating, and Alfred I. Tauber. Introduction: Immunology as a historical object. Journal of the History of Biology, 27 (3):375–378, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058990.

Clause:1993:WRR

[Cla93] Bonnie Tocher Clause. The Wistar rat as a right choice: Establishing mammalian standards and the ideal of a standardized mammal. Journal of the History of Biology, 26(2): 329–349, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061973.

Cameron:2011:TEN

[CM11] Laura Cameron and David Matless. Translocal ecologies: The Norfolk Broads, the “natural,” and the International Phytogeographical Excursion, 1911. Journal of the History of Biology, 44(1): 15–41, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9245-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
010-9245-5.pdf.

Cobb:2025:CIC

[Cob25] Matthew Cobb. The curious incident of Crick in the night-time and other Asilomar enigmas. Journal of the History of Biology, 58 (1):21–47, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09805-y.

Cock:1973:WBM

[Coc73] A. G. Cock. William Bateson, Mendelism and biometry. Journal of the History of Biology, 6(1):1–36, ???? 1973. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137297.

Coen:2006:LPF

[Coe06] Deborah R. Coen. Living precisely in Fin-de-Siècle Vienna. Journal of the History of Biology, 39(3):493– 523, September 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9101-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9101-9.pdf.

Coggon:2002:QAD

[Cog02] Jennifer Coggon. Quinarianism after Darwin’s Origin: The circular system of William Hincks. Journal of the History of Biology, 35 (1):5–42, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014582710287;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1014582710287.
                                        pdf.

Coggon:2022:SFE

[Cog22]  Jennifer Coggon.  Sperm-

force: Naturphilosophie and George Newport’s quest to discover the secret of fertilization. Journal of the History of Biology, 55(4):615– 687, December 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09696-3.

Colp:1985:NCD

[Col85] Ralph Colp, Jr., M.D. Notes on Charles Darwin’s autobiography. Journal of the History of Biology, 18(3): 357–401, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138930.

Coleman:1986:EES

[Col86a] William Coleman. Evolution into ecology? The strategy of warming’s ecological plant geography. Journal of the History of Biology, 19(2): 181–196, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138875.

Collins:1986:EEU

[Col86b] James P. Collins. Evolutionary ecology and the use of natural selection in ecological theory. Journal of the History of Biology, 19(2): 257–288, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138879.

Colp:1998:IR

[Col98] Ralph Colp, Jr. To be an invalid, redux. Journal of the History of Biology, 31(2): 211–240, June 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004331503063;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004331503063.
                                        pdf.

Comfort:1999:RPC

[Com99] Nathaniel C. Comfort. “The real point is control”: The reception of Barbara Mc-Clintock’s controlling elements. Journal of the History of Biology, 32(1):133– 162, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004468625863;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004468625863.
                                        pdf.

Comfort:2011:WYS

[Com11] Nathaniel Comfort. When your sources talk back: Toward a multimodal approach to scientific biography. Journal of the History of Biology, 44(4):651– 669, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1007/s10739-011-9273-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9273-9.pdf.

Comfort:2012:BR

[Com12] Nathaniel Comfort. Book review. Journal of the History of Biology, 45(3):573– 574, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9329-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
012-9329-5.pdf.

Cooke:1990:DMO

[Coo90] Kathy J. Cooke. Darwin on man in the Origin of Species: An addendum to the Bajema– Bowler debate. Journal of the History of Biology, 23(3): 517–521, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136381.

Cooke:1998:LHN

[Coo98] Kathy J. Cooke. The limits of heredity: Nature and nurture in American eugenics before 1915. Journal of the History of Biology, 31(2): 263–278, June 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

1023/A:1004387603971;
http://link.springer.com/
content/pdf/10.1023/A:1004387603971.
pdf.

Cooke:2002:DDE

[Coo02] Kathy J. Cooke. Duty or dream? Edwin G. Conklin’s critique of eugenics and support for American individualism. Journal of the History of Biology, 35(2): 365–384, June 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016077829496;
http://link.springer.com/
content/pdf/10.1023/A:1016077829496.
pdf.

Cooper:2020:BRA

[Coo20] Austin R. Cooper. Book review: Abena Dove Osseo-Asare, Atomic Junction: Nuclear Power in Africa after Independence. Cambridge University Press, Cambridge, UK and New York, 2019, 296 pp., 28 b&w illus., $32.99 Paperback, ISBN 978-1-108-45737-8. Journal of the History of Biology, 53(2): 315–317, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09607-4.

Corcos:1971:FPG

[Cor71] Alain F. Corcos. Fontenelle and the problem of generation in the eighteenth century. Journal of the History of Biology, 4(2):363–372, ???? 1971. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138317.

Cornell:1983:CES

[Cor83] John F. Cornell. From creation to evlution: Sir William Dawson and the idea of design in the nineteenth century. Journal of the History of Biology, 16(1):137–170, ???? 1983. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00186678.

Cornell:1984:ATD

[Cor84] John F. Cornell. Analogy and technology in Darwin’s vision of nature. Journal of the History of Biology, 17(3): 303–344, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126367.

Cornell:1987:GML

[Cor87] John F. Cornell. God’s magnificent law: The bad influence of theistic metaphysics on Darwin’s estimation of natural selection. Journal of the History of Biology, 20(3): 381–412, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00139461.

Corliss:1989:PCB

[Cor89] John O. Corliss. The protozoon and the cell: a brief twentieth-century overview. Journal of the History of Biology, 22(2):307–323, ???? 1989. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00139516.

Corsi:2005:BDT

[Cor05] Pietro Corsi. Before Darwin: Transformist concepts in European natural history. Journal of the History of Biology, 38(1):67–83, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-6510-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-6510-5.pdf.

Corsi:2021:ELA

[Cor21]  Pietro Corsi.  Edinburgh

Lamarckians? The authorship of three anonymous papers (1826–1829). Journal of the History of Biology, 54 (3):345–374, September 2021. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09646-5.

Cosans:1997:GCR

[Cos97] Christopher E. Cosans. Galen’s critique of rationalist and empiricist anatomy. Journal of the History of Biology, 30 (1):35–54, March 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004266427468;
http://link.springer.com/
content/pdf/10.1023/A:1004266427468.
pdf.

Cowan:1972:FGC

[Cow72] Ruth Schwartz Cowan. Francis Galton’s contribution to genetics. Journal of the History of Biology, 5(2): 389–412, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346665.

Craver:2003:MMM

[Cra03] Carl F. Craver. The making of a memory mechanism. Journal of the History of Biology, 36(1):153–195, March 2003. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022596107834;
http://link.springer.com/
content/pdf/10.1023/A:1022596107834.
pdf.

Cittadino:1990:JHB

[CRBD90] Eugene Cittadino, Ronald Rainger, Kieth R. Benson, and Virginia P. Dawson. The J. H. B. bookshelf. Journal of the History of Biology, 23 (1):155–162, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00158158.

Creager:1996:WSD

[Cre96] Angela N. H. Creager. Wendell Stanley’s dream of a freestanding biochemistry department at the University of California, Berkeley. Journal of the History of Biology, 29(3): 331–360, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127379.

Creager:2003:ERB

[Cre03] Angela N. H. Creager. Essay review: Building biology across the Atlantic. Journal of the History of Biology, 36 (3):579–589, September 2003. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004574.
50758.19;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004574.
50758.19.pdf.

Creager:2006:NES

[Cre06] Angela N. H. Creager. Nuclear energy in the service of biomedicine: The U.S. Atomic Energy Commission’s radioisotope program, 1946– 1950. Journal of the History of Biology, 39(4):649– 684, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9108-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-9108-2.pdf.

Creager:2010:PPG

[Cre10a] Angela N. H. Creager. The paradox of the phage group: Essay review. Journal of the History of Biology, 43(1):183– 193, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9226-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
010-9226-8.pdf.

Creath:2010:RHS

[Cre10b] Richard Creath. The role of history in science. Journal of the History of Biology, 43 (2):207–214, May 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9208-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9208-x.pdf.

Creager:2019:BRM

[Cre19] Angela N. H. Creager. Book review: María Jesús Santesmases. The Circulation of Penicillin in Spain: Health, Wealth and Authority (London: Palgrave Macmillan, 2018), XI, 239 pp., 8 b/w 1 color illus., $99.99 Hardcover, ISBN: 978-3-319-69717-8. Journal of the History of Biology, 52(1):199– 201, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9533-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9533-z.pdf.

Creager:2022:MOU

[Cre22] Angela N. H. Creager. Model organisms unbound. Journal of the History of Biology, 55 (1):21–28, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09675-8.

Crow:1987:MDO

[Cro87]  James F. Crow.  Muller,

Dobzhansky, and overdominance. Journal of the History of Biology, 20(3):351– 380, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF00139460.

Crook:1989:PCM

[Cro89] D. P. Crook. Peter Chalmers Mitchell and antiwar evolutionism in Britain during the Great War. Journal of the History of Biology, 22(2): 325–356, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139517.

Crow:1990:SWP

[Cro90]  James F. Crow.  Sewall

Wright’s place in twentiethcentury biology. Journal of the History of Biology, 23 (1):57–89, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00158155.

Crowe:2005:BRS

[Cro05] Michael J. Crowe. Book reviews: Steven J. Dick and James E. Strick, The Living Universe: NASA and the Development of Astrobiology (New Brunswick, N.J.: Rutgers University Press, 2004), xiii + 308 pp., illus., $49.95. Journal of the History of Biology, 38(2):387, Summer 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4228-7;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-4228-7.pdf.

Crowe:2012:BR

[Cro12] Nathan Crowe. Book review. Journal of the History of Biology, 45(4):721– 724, November 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9340-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9340-x.pdf.

Crowe:2014:CCD

[Cro14] Nathan Crowe. Cancer, conflict, and the development of nuclear transplantation techniques. Journal of the History of Biology, 47(1):63– 105, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9359-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9359-7.pdf.

Crowe:2015:BR

[Cro15] Nathan Crowe. Book review. Journal of the History of Biology, 48(3):495– 497, August 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1007/s10739-015-9412-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9412-9.pdf.

Creager:2006:RAA

[CS06] Angela N. H. Creager and María Jesús Santesmases. Radiobiology in the atomic age: Changing research practices and policies in comparative perspective. Journal of the History of Biology, 39(4):637– 647, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9107-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9107-3.pdf.

Chernyak:1990:IIM

[CT90] Leon Chernyak and Alfred I.

Tauber. The idea of immunity: Metchnikoff’s metaphysics and science. Journal of the History of Biology, 23 (2):187–249, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141470.

Crist:1997:DHI

[CT97] Eileen Crist and Alfred I.

Tauber. Debating humoral immunity and epistemology: The rivalry of the immunochemists Jules Bordet and Paul Ehrlich. Journal of the History of Biology, 30(3):321– 356, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004269605715;
http://link.springer.com/
content/pdf/10.1023/A:1004269605715.
pdf.

Dannenfeldt:1985:EDC

[Dan85] Karl H. Dannenfeldt. Europe discovers civet cats and civet. Journal of the History of Biology, 18(3):403–431, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138931.

Dande:2021:CSA

[Dan21] Innocent Dande. The colonial state, African dog-owners, and the political economy of rabies vaccination campaigns in Southern Rhodesia in the 1950s and 1960s. Journal of the History of Biology, 54(4): 689–717, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09661-6.

Darden:1977:WBP

[Dar77]  Lindley Darden.  William

Bateson and the promise of Mendelism. Journal of the History of Biology, 10(1): 87–106, ???? 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126096.

Davis:1997:BRW

[Dav97a] Frederick R. Davis. Book review: William E. Davis, Jr., and Jerome A. Jackson, eds., Contributions to the History of North American Ornithology. Journal of the History of Biology, 30(3):488– 489, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004310813846;
http://link.springer.com/
content/pdf/10.1023/A:1004310813846.
pdf.

Davis:1997:PLF

[Dav97b] Frederick R. Davis. Paul Lawrence Farber, discovering birds: The emergence of ornithology as a scientific discipline. Journal of the History of Biology, 30(3):487– 488, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004358729776;
http://link.springer.com/
content/pdf/10.1023/A:1004358729776.
pdf.

Davis:2012:BR

[Dav12] Frederick R. Davis. Book review. Journal of the History of Biology, 45(4):719– 720, November 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9328-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9328-6.pdf.

Davis:2015:BR

[Dav15] Frederick R. Davis. Book review. Journal of the History of Biology, 48(3):491– 493, August 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9413-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9413-8.pdf.

deBont:2011:PPJ

[dB11] Raf de Bont. Poetry and precision: Johannes Thienemann, the bird observatory in Rossitten and civic ornithology, 1900–1930. Journal of the History of Biology, 44(2): 171–203, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9209-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9209-9.pdf.

Devlin:1996:PSG

[DC96] C. Leah Devlin and P. J.

Capelotti. Proximity to seacoast: G. W. Field and the marine laboratory at Point Judith Pond, Rhode Island, 1896–1900. Journal of the History of Biology, 29(2):251– 265, Summer 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00571084.

deChadarevian:2006:MRG

[dC06] Soraya de Chadarevian. Mice and the reactor: The “genetics experiment” in 1950s Britain. Journal of the History of Biology, 39(4):707– 735, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9110-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-9110-8.pdf.

deChadarevian:2014:BRH

[dC14] Soraya de Chadarevian. Book review: Hallam Stevens, Life Out of Sequence: A Data-Driven History of Bioinformatics (Chicago: Chicago University Press, 2013), 304 pp., 19halftones, 3 line drawings, $90 (cloth), $30 (paper). Journal of the History of Biology, 47(3):495– 496, August 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9389-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9389-9.pdf.

deChadarevian:2018:WTC

[dC18] Soraya de Chadarevian. Whose turn? Chromosome research and the study of the human genome. Journal of the History of Biology, 51(4):631– 655, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9486-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9486-7.pdf.

deCambiaire:2023:KUN

[dC23] Elisabeth de Cambiaire. “From the known to the unknown:” Nature’s diversity, Materia Medica, and analogy in 18th century botany, through the work of Tournefort, the Jussieu Brothers, and Linnaeus. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09741-9.

Diogo:2001:PNC

[DCS01] Maria Paula Diogo, Ana Carneiro, and Ana Simões. The Portuguese naturalist Correia da Serra (1751–1823) and his impact on early nineteenth-century botany. Journal of the History of Biology, 34(2):353–393, June 2001. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010350218005;
http://link.springer.com/
content/pdf/10.1023/A:1010350218005.
pdf.

deLimaNavarro:2022:COC

[dd22] Pedro de Lima Navarro and Cristina de Amorim Machado. Correction to: An Origin of Citations: Darwin’s Collaborators and Their Contributions to the Origin of Species. Journal of the History of Biology, 55(1):205– 206, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09640-x.

See [ddAM20].

deLimaNavarro:2020:OCD

[ddAM20] Pedro de Lima Navarro and Cristina de Amorim Machado. An origin of citations: Darwin’s collaborators and their contributions to the Origin of Species. Journal of the History of Biology, 53(1):

45–79, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09592-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        020-09592-8.pdf. See correc-

tion [dd22].

DeMarrais:1974:DEE

[De 74] Robert De Marrais. The double-edged effect of Sir Francis Galton: a search for the motives in the biometrician-Mendelian debate. Journal of the History of Biology, 7(1):141– 174, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00179297.

DeChadarevian:1996:SCI

[De 96] Soraya De Chadarevian. Sequences, conformation, information: Biochemists and molecular biologists in the 1950s. Journal of the History of Biology, 29(3):361– 386, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127380.

DeBont:2008:EMB

[De 08] Raf De Bont. Evolutionary morphology in Belgium. Journal of the History of Biology, 41(1):81–118, March 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9129-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9129-5.pdf.

DeRibera-Martin:2019:SSK

[De 19] Ignacio De Ribera-Martin.

Seed ( sperma ) and Kuêma in Aristotle’s generation of animals. Journal of the History of Biology, 52(1): 87–124, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9508-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
018-9508-0.pdf.

Debus:1970:HFI

[Deb70] Allen G. Debus. Harvey and Fludd: The irrational factor in the rational science of the seventeenth century. Journal of the History of Biology, 3 (1):81–105, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00569307.

Delaporte:1983:TOE

[Del83] François Delaporte. Theories of osteogenesis in the eighteenth century. Journal of the History of Biology, 16 (3):343–360, Fall 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00582406.

Delaporte:1997:RS

[Del97] François Delaporte. Romana’s sign. Journal of the History of Biology, 30(3):357– 366, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004221722554;
http://link.springer.com/
content/pdf/10.1023/A:1004221722554.
pdf.

Dentinger:2016:PIP

[Den16] Rachel Mason Dentinger. Patterns of infection and patterns of evolution: How a malaria parasite brought “monkeys and man” closer together in the 1960s. Journal of the History of Biology, 49(2): 359–395, April 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9421-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9421-8.pdf.

Desmond:1984:RGS

[Des84] Adrian Desmond. Robert E. Grant: The social predicament of a pre-Darwinian transmutationist. Journal of the History of Biology, 17(2): 189–223, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00143732.

Desmond:2001:RAP

[Des01] Adrian Desmond. Redefining the X axis: “professionals,” “amateurs” and the making of mid–Victorian biology

— a progress report. Journal of the History of Biology, 34 (1):3–50, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010346828270;
http://link.springer.com/
content/pdf/10.1023/A:1010346828270.
pdf.

Deslauriers:2005:BRR

[Des05] Marguerite Deslauriers. Book reviews: Robert Mayhew, The Female in Aristotle’s Biology: Reason or Rationalization (Chicago and London: The University of Chicago Press, 2004), xi + 136 pp., $28.00. Journal of the History of Biology, 38(2):400– 402, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4236-7;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-4236-7.pdf.

Dexter:1970:HAF

[Dex70] Ralph W. Dexter. Historical aspects of F. W. Putnam’s systematic studies on fishes. Journal of the History of Biology, 3(1):131– 135, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00569309.

DeChadarevian:1996:TDB

[DG96] Soraya De Chadarevian and Jean-Paul Gaudillière. The tools of the discipline: Biochemists and molecular biologists. Journal of the History of Biology, 29(3):327– 330, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127378.

diGregorio:2002:RNN

[dG02] Mario A. di Gregorio. Reflections of a nonpolitical naturalist: Ernst Haeckel, Wilhelm Bleek, Friedrich Müller and the meaning of language. Journal of the History of Biology, 35(1):79– 109, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1023/A:1014526629425;
http://link.springer.com/
content/pdf/10.1023/A:1014526629425.
pdf.

DHombres:2012:DPL

[D’H12] Emmanuel D’Hombres. The ‘division of physiological labour’: The birth, life and death of a concept. Journal of the History of Biology, 45 (1):3–31, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9256-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
010-9256-2.pdf.

Dhein:2021:KFD

[Dhe21] Kelle Dhein. Karl von Frisch and the discipline of ethology. Journal of the History of Biology, 54(4):739– 767, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09660-7.

DiTrocchio:1991:MER

[Di 91] Federico Di Trocchio. Mendel’s experiments: a reinterpretation. Journal of the History of Biology, 24(3):485– 519, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF00156322.

DiGregorio:1995:WSC

[Di 95] Mario A. Di Gregorio. A wolf in sheep’s clothing: Carl Gegenbaur, Ernst Haeckel, the vertebral theory of the skull, and the survival of Richard Owen. Journal of the History of Biology, 28(2): 247–280, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01059190.

Dietrich:1994:ONT

[Die94] Michael R. Dietrich. The origins of the neutral theory of molecular evolution. Journal of the History of Biology, 27 (1):21–59, Spring 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058626.

Dietrich:1995:RGH

[Die95] Michael R. Dietrich. Richard Goldschmidt’s “heresies” and the evolutionary synthesis. Journal of the History of Biology, 28(3):431–461, Fall 1995. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01059388.

Dietrich:1998:PPN

[Die98] Michael R. Dietrich. Paradox and persuasion: Negotiating the place of molecular evolution within evolutionary biology. Journal of the History of Biology, 31(1): 85–111, March 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004257523100;
http://link.springer.com/
content/pdf/10.1023/A:1004257523100.
pdf.

Dietrich:2004:BRE

[Die04] Michael R. Dietrich. Book review: Elof Axel Carlson, Mendel’s Legacy: The Origin of Classical Genetics. Journal of the History of Biology, 37(3):590–591, October 2004. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2088-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2088-1.pdf.

Dietrich:2010:BR

[Die10] Michael Dietrich. Book review. Journal of the History of Biology, 43(3):621– 622, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9254-4;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9254-4.pdf.

Dietrich:2011:RRG

[Die11] Michael R. Dietrich. Reinventing Richard Goldschmidt: Reputation, memory, and biography. Journal of the History of Biology, 44(4):693– 712, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9271-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9271-y.pdf.

Dietrich:2012:EI

[Die12] Michael R. Dietrich. Editorial introduction. Journal of the History of Biology, 45 (1):1, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9318-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9318-0.pdf.

Dietrich:2016:IEM

[Die16a] Michael R. Dietrich. Inaugurating the Everett Mendelsohn Prize. Journal of the History of Biology, 49(1): 1–2, February 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10739-016-9435-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9435-x.pdf.

Dietrich:2016:IRG

[Die16b] Michael R. Dietrich. Introduction: Revisiting Garland Allen’s views on the history of the life sciences in the Twentieth Century. Journal of the History of Biology, 49(4):581– 582, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9457-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9457-4.pdf.

Dietrich:2017:JHB

[Die17a] Michael R. Dietrich. The Journal of the History of Biology at 50. Journal of the History of Biology, 50(1): 1–2, February 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9463-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9463-6.pdf.

Dietrich:2017:FEM

[Die17b] Michael R. Dietrich. The First Everett Mendelsohn Prize. Journal of the History of Biology, 50(1):3–4, February 2017.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9464-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9464-5.pdf.

Dietrich:2017:LTN

[Die17c] Michael R. Dietrich. Looking toward the next fifty years at the journal of the history of biology. Journal of the History of Biology, 50(4):689– 690, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9498-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9498-3.pdf.

deJong-Lambert:2012:LP

[dL12] William deJong Lambert. Lysenkoism in Poland. Journal of the History of Biology, 45(3):499–524, August 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9289-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9289-1.pdf.

deJong-Lambert:2012:LIL

[dLK12] William deJong Lambert and Nikolai Krementsov. On labels and issues: The Lysenko Controversy and the Cold War. Journal of the History of Biology, 45(3): 373–388, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9292-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9292-6.pdf.

Dann:1997:ERE

[DM97] Kevin Dann and Gregg Mitman. Essay review: Exploring the borders of environmental history and the history of ecology. Journal of the History of Biology, 30(2): 291–302, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004291821562;
http://link.springer.com/
content/pdf/10.1023/A:1004291821562.
pdf.

Penteado:2023:TEM

[dMP23] David Francisco de Moura Penteado. A tale of enduring myths: Buffon’s theory of animal degeneration and the regeneration of domesticated animals in mid-19th century Brazil. Journal of the History of Biology, 56 (4):??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10739-023-09742-8.

deRooy:2023:SLS

[dR23] Laurens de Rooy. The shelf life of skulls: Anthropology and ‘race’ in the Vrolik Craniological Collection. Journal of the History of Biology, 56(2): 309–337, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09716-w.

Dresow:2017:BR

[Dre17]  Max Dresow.  Book re-

view. Journal of the History of Biology, 50(3):673– 675, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9481-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9481-z.pdf.

Dresow:2020:RFM

[Dre20] Max Dresow. Re-forming morphology: Two attempts to rehabilitate the problem of form in the first half of the Twentieth Century. Journal of the History of Biology, 53 (2):231–248, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09603-8.

Dresow:2023:BSR

[Dre23] Max Dresow. Biased, spasmodic, and ridiculously incomplete: Sequence stratigraphy and the emergence of a new approach to stratigraphic complexity in paleobiology, 1973–1995. Journal of the History of Biology, 56(3): 419–454, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09720-0.

Druglitro:2023:NPP

[Dru23] Tone Druglitrø. Nonhuman primates in public health: Between biological standardization, conservation and care. Journal of the History of Biology, 56(3):455– 477, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09721-z.

Dietrich:2007:BBB

[DT07] Michael R. Dietrich and Brandi H. Tambasco. Beyond the boss and the boys: Women and the division of labor in Drosophila genetics in the United States, 1934– 1970. Journal of the History of Biology, 40(3):509– 528, September 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-007-9131-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9131-y.pdf.

Duancic:2020:LYH

[Dua20] Vedran Duancić. Lysenko in Yugoslavia, 1945–1950s: How to de-Stalinize Stalinist science. Journal of the History of Biology, 53(1): 159–194, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09598-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        020-09598-2.pdf.

deVos:2009:RAP

[dV09] John de Vos. Receiving an ancestor in the phylogenetic tree. Journal of the History of Biology, 42(2):361–379, May 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9165-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9165-9.pdf.

Eddy:2023:BSF

[Edd23] John H. Eddy. Buffon, species and the forces of reproduction. Journal of the History of Biology, 56(3):479– 493, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09722-y.

Edsall:1972:BHE

[Eds72]  John T. Edsall.  Blood

and hemoglobin: The evolution of knowledge of functional adaptation in a biochemical system. Journal of the History of Biology, 5 (2):205–257, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346659.

Edsall:1980:ERH

[Eds80] John T. Edsall. Essay review: Horace Judson and the molecular biologists. Journal of the History of Biology, 13 (1):141–158, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125357.

Egerton:1968:LFA

[Ege68a] Frank N. Egerton. Leeuwenhoek as a founder of animal demography. Journal of the History of Biology, 1 (1):1–22, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00149773.

Egerton:1968:SAP

[Ege68b] Frank N. Egerton. Studies of animal populations from Lamarck to Darwin. Journal of the History of Biology, 1 (2):225–259, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00351922.

Egerton:1969:RBU

[Ege69] Frank N. Egerton. Richard Bradley’s understanding of biological productivity: a study of eighteenth-century ecological ideas. Journal of the History of Biology, 2(2): 391–410, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125025.

Egerton:1970:HDP

[Ege70] Frank N. Egerton. Humboldt, Darwin, and population. Journal of the History of Biology, 3(2):325– 360, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137357.

Egerton:1983:ERW

[Ege83a] Frank N. Egerton. Essay review: a worldwide inventory of the history of ecology. Journal of the History of Biology, 16(1):171–175, ???? 1983. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00186679.

Egerton:1983:HEA

[Ege83b] Frank N. Egerton. The history of ecology: Achievements and opportunities, part one. Journal of the History of Biology, 16(2):259–310, ???? 1983. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00124700.

Egerton:1985:HEA

[Ege85] Frank N. Egerton. The history of ecology: Achievements and opportunities, Part two. Journal of the History of Biology, 18(1):103–143, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127959.

Eigen:1997:OFI

[Eig97] Edward A. Eigen. Overcoming first impresions: Georges Cuvier’s types. Journal of the History of Biology, 30(2): 179–209, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004227517928;
http://link.springer.com/

                                        content/pdf/10.1023/A:1004227517928.
                                        pdf.

Elias:2014:BHC

[Eli14] Scott A. Elias. A brief history of the changing occupations and demographics of coleopterists from the 18th through the 20th Century. Journal of the History of Biology, 47(2):213–242, May 2014. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9365-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9365-9.pdf.

Ellis:2007:PEE

[Ell07]  Erik Ellis.  Politics and

the environment: Essay review. Journal of the History of Biology, 40(4):761– 787, December 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9146-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9146-4.pdf.

Ellis:2009:BR

[Ell09] Erik Ellis. Book review. Journal of the History of Biology, 42(2):391–393, May 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1007/s10739-009-9191-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9191-2.pdf.

Ellis:2011:BR

[Ell11] Erik Ellis. Book review. Journal of the History of Biology, 44(4):791–792, November 2011. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9307-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9307-3.pdf.

Elwick:2007:SRE

[Elw07]  James Elwick.  Styles of

reasoning in early to mid-Victorian life research: Analysis:synthesis and palaetiology. Journal of the History of Biology, 40(1):35–69, March 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9106-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9106-4.pdf.

Endersby:2003:EDS

[End03] Jim Endersby. Escaping Darwin’s shadow. Journal of the History of Biology, 36(2): 385–403, June 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1023/A:1024442921798;
http://link.springer.com/
content/pdf/10.1023/A:1024442921798.
pdf.

Endersby:2011:LMO

[End11] Jim Endersby. A life more ordinary: The dull life but interesting times of Joseph Dalton Hooker. Journal of the History of Biology, 44(4):611– 631, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9270-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9270-z.pdf.

Endersby:2023:BRL

[End23]  Jim Endersby.  Book re-

view: Luke Keogh, The Wardian Case: How a Simple Box Moved Plants and Changed the World, Chicago: University of Chicago Press, ISBN: 978-0-226-71361-8, 288 pp. Journal of the History of Biology, 56(3):567– 569, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09736-6.

England:1997:NSB

[Eng97] Richard England. Natural selection before the origin: Public reactions of some naturalists to the Darwin–Wallace papers (Thomas Boyd, Arthur Hussey, and Henry Baker Tristram). Journal of the History of Biology, 30(2): 267–290, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004287720654;
http://link.springer.com/
content/pdf/10.1023/A:1004287720654.
pdf.

Ericson:2017:MSR

[Eri17] Kjell Ericson. Making space for red tide: Discolored water and the early Twentieth Century bayscape of Japanese pearl cultivation. Journal of the History of Biology, 50(2): 393–423, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9443-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9443-x.pdf.

Erlingsson:2002:HMA

[Erl02] Steindór J. Erlingsson. From Haeckelian monist to anti-Haeckelian vitalist: The transformation of the Icelandic naturalist Thorvaldur Thoroddsen (1855–1921). Journal of the History of Biology, 35(3):443–470, September 2002. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1021115027965;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1021115027965.
                                        pdf.

Erlingsson:2009:PLI

[Erl09] Steindór J. Erlingsson. The Plymouth Laboratory and the institutionalization of experimental zoology in Britain in the 1920s. Journal of the History of Biology, 42(1):151– 183, February 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9157-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9157-9.pdf.

Erlingsson:2016:ETL

[Erl16] Steindór J. Erlingsson. “Enfant terrible”: Lancelot Hogben’s life and work in the 1920s. Journal of the History of Biology, 49(3):495– 526, August 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9427-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9427-2.pdf.

Esposito:2013:WVM

[Esp13] Maurizio Esposito. Weismann versus Morgan revisited: Clashing interpretations on animal regeneration. Journal of the History of Biology, 46(3):511–541, August 2013. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9341-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9341-9.pdf.

Esposito:2020:BHL

[Esp20]  Maurizio Esposito.  Be-

yond Haeckel’s Law: Walter Garstang and the evolutionary biology that might have been. Journal of the History of Biology, 53(2): 249–268, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09602-9.

Evans:1984:DUA

[Eva84] L. T. Evans. Darwin’s use of the analogy between artificial and natural selection. Journal of the History of Biology, 17 (1):113–140, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00397504.

Fagan:2007:WDP

[Fag07] Melinda B. Fagan. Wallace, Darwin, and the practice of natural history. Journal of the History of Biology, 40(4): 601–635, December 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9126-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9126-8.pdf.

Fagan:2010:SSS

[Fag10] Melinda Bonnie Fagan. Stems and standards: Social interaction in the search for blood stem cells. Journal of the History of Biology, 43(1):67– 109, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9174-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9174-8.pdf.

Falk:1991:DTG

[Fal91] Raphael Falk. The dominance of traits in genetic analysis. Journal of the History of Biology, 24(3):457–484, ???? 1991. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00156321.

Falk:1995:SGI

[Fal95] Raphael Falk. The struggle of genetics for independence. Journal of the History of Biology, 28(2):219– 246, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059189.

Falk:2003:LPI

[Fal03] Raphael Falk. Linkage: From particulate to interactive genetics. Journal of the History of Biology, 36(1):87– 117, March 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022548108762;
http://link.springer.com/
content/pdf/10.1023/A:1022548108762.
pdf.

Farber:1972:BCS

[Far72a] Paul L. Farber. Buffon and the concept of species. Journal of the History of Biology, 5(2):259–284, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346660.

Farley:1972:SGCa

[Far72b] John Farley. The spontaneous generation controversy (1700–1860): The origin of parasitic worms. Journal of the History of Biology, 5(1): 95–125, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113487.

Farley:1972:SGCb

[Far72c] John Farley. The spontaneous generation controversy (1859–1880): British and German reactions to the problem of abiogenesis. Journal of the History of Biology, 5 (2):285–319, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346661.

Farley:1974:IRF

[Far74] John Farley. The initial reactions of French biologists to Darwin’s Origin of Species. Journal of the History of Biology, 7(2):275–300, Fall 1974. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351206.

Farber:2005:BRR

[Far05] Paul Lawrence Farber. Book reviews: Richard Weikart, From Darwin to Hitler: Evolutionary Ethics, Eugenics,and Racism in Germany (New York: Palgrave Macmillan, 2004), xi + 312 pp., $59.95. Journal of the History of Biology, 38(2):390– 391, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-005-4230-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4230-0.pdf.

Farber:2006:E

[Far06] Paul Farber. Editorial. Journal of the History of Biology, 39(2):235–236, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0002-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-0002-8.pdf.

Farber:2007:E

[Far07] Paul Lawrence Farber. Editorial. Journal of the History of Biology, 40(3): 387, September 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9133-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
007-9133-9.pdf.

Farber:2009:EDY

[Far09a] Paul Farber. Editorial: Darwin year. Journal of the History of Biology, 42(1): 1–2, February 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9172-x;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9172-x.pdf.

Farber:2009:JUT

[Far09b] Paul Farber. Journals under threat: a joint response from HSTM Editors. Journal of the History of Biology, 42(1): 185–187, February 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9162-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9162-z.pdf.

Farber:2011:FE

[Far11a] Paul Farber. Farewell editorial. Journal of the History of Biology, 44(4):603– 605, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9309-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9309-1.pdf.

Farber:2011:BR

[Far11b] Paul Lawrence Farber. Book review. Journal of the History of Biology, 44(3): ??, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9304-6;

http://link.springer.com/
content/pdf/10.1007/s10739-
011-9304-6.pdf.

Farber:2016:DMD

[Far16] Paul Lawrence Farber. Dobzhansky and Montagu’s debate on race: The aftermath. Journal of the History of Biology, 49 (4):625–639, December 2016. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9428-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9428-1.pdf.

Farber:2017:R

[Far17] Paul Lawrence Farber. Reflections. Journal of the History of Biology, 50(2): 231–233, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9470-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
017-9470-2.pdf.

Fattorini:2017:WFB

[Fat17]  Simone Fattorini.  The

Watson–Forbes biogeographical controversy untangled 170 years later. Journal of the History of Biology, 50(3): 473–496, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-016-9454-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9454-7.pdf.

Ferguson-Cradler:2021:OPN

[FC21] Gregory Ferguson-Cradler. The overfishing problem: Natural and social categories in early Twentieth-Century fisheries science. Journal of the History of Biology, 54(4):719– 738, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09655-4.

Friedman:2020:AMS

[FE20] William E. Friedman and Peter K. Endress. Alexander Moritzi, a Swiss pre-Darwinian evolutionist: Insights into the creationist– transmutationist debates of the 1830s and 1840s. Journal of the History of Biology, 53 (4):549–585, December 2020. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09619-0.

Frank:1978:JHBa

[FG78] Robert G. Frank, Jr. and Judith H. Goetzl. The J. H. B. Archive report: the Alexander Forbes papers. Journal of the History of Biology, 11 (2):387–393, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00389305.

Fitzpatrick:2017:BSA

[FG17] Simon Fitzpatrick and Grant Goodrich. Building a science of animal minds: Lloyd Morgan, experimentation, and Morgan’s canon. Journal of the History of Biology, 50(3): 525–569, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9451-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9451-x.pdf.

Frank:1977:JHB

[FH77] Robert C. Frank, Jr. and Diana Long Hall. The J. H. B. bookshelf. Journal of the History of Biology, 10 (2):365–371, Fall 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00572647.

Frank:1978:JHBb

[FH78] Robert G. Frank, Jr. and Diana L. Hall. The J. H. B. bookshelf. Journal of the History of Biology, 11 (2):395–400, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00389306.

Frank:1979:JHBb

[FH79] Robert G. Frank, Jr. and Diana Long Hall. The J. H. B. bookshelf. Journal of the History of Biology, 12(1): 203–210, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00128141.

Findlen:1991:JHB

[FHP^{+}91] Paula Findlen, Anne Harrington, Dorothy Porter, M. Susan Lindee, and Pnina G. Abir-Am. The J. H. B. bookshelf. Journal of the History of Biology, 24(3):537–548, ???? 1991. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00156325.

Fichman:1977:WZP

[Fic77] Martin Fichman. Wallace: Zoogeography and the problem of land bridges. Journal of the History of Biology, 10 (1):45–63, ???? 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00126094.

Fichman:2024:JCE

[Fic24] Martin Fichman. James T. Costa, Radical by Nature: The Revolutionary Life of Alfred Russel Wallace, Princeton: Princeton University Press, 2023, ISBN: 978-0-691-23379-6, 515 pp. Journal of the History of Biology, 57(4):603– 605, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09794-4.

Findlen:1991:ERG

[Fin91] Paula Findlen. Essay review:

Gender and the scientific “civilizing process”. Journal of the History of Biology, 24(2): 331–338, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209434.

Finnegan:2008:STG

[Fin08] Diarmid A. Finnegan. The spatial turn: Geographical approaches in the history of science. Journal of the History of Biology, 41(2): 369–388, June 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9136-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9136-6.pdf.

Fischedick:2000:SER

[Fis00] Kaat Schulte Fischedick.

From survey to ecology:

The role of the British vegetation committee, 1904– 1913. Journal of the History of Biology, 33(2):291– 314, September 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004829916824;
http://link.springer.com/
content/pdf/10.1023/A:1004829916824.
pdf.

Fisher:2010:BRD

[Fis10] Susie Fisher. Not beyond reasonable doubt: Howard Temin’s provirus hypothesis revisited. Journal of the History of Biology, 43(4):661– 696, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9202-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9202-3.pdf.

Fite:2019:BRC

[Fit19] Chris Fite. Book review: Courtney Fullilove, The Profit of the Earth: The Global Seeds of American Agriculture. Journal of the History of Biology, 52(3):485– 487, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09573-6;

http://link.springer.com/
content/pdf/10.1007/s10739-
019-09573-6.pdf.

Flannery:2024:BRJ

[Fla24] Maura C. Flannery. Book review: Joel Schwartz, Robert Brown and Mungo Park: Travels and Explorations in Natural History for the Royal Society, Memoirs of the New York Botanical Garden no. 122, Cham: Springer, 2021, ISBN: 978-3-030-74861-6, 217 pp. Journal of the History of Biology, 57(1):169– 171, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09763-x.

Fortun:1998:JHB

[FMR^{+}98] Michael Fortun, Mark Madison, Edmund Russell, Frederick R. Davis, Ann F. La Berge, and Sally G. Kohlstedt. The J. H. B. bookshelf. Journal of the History of Biology, 31 (1):143–154, March 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004213724009;
http://link.springer.com/
content/pdf/10.1023/A:1004213724009.
pdf.

Folkers:2021:DIR

[Fol21] Cynthia Folkers. Disproportionate impacts of radiation exposure on women, children, and pregnancy: Taking back our narrative. Journal of the History of Biology, 54 (1):31–66, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09630-z.

Forsdyke:2012:INS

[For12] Donald R. Forsdyke. Immunology (1955–1975): The natural selection theory, the two signal hypothesis and positive repertoire selection. Journal of the History of Biology, 45(1):139– 161, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9274-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9274-8.pdf.

Forgione:2020:ESF

[For20]  Fabio Forgione.  Evolu-

tion as a solution: Franco Andrea Bonelli, Lamarck, and the origin of man in early-Nineteenth-Century Italy. Journal of the History of Biology, 53(4):521– 548, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09617-2.

Frank:1972:HR

[Fra72] Robert G. Frank, Jr. Harvey redux. Journal of the History of Biology, 5(1):189– 204, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113491.

Frank:1979:JHBa

[Fra79] Robert G. Frank, Jr. The J. H.

B. archive report the Joseph Erlanger collection at Washington University School of Medicine, St. Louis. Journal of the History of Biology, 12 (1):193–201, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00128140.

Frampton:1991:ACM

[Fra91] Michael F. Frampton. Aristotle’s cardiocentric model of animal locomotion. Journal of the History of Biology, 24 (2):291–330, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209433.

Francis:2010:BR

[Fra10] Kevin Francis. Book review. Journal of the History of Biology, 43(1):203– 206, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9225-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
010-9225-9.pdf.

French:1970:DPM

[Fre70] Richard D. French. Darwin and the physiologists, or the Medusa and modern cardiology. Journal of the History of Biology, 3(2): 253–274, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137354.

Freund:2025:BRC

[Fre25] Daniel Freund. Book review: Christian Warren, Starved for Light: The long Shadow of Rickets and Vitamin D Deficiency, Chicago: University of Chicago Press, 2024, ISBN: 978-0-226-15193-9, 288 pp. Journal of the History of Biology, 58(1):133– 139, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09816-9.

Findlen:1995:JHB

[FRS^{+}95] Paula Findlen, Ronald Rainger, Vassiliki Betty Smocovitis, Richard W. Burkhardt, and Diane Paul. The J. H. B. bookshelf. Journal of the History of Biology, 28(2):369– 379, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059194.

Fry:1996:BSP

[Fry96] Iris Fry. On the biological significance of the properties of matter: L. J. Henderson’s theory of the fitness of the environment. Journal of the History of Biology, 29(2): 155–196, Summer 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00571081.

Fausto-Sterling:1990:ERP

[FS90] Anne Fausto-Sterling. Essay review: Primate visions, a model for historians of science? Journal of the History of Biology, 23(2):329– 333, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141475.

Funk:2012:RJG

[Fun12] Holger Funk. R. J. Gordon’s discovery of the spotted hyena’s extraordinary genitalia in 1777. Journal of the History of Biology, 45(2): 301–328, May 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9285-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9285-5.pdf.

Futuyma:1986:RRE

[Fut86] Douglas J. Futuyma. Reflections on reflections: Ecology and evolutionary biology. Journal of the History of Biology, 19(2):303–312, ???? 1986. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138881.

Gaissinovitch:1973:PVH

[Gai73] A. E. Gaissinovitch. Problems of variation and heredity in Russian biology in the late nineteenth century. Journal of the History of Biology, 6(1):97–123, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137300.

Gaissinovitch:1980:OSG

[Gai80] A. E. Gaissinovitch. The origins of Soviet genetics and the struggle with Lamarckism, 1922–1929. Journal of the History of Biology, 13 (1):1–51, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF00125353.

Galera:2017:ILT

[Gal17] Andrés Galera. The impact of Lamarck’s theory of evolution before Darwin’s theory. Journal of the History of Biology, 50(1):53–70, February 2017. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9432-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9432-5.pdf.

Galera:2021:EGS

[Gal21] Andrés Galera. Étienne Geoffroy Saint-Hilaire and the first embryological evolutionary model on the origin of vertebrates. Journal of the History of Biology, 54(2): 229–245, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09638-5.

Galka:2022:ELL

[Gal22] Jonathan M. Galka. Liguus landscapes: Amateur liggers, professional malacology, and the social lives of snail sciences. Journal of the History of Biology, 55(4):689– 723, December 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10739-022-09695-4.

Gamini:2021:CDD

[Gam21] Amir-Mohammad Gamini. A critique of Darwin’s The Descent of Man by a Muslim scholar in 1912: Muh.ammadrid.a is.fahānī’s examination of the anatomical and embryological similarities between human and other animals. Journal of the History of Biology, 54(3):485– 511, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09641-w.

Gascoigne:1991:JHB

[Gas91] Robert M. Gascoigne. Julian Huxley and biological progress. Journal of the History of Biology, 24(3): 433–455, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00156320.

Gaudilliere:1993:MBF

[Gau93] Jean-Paul Gaudillière. Molecular biology in the French tradition? Redefining local traditions and disciplinary patterns. Journal of the History of Biology, 26(3): 473–498, Fall 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF01062058.

Gaudilliere:1996:MBB

[Gau96] Jean-Paul Gaudillière. Molecular biologists, biochemists, and messenger RNA: The birth of a scientific network. Journal of the History of Biology, 29(3):417–445, ???? 1996. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127382.

Gaudilliere:1998:ERC

[Gau98] Jean-Paul Gaudillière. Essay review: Cancer and science: The hundred years war. Journal of the History of Biology, 31(2):279–288, June 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004339720810;
http://link.springer.com/
content/pdf/10.1023/A:1004339720810.
pdf.

Gaudilliere:2006:NPC

[Gau06] Jean-Paul Gaudillière. Normal pathways: Controlling isotopes and building biomedical research in postwar France. Journal of the History of Biology, 39(4):737– 764, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9111-7;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9111-7.pdf.

Gayon:1998:CIC

[Gay98] Jean Gayon. The concept of individuality in Canguilhem’s philosophy of biology. Journal of the History of Biology, 31 (3):305–325, September 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004312823993;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004312823993.
                                        pdf.

Gaziel:2012:QMA

[Gaz12] Ahuva Gaziel. Questions of methodology in Aristotle’s zoology: a medieval perspective. Journal of the History of Biology, 45(2):329–352, May 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9284-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9284-6.pdf.

Green:2018:SIE

[GD18] Eric D. Green and Christopher R. Donohue. Special Issue Editors’ introduction: “Genomics and the Human Genome Project”. Journal of the History of Biology, 51(4): 625–629, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9548-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9548-5.pdf.

Gerstner:1970:VPE

[Ger70] Patsy A. Gerstner. Vertebrate paleontology, an early nineteenth-century transatlantic science. Journal of the History of Biology, 3(1): 137–148, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00569310.

Getz:1991:BDS

[Get91] Faye Marie Getz. Black Death and the silver lining: Meaning, continuity, and revolutionary change in histories of medieval plague. Journal of the History of Biology, 24(2): 265–289, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209432.

Griesemer:1993:CMV

[GG93] James R. Griesemer and Elihu M. Gerson. Collaboration in the museum of vertebrate zoology. Journal of the History of Biology, 26(2): 185–203, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01061965.

Ghiselin:1997:EMA

[GG97] Michael T. Ghiselin and Christiane Groeben. Elias Metschnikoff, Anton Dohrn, and the metazoan common ancestor. Journal of the History of Biology, 30(2): 211–228, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004279501998;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004279501998.
                                        pdf.

Gayon:2019:MST

[GH19] Jean Gayon and Philippe Huneman. The modern synthesis: Theoretical or institutional event? Journal of the History of Biology, 52(4): 519–535, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09569-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09569-2.pdf.

Ghiselin:1973:MDC

[Ghi73] Michael T. Ghiselin. Mr. Darwin’s critics, old and new. Journal of the History of Biology, 6(1):155–165, ???? 1973.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137302.

Ghiselin:2005:DRV

[Ghi05] Michael T. Ghiselin. The Darwinian revolution as viewed by a philosophical biologist. Journal of the History of Biology, 38(1):123–136, ???? 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-6513-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-6513-2.pdf.

Giacomoni:1993:ODR

[Gia93] Dario Giacomoni. The origin of DNA:RNA hybridization. Journal of the History of Biology, 26(1):89– 107, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01060681.

Gianquitto:2018:BRE

[Gia18] Tina Gianquitto. Book review: Elizabeth Gilbert, The Signature of All Things: a Novel (Riverhead Books, 2014), 528 pp., $17.00 Paperback, ISBN: 978-0-14-312584-6. Journal of the History of Biology, 51(3):621– 624, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9534-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9534-y.pdf.

Gibson:2013:EWO

[Gib13] Abraham H. Gibson. Edward O. Wilson and the organicist tradition. Journal of the History of Biology, 46(4): 599–630, November 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9347-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9347-3.pdf.

Gibson:2018:BR

[Gib18] Abraham Gibson. Book review. Journal of the History of Biology, 51(1):183– 185, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9506-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9506-2.pdf.

Gibson:2019:BRB

[Gib19] Abraham Gibson. Book review: Biology in the Marvel cinematic universe. Journal of the History of Biology, 52 (2):365–369, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9564-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9564-0.pdf.

Gibson:2024:NKE

[Gib24] Abraham Gibson. Neal A.

Knapp, Making Machines of Animals: The International Livestock Exposition, Baltimore: Johns Hopkins University Press, 2023, ISBN: 978-1-4214-4655-4, 216 pp. Journal of the History of Biology, 57 (4):607–608, December 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09797-1.

Gilbert:1978:EOG

[Gil78] Scott F. Gilbert. The embryological origins of the gene theory. Journal of the History of Biology, 11(2): 307–351, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00389303.

Gillespie:1987:NHN

[Gil87] Neal C. Gillespie. Natural history, natural theology, and social order: John Ray and the “Newtonian ideology”. Journal of the History of Biology, 20(1):1–49, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125257.

Gillespie:1990:INT

[Gil90] Neal C. Gillespie. The interface of natural theology and science in the ethology of W. H. Thorpe. Journal of the History of Biology, 23 (1):1–38, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00158153.

Gingras:2010:RQD

[Gin10] Yves Gingras. Revisiting the “quiet debut” of the double helix: a bibliometric and methodological note on the “impact” of scientific publications. Journal of the History of Biology, 43(1):159– 181, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9183-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9183-2.pdf.

Gissis:2023:NLT

[Gis23] Snait B. Gissis. The neo-Lamarckian tools deployed by the young Durkheim: 1882– 1892. Journal of the History of Biology, 56(1):153– 190, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09708-w.

Glass:1969:EB

[Gla69]  Bentley Glass.  Explana-

tion in biology. Journal of the History of Biology, 2(1): 47–53, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137267.

Glass:1974:LNG

[Gla74] Bentley Glass. The long neglect of genetic discoveries and the criterion of prematurity. Journal of the History of Biology, 7(1):101– 110, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00179295.

Grene:1994:JHB

[GLB^{+}94] Marjorie Grene, Sherrie L.

Lyons, Mark V. Barrow, Jr., Ronald Rainger, Susan Lindee, Jane Maienschein, Michael Fortun, and Joel B. Hagen. The J. H. B. bookshelf. Journal of the History of Biology, 27(1):161– 175, Spring 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058630.

Glick:1976:DNS

[Gli76] Thomas F. Glick. On the diffusion of a new specialty: Marañón and the “crisis” of endocrinology in Spain. Journal of the History of Biology, 9(2):287–300, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00209886.

Gliboff:1998:ERR

[Gli98] Sander Gliboff. Evolution, revolution, and reform in Vienna: Franz Unger’s ideas on descent and their post-1848 reception. Journal of the History of Biology, 31(2): 179–209, June 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004379402154;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004379402154.
                                        pdf.

Gliboff:2004:BRT

[Gli04] Sander Gliboff. Book review: Thomas junker,die zweite darwinsche revolution. Journal of the History of Biology, 37(3): 601–602, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10739-004-2094-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
004-2094-3.pdf.

Gliboff:2006:CPK

[Gli06] Sander Gliboff. The case of Paul Kammerer: Evolution and experimentation in the early 20th Century. Journal of the History of Biology, 39 (3):525–563, September 2006. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-3051-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
005-3051-5.pdf.

Gliboff:2007:HGB

[Gli07] Sander Gliboff. H. G. Bronn and the history of nature. Journal of the History of Biology, 40(2):259–294, June 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9114-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9114-4.pdf.

Gliboff:2024:BRG

[Gli24] Sander Gliboff. Book review: Gregory Radick, Disputed Inheritance: The Battle over Mendel and the Future of Biology, Chicago: University of Chicago Press, 2023, ISBN: 978-0-226-82272-3, 630 pp. Journal of the History of Biology, 57(1):153– 156, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09761-z.

Guerrini:2013:IEA

[GM13] Anita Guerrini and Domenico Bertoloni Meli. Introduction: Experimenting with animals in the early modern era. Journal of the History of Biology, 46 (2):167–170, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9337-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9337-5.pdf.

Gamito-Marques:2018:SOO

[GM18] Daniel Gamito-Marques. A space of one’s own: Barbosa du Bocage, the foundation of the National Museum of Lisbon, and the construction of a career in zoology (1851–1907). Journal of the History of Biology, 51(2):223–257, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9487-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9487-6.pdf.

Gokyigit:1994:RFG

[Gök94] Emel Aileen Gökyiḡit. The reception of Francis Galton’s hereditary genius in the Victorian periodical press. Journal of the History of Biology, 27(2):215–240, Summer 1994. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062563.

Goodrum:2017:BR

[Goo17] Matthew R. Goodrum. Book review. Journal of the History of Biology, 50(2): 457–459, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9473-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9473-z.pdf.

Goodrum:2019:BRC

[Goo19] Matthew R. Goodrum. Book review: Christa Kuljian, Darwin’s Hunch: Science, Race and the Search for Human Origins (Auckland Park: Jacana Media, 2016), 1 + 352 pp., illus., $23.40 paperback, ISBN: 978-1-4314-2425-2. Journal of the History of Biology, 52(2): 357–358, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-019-9560-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-9560-4.pdf.

Gordon:1989:DPE

[Gor89] Scott Gordon. Darwin and political economy: The connection reconsidered. Journal of the History of Biology, 22 (3):437–459, Fall 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00401577.

Gordin:1997:ASS

[Gor97] Michael D. Gordin. The anthrax solution: The Sverdlovsk Incident and the resolution of a biological weapons controversy. Journal of the History of Biology, 30(3):441– 480, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004299907049;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004299907049.
                                        pdf.

Gormley:2009:SDA

[Gor09] Melinda Gormley. Scientific discrimination and the activist scientist: L. C. Dunn and the professionalization of genetics and human genetics in the United States. Journal of the History of Biology, 42(1):33–72, February 2009.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9170-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9170-z.pdf.

Gordin:2012:HLB

[Gor12] Michael D. Gordin. How Lysenkoism became pseudoscience: Dobzhansky to Velikovsky. Journal of the History of Biology, 45(3):443– 468, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9287-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9287-3.pdf.

Gormley:2019:BRR

[Gor19] Melinda Gormley. Book review: Robert E. Kohler, Inside Science: Stories from the Field in Human and Animal Science (Chicago: University of Chicago Press, 2019), 264 40pp., $35.00 Cloth, ISBN: 978-0-226-61798-5. Journal of the History of Biology, 52(4): 737–739, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09586-1;
 http://link.springer.com/

                                        content/pdf/10.1007/s10739-
                                        019-09586-1.pdf.

Goss:2023:DBI

[Gos23] Andrew Goss. Decolonizing botany: Indonesia, UN-ESCO, and the making of a global science. Journal of the History of Biology, 56(3): 495–523, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09734-8.

Gotthelf:1999:DA

[Got99] Allan Gotthelf. Darwin on Aristotle. Journal of the History of Biology, 32(1): 3–30, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004404206298;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004404206298.
                                        pdf.

Gould:1968:TOS

[Gou68] Stephen Jay Gould. Trigonia and the origin of species. Journal of the History of Biology, 1(1):41–56, ???? 1968. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00149775.

Gould:1970:DDL

[Gou70] Stephen Jay Gould. Dollo on Dollo’s law: Irreversibility and the status of evolutionary laws. Journal of the History of Biology, 3(2): 189–212, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137351.

Graham:1992:I

[Gra92] Loren Graham. Introduction.

Journal of the History of Biology, 25(3):363–368, Fall 1992. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351997.

Gramelsberger:2020:SMV

[Gra20] Gabriele Gramelsberger. Synthetic morphology: a vision of engineering biological form. Journal of the History of Biology, 53(2):295–309, June 2020. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09601-w.

Greene:1977:DSE

[Gre77] John C. Greene. Darwin as a social evolutionist. Journal of the History of Biology, 10(1):1–27, ???? 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126092.

Greene:1990:ISW

[Gre90] John C. Greene. The interaction of science and world view in Sir Julian Huxley’s evolutionary biology. Journal of the History of Biology, 23 (1):39–55, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00158154.

Greene:1992:ADR

[Gre92] John C. Greene. From Aristotle to Darwin: Reflections on Ernst Mayr’s interpretation in the growth of biological thought. Journal of the History of Biology, 25(2): 257–284, ???? 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00162842.

Greene:1994:SPM

[Gre94] John C. Greene. Science, philosophy, and metaphor in Ernst Mayr’s writings. Journal of the History of Biology, 27(2):311–347, Summer 1994. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01062565.

Greene:2004:BRM

[Gre04] John C. Greene. Book review: Martin Fichman, An Elusive Victorian: The Evolution of Alfred Russel Wallace. Journal of the History of Biology, 37(3):598–600, October 2004. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2093-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2093-4.pdf.

Greene:2007:WSB

[Gre07] Mott T. Greene. Writing scientific biography. Journal of the History of Biology, 40(4): 727–759, December 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9124-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
007-9124-x.pdf.

Gregory:2025:BRA

[Gre25] Frederick Gregory. Book review: Andreas Daum, Alexander Von Humboldt: a Concise Biography, Trans. Robert Savage, Princeton: Princeton University Press, 2024, ISBN: 978-0-691-24736-6, 208 pp. Journal of the History of Biology, 58(1):141– 143, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09815-w.

Grinnell:1974:RFD

[Gri74] George Grinnell. The rise and fall of Darwin’s first theory of transmutation. Journal of the History of Biology, 7(2):259–273, Fall 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00351205.

Grinnell:1985:RFD

[Gri85] George James Grinnell. The rise and fall of Darwin’s second theory. Journal of the History of Biology, 18(1): 51–70, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127957.

Grmek:1968:FSC

[Grm68] M. D. Grmek. First steps in Claude Bernard’s discovery of the glycogenic function of the liver. Journal of the History of Biology, 1(1): 141–154, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00149779.

Grobstein:1969:OLE

[Gro69] Clifford Grobstein. Organizational levels and explanation. Journal of the History of Biology, 2(1):199–206, ???? 1969.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137274.

Gross:1979:LLF

[Gro79] Michael Gross. The lessened locus of feelings: a transformation in French physiology in the early nineteenth century. Journal of the History of Biology, 12(2):231– 271, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00124193.

Gronim:2009:BR

[Gro09] Sara Stidstone Gronim. Book review. Journal of the History of Biology, 42(3):605– 607, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9197-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9197-9.pdf.

Grote:2013:PMM

[Gro13] Mathias Grote. Purple matter, membranes and ‘molecular pumps’ in rhodopsin research (1960s–1980s). Journal of the History of Biology, 46(3):331–368, August 2013. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9333-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9333-9.pdf.

Grodwohl:2017:NSA

[Gro17a] Jean-Baptiste Grodwohl. Natural selection, adaptive topographies and the problem of statistical inference: The Moraba scurra controversy under the microscope. Journal of the History of Biology, 50 (4):753–796, November 2017. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9448-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9448-5.pdf.

Grodwohl:2017:TWB

[Gro17b] Jean-Baptiste Grodwohl. “The theory was beautiful indeed”: Rise, fall and circulation of maximizing methods in population genetics (1930– 1980). Journal of the History of Biology, 50(3):571– 608, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9449-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9449-4.pdf.

Grodwohl:2019:ABP

[Gro19] Jean-Baptiste Grodwohl. Animal behavior, population biology and the modern synthesis (1955–1985). Journal of the History of Biology, 52 (4):597–633, December 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9553-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
018-9553-8.pdf.

Grunert:2018:BRL

[Gru18] Jonathan Grunert. Book review: Lance Grande, Curators: Behind the Scenes of Natural History Museums (Chicago: University of Chicago Press, 2017), 432 pp., 146 color plates, $35.00 Cloth ISBN: 978-0-226-19275-8. Journal of the History of Biology, 51(2): 403–405, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9517-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
018-9517-z.pdf.

Gruevska:2022:AIN

[Gru22] Julia Gruevska. Analysis and/ or interpretation in neurophysiology? A transatlantic discussion between F. J. J.

Buytendijk and K. S. Lashley, 1929–1932. Journal of the History of Biology, 55(2): 321–347, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09680-x.

Gerovitch:1992:ERR

[GS92] Vyacheslav Gerovitch and Anton Struchkov. Epilogue: Russian reflections. Journal of the History of Biology, 25 (3):487–495, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00352004.

Garcia-Sancho:2010:NIS

[GS10] Miguel García-Sancho. A new insight into Sanger’s development of sequencing: From proteins to DNA, 1943–1977. Journal of the History of Biology, 43(2):265–323, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9184-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9184-1.pdf.

Goldstein:2015:JVN

[GS15] Donna M. Goldstein and Magdalena E. Stawkowski. James V. Neel and Yuri E. Dubrova: Cold War debates and the genetic effects of low-dose radiation. Journal of the History of Biology, 48(1):67– 98, February 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9385-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
014-9385-0.pdf.

Goodfield-Toulmin:1969:SAE

[GT69] June Goodfield-Toulmin. Some aspects of English physiology: 1780–1840. Journal of the History of Biology, 2(2): 283–320, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125021.

Guerrini:1985:JKG

[Gue85] Anita Guerrini. James Keill, George Cheyne, and Newtonian physiology, 1690–1740. Journal of the History of Biology, 18(2):247–266, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00120111.

Guerrini:2013:ECU

[Gue13] Anita Guerrini. Experiments, causation, and the uses of vivisection in the first half of the seventeenth century. Journal of the History of Biology, 46 (2):227–254, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9319-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9319-7.pdf.

Guerrini:2021:EPL

[Gue21]  Anita Guerrini.  Eloge:

Paul Lawrence Farber (1944– 2021). Journal of the History of Biology, 54(4):551– 553, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09665-2.

Guerrini:2023:BRA

[Gue23] Anita Guerrini. Book review: Alisha Rankin. The Poison Trials: Wonder Drugs, Experiment, and the Battle for Authority in Renaissance Science, Chicago: University of Chicago Press, 2021, ISBN 978-0-226-74485-8, 329 pp. Journal of the History of Biology, 56(1):207– 209, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09710-2.

Guerrini:2024:ACE

[Gue24] Anita Guerrini. Andrew Cunningham, “I Follow Aristotle:” How William Harvey Discovered the Circulation of the Blood, Routledge: London and New York, 2022, ISBN: 978-1-032-16223-2, xii + 180 pp. Journal of the History of Biology, 57(4):609– 611, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09795-3.

Guralnick:2002:RRF

[Gur02] Robert Guralnick. A recapitulation of the rise and fall of the cell lineage research program: The evolutionary– developmental relationship of cleavage to homology, body plans and life history. Journal of the History of Biology, 35 (3):537–567, September 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021119112943;
http://link.springer.com/
content/pdf/10.1023/A:1021119112943.
pdf.

Guttmann:1997:TCS

[Gut97] Ronald D. Guttmann. Technology, clinical studies, and control in the field of organ transplantation. Journal of the History of Biology, 30 (3):367–379, September 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004273805232;

                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004273805232.
                                        pdf.

Gayon:1998:RVC

[GZ98] Jean Gayon and Doris T. Zallen. The role of the Vilmorin Company in the promotion and diffusion of the experimental science of heredity in France, 1840–1920. Journal of the History of Biology, 31 (2):241–262, June 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004335619901;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004335619901.
                                        pdf.

Ha:2011:RSB

[Ha11] Nathan Q. Ha. The riddle of sex: Biological theories of sexual difference in the early Twentieth Century. Journal of the History of Biology, 44(3): 505–546, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9257-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9257-1.pdf.

Hill-Andrews:2019:LOM

[HA19] Oliver Hill-Andrews. Lamarckism by other means: Interpreting Pavlov’s conditioned reflexes in Twentieth-Century Britain. Journal of the History of Biology, 52 (1):3–43, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9540-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
018-9540-0.pdf.

Hagen:1984:ENT

[Hag84]  Joel B. Hagen.  Exper-

imentalists and naturalists in twentieth-century botany: Experimental taxonomy, 1920– 1950. Journal of the History of Biology, 17(2):249– 270, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00143734.

Hagen:1986:ETD

[Hag86] Joel B. Hagen. Ecologists and taxonomists: Divergent traditions in twentieth-century plant geography. Journal of the History of Biology, 19(2): 197–214, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138876.

Hagen:1999:NMB

[Hag99] Joel B. Hagen. Naturalists, molecular biologists, and the challenges of molecular evolution. Journal of the History of Biology, 32(2):321– 341, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004660202226;
http://link.springer.com/
content/pdf/10.1023/A:1004660202226.
pdf.

Hagen:2003:SFM

[Hag03] Joel Hagen. The statistical frame of mind in systematic biology from quantitative zoology to biometry. Journal of the History of Biology, 36 (2):353–384, June 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1024479322226;
http://link.springer.com/
content/pdf/10.1023/A:1024479322226.
pdf.

Hagen:2010:WSM

[Hag10] Joel B. Hagen. Waiting for sequences: Morris Goodman, immunodiffusion experiments, and the origins of molecular anthropology. Journal of the History of Biology, 43(4):697– 725, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9219-7;
 http://link.springer.com/

content/pdf/10.1007/s10739-
009-9219-7.pdf.

Hagen:2015:CCK

[Hag15]  Joel B. Hagen.  Camels,

cormorants, and kangaroo rats: Integration and synthesis in organismal biology after World War II. Journal of the History of Biology, 48 (2):169–199, May 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9399-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9399-2.pdf.

Hagen:2017:BRA

[Hag17] Joel B. Hagen. Bergmann’s rule, adaptation, and thermoregulation in Arctic animals: Conflicting perspectives from physiology, evolutionary biology, and physical anthropology after World War II. Journal of the History of Biology, 50(2):235–265, May 2017. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9446-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
016-9446-7.pdf.

Hagen:2024:BRS

[Hag24] Joel B. Hagen. Book review:

Sharon E. Kingsland, A Lab for All Seasons: The Laboratory Revolution in Modern Botany and the Rise of Physiological Plant Ecology, 2023, New Haven: Yale University Press, ISBN: 978-0-300-26722-8, 385 pp. Journal of the History of Biology, 57 (1):165–167, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09765-9.

Haines:1991:SDQ

[Hai91] Valerie A. Haines. Spencer, Darwin, and the question of reciprocal influence. Journal of the History of Biology, 24 (3):409–431, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00156319.

Hall:1970:DPM

[Hal70] Thomas S. Hall. Descartes’ physiological method: Position, principles, examples. Journal of the History of Biology, 3(1):53–79, Spring 1970. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00569306.

Hall:1971:IBL

[Hal71] Diana Long Hall. The iatromechanical background of Lagrange’s theory of animal heat. Journal of the History of Biology, 4(2):245–

248, ????  1971.  CO-

DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138312.

Hall:1974:JHBb

[Hal74] Diana Long Hall. The J.

H. B. bookshelf. Journal of the History of Biology, 7 (2):339–345, Fall 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351209.

Hall:1976:CAC

[Hal76] Diana Long Hall. The critic and the advocate: Contrasting British views on the state of endocrinology in the early 1920s. Journal of the History of Biology, 9(2):269– 285, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209885.

Hall:1998:MKP

[Hal98] Nancy S. Hall. Metagons in killer paramecia: Problems of reproducibility and alternative hypotheses. Journal of the History of Biology, 31 (3):425–446, September 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004345725810;

                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004345725810.
                                        pdf.

Hale:2003:LHR

[Hal03] Piers J. Hale. Labor and the human relationship with nature: The naturalization of politics in the work of Thomas Henry Huxley, Herbert George Wells, and William Morris. Journal of the History of Biology, 36(2): 249–284, June 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1024486021318;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1024486021318.
                                        pdf.

Hall:2007:RFH

[Hal07] Nancy S. Hall. R. A. Fisher and his advocacy of randomization. Journal of the History of Biology, 40(2): 295–325, June 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9119-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9119-z.pdf.

Hale:2009:BR

[Hal09] Piers J. Hale. Book review. Journal of the History of Biology, 42(2):381–385, May 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9193-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9193-0.pdf.

Hale:2010:MME

[Hal10] Piers J. Hale. Of mice and men: Evolution and the socialist utopia. William Morris, H. G. Wells, and George Bernard Shaw. Journal of the History of Biology, 43(1): 17–66, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9177-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9177-0.pdf.

Hale:2012:BRa

[Hal12a] Piers J. Hale. Book review. Journal of the History of Biology, 45(1):171– 173, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9315-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9315-3.pdf.

Hale:2012:BRb

[Hal12b] Piers J. Hale. Book review.

Journal of the History of Biology, 45(2):365–366, May 2012.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9323-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9323-y.pdf.

Hale:2013:MMM

[Hal13] Piers J. Hale. Monkeys into men and men into monkeys: Chance and contingency in the evolution of man, mind and morals in Charles Kingsley’s water babies. Journal of the History of Biology, 46 (4):551–597, November 2013. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9345-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9345-5.pdf.

Hale:2017:BR

[Hal17] Piers J. Hale. Book review. Journal of the History of Biology, 50(3):681– 683, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9483-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9483-x.pdf.

Hale:2025:BRM

[Hal25] Piers J. Hale. Book review:

Maria Elice Brzezinski Prestes (ed.), Understanding Evolution in Darwin’s Origin: The Emerging Context of Evolutionary Theory, (History, Philosophy, and Theory of the Life Sciences, volume 34), Cham: Springer, 2023, ISBN: 978-3-031-40164-0, 423 pp. Journal of the History of Biology, 58(2):305–308, June 2025. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09823-w.

Hamburger:1984:HMC

[Ham84] Viktor Hamburger. Hilde Mangold, co-discoverer of the organizer. Journal of the History of Biology, 17(1): 1–11, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00397500.

Hamlin:1986:RWM

[Ham86] Christopher Hamlin. Robert Warington and the moral economy of the aquarium. Journal of the History of Biology, 19(1):131–153, Spring 1986. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http://

link.springer.com/article/
10.1007/BF00346619.

Hamburger:1997:WRV

[Ham97a]  Viktor Hamburger.  Wil-

helm Roux: Visionary with a blind spot. Journal of the History of Biology, 30(2): 229–238, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004231618837;
http://link.springer.com/
content/pdf/10.1023/A:1004231618837.
pdf.

Hammonds:1997:BRD

[Ham97b] Evelynn M. Hammonds. Book review: Donna J. Harway, Modest Witness@Second Millennium.FemaleMan\begin{matrix}c \\ ◯ \end{matrix} MeetsOncoMouse^{T M}: Feminism and Technoscience. Journal of the History of Biology, 30(3):494–497, September 1997. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004349615837;
http://link.springer.com/
content/pdf/10.1023/A:1004349615837.
pdf.

Hamburger:1999:HSV

[Ham99] Viktor Hamburger. Hans Spemann on vitalism in biology: Translation of a portion of Spemann’s autobiography. Journal of the History of Biology, 32(2):231– 243, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004659911795;
http://link.springer.com/

content/pdf/10.1023/A:1004659911795.
pdf.

Hamblin:2007:DOE

[Ham07] Jacob Darwin Hamblin. ‘A dispassionate and objective effort:’ negotiating the first study on the biological effects of atomic radiation. Journal of the History of Biology, 40 (1):147–177, March 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-6531-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
005-6531-8.pdf.

Hansen:1993:ENM

[Han93] Leeann Hansen. From enlightenment to Naturphilosophie: Marcus Herz, Johann Christian Reil, and the problem of border crossings. Journal of the History of Biology, 26 (1):39–64, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01060679.

Harrington:1989:ERR

[Har89]  Anne Harrington.  Essay

review: Race hygiene and Nazi medicine. Journal of the History of Biology, 22 (3):501–505, Fall 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00401580.

Harvey:1990:ERR

[Har90] Joy Harvey. Essay review: Russian Darwinism. Journal of the History of Biology, 23 (3):523–527, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136382.

Harwood:1994:MFE

[Har94] Jonathan Harwood. Metaphysical foundations of the evolutionary synthesis: a historiographical note. Journal of the History of Biology, 27 (1):1–20, Spring 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058625.

Harman:2003:CDD

[Har03] Oren Solomon Harman. C. D. Darlington and the British and American reaction to Lysenko and the Soviet conception of science. Journal of the History of Biology, 36(2): 309–352, June 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1024483131660;
http://link.springer.com/
content/pdf/10.1023/A:1024483131660.
pdf.

Harman:2006:MFD

[Har06a] Oren Solomon Harman. Method as a function of “disciplinary landscape”: C. D. Darlington and cytology, genetics and evolution, 1932–1950. Journal of the History of Biology, 39 (1):165–197, Spring 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2085-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2085-4.pdf. See correc-

tion [Har22b].

Harwood:2006:ISI

[Har06b] Jonathan Harwood. Introduction to the special issue on biology and agriculture. Journal of the History of Biology, 39 (2):237–239, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0003-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-0003-7.pdf.

Harman:2008:PII

[Har08] Oren Harman. Pure intelligence: On intelligence testing, Puritanism, and the methods and burdens of history. Journal of the History of Biology, 41(1):167–202, March 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-008-9151-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9151-2.pdf.

Harman:2011:HBH

[Har11a] Oren Harman. Helical biography and the historical craft: The case of altruism and George price. Journal of the History of Biology, 44(4): 671–691, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9269-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9269-5.pdf.

Harman:2011:ISI

[Har11b] Oren Harman. Introduction to the special issue — “Scientific biography: a many faced art form”. Journal of the History of Biology, 44(4):607– 609, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9275-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9275-7.pdf.

Harman:2011:MDS

[Har11c] Oren Harman. The myth of disinterested science. Journal of the History of Biology, 44 (2):345–355, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9293-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9293-5.pdf.

Harman:2012:NFD

[Har12a] Oren Harman. Is the naturalistic fallacy dead (and if so, ought it be?). Journal of the History of Biology, 45(3): 557–572, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9332-x;
http://link.springer.com/
content/pdf/10.1007/s10739-
012-9332-x.pdf.

Hart:2012:WEE

[Har12b] Bradley W. Hart. Watching the ‘eugenic experiment’ unfold: The mixed views of British eugenicists toward Nazi Germany in the early 1930s. Journal of the History of Biology, 45(1):33– 63, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9259-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9259-z.pdf.

Harman:2013:EGH

[Har13] Oren Harman. Explaining the gap: On humans and other animals — essay review. Journal of the History of Biology, 46(4):739– 755, November 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9372-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9372-x.pdf.

Harman:2014:CNR

[Har14] Oren Harman. Chance and necessity revisited. Journal of the History of Biology, 47(3): 479–493, August 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9390-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9390-3.pdf. A version

of this article appeared as “Two Road[s] to the Meaning of Life” in the June 8, 2014 edition of the Los Angeles Review of Books.

Harman:2015:EOF

[Har15]  Oren Harman.  Evolution

on one foot. Journal of the History of Biology, 48(2): 337–351, May 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-015-9406-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9406-7.pdf.

Harman:2022:CDM

[Har22a] Oren Harman. A conversation with Darwin on Man revisited: 150 years to The Descent of Man. Journal of the History of Biology, 55(1): 185–201, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09673-w.

Harman:2022:CMF

[Har22b] Oren Soloman Harman. Correction to: Method as a Function of “Disciplinary Landscape”: C. D. Darlington and Cytology, Genetics and Evolution, 1932–1950. Journal of the History of Biology, 55 (1):203, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-018-9551-x. See

[Har06a].

Harman:2023:KFM

[Har23] Oren Harman. “Keep the faith:” memories of Everett Mendelson. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09750-8.

Hatakeyama:2021:LCS

[Hat21] Sumiko Hatakeyama. Let chromosomes speak: The cytogenetics project at the Atomic Bomb Casualty Commission (ABCC). Journal of the History of Biology, 54(1): 107–126, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09628-7.

Haygood:1988:HRV

[Hay88] Tamara Miner Haygood. Henry Ravenel (1814–1887): Views on evolution in social context. Journal of the History of Biology, 21(3): 457–472, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144091.

Hay:2017:BR

[Hay17] Amy M. Hay. Book review. Journal of the History of Biology, 50(1):219– 222, February 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9465-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9465-z.pdf.

Hearty:2020:RBR

[Hea20]  Ryan Hearty.  Redefining

boundaries: Ruth Myrtle Patrick’s ecological program at the Academy of Natural Sciences of Philadelphia, 1947–1975. Journal of the History of Biology, 53(4):587– 630, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09622-5.

Hecht:2021:EMR

[Hec21] David K. Hecht. Embracing mystery: Radiation risks and popular science writing in the Early Cold War. Journal of the History of Biology, 54(1): 127–141, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09629-6.

Heggie:2015:ISS

[Heg15] Vanessa Heggie. Introduction — Special section: Harvard Fatigue Laboratory. Journal of the History of Biology, 48(3):361–364, August 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9411-x;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9411-x.pdf.

Hein:1972:EMV

[Hei72] Hilde Hein. The endurance of the mechanism-vitalism controversy. Journal of the History of Biology, 5(1):159– 188, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113490.

Heinrich:2019:BRG

[Hei19] Ari Larissa Heinrich. Book review: Giovanni Aloi, Speculative Taxidermy: Natural History, Animal Surfaces, and Art in the Anthropocene (New York: Columbia University Press, 2018), 328 pp., ISBN: 978-0-231-18070-2 (hard cover), ISBN: 978-0-231-18071-9 (Paperback). Journal of the History of Biology, 52(2):353–355, June 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9556-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9556-0.pdf.

Henson:1997:PPE

[Hen97] Pamela M. Henson. Paolo Palladino, entomology, ecology and agriculture: The making of scientific careers in North America. Journal of the History of Biology, 30(3):485– 487, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004306712938;
http://link.springer.com/
content/pdf/10.1023/A:1004306712938.
pdf.

Henson:2003:WHE

[Hen03] Pamela M. Henson. ‘What holds the Earth together’: Agnes Chase and American agrostology. Journal of the History of Biology, 36(3):437– 460, September 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004568.
11609.2d;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004568.
11609.2d.pdf.

Henry:2006:AMI

[Hen06] Devin Henry. Aristotle on the mechanism of inheritance. Journal of the History of Biology, 39(3):425– 455, September 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-3058-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-3058-y.pdf.

Henderson:2010:BR

[Hen10] Gabriel Henderson. Book review. Journal of the History of Biology, 43(4):813– 815, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9263-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9263-3.pdf.

Herbert:1971:DMS

[Her71]  Sandra Herbert.  Darwin,

Malthus, and selection. Journal of the History of Biology, 4 (1):209–217, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00356983.

Herbert:1974:PMD

[Her74] Sandra Herbert. The place of man in the development of Darwin’s theory of transmutation. Journal of the History of Biology, 7(2): 217–258, Fall 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00351204.

Herbert:1977:PMD

[Her77] Sandra Herbert. The place of man in the development of Darwin’s theory of transmutation. Part II. Journal of the History of Biology, 10 (2):155–227, Fall 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00572643.

Herbert:2005:DRR

[Her05] Sandra Herbert. The Darwinian revolution revisited. Journal of the History of Biology, 38(1):51–66, ???? 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-6509-y;
http://link.springer.com/
content/pdf/10.1007/s10739-
004-6509-y.pdf.

Herbert:2010:BR

[Her10] Sandra Herbert. Book review. Journal of the History of Biology, 43(1):199– 202, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9224-x;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9224-x.pdf.

Hall:1973:JHB

[HF73] Diana Long Hall and Robert G.

Frank, Jr. The J. H. B. bookshelf. Journal of the History of Biology, 6(1):167– 170, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00137303.

Hall:1974:JHBa

[HF74] Diana Long Hall and Robert G. Frank, Jr. The J. H. B. bookshelf. Journal of the History of Biology, 7(1):175– 177, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00179298.

Hall:1976:EBI

[HG76] Diana Long Hall and Thomas F. Glick. Endocrinology: a brief introduction. Journal of the History of Biology, 9(2): 229–233, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209883.

Hickling:2021:VCE

[Hic21] James Hickling. The vera causa of endangered species legislation: Alfred Newton and the Wild Bird Preservation Acts, 1869–1894. Journal of the History of Biology, 54 (2):275–309, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09633-w.

Harwood:1995:JHB

[HLM^{+}95] Jonathan Harwood, M. Susan Lindee, David Magnus, Angela Creager, Mark V. Barrow, Jr., and Myles W. Jackson. The J. H. B. bookshelf. Journal of the History of Biology, 28(1):167– 179, Spring 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061251.

Hobbins:2015:SAP

[Hob15] Peter Hobbins. A spur to atavism: Placing platypus poison. Journal of the History of Biology, 48(4):499– 537, November 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9409-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9409-4.pdf.

Hodge:1972:UGN

[Hod72] M. J. S. Hodge. The universal gestation of nature: Chambers’ vestiges and explanations. Journal of the History of Biology, 5(1):127– 151, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113488.

Hodge:2005:ARE

[Hod05] Jonathan Hodge. Against “revolution” and “evolution”.

Journal of the History of Biology, 38(1):101–121, ???? 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-6512-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-6512-3.pdf.

Hodge:2009:CCD

[Hod09] M. J. S. Hodge. Capitalist contexts for Darwinian theory: Land, finance, industry and empire. Journal of the History of Biology, 42(3): 399–416, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9187-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9187-y.pdf.

Holmes:1969:JBF

[Hol69] Frederic L. Holmes. Joseph Barcroft and the fixity of the internal environment. Journal of the History of Biology, 2(1):89–122, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137269.

Holmes:1993:OMS

[Hol93] Frederic L. Holmes. The old martyr of science: The frog in experimental physiology. Journal of the History of Biology, 26(2):311– 328, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061972.

Holterhoff:2014:HRC

[Hol14] Kate Holterhoff. The history and reception of Charles Darwin’s hypothesis of pangenesis. Journal of the History of Biology, 47(4):661– 695, November 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9377-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
014-9377-0.pdf.

Holmes:2017:SQS

[Hol17] Matthew Holmes. The sparrow question: Social and scientific accord in Britain, 1850–1900. Journal of the History of Biology, 50(3): 645–671, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9455-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
016-9455-6.pdf.

Home:1970:ENF

[Hom70] Roderick W. Home. Electricity and the nervous fluid. Journal of the History of Biology, 3(2):235–251, ???? 1970. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137353.

Honigsbaum:2016:TBK

[Hon16] Mark Honigsbaum. ‘Tipping the balance’: Karl Friedrich Meyer, latent infections, and the birth of modern ideas of disease ecology. Journal of the History of Biology, 49(2): 261–309, April 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9430-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9430-7.pdf.

Hook:1997:EMM

[Hoo97] Ernest B. Hook. The exclusion of minor malformations in the study of mutation in the offspring of survivors of atomic bombs: Methodological, not sociopolitical, rationale. Journal of the History of Biology, 30(2):239–242, June 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004283602907;
http://link.springer.com/
content/pdf/10.1023/A:1004283602907.
pdf.

Hopwood:2024:SCM

[Hop24]  Nick Hopwood.  Species

choice and model use: Reviving research on human development. Journal of the History of Biology, 57(2): 231–279, June 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09775-7.

Hoquet:2016:WDI

[Hoq16] Thierry Hoquet. What does it mean to be central? A botanical geography of Paris 1830–1848. Journal of the History of Biology, 49(1):191– 230, February 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9416-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9416-5.pdf.

Hoquet:2024:DWS

[Hoq24] Thierry Hoquet. Darwin and the white shipwrecked sailor: Beyond blending inheritance and the Jenkin Myth. Journal of the History of Biology, 57 (1):17–49, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09770-y.

Horowitz:1976:AW

[Hor76] Mary Anne Cline Horowitz. Aristotle and woman. Journal of the History of Biology, 9(2):183–213, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209881.

Hoffman:2011:RTT

[HR11] M. Timm Hoffman and Richard Frederick Rohde. Rivers through time: Historical changes in the riparian vegetation of the semi-arid, winter rainfall region of South Africa in response to climate and land use. Journal of the History of Biology, 44(1): 59–80, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9246-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
010-9246-4.pdf.

Hamblin:2021:CLH

[HR21] Jacob Hamblin and Linda M. Richards. Connecting to the living history of radiation exposure. Journal of the History of Biology, 54 (1):1–6, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09634-9.

Hall:1972:JHB

[HS72] Diana Long Hall and Judith P. Swazey. The J. H. B. bookshelf. Journal of the History of Biology, 5(2): 439–444, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346667.

Hepler-Smith:2018:BRJ

[HS18] Evan Hepler-Smith. Book review: Jutta Schickore, About Method: Experimenters, Snake Venom, and the History of Writing Scientifically (Chicago: University of Chicago Press, 2017), 316 pp., $50.00 Cloth, ISBN: 978-0-226-44998-2. Journal of the History of Biology, 51(3):601–603, September 2018. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9523-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9523-1.pdf.

Hubbard:2020:BRA

[Hub20] Joshua A. Hubbard. Book review: Ayo Wahlberg, Good Quality: The Routinization of Sperm Banking in China, Oakland: University of California Press, 2018, 248 pp., 14 b&w illus., 3 tables, $34.95 Paperback, ISBN: 978-0-520-29778-4. Journal of the History of Biology, 53(4):663– 665, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09621-6.

Hughes:2011:ADR

[Hug11] J. Donald Hughes. Ancient deforestation revisited. Journal of the History of Biology, 44(1):43–57, February 2011. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9247-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9247-3.pdf.

Hull:1969:WPB

[Hul69] David Hull. What philosophy of biology is not. Journal of the History of Biology, 2(1):241–268, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137277.

Hull:2005:DDE

[Hul05] David L. Hull. Deconstructing Darwin: Evolutionary theory in context. Journal of the History of Biology, 38(1): 137–152, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-6514-1;

http://link.springer.com/
content/pdf/10.1007/s10739-
004-6514-1.pdf.

Hume:2008:QCP

[Hum08] Brad D. Hume. Quantifying characters: Polygenist anthropologists and the hardening of heredity. Journal of the History of Biology, 41(1): 119–158, March 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9143-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9143-7.pdf.

Huneman:2019:HMS

[Hun19a] Philippe Huneman. How the modern synthesis came to ecology. Journal of the History of Biology, 52(4):635– 686, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09570-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-09570-9.pdf.

Huneman:2019:SIE

[Hun19b] Philippe Huneman. Special Issue Editor’s introduction: “revisiting the modern synthesis”. Journal of the History of Biology, 52(4):509– 518, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09585-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-09585-2.pdf.

Hung:2019:SSC

[Hun19c] Kuang-Chi Hung. Subscribing to specimens, cataloging subscribed specimens, and assembling the first phytogeographical survey in the United States. Journal of the History of Biology, 52(3):391– 431, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9565-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9565-z.pdf.

Hutcheson:2022:CBF

[Hut22] Emily S. Hutcheson. A “Central Bureau of Feminine Algology:” algae, mutualism, and gendered ecological perspectives, 1880–1910. Journal of the History of Biology, 55(4): 791–825, December 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09698-1.

Ihde:1971:CCE

[IB71] Aaron J. Ihde and Stanley L. Becker. Conflict of concepts in early vitamin studies. Journal of the History of Biology, 4(1):1–33, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356975.

Iida:2010:PPJ

[Iid10] Kaori Iida. Practice and politics in Japanese science: Hitoshi Kihara and the formation of a genetics discipline. Journal of the History of Biology, 43(3):529– 570, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9194-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9194-z.pdf.

Ilerbaig:1999:ASF

[Ile99] Juan Ilerbaig. Allied sciences and fundamental problems: C. C. Adams and the search for method in early American ecology. Journal of the History of Biology, 32(3):439– 463, December 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004737021541;
http://link.springer.com/
content/pdf/10.1023/A:1004737021541.
pdf.

Ilerbaig:2005:BRD

[Ile05] Juan Ilerbaig. Book reviews: David N. Livingstone, Putting Science in Its Place: Geographies of Scientific Knowledge, science * culture Series (Chicago and London: University of Chicago Press, 2003), xii + 234 pp., illus., $27.50. Journal of the History of Biology, 38(2):388– 389, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4229-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4229-6.pdf.

Irmscher:2005:BRL

[Irm05] Christoph Irmscher. Book reviews: Louis Agassiz, Essay on Classification, with an introduction by Edward Lurie (Mineola, N.Y.: Dover, 2004), xxxiii + 268 pp., illus., $22.95 (paper). Elizabeth Higgins Gladfelter, Agassiz’s Legacy: Scientists’ Reflections on the Value of the Field Experience (New York: Oxford University Press, 2002), ix + 437 pp., illus., $17.95 (paper). Journal of the History of Biology, 38(2):396–397, Summer 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4233-x;
 http://link.springer.com/

content/pdf/10.1007/s10739-
005-4233-x.pdf.

Iverson:2010:BR

[Ive10] Margot Iverson. Book review. Journal of the History of Biology, 43(4):801– 803, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9260-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9260-6.pdf.

Jacyna:1983:JGM

[Jac83] L. S. Jacyna. John Goodsir and the making of cellular reality. Journal of the History of Biology, 16(1): 75–99, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00186676.

Jacyna:1984:RPR

[Jac84] L. S. Jacyna. The romantic programme and the reception of cell theory in Britain. Journal of the History of Biology, 17(1):13–48, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00397501.

Jacobs:1989:UUP

[Jac89] Natasha X. Jacobs. From unit to unity: Protozoology, cell theory, and the new concept of life. Journal of the History of Biology, 22(2): 215–242, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139513.

Jackson:2001:WUR

[Jac01] John P. Jackson, Jr. “In ways unacademical”: The reception of Carleton S. Coon’s the origin of races. Journal of the History of Biology, 34 (2):247–285, June 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010366015968;
http://link.springer.com/
content/pdf/10.1023/A:1010366015968.
pdf.

Jacyna:2003:MFN

[Jac03] L. S. Jacyna. Moral fibre: The negotiation of microscopic facts in Victorian Britain. Journal of the History of Biology, 36(1):39–85, March 2003. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022532313285;
http://link.springer.com/
content/pdf/10.1023/A:1022532313285.
pdf.

Jackson:2009:BR

[Jac09] John P. Jackson, Jr. Book review. Journal of the History of Biology, 42(2): 395–397, May 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9192-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9192-1.pdf.

Jackson:2010:BR

[Jac10] John P. Jackson, Jr. Book review. Journal of the History of Biology, 43(4):809– 811, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9262-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9262-4.pdf.

Jones:2018:BTP

[JACD18] Kathryn Maxson Jones, Rachel A.

Ankeny, and Robert Cook-Deegan. The Bermuda Triangle: The pragmatics, policies, and principles for data sharing in the history of the Human Genome Project. Journal of the History of Biology, 51 (4):693–805, December 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9538-7;
http://link.springer.com/

content/pdf/10.1007/s10739-
018-9538-7.pdf.

Janssen:2024:DSR

[Jan24] Diederik F. Janssen. Developing sex: From recremental semen to developmental endocrinology. Journal of the History of Biology, 57(1): 113–151, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09766-8.

Jax:1998:HEO

[Jax98] Kurt Jax. Holocoen and ecosystem — on the origin and historical consequences of two concepts. Journal of the History of Biology, 31(1): 113–142, March 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004261607170;
http://link.springer.com/
content/pdf/10.1023/A:1004261607170.
pdf.

Jenkins:2016:NTR

[Jen16] Bill Jenkins. Neptunism and transformism: Robert Jameson and other evolutionary theorists in Early Nineteenth-Century Scotland. Journal of the History of Biology, 49(3): 527–557, August 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10739-015-9425-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9425-4.pdf.

Jianhui:2003:SIR

[JF03] Li Jianhui and Hong Fan. Science as ideology: The rejection and reception of sociobiology in China. Journal of the History of Biology, 36(3):567– 578, September 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004580.
34274.78;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004580.
34274.78.pdf.

Jiang:2014:CAL

[Jia14] Lijing Jiang. Causes of aging are likely to be many: Robin Holliday and changing molecular approaches to cell aging, 1963–1988. Journal of the History of Biology, 47(4): 547–584, November 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9382-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
014-9382-3.pdf.

Jiang:2024:BRS

[Jia24] Lijing Jiang. [Book review:] Samantha Muka, Oceans under Glass: Tank Craft and the Sciences of the Sea, Chicago: University of Chicago Press, 2023, ISBN 978-0-226-82413-0, 242 pp. Journal of the History of Biology, 57(2): 331–334, June 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09776-6.

Jin:2025:BRH

[Jin25] Xiaoxing Jin. Book review: Hyung Wook Park, Creationism in South Korean Culture: Science, Religion, and the Struggle against Evolution, London: Routledge, 2024, ISBN: 978-1-032-75714-8, 222 pp. Journal of the History of Biology, 58(1): 145–146, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09812-z.

Jackson:1999:TPE

[JK99] James R. Jackson and William C. Kimler. Taxonomy and the personal equation: The historical fates of Charles Girard and Louis Agassiz. Journal of the History of Biology, 32 (3):509–555, December 1999. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004784904703;
http://link.springer.com/

content/pdf/10.1023/A:1004784904703.
pdf.

Johnson:2004:ITT

[Joh04] Kristin Johnson. The Ibis:

Transformations in a Twentieth Century British natural history journal. Journal of the History of Biology, 37(3): 515–555, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-1499-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
004-1499-3.pdf.

Johnson:2007:PDO

[Joh07]  Curtis N. Johnson.  The

preface to Darwin’s Origin of Species: The curious history of the “historical sketch” *. Journal of the History of Biology, 40(3):529– 556, September 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9118-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9118-0.pdf.

Johnson:2009:RPI

[Joh09] Kristin Johnson. The return of the Phoenix: The 1963 International Congress of Zoology and American Zoologists in the Twentieth Century. Journal of the History of Biology, 42(3):417– 456, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9160-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9160-1.pdf.

Johnson:2015:TSS

[Joh15] Andi Johnson. “They sweat for science”: The Harvard Fatigue Laboratory and selfexperimentation in American exercise physiology. Journal of the History of Biology, 48(3): 425–454, August 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9387-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9387-y.pdf.

Johnson:2019:CDR

[Joh19a] Curtis N. Johnson. Charles Darwin, Richard Owen, and natural selection: a question of priority. Journal of the History of Biology, 52 (1):45–85, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9514-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9514-2.pdf.

Johnson:2019:BRA

[Joh19b] Tina Phillips Johnson. Book review: Angela Ki Che Leung and Izumi Nakayama, eds.: Gender, Health, and History in Modern East Asia. Journal of the History of Biology, 52 (3):501–503, September 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9562-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-9562-2.pdf.

Johnson:2022:RGT

[Joh22] Kristin Johnson. The return of the geneticist: Theodosius Dobzhansky, Edward Chapin, and museum taxonomy. Journal of the History of Biology, 55(3):443–463, October 2022. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09694-5.

Jones:2017:PCD

[Jon17] Susan D. Jones. Population cycles, disease, and networks of ecological knowledge. Journal of the History of Biology, 50(2):357–391, May 2017. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9441-z;
http://link.springer.com/

content/pdf/10.1007/s10739-
016-9441-z.pdf.

Jiang:2018:EIB

[JRR18] Lijing Jiang, Karen Rader, and Marsha Richmond. Editorial: Introducing “biology in culture” reviews. Journal of the History of Biology, 51 (2):407–409, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9521-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9521-3.pdf.

Jumonville:2002:CPS

[Jum02] Neil Jumonville. The cultural politics of the sociobiology debate. Journal of the History of Biology, 35(3):569– 593, September 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021190227056;
http://link.springer.com/
content/pdf/10.1023/A:1021190227056.
pdf.

Junker:1996:FSE

[Jun96] Thomas Junker. Factors shaping Ernst Mayr’s concepts in the history of biology. Journal of the History of Biology, 29(1):29–77, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00129696.

Joseph:2013:ERH

[JW13] Jay Joseph and Norbert A.

Wetzel. Ernst Rüdin: Hitler’s racial hygiene mastermind. Journal of the History of Biology, 46(1):1– 30, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9344-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9344-6.pdf.

Kalikow:1983:KLE

[Kal83] Theodora J. Kalikow. Konrad Lorenz’s ethological theory: Explanation and ideology, 1938–1943. Journal of the History of Biology, 16 (1):39–73, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00186675.

Kay:1985:CMA

[Kay85] Lily E. Kay. Conceptual models and analytical tools: The biology of physicist Max Delbrück. Journal of the History of Biology, 18(2): 207–246, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00120110.

Kay:1989:SPS

[Kay89] Lily E. Kay. Selling pure science in wartime: The biochemical genetics of G. W. Beadle. Journal of the History of Biology, 22(1):73– 101, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00209604.

Kay:1996:BMB

[Kay96] Lily E. Kay. “Biochemists and molecular biologists: Laboratories, networks, disciplines”: Comments. Journal of the History of Biology, 29(3): 447–450, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127383.

Keating:1997:IHI

[KBCT97] Peter Keating, Miriam Balaban, Alberto Cambrosio, and Alfred I. Tauber. Introduction: Historiographic issues. Journal of the History of Biology, 30(3):317– 320, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004241721645;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004241721645.
                                        pdf.

Keating:1994:OEA

[KC94] Peter Keating and Alberto Cambrosio. “Ours is an engineering approach”: Flow cytometry and the constitution of human T-cell subsets. Journal of the History of Biology, 27(3):449–479, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058994.

Keating:1997:HSF

[KC97] Peter Keating and Alberto Cambrosio. Helpers and suppressors: On fictional characters in immunology. Journal of the History of Biology, 30 (3):381–396, September 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004243822070;
http://link.springer.com/
content/pdf/10.1023/A:1004243822070.
pdf.

Keiner:1998:WKB

[Kei98] Christine Keiner. W. K.

Brooks and the oyster question: Science, politics, and resource management in Maryland, 1880–1930. Journal of the History of Biology, 31 (3):383–424, September 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004393608972;

                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004393608972.
                                        pdf.

Keirns:1999:SPM

[Kei99] Carla Keirns. Seeing patterns: Models, visual evidence and pictorial communication in the work of Barbara McClintock. Journal of the History of Biology, 32(1): 163–195, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004420726771;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004420726771.
                                        pdf.

Keiner:2001:SKH

[Kei01] Christine Keiner. Sheldon Krimsky, hormonal chaos: The scientific and social origins of the environmental endocrine hypothesis. Journal of the History of Biology, 34 (1):195–226, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1010358710108;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1010358710108.
                                        pdf.

Keiner:2011:BR

[Kei11a]  Christine Keiner.  Book

review. Journal of the History of Biology, 44(2): 361–364, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9295-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9295-3.pdf.

Keiner:2011:BRN

[Kei11b] Christine Keiner. Book review: Nancy Langston, Toxic Bodies: Hormone Disruptors and the Legacy of DES (New Haven: Yale University Press, 2010), xvii + 233 pp., illus., $30.00. Journal of the History of Biology, 44(1):167– 169, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9281-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9281-9.pdf.

Keiner:2017:TOB

[Kei17] Christine Keiner. A two-ocean bouillabaisse: Science, politics, and the Central American sea-level canal controversy. Journal of the History of Biology, 50(4):835– 887, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9461-8;
http://link.springer.com/

content/pdf/10.1007/s10739-
016-9461-8.pdf.

Keiner:2023:BRL

[Kei23] Christine Keiner. Book review: Laura J. Martin, Wild by Design: The Rise of Ecological Restoration, Cambridge, USA: Harvard University Press, 2022, ISBN: 978-0-674-97942-0, 336 pp. Journal of the History of Biology, 56(2):407–409, August 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09729-5.

Keller:1988:DPP

[Kel88] Evelyn Fox Keller. Demarcating public from private values in evolutionary discourse. Journal of the History of Biology, 21(2):195–211, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146986.

Keller:1990:PEM

[Kel90] Evelyn Fox Keller. Physics and the emergence of molecular biology: a history of cognitive and political synergy. Journal of the History of Biology, 23(3):389–409, ???? 1990. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136376.

Kevles:1993:RDN

[Kev93] Daniel J. Kevles. Renato Dulbecco and the new animal virology: Medicine, methods, and molecules. Journal of the History of Biology, 26 (3):409–442, Fall 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062056.

Keveles:2005:BRP

[Kev05] Daniel Keveles. Book reviews:

Paul Berg and Maxine Singer, George Beadle, An Uncommon Farmer: The Emergence of Genetics in the 20th Century (Cold Spring Harbor, N.Y.: Cold Spring Harbor Laboratory Press, 2003), ix + 383 pp., illus., $35.00. Journal of the History of Biology, 38(2):381–382, Summer 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4211-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4211-3.pdf.

Koehler:2009:ESA

[KFP09] Peter J. Koehler, Stanley Finger, and Marco Piccolino. The “eels” of South America: Mid-18th-century Dutch contributions to the theory of animal electricity. Journal of the History of Biology, 42(4):715– 763, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9186-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9186-z.pdf.

Kirsh:2021:FHS

[KG21] Nurit Kirsh and L. Joanne Green. A feeling for the human subject: Margaret Lasker and the genetic puzzle of pentosuria. Journal of the History of Biology, 54 (2):247–274, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09642-9.

Kimler:1986:AAE

[Kim86] William C. Kimler. Advantage, adaptiveness, and evolutionary ecology. Journal of the History of Biology, 19(2): 215–233, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138877.

Kimmelman:2006:MBB

[Kim06] Barbara A. Kimmelman. Mr. Blakeslee builds his dream house: Agricultural institutions, genetics, and careers 1900–1915. Journal of the History of Biology, 39(2): 241–280, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0004-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-0004-6.pdf.

Kimmelman:2018:BRH

[Kim18] Barbara Kimmelman. Book review: Helen Anne Curry, Evolution Made to Order: Plant Breeding and Technological Innovation in Twentieth-Century America (Chicago and London: The University of Chicago Press, 2016), x + 285 pp., illus., index, bibl. $45.00, cloth, ISBN: 978-0-226-39008-6. Journal of the History of Biology, 51(2): 387–390, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9518-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9518-y.pdf.

Kimler:2023:BS

[Kim23a] William Kimler. Balancing the synthesis. Journal of the History of Biology, 56(1): 191–194, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09714-y.

Kimler:2023:CJH

[Kim23b] William Kimler. Correction: Joel Hagen. Life out of Balance: Homeostasis and Adaptation in a Darwinian World, Tuscaloosa: University of Alabama Press, 2021, ISBN 978-0-8173-2089-8, 360 pp. Journal of the History of Biology, 56(1):195, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09718-8.

Kingsland:1978:ATP

[Kin78] Sharon Kingsland. Abbott Thayer and the protective coloration debate. Journal of the History of Biology, 11 (2):223–244, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00389300.

Kinch:1980:GDO

[Kin80] Michael Paul Kinch. Geographical distribution and the origin of life: The development of early nineteenthcentury British explanations. Journal of the History of Biology, 13(1):91–119, ???? 1980. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125355.

Kingsland:1986:MFB

[Kin86] Sharon E. Kingsland. Mathematical figments, biological facts: Population ecology in the thirties. Journal of the History of Biology, 19(2): 235–256, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138878.

Kingsland:1994:ERH

[Kin94] Sharon E. Kingsland. Essay review: The history of ecology. Journal of the History of Biology, 27(2):349– 357, Summer 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062566.

Kingsland:2009:IRP

[Kin09a] Sharon Kingsland. Introduction: Remembering Phil. Journal of the History of Biology, 42(2):205–214, May 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9182-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9182-3.pdf.

Kingsland:2009:FWA

[Kin09b] Sharon E. Kingsland. Frits Went’s atomic age greenhouse: The changing labscape on the lab–field border. Journal of the History of Biology, 42(2):289–324, May 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9179-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9179-y.pdf.

Kinzelman:2009:BR

[Kin09c] Cara Kinzelman. Book review. Journal of the History of Biology, 42(3):609– 611, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9198-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9198-8.pdf.

Kingsland:2020:BRE

[Kin20] Sharon E. Kingsland. Book review: Etienne S. Benson, Surroundings: a History of Environments and Environmentalism (Chicago and London: University of Chicago Press, 2020), 278 pp., 16 b&w illus., $27.50 Paper, ISBN: 978-0-226-70629-0. Journal of the History of Biology, 53(4):653– 655, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09620-7.

Kirsh:2013:BR

[Kir13]  Nurit Kirsh.  Book re-

view. Journal of the History of Biology, 46(3):543– 545, August 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9360-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9360-1.pdf.

Kocandrle:2013:EBM

[KK13] Radim Kocandrle and Karel Kleisner. Evolution born of moisture: Analogies and parallels between Anaximander’s ideas on origin of life and man and later pre–Darwinian and Darwinian evolutionary concepts. Journal of the History of Biology, 46(1):103– 124, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9334-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9334-8.pdf.

Kleinman:1999:HOS

[Kle99] Kim Kleinman. His own synthesis: Corn, Edgar Anderson, and evolutionary theory in the 1940s. Journal of the History of Biology, 32(2):293– 320, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004424810841;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004424810841.
                                        pdf.

Kleinman:2004:BRA

[Kle04] Kim Kleinman. Book review: Arturo Warman, Corn and Capitalism: How a Botanical Bastard Grew to Global Dominance. Journal of the History of Biology, 37(3):594– 595, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-2090-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-2090-7.pdf.

Kleinman:2010:BR

[Kle10] Kim Kleinman. Book review. Journal of the History of Biology, 43(3):601– 603, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9249-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9249-1.pdf.

Kleinman:2011:BRN

[Kle11] Kim Kleinman. Book review: Noel Kingsbury, Hybrid: The History and Science of Plant Breeding (Chicago and London: University of Chicago Press, 2009), xv + 493 pp., $35.00. Journal of the History of Biology, 44(1):153– 154, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9277-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9277-5.pdf.

Kleinman:2013:SOS

[Kle13] Kim Kleinman. Systematics and the Origin of Species from the viewpoint of a botanist: Edgar Anderson prepares the 1941 Jesup Lectures with Ernst Mayr. Journal of the History of Biology, 46(1):73– 101, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9325-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9325-9.pdf.

Kleinman:2016:BTS

[Kle16] Kim Kleinman. “Bringing taxonomy to the service of genetics”: Edgar Anderson and introgressive hybridization. Journal of the History of Biology, 49(4):603– 624, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-016-9436-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9436-9.pdf.

Kleinman:2023:GAM

[Kle23] Kim Kleinman. Garland Allen and Marxism: an appreciation. Journal of the History of Biology, 56(2):227– 238, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09725-9.

Kleinman:2024:LBK

[Kle24] Kim Kleinman. Lee B. Kass, From Chromosomes to Mobile Genetic Elements: The Life and Work of Nobel Laureate Barbara McClintock, Boca Raton: CRC Press, 2024, ISBN: 978-1-032-36532-9, 265 pp. Journal of the History of Biology, 57(4):613– 615, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09796-2.

Klingle:1998:PAW

[Kli98] Matthew W. Klingle. Plying atomic waters: Lauren Donaldson and the “Fern Lake concept” of fisheries management. Journal of the History of Biology, 31(1): 1–32, March 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004246421283;
http://link.springer.com/
content/pdf/10.1023/A:1004246421283.
pdf.

Klyve:2014:DMS

[Kly14]  Dominic Klyve.  Darwin,

Malthus, Süssmilch, and Euler: The ultimate origin of the motivation for the theory of natural selection. Journal of the History of Biology, 47 (2):189–212, May 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9366-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
013-9366-8.pdf.

Kay:1993:JHB

[KNM^{+}93] Lily E. Kay, Lynn K. Nyhart, James Moore, Ronald Rainger, and Kristie Macrakis. The J. H. B. bookshelf. Journal of the History of Biology, 26(2):369–381, Summer 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061975.

Keating:1991:PNA

[KO91] Peter Keating and Abdelkérim Ousman. The problem of natural antibodies, 1894–1905. Journal of the History of Biology, 24(2): 245–263, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209431.

Kohler:1971:BEB

[Koh71] Robert Kohler. The background to Eduard Buchner’s discovery of cell-free fermentation. Journal of the History of Biology, 4(1): 35–61, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356976.

Kohler:1972:REB

[Koh72]  Robert E. Kohler.  The

reception of Eduard Buchner’s discovery of cell-free fermentation. Journal of the History of Biology, 5(2): 327–353, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346663.

Kohler:1973:BOW

[Koh73]  Robert E. Kohler.  The

background to Otto Warburg’s conception of the Atmungsferment. Journal of the History of Biology, 6(2): 171–192, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127607.

Kohler:1993:DLL

[Koh93] Robert E. Kohler. Drosophila:

a life in the laboratory. Journal of the History of Biology, 26(2):281–310, Summer 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061971.

Kohlstedt:1995:ERM

[Koh95] Sally Gregory Kohlstedt. Essay review: Museums: Revisiting sites in the history of the natural sciences. Journal of the History of Biology, 28(1): 151–166, Spring 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061250.

Kohler:2004:BRF

[Koh04] Robert E. Kohler. Book review: Frederic Lawrence Holmes, Investigative Pathways: Patterns and Stages in the Careers of Experimental Scientists. Journal of the History of Biology, 37(3): 585–588, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2086-3;
 http://link.springer.com/

                                        content/pdf/10.1007/s10739-
                                        004-2086-3.pdf.

Kohler:2011:LAP

[Köh11] Piotr Köhler. Lysenko Affair and Polish botany. Journal of the History of Biology, 44 (2):305–343, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9238-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9238-4.pdf.

Kohler:2012:PPT

[Koh12] Robert E. Kohler. Practice and place in Twentieth-Century field biology: a comment. Journal of the History of Biology, 45(4):579– 586, November 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9300-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9300-x.pdf.

Kohler:2022:ELF

[Koh22a] Robert Kohler. Lords of the Fly revisited. Journal of the History of Biology, 55 (1):15–19, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09671-y.

Kohlstedt:2022:CGL

[Koh22b] Sally Gregory Kohlstedt. Collaboration, gender, and leadership at the Minnesota Seaside Station, 1901–1907. Journal of the History of Biology, 55(4):751–790, December 2022. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09679-4.

Kohlstedt:2024:BRK

[Koh24] Sally Gregory Kohlstedt.

Book review: Kathleen S. Murphy, Captivity’s Collections: Science, Natural History, and the British Transatlantic Slave Trade, Chapel Hill: University of North Carolina Press, 2023, ISBN: 978-1-4696-7590-9, 256 pp. Journal of the History of Biology, 57(3):489– 491, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09788-2.

Kollmer:2022:ICC

[Kol22] Charles A. Kollmer. International culture collections and the value of microbial life: Johanna Westerdijk’s fungi and Ernst Georg Pringsheim’s algae. Journal of the History of Biology, 55 (1):59–87, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09669-6.

Konashev:2023:RBD

[Kon23] Mikhail B. Konashev. The Russian backdrop to Dobzhansky’s Genetics and the Origin of Species. Journal of the History of Biology, 56(2): 285–307, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09709-9.

Kottek:1981:ETM

[Kot81]  Samuel S. Kottek.  Em-

bryology in Talmudic and Midrashic literature. Journal of the History of Biology, 14 (2):299–315, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00141096.

Kraft:2004:PPP

[Kra04] Alison Kraft. Pragmatism, patronage and politics in English biology: The rise and fall of economic biology 1904–1920. Journal of the History of Biology, 37(2): 213–258, June 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/B:HIST.0000038268.
                                        39357.28;   http://link.

springer.com/content/pdf/
10.1023/B:HIST.0000038268.
39357.28.pdf.

Kranke:2024:HPC

[Kra24] Nina Kranke. How phenograms and cladograms became molecular phylogenetic trees. Journal of the History of Biology, 57(3):423–443, September 2024. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09782-8.

Krementsov:1996:SFS

[Kre96] Nikolai Krementsov. A “second front” in Soviet genetics: The international dimension of the Lysenko controversy, 1944–1947. Journal of the History of Biology, 29(2): 229–250, Summer 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00571083.

Krementsov:2007:PSA

[Kre07] Nikolai Krementsov. A particular synthesis: Aleksandr Promptov and speciation in birds. Journal of the History of Biology, 40(4):637– 682, December 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9123-y;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9123-y.pdf.

Kroll:1997:BRH

[Kro97] Gary Kroll. Book review: Henrika Kuklick and Robert E. Kohler, eds., Science in the Field, Osiris. Journal of the History of Biology, 30(3):481– 484, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004326428867;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004326428867.
                                        pdf.

LaBonte:2017:BSB

[LaB17] Michelle Lynne LaBonte. Blobel and Sabatini’s “beautiful idea”: Visual representations of the conception and refinement of the signal hypothesis. Journal of the History of Biology, 50(4):797– 833, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9462-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9462-7.pdf.

Ladouceur:2008:ETS

[Lad08] Ronald P. Ladouceur. Ella Thea Smith and the lost history of American high school biology textbooks. Journal of the History of Biology, 41 (3):435–471, September 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9139-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
007-9139-3.pdf.

Lagueux:2003:GGH

[Lag03] Olivier Lagueux. Geoffroy’s giraffe: The hagiography of a charismatic mammal. Journal of the History of Biology, 36 (2):225–247, June 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1024435232569;
http://link.springer.com/
content/pdf/10.1023/A:1024435232569.
pdf.

Lamm:2015:STV

[Lam15] Ehud Lamm. Systems thinking versus population thinking: Genotype integration and chromosomal organization 1930s–1950s. Journal of the History of Biology, 48(4): 641–677, November 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9400-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9400-0.pdf.

Lanzarotta:2020:BRV

[Lan20] Tess Lanzarotta. Book review: Vanessa Heggie, Higher and Colder: A History of Extreme Physiology and Exploration (Chicago: University of Chicago Press, 2019), 253 pp., 9 b&w illus., $40.00 Cloth, ISBN: 978-0-226-65088-3. Journal of the History of Biology, 53(1): 199–201, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09596-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        020-09596-4.pdf.

Laporte:1990:WWD

[Lap90] Léo F. Laporte. The world into which Darwin led Simpson. Journal of the History of Biology, 23(3):499– 516, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00136380.

Laporte:1994:SS

[Lap94] Léo F. Laporte. Simpson on species. Journal of the History of Biology, 27(1): 141–159, Spring 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058629.

Larson:1986:PGN

[Lar86] James Larson. Not without a plan: Geography and natural history in the late eighteenth century. Journal of the History of Biology, 19(3): 447–488, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138287.

Larson:1987:BCE

[Lar87] Edward J. Larson. Before the crusade: Evolution in American secondary education before 1920. Journal of the History of Biology, 20(1): 89–114, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125259.

Largent:1999:BVL

[Lar99]  Mark A. Largent.  Bio-

nomics: Vernon Lyman Kellogg and the defense of Darwinism. Journal of the History of Biology, 32(3):465– 488, December 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004732920632;
http://link.springer.com/
content/pdf/10.1023/A:1004732920632.
pdf.

Largent:2006:BR

[Lar06] Mark Largent. Book reviews. Journal of the History of Biology, 39(4):795– 820, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9115-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9115-3.pdf.

Largent:2007:BRb

[Lar07a] Mark Largent. Book reviews. Journal of the History of Biology, 40(2):369–386, June 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9128-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9128-6.pdf.

Largent:2007:BRc

[Lar07b] Mark Largent. Book reviews. Journal of the History of Biology, 40(3):577– 600, September 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9141-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9141-9.pdf.

Largent:2007:BRa

[Lar07c] Mark A. Largent. Book reviews. Journal of the History of Biology, 40(1):185– 206, March 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9121-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9121-5.pdf.

Largant:2008:BR

[Lar08a] Mark Largant. Book reviews. Journal of the History of Biology, 41(3):573– 587, September 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9159-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
008-9159-7.pdf.

Largent:2008:BRa

[Lar08b] Mark Largent. Book reviews.

Journal of the History of Biology, 41(2):389–408, June 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9154-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
008-9154-z.pdf.

Largent:2008:BRb

[Lar08c] Mark Largent. Book reviews. Journal of the History of Biology, 41(4):763– 783, December 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9168-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9168-6.pdf.

Largent:2009:BR

[Lar09] Mark Largent. Book reviews. Journal of the History of Biology, 42(1):189– 203, February 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9185-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9185-0.pdf.

Laurence:2019:BRD

[Lau19] Alison Laurence. Book review: Deborah Bird Rose, Thom van Dooren, and Matthew Chrulew, eds.: Extinction Studies: Stories of Time, Death, and Generations. Journal of the History of Biology, 52(2):361–363, June 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9567-x;

 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9567-x.pdf.

Lederman:1993:I

[LB93] Muriel Lederman and Richard M.

Burian. Introduction. Journal of the History of Biology, 26(2):235–237, Summer 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061967.

Lopez-Beltran:2004:CHF

[LB04] Carlos López-Beltrán. In the cradle of heredity; French physicians and L’Hérédité naturelle in the early 19th Century. Journal of the History of Biology, 37(1): 39–72, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020385.
44164.e2;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020385.
44164.e2.pdf.

Love:2012:BR

[LD12] Alan C. Love and Thomas J.

Doyle. Book review. Journal of the History of Biology, 45 (2):367–371, May 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9324-x;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9324-x.pdf.

Lidwell-Durnin:2019:IDD

[LD19] John Lidwell-Durnin. Inevitable decay: Debates over climate, food security, and plant heredity in Nineteenth-Century Britain. Journal of the History of Biology, 52(2): 271–292, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9550-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9550-y.pdf.

Lidwell-Durnin:2020:WBC

[LD20] John Lidwell-Durnin. William Benjamin Carpenter and the emerging science of heredity. Journal of the History of Biology, 53(1):81– 103, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09568-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09568-3.pdf.

Lederman:1989:RNG

[Led89] Muriel Lederman. Research note: Genes on chromosomes: The conversion of Thomas Hunt Morgan. Journal of the History of Biology, 22(1):

163–176, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209607.

Lederer:2004:BRK

[Led04] Susan E. Lederer. Book review: Karen Rader, Making Mice: Standardizing Animals for American Biomedical Research. Journal of the History of Biology, 37(3): 588–590, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2087-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2087-2.pdf.

Lennox:1980:AGS

[Len80] James G. Lennox. Aristotle on genera, species, and “the more and the less”. Journal of the History of Biology, 13 (2):321–346, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125746.

Lenoir:1987:ERD

[Len87] Timothy Lenoir. Essay review: The Darwin industry. Journal of the History of Biology, 20(1):115–130, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125260.

Lennox:2005:DME

[Len05]  James G. Lennox.  Dar-

win’s methodological evolution. Journal of the History of Biology, 38(1):85– 99, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-6511-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-6511-4.pdf.

Lepisto:2015:RLW

[Lep15] Antti Lepistö. Revisiting the left-wing response to sociobiology: The case of Finland in a European context. Journal of the History of Biology, 48(1):99–136, February 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9386-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9386-z.pdf.

Levallois:2018:DSR

[Lev18] Clement Levallois. The development of sociobiology in relation to animal behavior studies, 1946–1975. Journal of the History of Biology, 51 (3):419–444, September 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9491-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9491-x.pdf.

Lewontin:1969:BCB

[Lew69] Richard C. Lewontin. The bases of conflict in biological explanation. Journal of the History of Biology, 2(1): 35–45, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137266.

Lewontin:1987:PHO

[Lew87] R. C. Lewontin. Polymorphism and heterosis: Old wine in new bottles and vice versa. Journal of the History of Biology, 20(3):337–349, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139459.

Lewis:2004:VRM

[Lew04] Jeffrey Lewis. From virus research to molecular biology: Tobacco mosaic virus in Germany, 1936–1956. Journal of the History of Biology, 37 (2):259–301, June 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1023/B:HIST.0000038206.
                                        09052.7a;   http://link.
                                        springer.com/content/pdf/
                                        10.1023/B:HIST.0000038206.
                                        09052.7a.pdf.

Lewis:2005:ISI

[Lew05] Michael Lewis. Indian science for Indian tigers?: Conservation biology and the question of cultural values. Journal of the History of Biology, 38(2): 185–207, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-1486-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-1486-8.pdf.

Lewis:2018:ASP

[Lew18] Amanda E. Lewis. Africanizing science in post-colonial Kenya: Long–term field research in the Amboseli ecosystem, 1963–1989. Journal of the History of Biology, 51 (3):535–562, September 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9502-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9502-y.pdf.

Loison:2017:CCL

[LGB17] Laurent Loison, Jean Gayon, and Richard M. Burian. The contributions — and collapse — of Lamarckian heredity in Pasteurian molecular biology:

1. Lysogeny, 1900–1960. Journal of the History of Biology, 50(1):5–52, February 2017. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9434-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9434-3.pdf.

Li:2004:NPG

[Li04] Shang-Jen Li. The nurse of parasites: Gender concepts in Patrick Manson’s parasitological research. Journal of the History of Biology, 37(1): 103–130, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020280.
93881.48;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020280.
93881.48.pdf.

Lightman:2018:BRE

[Lig18] Bernard Lightman. Book review: Evelleen Richards, Darwin and the Making of Sexual Selection (Chicago: The University of Chicago Press, 2017), 672 pp., 48 halftones, $47.50 Cloth, ISBN 978-0-226-43690-6. Journal of the History of Biology, 51(3):597– 600, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9524-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9524-0.pdf.

Lightman:2022:DCP

[Lig22] Bernard Lightman. The Darwin Correspondence Project and pedagogy. Journal of the History of Biology, 55(2): 403–409, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09684-7.

Lindee:1992:WMI

[Lin92] M. Susan Lindee. What is a mutation? Identifying heritable change in the offspring of survivors at Hiroshima and Nagasaki. Journal of the History of Biology, 25(2): 231–255, ???? 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00162841.

Linquist:2015:BR

[Lin15] Stefan Linquist. Book review. Journal of the History of Biology, 48(2):357–360, May 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9405-8;

http://link.springer.com/
content/pdf/10.1007/s10739-
015-9405-8.pdf.

Link:2018:BRS

[Lin18] Adrianna Link. Book review: Sarah S. Richardson and Hallam Stevens, eds., Postgenomics: Perspectives on Biology after the Genome (Durham: Duke University Press, 2015), 294 pp., $99.95 Cloth, ISBN: 978-0-8223-5894-7. Journal of the History of Biology, 51(4):891– 895, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9545-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
018-9545-8.pdf.

Lindee:2023:BRR

[Lin23] M. Susan Lindee. Book review: Rena Selya, Salvador Luria: An Immigrant Biologist in Cold War America, Cambridge, USA: MIT Press, 2022, ISBN: 978-0-262-04646-6, 248 pp. Journal of the History of Biology, 56 (4):??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09758-0.

Lindee:2024:HCE

[Lin24] M. Susan Lindee. How to civilize elites: Controlling “Foreign scientists” at a field station in the Galápagos Islands. Journal of the History of Biology, 57(4):581– 602, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09801-8.

Liu:2017:CPC

[Liu17] Daniel Liu. The cell and protoplasm as container, object, and substance, 1835– 1861. Journal of the History of Biology, 50(4):889– 925, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9460-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9460-9.pdf.

Liu:2023:BRK

[Liu23]  Daniel Liu.  Book re-

view: Karl S. Matlin, Crossing the Boundaries of Life: Günter Blobel and the Origins of Molecular Cell Biology, Chicago: University of Chicago Press, 2022, ISBN: 978-0-226-81923-5, 368 pp. Journal of the History of Biology, 56(2):411– 414, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10739-023-09730-y.

Liu:2024:SOC

[Liu24] Daniel Liu. The schema and organization of the cell: An introduction to Ernst Brücke’s Die Elementarorganismen (1861). Journal of the History of Biology, 57(2): 281–304, June 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09774-8.

Lidgard:2025:LFM

[LK25] Scott Lidgard and Emma Kitchen. Living fossil: A metaphor’s travels across popular culture and the foundations of Darwinian evolution and anthropology. Journal of the History of Biology, 58 (2):157–162, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09807-w.

Llana:2000:NHE

[Lla00] James Llana. Natural history and the Encyclopédie. Journal of the History of Biology, 33(1):1–25, March 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004736103768;
http://link.springer.com/

                                        content/pdf/10.1023/A:1004736103768.
                                        pdf.

Lachapelle:2014:WET

[LM14] Sofie Lachapelle and Heena Mistry. From the waters of the empire to the tanks of Paris: The creation and early years of the aquarium tropical, Palais de la Porte Dorée. Journal of the History of Biology, 47(1):1– 27, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9362-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9362-z.pdf.

Ledesma-Mateos:2003:IBM

[LMB03] Ismael Ledesma-Mateos and Ana Barahona. The institutionalization of biology in Mexico in the early 20th Century. The conflict between Alfonso Luis Herrera (1868– 1942) and Isaac Ochoterena (1885–1950). Journal of the History of Biology, 36(2): 285–307, June 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1024443517547;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1024443517547.
                                        pdf.

Logan:2001:NRL

[Log01] Cheryl A. Logan. “[A]re Norway rats … things?”: Diversity versus generality in the use of albino rats in experiments on development and sexuality. Journal of the History of Biology, 34(2): 287–314, June 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010398116188;
http://link.springer.com/
content/pdf/10.1023/A:1010398116188.
pdf.

Logan:2002:BTW

[Log02] Cheryl A. Logan. Before there were standards: The role of test animals in the production of empirical generality in physiology. Journal of the History of Biology, 35 (2):329–363, June 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016036223348;
http://link.springer.com/
content/pdf/10.1023/A:1016036223348.
pdf.

Logan:2007:ORR

[Log07] Cheryl A. Logan. Overheated rats, race, and the double gland: Paul Kammerer, endocrinology and the problem of somatic induction. Journal of the History of Biology, 40 (4):683–725, December 2007.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9130-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9130-z.pdf.

Loison:2011:FRF

[Loi11] Laurent Loison. French roots of French neo–Lamarckisms, 1879–1985. Journal of the History of Biology, 44(4):713– 744, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9240-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9240-x.pdf.

Loison:2025:CCL

[Loi25] Laurent Loison. The contributionsand collapse-of Lamarckism in Pasteurian molecular biology: 2. Enzymatic adaptation, 1920–1965. Journal of the History of Biology, 58 (2):215–246, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-025-09818-7.

Long:2009:WMM

[Lon09] Tulley Long. William McElroy, the McCollum–Pratt institute, and the transformation of biology at Johns Hopkins, 1945–1960. Journal of the History of Biology, 42(4): 765–809, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9188-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9188-x.pdf.

Louca:2009:ETI

[Lou09]  Francisco Louçã.  Eman-

cipation through interaction — how eugenics and statistics converged and diverged. Journal of the History of Biology, 42(4):649– 684, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9167-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9167-7.pdf.

Love:2002:DCP

[Lov02] Alan C. Love. Darwin and Cirripedia prior to 1846: Exploring the origins of the barnacle research. Journal of the History of Biology, 35(2): 251–289, June 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016020816265;
http://link.springer.com/

                                        content/pdf/10.1023/A:1016020816265.
                                        pdf.

Loveland:2006:ADA

[Lov06]  Jeff Loveland.  Another

Daubenton, another Histoire naturelle. Journal of the History of Biology, 39(3):457– 491, September 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-6533-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-6533-6.pdf.

Lowy:1994:ESC

[Löw94] Ilana Löwy. Experimental systems and clinical practices: Tumor immunology and cancer immunotherapy, 1895–1980. Journal of the History of Biology, 27(3): 403–435, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058992.

Lowy:1997:EIY

[Löw97] Ilana Löwy. Epidemiology, immunology, and yellow fever: The Rockefeller Foundation in Brazil, 1923–1939. Journal of the History of Biology, 30 (3):397–417, September 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1023/A:1004295806141;
http://link.springer.com/
content/pdf/10.1023/A:1004295806141.
pdf.

Lozano:2019:BRS

[Loz19] Cristina Moreno Lozano.

Book review: Sara Kenney and John Watkiss, Surgeon X, Vol. 1–6: The Path of Most Resistance (London: Image Comics, 2017), 218 pp., $14.99, ISBN-10: 1-5343-0154-2. Journal of the History of Biology, 52(1): 219–222, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9557-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9557-z.pdf.

Lederman:1993:OOV

[LT93] Muriel Lederman and Sue A.

Tolin. OVATOOMB: Other viruses and the origins of molecular biology. Journal of the History of Biology, 26(2): 239–254, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061968.

Ludmerer:1969:AGE

[Lud69] Kenneth M. Ludmerer. American geneticists and the eugenics movement: 1905–1935. Journal of the History of Biology, 2(2):337–362, ???? 1969. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00125023.

Luedee:2021:LBN

[Lue21] Jonathan Luedee. Locating the boundaries of the Nuclear North: Arctic biology, contaminated caribou, and the problem of the threshold. Journal of the History of Biology, 54(1):67–93, April 2021. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09631-y.

Luk:2015:BBM

[Luk15] Yi Lai Christine Luk. Building biophysics in mid-century China: The University of Science and Technology of China. Journal of the History of Biology, 48(2):201–235, May 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9398-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9398-8.pdf.

Luk:2018:BR

[Luk18] Christine Y. L. Luk. Book review. Journal of the History of Biology, 51(1):179– 182, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9505-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
018-9505-3.pdf.

Luk:2021:CZW

[Luk21] Christine Y. L. Luk. Chen Ziying and Woods Hole: Bringing the marine biological laboratory to Amoy, China, 1930–1936. Journal of the History of Biology, 54(2): 151–173, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09636-7.

See [Luk22].

Luk:2022:CCZ

[Luk22] Christine Y. L. Luk. Correction to: [Chen] Ziying and Woods Hole: Bringing the Marine Biological Laboratory to Amoy, China, 1930–1936. Journal of the History of Biology, 55(1):207, March 2022. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09653-6.

See [Luk21].

Lussenhop:1974:VHD

[Lus74]  John Lussenhop.  Victor

Hensen and the development of sampling methods in ecology. Journal of the History of Biology, 7(2): 319–337, Fall 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351208.

Lustig:2000:SDE

[Lus00] A. J. Lustig. Sex, death, and evolution in proto- and metazoa, 1876–1913. Journal of the History of Biology, 33(2):221– 246, September 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004740325150;
http://link.springer.com/
content/pdf/10.1023/A:1004740325150.
pdf.

Lustig:2001:ERB

[Lus01]  A. J. Lustig.  Essay re-

view: Botanists sow, historians reap. Journal of the History of Biology, 34(3):581– 591, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012982808852;
http://link.springer.com/
content/pdf/10.1023/A:1012982808852.
pdf.

Lynch:2009:BR

[Lyn09] John M. Lynch. Book review. Journal of the History of Biology, 42(3):617– 619, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9200-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9200-5.pdf.

Lynch:2010:BR

[Lyn10] John M. Lynch. Book review. Journal of the History of Biology, 43(3):609– 611, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9251-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9251-7.pdf.

Lynch:2019:DAR

[Lyn19] William T. Lynch. The domestication of animals and the roots of the Anthropocene. Journal of the History of Biology, 52(1):209– 217, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9555-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-9555-1.pdf.

Lyons:1993:THF

[Lyo93] Sherrie L. Lyons. Thomas Huxley: Fossils, persistence, and the argument from design.

Journal of the History of Biology, 26(3):545–569, Fall 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062061.

Lyons:1995:OHH

[Lyo95] Sherrie L. Lyons. The origins of T. H. Huxley’s saltationism: History in Darwin’s shadow. Journal of the History of Biology, 28(3):463–494, Fall 1995. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059389.

Macrakis:1993:SBB

[Mac93] Kristie Macrakis. The survival of basic biological research in National Socialist Germany. Journal of the History of Biology, 26(3):519–543, Fall 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062060.

Mitman:1994:JHB

[MAC^{+}94] Gregg Mitman, Garland E. Allen, Joseph Cain, Nancy G. Slack, Keith R. Benson, Lily E. Kay, and Alix Cooper. The J. H. B. bookshelf. Journal of the History of Biology, 27(2):359–373, Summer 1994. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF01062567.

MacLeod:2001:SBS

[Mac01a] Roy MacLeod. “Strictly for the birds”: Science, the military and the Smithsonian’s Pacific Ocean Biological Survey Program, 1963–1970. Journal of the History of Biology, 34(2):315–352, June 2001. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010371321083;
http://link.springer.com/
content/pdf/10.1023/A:1010371321083.
pdf.

MacPherson:2001:ERW

[Mac01b] Ryan Cameron MacPherson.

Essay review: When evolution became conversation: Vestiges of creation, its readers, and its respondents in Victorian Britain. Journal of the History of Biology, 34(3):565– 579, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012951807944;
http://link.springer.com/
content/pdf/10.1023/A:1012951807944.
pdf.

MacLea:2012:BR

[Mac12] Kyle S. MacLea. Book review.

Journal of the History of Biology, 45(2):357–360, May 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9320-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9320-1.pdf.

Macfarlane:2021:CWL

[Mac21a] R. Ashton Macfarlane. Correction to: Wild laboratories of climate change: Plants, phenology, and global warming, 1955–1980. Journal of the History of Biology, 54(2): 341–342, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09648-3.

See [Mac21b].

Macfarlane:2021:WLC

[Mac21b] R. Ashton Macfarlane. Wild laboratories of climate change: Plants, phenology, and global warming, 1955–1980. Journal of the History of Biology, 54 (2):311–340, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09643-8.

See correction [Mac21a].

Madison:2017:BR

[Mad17] Paige Madison. Book review. Journal of the History of Biology, 50(3):677– 679, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9482-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9482-y.pdf.

Madison:2020:CDR

[Mad20] Paige Madison. Characterized by darkness: Reconsidering the origins of the brutish Neanderthal. Journal of the History of Biology, 53(4):493– 519, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09623-4.

Maienschein:1993:WC

[Mai93] Jane Maienschein. Why collaborate? Journal of the History of Biology, 26(2):167– 183, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061964.

Maienschein:2001:CAH

[Mai01] Jane Maienschein. On cloning:

Advocating history of biology in the public interest. Journal of the History of Biology, 34 (3):423–432, December 2001. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012994614765;

                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1012994614765.
                                        pdf.

Maienschein:2009:CLJ

[Mai09] Jane Maienschein. Controlling life: From Jacques Loeb to regenerative medicine. Journal of the History of Biology, 42 (2):215–230, May 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9180-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9180-5.pdf.

Maienschein:2016:GAT

[Mai16] Jane Maienschein. Garland Allen, Thomas Hunt Morgan, and development. Journal of the History of Biology, 49(4): 587–601, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9426-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9426-3.pdf.

Maienschein:2017:BR

[Mai17] Jane Maienschein. Book review. Journal of the History of Biology, 50(1):223– 225, February 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1007/s10739-017-9466-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9466-y.pdf.

Maienschein:2023:EMM

[Mai23a] Jane Maienschein. Everett Mendelsohn at the MBL. Journal of the History of Biology, 56(4):??, December 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09754-4.

Maienschein:2023:GAL

[Mai23b] Jane Maienschein. Garland Allen’s last book project. Journal of the History of Biology, 56(2):239–250, August 2023. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09726-8.

Manganiello:2009:HWH

[Man09] Christopher J. Manganiello.

From a howling wilderness to howling safaris: Science, policy and red wolves in the American South. Journal of the History of Biology, 42(2): 325–359, May 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9176-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9176-6.pdf.

Manganaro:2013:BRT

[Man13] Christine Manganaro. Book review: Theodore W. Pietsch, Trees of Life: a Visual History of Evolution (Baltimore: The Johns Hopkins University Press, 2012). Journal of the History of Biology, 46(4): 761–763, November 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9369-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9369-5.pdf.

Manias:2015:BBI

[Man15]  Chris Manias.  Building

Baluchitherium and Indricotherium: Imperial and international networks in early-Twentieth Century paleontology. Journal of the History of Biology, 48(2): 237–278, May 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9395-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9395-y.pdf.

Marx:1992:EDA

[Mar92] Leo Marx. Environmental degradation and the ambiguous social role of science and technology. Journal of the History of Biology, 25(3):

449–468, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00352002.

Marks:2004:BRO

[Mar04] Jonathan Marks. Book review: Orin Starn, Ishi’s Brain: In Search of American’s Last “Wild Indian”. Journal of the History of Biology, 37(3):610–611, October 2004. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2099-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2099-y.pdf.

Marti:2009:BRF

[Mar09] Oscar R. Marti. Book review: Francesca Bordogna, William James at the Boundaries: Philosophy, Science, and the Geography of Knowledge (Chicago: University of Chicago Press, 2008), x + 382 pp., illus., $39.00. Journal of the History of Biology, 42 (4):833–836, November 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9215-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9215-y.pdf.

Markus:2011:MOT

[Mar11] Miles B. Markus. Malaria: Origin of the term “hypnozoite”. Journal of the History of Biology, 44(4):781– 786, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9239-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9239-3.pdf.

Maienschein:2021:MPF

[MAR21] Jane Maienschein, Garland E. Allen, and Karen Rader. In memory of Paul Farber (1944– 2021), third Editor of the Journal of the History of Biology. Journal of the History of Biology, 54(4):549– 550, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09664-3.

Martin:2023:YGL

[Mar23] Laura J. Martin. The Yale Geochronometric Laboratory and the rewriting of global environmental history. Journal of the History of Biology, 56 (1):35–63, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09704-0.

Marks:2024:MSE

[Mar24] Jonathan Marks. Marianne Sommer, The Diagrammatics of ‘Race:’ Visualizing Human Relatedness in the History of Physical, Evolutionary, and Genetic Anthropology, ca. 1770–2020, Cambridge, UK: Open Book Publishers, 2024, ISBN: 978-1-80511-265-5. Journal of the History of Biology, 57(4):617– 619, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09798-0.

Matijasic:1987:SRF

[Mat87] Thomas D. Matijasic. Science, religion, and the fossils at Big Bone Lick. Journal of the History of Biology, 20 (3):413–421, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139462.

Matchett:2006:OIA

[Mat06] Karin Matchett. At odds over inbreeding: An abandoned attempt at Mexico/ United States collaboration to “improve” Mexican corn, 1940–1950. Journal of the History of Biology, 39(2): 345–372, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-006-0007-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-0007-3.pdf.

Matta:2010:SGD

[Mat10] Christina Matta. Spontaneous generation and disease causation: Anton de Bary’s experiments with Phytophthora infestans and late blight of potato. Journal of the History of Biology, 43(3):459– 491, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9220-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9220-1.pdf.

Matlin:2016:HFM

[Mat16] Karl S. Matlin. The heuristic of form: Mitochondrial morphology and the explanation of oxidative phosphorylation. Journal of the History of Biology, 49(1):37– 94, February 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9418-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9418-3.pdf.

Matlin:2021:BRA

[Mat21] Karl S. Matlin. Book review:

Andrew S. Reynolds, The Third Lens: Metaphor and the Creation of Modern Cell Biology (Chicago: University of Chicago Press, 2018), 272 pp., $30.00 Paper, ISBN: 978-0-226-56312-1. Journal of the History of Biology, 54(3):545– 547, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-019-9558-y.

Mayr:1968:IBS

[May68] Ernst Mayr. Illiger and the biological species concept. Journal of the History of Biology, 1(2):163–178, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351918.

Mayr:1972:LR

[May72] Ernst Mayr. Lamarck revisited. Journal of the History of Biology, 5(1): 55–94, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113486.

Mayr:1973:RHG

[May73] Ernst Mayr. The recent historiography of genetics. Journal of the History of Biology, 6(1):125–154, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137301.

Mayr:1985:WE

[May85] Ernst Mayr. Weismann and evolution. Journal of the History of Biology, 18(3): 295–329, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138928.

Mayr:1992:DPD

[May92] Ernst Mayr. Darwin’s principle of divergence. Journal of the History of Biology, 25 (3):343–359, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00351996.

Mayr:1997:GES

[May97] Ernst Mayr. Goldschmidt and the evolutionary synthesis: a response. Journal of the History of Biology, 30 (1):31–33, March 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004262310629;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004262310629.
                                        pdf.

Maydom:2023:BRS

[May23]  Katrina Maydom.  Book

review: Stefanie Gänger, A Singular Remedy: Cinchona across the Atlantic World, 1751–1820, Cambridge: Cambridge University Press, 2020, ISBN: 978-1-108-84216-7, 300 pp. Journal of the History of Biology, 56 (4):??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09757-1.

Mazumdar:1975:PIL

[Maz75] Pauline M. H. Mazumdar.

The purpose of immunity: Landsteiner’s interpretation of the human isoantibodies. Journal of the History of Biology, 8(1):115–133, ???? 1975. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129598.

Maienschein:1986:P

[MCB86] Jane Maienschein, James P.

Collins, and John Beatty. Preface. Journal of the History of Biology, 19(2):167– 168, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138873.

McCullough:1969:WKB

[McC69] Dennis M. McCullough. W. K. Brooks’s role in the history of American biology. Journal of the History of Biology, 2 (2):411–438, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125026.

McEvoy:1992:SCP

[McE92]  Arthur F. McEvoy.  Sci-

ence, culture, and politics in U.S. natural resources management. Journal of the History of Biology, 25(3): 469–486, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00352003.

McGovern:2018:BRM

[McG18] Michael F. McGovern. Book review: Michelle Murphy, The Economization of Life (Durham: Duke University Press, 2017), 232 pp., 25 b&w illus., $24.95 pbk., ISBN: 978-0-8223-6345-3. Journal of the History of Biology, 51(3):613– 615, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9536-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9536-9.pdf.

McGuffie:2019:PEI

[McG19]  Joshua McGuffie.  Planet

Earth II: BBC (November 2016–January 2017) Television. Journal of the History of Biology, 52(3):505– 507, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09575-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-09575-4.pdf.

McGonigle:2021:BRD

[McG21] Ian McGonigle. Book review: Daniel Navon, Mobilizing Mutations: Human Genetics in the Age of Patient Advocacy (Chicago: University of Chicago Press, 2019), 348 pp. 11 halftones, $40.00 Cloth, ISBN: 978-0-226-63809-6. Journal of the History of Biology, 54(4):769– 770, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09659-0.

McLaughlin:2002:NB

[McL02]  Peter McLaughlin.  Nam-

ing biology. Journal of the History of Biology, 35 (1):1–4, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1023/A:1014535811678;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1014535811678.
                                        pdf.

Methot:2016:EIS

[MD16] Pierre-Olivier Méthot and Rachel Mason Dentinger. Ecology and infection: Studying host–parasite interactions at the interface of biology and medicine. Journal of the History of Biology, 49(2): 231–240, April 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9440-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9440-0.pdf.

Meli:2013:EME

[Mel13] Domenico Bertoloni Meli. Early modern experimentation on live animals. Journal of the History of Biology, 46 (2):199–226, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9327-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9327-7.pdf.

Mendelsohn:1968:EF

[Men68] Everett Mendelsohn. Editorial foreword. Journal of the History of Biology, 1(1): iii–iv, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00149772.

Mendelsohn:1998:NR

[Men98] Everett Mendelsohn. Notes to the reader. Journal of the History of Biology, 31(3):303– 304, September 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004340722396;
http://link.springer.com/
content/pdf/10.1023/A:1004340722396.
pdf.

Mendelsohn:2003:LC

[Men03] J. Andrew Mendelsohn. Lives of the cell. Journal of the History of Biology, 36(1): 1–37, March 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022591924692;
http://link.springer.com/
content/pdf/10.1023/A:1022591924692.
pdf.

Mendelsohn:2016:GAA

[Men16] Everett Mendelsohn. Garland Allen: An appreciation. Journal of the History of Biology, 49(4):583– 586, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-016-9459-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9459-2.pdf.

Mendelsohn:2017:RYJ

[Men17] Everett Mendelsohn. Reflections on 50 years of the journal of the history of biology. Journal of the History of Biology, 50(4):691– 693, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9492-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9492-9.pdf.

Messbarger:2019:BRD

[Mes19] Rebecca Messbarger. Book review: Domenico Bertoloni Meli, Visualizing Disease: The Art and History of Pathological Illustration (Chicago: The University of Chicago Press, 2018), 288 pp., 36 color plates, 36 halftones, $55.00 Cloth, ISBN: 978-0-226-11029-5. Journal of the History of Biology, 52(1): 207–208, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9546-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9546-7.pdf.

Methot:2014:BRS

[Mét14] Pierre-Olivier Méthot. Book review: Sebastian Normandin and Charles T. Wolfe, eds., Vitalism and the Scientific Image in Post-Enlightenment Life Science, 1800–2010. History, Philosophy & Theory of the Life Sciences, Volume 2 (Dordrecht, Heidelberg, New York, London: Springer, 2013), iv + 377 pp., illus., $179.00. Journal of the History of Biology, 47(3): 497–499, August 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9391-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9391-2.pdf.

Methot:2016:BTO

[Mét16] Pierre-Olivier Méthot. Bacterial transformation and the origins of epidemics in the interwar period: The epidemiological significance of Fred Griffith’s “transforming experiment”. Journal of the History of Biology, 49(2): 311–358, April 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9415-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9415-6.pdf.

Meulendijks:2021:CEE

[Meu21a] Max Meulendijks. Correction to: Eclipsing the Eclipse?: a neo-Darwinian historiography revisited. Journal of the History of Biology, 54(3): 445, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09654-5.

See [Meu21b].

Meulendijks:2021:EEN

[Meu21b] Max Meulendijks. Eclipsing the Eclipse?: a neo-Darwinian historiography revisited. Journal of the History of Biology, 54(3):403– 443, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09650-9.

See correction [Meu21a].

Mitman:1996:JHB

[MFM^{+}96] Gregg Mitman, Michael Fortun, Jordan D. Marché, Joseph E. Taylor, Mark V. Barrow, Jr., and Barbara Gutmann Rosenkrantz. The J. H. B. bookshelf. Journal of the History of Biology, 29(2): 309–325, Summer 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00571086.

Michod:1986:FAT

[Mic86] Richard E. Michod. On fitness and adaptedness and their role in evolutionary explanation. Journal of the History of Biology, 19(2):289–302, ???? 1986. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138880.

Millstein:2008:DDS

[Mil08] Roberta L. Millstein. Distinguishing drift and selection empirically: “The Great Snail Debate” of the 1950s. Journal of the History of Biology, 41 (2):339–367, June 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9145-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
007-9145-5.pdf.

Milam:2011:BRR

[Mil11] Erika Lorraine Milam. Book review: Rebecca M. Jordan-Young, Brain Storm: The Flaws in the Science of Sex Differences (Cambridge, MA: Harvard University Press, 2010), xiv + 394 pp., illus., $35.00. Journal of the History of Biology, 44(1):163– 165, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-011-9280-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9280-x.pdf.

Milam:2012:BR

[Mil12] Erika Lorraine Milam. Book review. Journal of the History of Biology, 45(2): 361–363, May 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9321-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9321-0.pdf.

Milam:2014:BR

[Mil14] Erika Lorraine Milam. Book review. Journal of the History of Biology, 47(2): 329–331, May 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9380-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9380-5.pdf.

Miracle:2008:STW

[Mir08] M. Eulàlia Gassó Miracle. The significance of Temminck’s work on biogeography: Early Nineteenth Century natural history in Leiden, The Netherlands. Journal of the History of Biology, 41(4):677– 716, December 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9158-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
008-9158-8.pdf.

Miracle:2011:WAT

[Mir11] M. Eulàlia Gassó Miracle.

On whose authority? Temminck’s debates on zoological classification and nomenclature: 1820–1850. Journal of the History of Biology, 44(3): 445–481, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9265-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9265-1.pdf.

Mirza:2022:ADU

[Mir22] Ali Mirza. Alexander Dalrymple, the utility of coral reefs, and Charles Darwin’s Structure and Distribution of Coral Reefs. Journal of the History of Biology, 55(4):827– 864, December 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09699-0.

Mitman:1988:PSC

[Mit88] Gregg Mitman. From the population to society: The cooperative metaphors of W.

C. Allee and A. E. Emerson. Journal of the History of Biology, 21(2):173–194, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146985.

Mitman:2003:RKL

[Mit03]  Gregg Mitman.  Robert

E. Kohler, landscapes and labscapes: Exploring the lab–field border in biology. Journal of the History of Biology, 36(3):599– 629, September 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004701.
91491.c9;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004701.
91491.c9.pdf.

Mitchell:2017:SCH

[Mit17] M. X. Mitchell. Screening out controversy: Human genetics, emerging techniques of diagnosis, and the origins of the social issues committee of the American society of human genetics, 1964–1973. Journal of the History of Biology, 50 (2):425–456, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9437-8;
 http://link.springer.com/

content/pdf/10.1007/s10739-
016-9437-8.pdf.

Mitchell:2021:CES

[Mit21] M. X. Mitchell. The cosmology of evidence: Suffering, science, and biological witness after Three Mile Island. Journal of the History of Biology, 54(1):7–29, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09627-8.

Manier:1969:CPI

[MKC69] Edward Manier, David B.

Kitts, and William Coleman. Commentary — Part II. Journal of the History of Biology, 2(1):207–221, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137275.

Maienschein:2010:EPI

[ML10] Jane Maienschein and Manfred D. Laubichler. The embryo project: An integrated approach to history, practices, and social contexts of embryo research. Journal of the History of Biology, 43(1): 1–16, February 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9204-1;
 http://link.springer.com/

                                        content/pdf/10.1007/s10739-
                                        009-9204-1.pdf.

Marcattilio-McCracken:2017:CCS

[MM17] Ry Marcattilio-McCracken. Cacogenic cartographies: Space and place in the eugenic family study. Journal of the History of Biology, 50(3): 497–524, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-016-9452-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        016-9452-9.pdf.

MacCord:2024:SRT

[MM24] Kate MacCord and Jane Maienschein. Studying regeneration through history as a way of looking forward. Journal of the History of Biology, 57(1):5–15, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09769-5.

Mayr:1969:CP

[MN69] Ernst Mayr and Ernest Nagel. Commentary — Part I. Journal of the History of Biology, 2(1):123–147, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137270.

Mohaupt:2018:BRK

[Moh18] Hillary Moses Mohaupt. Book review: Karen Joy Fowler, We Are All Completely Beside Ourselves: a Novel. (Marian Wood Books/Putnam, 2013 & 2014), 320 pp., Hardcover & Paperback, ISBN: 978-0-14-218082-2. Journal of the History of Biology, 51(2): 411–413, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9520-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9520-4.pdf.

Montgomery:1970:OST

[Mon70] William M. Montgomery. The origins of the spiral theory of phyllotaxis. Journal of the History of Biology, 3(2): 299–323, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137356.

Montgomery:2005:PPP

[Mon05] Georgina M. Montgomery.

Place, practice and primatology: Clarence Ray Carpenter, primate communication and the development of field methodology, 1931– 1945. Journal of the History of Biology, 38(3):495– 533, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-0553-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-0553-0.pdf.

Montgomery:2011:BR

[Mon11] Georgina M. Montgomery. Book review. Journal of the History of Biology, 44(4):793– 795, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9311-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9311-7.pdf.

Montgomery:2018:BRA

[Mon18] Georgina M. Montgomery. Book review: Abigail Woods, Michael Bresalier, Angela Cassidy, and Rachel Mason Dentinger, Animals and the Shaping of Modern Medicine: One Heath and Its Histories (Palgrave 2018), 288 pp., $40.00 Hardcover, ISBN 978-3-319-64336-6. Journal of the History of Biology, 51(3):605– 607, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9527-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9527-x.pdf.

Moore:1991:DDP

[Moo91] James Moore. Deconstructing Darwinism: The politics of evolution in the 1860s. Journal of the History of Biology, 24(3):353–408, ???? 1991. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00156318.

Mooney:1995:HJM

[Moo95] Susan M. Mooney. H. J.

Muller and R. A. Fisher on the evolutionary significance of sex. Journal of the History of Biology, 28(1): 133–149, Spring 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061249.

Morsink:1979:WAB

[Mor79] Johannes Morsink. Was Aristotle’s biology sexist? Journal of the History of Biology, 12 (1):83–112, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00128136.

Morange:1997:RVC

[Mor97a] Michel Morange. From the regulatory vision of cancer to the Oncogene paradigm, 1975–1985. Journal of the History of Biology, 30(1):

1–29, March 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004255309721;
http://link.springer.com/
content/pdf/10.1023/A:1004255309721.
pdf.

Morris:1997:LAA

[Mor97b] Paul J. Morris. Louis Agassiz’s arguments against Darwinism in his additions to the French translation of the essay on classification. Journal of the History of Biology, 30 (1):121–134, March 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004274730193;
http://link.springer.com/
content/pdf/10.1023/A:1004274730193.
pdf.

Morgan:1998:EZL

[Mor98] Gregory J. Morgan. Emile Zuckerkandl, Linus Pauling, and the molecular evolutionary clock, 1959–1965. Journal of the History of Biology, 31 (2):155–178, June 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004394418084;
http://link.springer.com/
content/pdf/10.1023/A:1004394418084.
pdf.

Morange:2012:BR

[Mor12] Michel Morange. Book review. Journal of the History of Biology, 45(4):717– 718, November 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9330-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9330-z.pdf.

Morris:2014:WKP

[Mor14] Matthew Morris. We know in part: James McCosh on evolution and Christian faith. Journal of the History of Biology, 47(3):363–410, August 2014. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9371-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9371-y.pdf.

Morange:2023:BRH

[Mor23] Michel Morange. Book review:

Hans-Jörg Rheinberger, Split & Splice: A Phenomenology of Experimentation, Chicago: University of Chicago Press, 2023, ISBN: 978-0-226-82532-8, 256 pp. Journal of the History of Biology, 56(3):575– 576, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10739-023-09737-5.

Moscoso:1998:MEU

[Mos98] Javier Moscoso. Monsters as evidence: The uses of the abnormal body during the early eighteenth century. Journal of the History of Biology, 31 (3):355–382, September 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004341624902;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004341624902.
                                        pdf.

Moore-Sheeley:2018:BR

[MS18] Kirsten Moore-Sheeley. Book review. Journal of the History of Biology, 51(1):187– 189, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9507-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9507-1.pdf.

Mendelsohn:1969:EF

[MSA69] Everett Mendelsohn, Dudley Shapere, and Garland E. Allen. Editors’ foreword. Journal of the History of Biology, 2(1):v–vi, ???? 1969. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF00137263.

Muka:2014:POC

[Muk14] Samantha K. Muka. Portrait of an outsider: Class, gender, and the scientific career of Ida M. Mellen. Journal of the History of Biology, 47(1): 29–61, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9354-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
013-9354-z.pdf.

Muka:2017:BR

[Muk17]  Samantha Muka.  Book

review. Journal of the History of Biology, 50(2): 461–463, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9474-y;
http://link.springer.com/
content/pdf/10.1007/s10739-
017-9474-y.pdf.

Muka:2019:BRM

[Muk19] Samantha Muka. Book review: Michel Anctil, Luminous Creatures: The History and Science of Light Production in Living Organisms (Montreal: McGill– Queen’s University Press, 2018), 486 pp., 56 photos, C$49.95 Cloth, ISBN: 978-0-7735-5312-5. Journal of the History of Biology, 52(1): 195–197, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9552-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9552-9.pdf.

Mullett:1969:MPG

[Mul69] Charles F. Mullett. Multum in Parvo: Gilbert White of Selborne. Journal of the History of Biology, 2(2): 363–389, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125024.

Munzer:1993:ABT

[Mun93] Stephen R. Munzer. Aristotle’s biology and the transplantation of organs. Journal of the History of Biology, 26 (1):109–129, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01060682.

Munz:2005:BBK

[Mun05] Tania Munz. The bee battles: Karl von Frisch, Adrian Wenner and the honey bee dance language controversy. Journal of the History of Biology, 38 (3):535–570, November 2005.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-0552-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
005-0552-1.pdf.

Munns:2023:DFE

[Mun23] David P. D. Munns. “Not by a decree of fate:” Ellen Richards, euthenics, and the environment in the progressive era. Journal of the History of Biology, 56(3):525– 557, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09733-9.

Mursinna:2023:MQS

[Mur23] Jordan Thomas Mursinna. A “Mean quarrelsome spirit:” controversy in British systematics, 1822–1836. Journal of the History of Biology, 56 (4):??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09743-7.

Nathan:2024:BRC

[Nat24] Geoffrey Nathan. Book review: Colin Webster, Tools and the Organism: Technology and the Body in Ancient Greek and Roman Medicine, Chicago: University of Chicago Press, 2023, ISBN: 978-0-226-82877-0, 320 pp. Journal of the History of Biology, 57(3):493– 496, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09789-1.

Negbi:1995:MFT

[Neg95] Moshe Negbi. Male and female in Theophrastus’s botanical works. Journal of the History of Biology, 28(2):317– 332, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059192.

Nelson:1978:CCC

[Nel78] Gareth Nelson. From Candolle to Croizat: Comments on the history of biogeography. Journal of the History of Biology, 11(2): 269–305, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00389302.

Nicholson:2014:RWL

[NG14] Daniel J. Nicholson and Richard Gawne. Rethinking Woodger’s legacy in the philosophy of biology. Journal of the History of Biology, 47 (2):243–292, May 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9364-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9364-x.pdf.

Nickrand:2011:BR

[Nic11] Jessica Nickrand. Book review. Journal of the History of Biology, 44(4):787– 789, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9305-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9305-5.pdf.

Nickelsen:2022:PBK

[Nic22] Kärin Nickelsen. Physicochemical biology and knowledge transfer: The study of the mechanism of photosynthesis between the two world wars. Journal of the History of Biology, 55(2):349– 377, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-019-9559-x.

Nyhart:2011:ICB

[NL11] Lynn K. Nyhart and Scott Lidgard. Individuals at the center of biology: Rudolf Leuckart’s Polymorphismus der Individuen and the ongoing narrative of parts and wholes. With an annotated translation. Journal of the History of Biology, 44(3): 373–443, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9268-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9268-6.pdf.

Norton:1973:BDD

[Nor73] B. J. Norton. The biometric defense of Darwinism. Journal of the History of Biology, 6(2):283–316, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00127611.

Novick:2016:OQS

[Nov16] Aaron Novick. On the origins of the quinarian system of classification. Journal of the History of Biology, 49(1): 95–133, February 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9419-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9419-2.pdf.

November:2018:MTM

[Nov18] Joseph November. More than Moore’s mores: Computers, genomics, and the embrace of innovation. Journal of the History of Biology, 51(4):807– 840, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9539-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9539-6.pdf.

Novick:2019:RCD

[Nov19] Aaron Novick. A reappraisal of Charles Darwin’s engagement with the work of William Sharp Macleay. Journal of the History of Biology, 52(2): 245–270, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9541-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9541-z.pdf.

Novick:2023:BRJ

[Nov23] Rose Novick. Book review:

James Elwick, Styles of Reasoning in the British Life Sciences: Shared Assumptions 1820–1858, Pittsburgh: University of Pittsburgh Press, 2020, ISBN: 978-0-8229-6634-0, 234 pp. Journal of the History of Biology, 56(3):563– 565, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10739-023-09738-4.

Nyhart:1992:JHB

[NSMB92] Lynn K. Nyhart, P. F. Stevens, Jane Maienschein, and Mark V. Barrow, Jr. The J. H. B. bookshelf. Journal of the History of Biology, 25 (3):497–504, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00352005.

Noguera-Solano:2009:DII

[NSRG09] Ricardo Noguera-Solano and Rosaura Ruiz-Gutiérrez. Darwin and inheritance: The influence of Prosper Lucas. Journal of the History of Biology, 42(4):685– 714, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9175-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9175-7.pdf.

Numbers:1979:WBE

[Num79] Ronald L. Numbers. William Beaumont and the ethics of human experimentation. Journal of the History of Biology, 12(1):113–135, ???? 1979. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00128137.

Nyhart:1995:ERB

[Nyh95]  Lynn Nyhart.  Essay re-

view: Biology and imperialism. Journal of the History of Biology, 28(3): 533–543, Fall 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059391.

Nyhart:2018:BRJ

[Nyh18] Lynn K. Nyhart. Book review: John H. Zammito, The Gestation of German Biology: Philosophy and Physiology from Stahl to Schelling (Chicago: University of Chicago Press, 2018), 523 pp., $45.00 Cloth, ISBN: 978-0-226-52079-7. Journal of the History of Biology, 51(3):593– 595, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9528-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9528-9.pdf.

Oakes:2015:AHB

[Oak15] Jason Oakes. Alliances in human biology: The Harvard Committee on Industrial Physiology, 1929–1939. Journal of the History of Biology, 48(3):365–390, August 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-014-9396-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9396-x.pdf.

Ogawa:2001:MMC

[Oga01] Mariko Ogawa. The mysterious Mr. Collins: Living for 140 years in Origin of Species. Journal of the History of Biology, 34(3):461– 479, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1012983831897;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1012983831897.
                                        pdf.

Ogilvie:2007:IEH

[Ogi07] Marilyn Bailey Ogilvie. Inbreeding, eugenics, and Helen Dean King (1869–1955). Journal of the History of Biology, 40(3):467–507, September 2007. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9117-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9117-1.pdf.

Olby:1971:SPW

[Olb71] Robert Olby. Schrödinger’s problem: What is life? Journal of the History of Biology, 4 (1):119–148, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356979.

Olby:1974:OMG

[Olb74] Robert Olby. The origins of molecular genetics. Journal of the History of Biology, 7 (1):93–100, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00179294.

OHara-May:1971:MMN

[OM71] Jane O’Hara-May. Measuring man’s needs. Journal of the History of Biology, 4(2): 249–273, ???? 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138313.

OMalley:2018:ESB

[O’M18] Maureen A. O’Malley. The experimental study of bacterial evolution and its implications for the modern synthesis of evolutionary biology. Journal of the History of Biology, 51(2):319–354, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9493-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9493-8.pdf.

Onaga:2010:BR

[Ona10a] Lisa Onaga. Book review. Journal of the History of Biology, 43(2):411–414, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9234-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9234-8.pdf.

Onaga:2010:TKV

[Ona10b] Lisa Onaga. Toyama Kametaro and Vernon Kellogg: Silkworm inheritance experiments in Japan, Siam, and the United States, 1900–1912. Journal of the History of Biology, 43(2):215–264, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9222-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9222-z.pdf.

Onaga:2023:BRJ

[Ona23] Lisa Onaga. Book review: Jeannie N. Shinozuka, Biotic Borders: Transpacific Plant and Insect Migration and the Rise of Anti-Asian Racism in America, 1890– 1950, Chicago: University of Chicago Press, 2022, 296 pp. Journal of the History of Biology, 56(4):??, December 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09759-z.

Ospovat:1977:LTC

[Osp77] Dov Ospovat. Lyell’s theory of climate. Journal of the History of Biology, 10 (2):317–339, Fall 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00572645.

Ospovat:1979:DAM

[Osp79] Dov Ospovat. Darwin after Malthus. Journal of the History of Biology, 12(2): 211–230, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00124192.

Ospovat:1980:GNS

[Osp80] Dov Ospovat. God and natural selection: The Darwinian idea of design. Journal of the History of Biology, 13(2): 169–194, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125743.

Oudshoorn:1990:ECS

[Oud90] Nelly Oudshoorn. Endocrinologists and the conceptualization of sex, 1920–1940. Journal of the History of Biology, 23(2):163–186, ???? 1990.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141469.

Outram:1986:ULG

[Out86] Dorinda Outram. Uncertain legislator: Georges Cuvier’s laws of nature in their intellectual context. Journal of the History of Biology, 19(3): 323–368, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138285.

Orel:1968:MPH

[OV68] V. Orel and M. Vávra. Mendel’s program for the hybrization of apple trees. Journal of the History of Biology, 1(2):219–224, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351921.

Padian:1995:MHL

[Pad95] Kevin Padian. A missing Hunterian lecture on vertebrae by Richard Owen, 1837. Journal of the History of Biology, 28(2):333– 368, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059193.

Palladino:1991:DEE

[Pal91] Paolo Palladino. Defining ecology: Ecological theories, mathematical models, and applied biology in the 1960s and 1970s. Journal of the History of Biology, 24(2): 223–243, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209430.

Palladino:2022:MSM

[Pal22] Paolo Palladino. The making of the Sambucana: On memory, the body, and the production of bioheritage. Journal of the History of Biology, 55 (4):725–749, December 2022. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09697-2.

Parascandola:1971:NSM

[Par71a] John Parascandola. Notes on source materials: The L. J. Henderson papers at Harvard. Journal of the History of Biology, 4(1):115– 118, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356978.

Parascandola:1971:OHC

[Par71b]  John Parascandola.  Or-

ganismic and holistic concepts in the thought of L. J. Henderson. Journal of the History of Biology, 4(1): 63–113, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356977.

Parshall:1982:VIS

[Par82] Karen Hunger Parshall. Varieties as incipient species: Darwin’s numerical analysis. Journal of the History of Biology, 15(2):191–214, ???? 1982. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00233014.

Park:2000:PAC

[Par00] Hee-Joo Park. The politics of anti-creationism: The committees of correspondence. Journal of the History of Biology, 33(2):349– 370, September 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004821731846;
http://link.springer.com/
content/pdf/10.1023/A:1004821731846.
pdf.

Park:2008:EVC

[Par08] Hyung Wook Park. Edmund Vincent Cowdry and the making of gerontology as a multidisciplinary scientific field in the United States. Journal of the History of Biology, 41 (3):529–572, September 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9152-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9152-1.pdf.

Park:2013:SGG

[Par13] Hyung Wook Park. Senescence, growth, and gerontology in the United States. Journal of the History of Biology, 46(4):631–667, November 2013. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9348-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9348-2.pdf.

Parolini:2015:EMS

[Par15] Giuditta Parolini. The emergence of modern statistics in agricultural science: Analysis of variance, experimental design and the reshaping of research at Rothamsted experimental station, 1919–1933. Journal of the History of Biology, 48(2):301–335, May 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9394-z;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9394-z.pdf.

Partridge:2018:DTT

[Par18]  Derek Partridge.  Dar-

win’s two theories, 1844 and 1859. Journal of the History of Biology, 51(3):563– 592, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9509-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9509-z.pdf.

Parry:2021:CBM

[Par21] Alexander Ian Parry. Catharine Beecher and the mechanical body: Physiology, evangelism, and American social reform from the Antebellum Period to the Gilded Age. Journal of the History of Biology, 54 (4):603–638, December 2021. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09658-1.

Passariello:2021:ERC

[Pas21] Alessandra Passariello. From entomological research to culturing tissues: Aron Moscona’s Investigative Pathway. Journal of the History of Biology, 54(4):555– 601, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09663-4.

Patterson:2009:GJC

[Pat09] Andrea Patterson. Germs and Jim Crow: The impact of microbiology on public health policies in progressive era American South. Journal of the History of Biology, 42(3): 529–559, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9164-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9164-x.pdf.

Paul:1983:WTF

[Pau83] Diane B. Paul. A war on two fronts: J. B. S. Haldane and the response to Lysenkoism in Britain. Journal of the History of Biology, 16 (1):1–37, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00186674.

Pauly:1984:AAB

[Pau84] Philip J. Pauly. The appearance of academic biology in late nineteenth-century America. Journal of the History of Biology, 17(3): 369–397, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00126369.

Paul:1987:OLM

[Pau87] Diane B. Paul. “Our load of mutations” revisited. Journal of the History of Biology, 20 (3):321–335, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00139458.

Paul:1988:SSF

[Pau88] Diane B. Paul. The selection of the “survival of the fittest”. Journal of the History of Biology, 21(3):411–424, ???? 1988. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00144089.

Pauly:1993:ERE

[Pau93] Philip J. Pauly. Essay review: The eugenics industry — growth or restructuring? Journal of the History of Biology, 26(1): 131–145, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01060683.

Pauly:1996:HDE

[Pau96] Philip J. Pauly. How did the effects of alcohol on reproduction become scientifically uninteresting? Journal of the History of Biology, 29 (1):1–28, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129695.

Paul:2016:RHA

[Pau16] Diane B. Paul. Reflections on the historiography of American eugenics: Trends, fractures, tensions. Journal of the History of Biology, 49(4): 641–658, December 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9442-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9442-y.pdf.

Paul:2023:MVE

[Pau23] Diane B. Paul. Book review: Marga Vicedo, Intelligent Love: The Story of Clara Park, Her Autistic Daughter, and the Myth of the Refrigerator Mother, Boston: Beacon Press, 2021, ISBN: 978-0-8070-5551-9, 272 pp. Journal of the History of Biology, 56(3):577–578, October 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09739-3.

Peirson:2017:QPF

[PBDL17] B. R. Erick Peirson, Erin Bottino, Julia L. Damerow, and Manfred D. Laubichler. Quantitative perspectives on fifty years of the journal of the history of biology. Journal of the History of Biology, 50(4): 695–751, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9499-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9499-2.pdf.

Pearce:2010:GCC

[Pea10] Trevor Pearce. “A great complication of circumstances”

— Darwin and the economy of nature. Journal of the History of Biology, 43(3): 493–528, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9205-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9205-0.pdf.

Peterson:2008:WBB

[Pet08] Erik L. Peterson. William Bateson from Balanoglossus to materials for the study of variation: The transatlantic roots of discontinuity and the (un)naturalness of selection. Journal of the History of Biology, 41(2):267–305, June 2008.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9137-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9137-5.pdf.

Peterson:2022:WDM

[PH22] Erik L. Peterson and Crystal Hall. “What is dead may not die”: Locating marginalized concepts among ordinary biologists. Journal of the History of Biology, 55(2): 219–251, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09618-1.

Pierrel:2012:RPL

[Pie12] Jérôme Pierrel. An RNA phage lab: MS2 in Walter Fiers’ Laboratory of Molecular Biology in Ghent, from genetic code to gene and genome, 1963–1976. Journal of the History of Biology, 45 (1):109–138, February 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9267-z;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9267-z.pdf.

Piel:2019:CSP

[Pie19] Helen Piel. Complicating the story of popular science: John Maynard Smith’s “Little Penguin” on the theory of evolution. Journal of the History of Biology, 52(3):371– 390, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9566-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-9566-y.pdf.

Pinero:1988:JBB

[Piñ88] José M. López Piñero. Juan Bautista Bru (1740–1799) and the description of the genus Megatherium. Journal of the History of Biology, 21(1): 147–163, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125797.

Pinar:2002:EMG

[Pin02] Susana Pinar. The emergence of modern genetics in Spain and the effects of the Spanish Civil War (1936–1939) on its development. Journal of the History of Biology, 35(1): 111–148, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014544520340;
http://link.springer.com/
content/pdf/10.1023/A:1014544520340.
pdf.

Pioero:1979:VMS

[Pió79] José M. López Pióero. The Vesalian movement in sixteenthcentury Spain. Journal of the History of Biology, 12 (1):45–81, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00128135.

Pireddu:2023:RBG

[Pir23] Pier Luigi Pireddu. The relationship between George Evelyn Hutchinson and Vladimir Ivanovic Vernadsky: Roots and consequences of a biogeochemical approach. Journal of the History of Biology, 56(2): 339–363, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09717-9.

Pireddu:2024:BDO

[Pir24] Pier Luigi Pireddu. A biogeographical debate at the origins of limnology in Switzerland and Italy: The issue over pelagic fauna between Pietro Pavesi and François–Alphonse Forel. Journal of the History of Biology, 57(4):507– 532, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09791-7.

Park:1995:JHB

[PKS^{+}95] Katharine Park, Elizabeth B. Kenney, Michael Seltzer, Joseph Cain, Mark V. Barrow, Jr., and Nancy Slack. The J. H. B. bookshelf. Journal of the History of Biology, 28 (3):551–563, Fall 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01059393.

Podani:2018:HFD

[PKS18] János Podani, Ádám Kun, and András Szilágyi. How fast does Darwin’s elephant population grow? Journal of the History of Biology, 51 (2):259–281, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9488-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9488-5.pdf.

Platt:1969:OEI

[Pla69] John Platt. Organism, environment, and intelligence as a system. Journal of the History of Biology, 2(1): 225–239, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137276.

Plutynski:2019:SPS

[Plu19]  Anya Plutynski.  Specia-

tion post synthesis: 1960– 2000. Journal of the History of Biology, 52(4):569– 596, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9512-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9512-4.pdf.

Podolsky:1996:RVO

[Pod96] Scott Podolsky. The role of the virus in origin-of-life theorizing. Journal of the History of Biology, 29(1): 79–126, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129697.

Porter:1979:CRP

[Por79] Charlotte M. Porter. The concussion of revolution: Publications and reform at the early Academy of Natural Sciences, Philadelphia, 1812– 1842. Journal of the History of Biology, 12(2):273– 292, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00124194.

Porter:1985:ERS

[Por85] Roy Porter. Essay review: Stone ages, old and new. Journal of the History of Biology, 18(3):433–438, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138932.

Porter:1993:ROD

[Por93] Duncan M. Porter. On the road to the origin with Darwin, Hooker, and Gray. Journal of the History of Biology, 26(1):1–38, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01060678.

Porter:1995:ERH

[Por95] Charlotte M. Porter. Essay review: The history of scientific illustration. Journal of the History of Biology, 28 (3):545–550, Fall 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01059392.

Porter:2004:BRJ

[Por04] Charlotte M. Porter. Book review: John C. Greene, American Science in the Age of Jefferson. Journal of the History of Biology, 37(3):604– 605, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2096-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2096-1.pdf.

Porter:2009:BRS

[Por09] Charlotte M. Porter. Book review: Shepard Krech III, Spirits of the Air: Birds & American Indians in the South (Athens, GA: University of Georgia Press, 2009) xvi + 245 pp. illus. Journal of the History of Biology, 42(4):837– 838, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9216-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9216-x.pdf.

Porter:2025:BRP

[Por25] Charlotte Porter. Book review: Patrick Spero, André Michaux and Thomas Jefferson and the Conspiracy of 1793, Charlottesville: University of Virginia Press, 2024, ISBN: 978-0-8139-5219-2, 334 pp. Journal of the History of Biology, 58(2): 309–312, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09824-9.

Powell:2018:BRB

[Pow18] Miles A. Powell. Book review: Ben A. Minteer, Jane Maienschein, and James P. Collins, The Ark and Beyond: The Evolution of Zoo and Aquarium Conservation (Chicago and London: University of Chicago Press, 2018), 528 pp., 51 halftones, 2 line drawings, 6 tables, $35.00 Paper, ISBN: 978-0-226-53832-7. Journal of the History of Biology, 51 (3):609–611, September 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9525-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9525-z.pdf.

Preus:1970:SPA

[Pre70] Anthony Preus. Science and philosophy in Aristotle’s generation of animals. Journal of the History of Biology, 3 (1):1–52, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00569305.

Preus:1977:GCA

[Pre77] Anthony Preus. Galen’s criticism of Aristotle’s conception theory. Journal of the History of Biology, 10(1): 65–85, ???? 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126095.

Prete:1990:CHB

[Pre90] Frederick R. Prete. The conundrum of the honey bees: One impediment to the publication of Darwin’s theory. Journal of the History of Biology, 23(2):271–290, ???? 1990. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141472.

Prebble:2001:POM

[Pre01] John N. Prebble. The philosophical origins of Mitchell’s chemiosmotic concepts. Journal of the History of Biology, 34(3):433–460, December 2001. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012946715673;
http://link.springer.com/
content/pdf/10.1023/A:1012946715673.
pdf.

Prebble:2013:CAB

[Pre13] John N. Prebble. Contrasting approaches to a biological problem: Paul Boyer, Peter Mitchell and the mechanism of the ATP synthase, 1961–1985. Journal of the History of Biology, 46(4):699– 737, November 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9343-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9343-7.pdf.

Priest:2024:HHW

[Pri24] Greg Priest. Hey hey we’re the monkeys! An essay review of Gowan Dawson’s Monkey to Man. Journal of the History of Biology, 57(3):477– 484, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09786-4.

Poczai:2022:MSS

[PSBS^{+}22] Péter Poczai, Jorge A. Santiago-Blay, Jiří Sekerák, István Bariska, and Attila T. Szabó. Mimush sheep and the spectre of inbreeding: Historical background for Festetics’s organic and genetic laws four decades before Mendel’s experiments in peas. Journal of the History of Biology, 55(3): 495–536, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09678-5.

Prete:1992:RSS

[PW92] Frederick R. Prete and M. Melissa Wolfe. Religious supplicant, seductive cannibal, or reflex machine? In search of the praying mantis. Journal of the History of Biology, 25(1):91– 136, Spring 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01947506.

Yen:2014:EAP

[pY14] Hsiao pei Yen. Evolutionary Asiacentrism, Peking Man, and the origins of Sinocentric ethno-nationalism. Journal of the History of Biology, 47(4): 585–625, November 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9381-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
014-9381-4.pdf.

Quinn:2017:CGR

[Qui17] Aleta Quinn. Charles Girard: Relationships and representation in Nineteenth Century systematics. Journal of the History of Biology, 50(3): 609–643, August 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9447-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
016-9447-6.pdf.

Raacke:1976:CGH

[Raa76] I. D. Raacke. “The die is cast” — “I am going home”: The appointment of Herbert McLean Evans as head of anatomy at Berkeley. Journal of the History of Biology, 9(2):301–322, ???? 1976. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209887.

Raby:2011:BRP

[Rab11] Megan Raby. Book review: Paul D. Brinkman, The Second Jurassic Dinosaur Rush: Museums and Paleontology in America at the Turn of the Twentieth Century (Chicago: University of Chicago Press, 2010), xiv + 312 pp., illus., $49.00. Journal of the History of Biology, 44(1):155– 157, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9278-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9278-4.pdf.

Rader:1998:MPM

[Rad98] Karen A. Rader. “The mouse people”: Murine genetics work at the Bussey Institution, 1909–1936. Journal of the History of Biology, 31 (3):327–354, September 1998. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004389408064;
http://link.springer.com/
content/pdf/10.1023/A:1004389408064.
pdf.

Rader:2004:BRT

[Rad04] Karen Rader. Book review:

Tim Birkhead, A Brand New Bird: How Two Amateur Geneticists Created the First Genetically Engineered Animal. Journal of the History of Biology, 37(3):591– 593, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2089-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2089-0.pdf.

Radick:2005:PLP

[Rad05] Gregory Radick. Primate language and the playback experiment, in 1890 and 1980. Journal of the History of Biology, 38(3):461– 493, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-0554-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-0554-z.pdf.

Rader:2006:AHP

[Rad06] Karen A. Rader. Alexander Hollaender’s postwar vision for biology: Oak Ridge and beyond. Journal of the History of Biology, 39(4):685– 706, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9109-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9109-1.pdf.

Radick:2007:ERE

[Rad07]  Gregory Radick.  “Essay

review: The Ethologist’s world”. Journal of the History of Biology, 40(3):565– 575, September 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9140-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9140-x.pdf.

Radick:2012:BR

[Rad12] Gregory Radick. Book review. Journal of the History of Biology, 45(3):575– 578, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9331-y;
                                         http://link.springer.com/

content/pdf/10.1007/s10739-
012-9331-y.pdf.

Rader:2022:REM

[Rad22] Karen A. Rader. Reflections on Making Mice (2004). Journal of the History of Biology, 55(1):29–33, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09672-x.

Radick:2023:BRN

[Rad23] Gregory Radick. Book review: Nathan Crowe, Forgotten Clones: The Birth of Cloning and the Biological Revolution, Pittsburgh: University of Pittsburgh Press, 2021, ISBN: 978-0-8229-4627-4, 299 pp. Journal of the History of Biology, 56(3):559– 561, October 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09740-w.

Rader:2025:JCS

[Rad25] Karen Rader. John Calhoun’s ‘strange’ rodent tale? Constructing and consuming interdisciplinary pasts and futures. Journal of the History of Biology, 58(1):95– 132, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09813-y.

Rainger:1985:PPC

[Rai85] Ronald Rainger. Paleontology and philosophy: a critique. Journal of the History of Biology, 18(2):267–287, ???? 1985. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00120112.

Rainger:1989:WUW

[Rai89] Ronald Rainger. What’s the use: William King Gregory and the functional morphology of fossil vertebrates. Journal of the History of Biology, 22(1):103–139, ???? 1989. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00209605.

Rainger:2001:ERI

[Rai01] Ronald Rainger. Essay review: Improving Americans. Journal of the History of Biology, 34(3):557– 564, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1012986909761;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1012986909761.
                                        pdf.

Rainger:2003:AIL

[Rai03] Ronald Rainger. Adaptation and the importance of local culture: Creating a research school at the Scripps Institution of Oceanography. Journal of the History of Biology, 36 (3):461–500, September 2003. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004570.
20057.93;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004570.
20057.93.pdf.

Rasmussen:1995:MSP

[Ras95] Nicolas Rasmussen. Mitochondrial structure and the practice of cell biology in the 1950s. Journal of the History of Biology, 28(3): 381–429, Fall 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059387.

Rasmussen:1999:FPT

[Ras99]  Nicolas Rasmussen.  The

forgotten promise of thiamin: Merck, Caltech biologists, and plant hormones in a 1930s biotechnology project. Journal of the History of Biology, 32(2):245– 261, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004635500408;
http://link.springer.com/
content/pdf/10.1023/A:1004635500408.
pdf.

Rasmussen:2022:RDA

[Ras22] Nicolas Rasmussen. René Dubos, the autochthonous flora, and the discovery of the microbiome. Journal of the History of Biology, 55(3):537– 558, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09692-7.

Rasmussen:2024:BRA

[Ras24a] Nicolas Rasmussen. [Book review:] Alfredo Morabia, The Public Health Approach: Population Thinking from the Black Death to COVID-19, Baltimore: Johns Hopkin University Press, 2023, ISBN: 978-1-4214-4678-3, 232 pp. Journal of the History of Biology, 57(2):343–344, June 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09779-3.

Rasmussen:2024:BRJ

[Ras24b] Nicolas Rasmussen. Book review: Janina Wellman, Biological Motion: A History of Life, New York: Zone Books, 2024, ISBN: 978-1-942130-81-9, pp. 336. Journal of the History of Biology, 57(3):497– 498, September 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10739-024-09790-8.

Rasmussen:2025:SCE

[Ras25] Nicolas Rasmussen. Situating commemoration: an editorial introduction. Journal of the History of Biology, 58 (1):1–3, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09811-0.

Raymer:2013:MHT

[Ray13] Emilie J. Raymer. A man of his time: Thorstein Veblen and the University of Chicago Darwinists. Journal of the History of Biology, 46(4):669– 698, November 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9342-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9342-8.pdf.

Ray:2019:BRS

[Ray19] Sara Ray. Book review: Stefanie Buchenau and Roberto Lo Presti, eds.: Human and Animal Cognition in Early Modern Philosophy and Medicine, University of Pittsburg Press, Pittsburgh, 2017, 354 pp., ISBN: 978-0-8229-4472-0. Journal of the History of Biology, 52(2): 359–360, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-9561-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-9561-3.pdf.

Ray:2022:MMA

[Ray22] Sara Ray. From monsters to malformations: Anatomical preparations as objects of evidence for a developmental paradigm of embryology, 1770–1850. Journal of the History of Biology, 55 (1):35–57, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09670-z.

Roe:1984:JHBb

[RBFC84] Shirley A. Roe, Keith R. Benson, Georgiana Feldberg, and Pietro Corsi. The J. H. B. bookshelf. Journal of the History of Biology, 17(2): 295–301, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00143737.

Roe:1987:JHBa

[RBHK87] Shirley A. Roe, Keith R. Benson, Joy Harvey, and Sharon E. Kingsland. The J. H. B. bookshelf. Journal of the History of Biology, 20 (1):131–135, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125261.

Roe:1986:JHBc

[RBK^{+}86] Shirley A. Roe, Keith R.

Benson, Sharon Kingsland, Eugene Cittadino, and Jane Maienschein. The J. H. B. bookshelf. Journal of the History of Biology, 19(3): 489–494, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138288.

Roe:1990:JHBb

[RBWB90] Shirley A. Roe, Keith R. Benson, Douglas R. Weiner, and Janet Browne. The J. H. B. bookshelf. Journal of the History of Biology, 23(3): 529–537, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136383.

Robin:2011:IEH

[RC11] Libby Robin and Jane Carruthers. Introduction: Environmental history and the history of biology. Journal of the History of Biology, 44 (1):1–14, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-010-9242-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9242-8.pdf.

Roe:1986:JHBb

[RCK^{+}86] Shirley A. Roe, Eugene Cittadino, Sharon E. Kingsland, Janet Browne, Ronald Rainger, S. A. R., and Keith R. Benson. The J. H. B. bookshelf. Journal of the History of Biology, 19(2): 313–322, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138882.

Rehbock:1979:EDN

[Reh79] Philip F. Rehbock. The early dredgers: “naturalizing” in British seas, 1830–1850. Journal of the History of Biology, 12(2):293–368, ???? 1979. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00124195.

Reif:1986:SMT

[Rei86] Wolf-Ernst Reif. The search for a macroevolutionary theory in German paleontology. Journal of the History of Biology, 19(1):79– 130, Spring 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF00346618.

Rensing:2012:BR

[Ren12] Susan Rensing. Book review. Journal of the History of Biology, 45(1):167– 169, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9314-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9314-4.pdf.

Reynolds:2008:AEC

[Rey08] Andrew Reynolds. Amoebae as exemplary cells: The protean nature of an elementary organism. Journal of the History of Biology, 41(2): 307–337, June 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9142-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9142-8.pdf.

Rajagopalan:2018:VCT

[RF18] Ramya M. Rajagopalan and Joan H. Fujimura. Variations on a chip: Technologies of difference in human genetics research. Journal of the History of Biology, 51(4):841– 873, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9543-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9543-x.pdf.

Richmond:1998:JHB

[RFL^{+}98] Marsha L. Richmond, Paul Lawrence Farber, Hannah Landecker, Vassiliki Betty Smocovitis, Eileen Crist, Chris Young, and Sara F. Tjossem. The J. H. B. bookshelf. Journal of the History of Biology, 31(3):447– 461, September 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004397709881;
http://link.springer.com/
content/pdf/10.1023/A:1004397709881.
pdf.

Reeds:1976:JHB

[RG76] Karen M. Reeds and Thomas F. Glick. The J. H. B. bookshelf. Journal of the History of Biology, 9(2):323–327, ???? 1976. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209888.

Roe:1981:JHB

[RGH^{+}81] Shirley A. Roe, Thomas F. Glick, Joy Harvey, F. Weiling, and John Scarborough. The J. H. B. bookshelf. Journal of the History of Biology, 14 (2):355–362, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141099.

Roll-Hansen:1984:RJH

[RH84]  Nils Roll-Hansen.  E. S.

Russell and J. H. Woodger: The failure of two twentiethcentury opponents of mechanistic biology. Journal of the History of Biology, 17(3): 399–428, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126370.

Roll-Hansen:2009:SWJ

[RH09] Nils Roll-Hansen. Sources of Wilhelm Johannsen’s genotype theory. Journal of the History of Biology, 42(3): 457–493, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9166-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
008-9166-8.pdf.

Rheinberger:1993:EOE

[Rhe93] Hans-Jörg Rheinberger. Experiment and orientation: Early systems of in vitro protein synthesis. Journal of the History of Biology, 26 (3):443–471, Fall 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062057.

Rheinberger:1995:MRS

[Rhe95] Hans-Jörg Rheinberger. From microsomes to ribosomes: “strategies” of “representation”. Journal of the History of Biology, 28(1):49– 89, Spring 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061246.

Rheinberger:1996:CES

[Rhe96] Hans-Jörg Rheinberger. Comparing experimental systems: Protein synthesis in microbes and in animal tissue at Cambridge (Ernest F. Gale) and at the Massachusetts General Hospital (Paul C. Zamecnik), 1945–1960. Journal of the History of Biology, 29(3): 387–416, ???? 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127381.

Rheinberger:2000:EED

[Rhe00] Hans-Jörg Rheinberger. Ephestia: The experimental design of Alfred Kühn’s physiological developmental genetics. Journal of the History of Biology, 33(3):535– 576, December 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004858314375;
http://link.springer.com/
content/pdf/10.1023/A:1004858314375.
pdf.

Rhizopoulou:2025:PPP

[Rhi25] Sophia Rhizopoulou. Plants, pictures, and people. Journal of the History of Biology, 58 (2):275–296, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09822-x.

Rhodes:1987:DGL

[Rho87] Frank H. T. Rhodes. Darwinian gradualism and its limits: The development of Darwin’s views on the rate and pattern of evolutionary change. Journal of the History of Biology, 20(2):139– 157, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138435.

Rainger:1997:JHB

[RHW^{+}97] Ronald Rainger, Joy Harvey, Mary P. Winsor, Joe Cain, and Keith R. Benson. The J. H. B. bookshelf. Journal of the History of Biology, 30 (2):303–315, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004208323283;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004208323283.
                                        pdf.

Richards:1981:IIB

[Ric81] Robert J. Richards. Instinct and intelligence in British natural theology: Some contributions to Darwin’s theory of the evolution of behavior. Journal of the History of Biology, 14 (2):193–230, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00141092.

Richards:1989:MAR

[Ric89a]  Evelleen Richards.  The

“moral anatomy” of Robert Knox: The interplay between biological and social thought in Victorian scientific naturalism. Journal of the History of Biology, 22 (3):373–436, Fall 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00401576.

Richmond:1989:PPM

[Ric89b] Marsha L. Richmond. Protozoa as precursors of metazoa: German cell theory and its critics at the turn of the century. Journal of the History of Biology, 22(2):

243–276, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139514.

Richmond:2000:HHC

[Ric00] Marsha L. Richmond. T. H.

Huxley’s criticism of German cell theory: An epigenetic and physiological interpretation of cell structure. Journal of the History of Biology, 33(2):247– 289, September 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004881730937;
http://link.springer.com/
content/pdf/10.1023/A:1004881730937.
pdf.

Richards:2004:MRD

[Ric04] Robert J. Richards. Michael Ruse’s design for living. Journal of the History of Biology, 37(1):25–38, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020388.
99933.5b;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020388.
99933.5b.pdf.

Richmond:2006:DHF

[Ric06] Marsha L. Richmond. “The ‘domestication’ of heredity: The familial organization of geneticists at Cambridge University, 1895–1910”. Journal of the History of Biology, 39 (3):565–605, September 2006. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-5431-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-5431-7.pdf.

Richmond:2007:MWO

[Ric07] Marsha L. Richmond. Muriel Wheldale Onslow and early biochemical genetics. Journal of the History of Biology, 40 (3):389–426, September 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9134-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9134-8.pdf.

Richmond:2009:BR

[Ric09] Marsha L. Richmond. Book review. Journal of the History of Biology, 42(3):613– 616, August 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9199-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9199-7.pdf.

Richmond:2010:BR

[Ric10] Marsha Richmond. Book review. Journal of the History of Biology, 43(3):617– 619, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9253-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9253-5.pdf.

Richmond:2011:BR

[Ric11] Marsha L. Richmond. Book review. Journal of the History of Biology, 44(2): 357–360, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9294-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9294-4.pdf.

Richmond:2023:RGE

[Ric23] Marsha L. Richmond. Remembering Garland Edward Allen, III (1936–2023), second Editor of Journal of the History of Biology. Journal of the History of Biology, 56(2): 219–226, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09724-w.

Richards:2024:WLT

[Ric24] Eric J. Richards. William Lawrence Tower’s beetles: Experimental evolution and the manipulation of inheritance. Journal of the History of Biology, 57(2):173–206, June 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09771-x.

Richmond:2025:BRS

[Ric25] Marsha L. Richmond. Book review: Savithri Preetha Nair, Chromosome Woman, Nomad Scientist: E. K. Janaki Ammal, A Life 1897–1984, London and New York: Routledge/Taylor and Francis, 2023, ISBN: 978-1-032-03548-2, 611 pp. Journal of the History of Biology, 58(2): 297–304, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09825-8.

Rieppel:1988:RLP

[Rie88] Olivier Rieppel. The reception of Leibniz’s philosophy in the writings of Charles Bonnet (1720–1793). Journal of the History of Biology, 21(1): 119–145, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125796.

Rieppel:1990:SFF

[Rie90] Olivier Rieppel. Structuralism, functionalism, and the four Aristotelian causes. Journal of the History of Biology, 23(2):291–320, ???? 1990. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141473.

Rieppel:2012:KBN

[Rie12] Olivier Rieppel. Karl Beurlen (1901–1985), nature mysticism, and Aryan paleontology. Journal of the History of Biology, 45(2):253–299, May 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9283-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9283-7.pdf.

Rieppel:2020:MP

[Rie20] Olivier Rieppel. Morphology and phylogeny. Journal of the History of Biology, 53 (2):217–230, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09600-x.

Rinard:1981:POI

[Rin81] Ruth G. Rinard. The problem of the organic individual: Ernst Haeckel and the development of the biogenetic law.

Journal of the History of Biology, 14(2):249–275, ???? 1981. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141094.

Rinard:1988:NLT

[Rin88]  R. G. Rinard.  Neo–

Lamarckism and technique: Hans Spemann and the development of experimental embryology. Journal of the History of Biology, 21(1): 95–118, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125795.

Rizhinashvili:2020:PHU

[Riz20] Alexandra Rizhinashvili. Production hydrobiology in the USSR under the pressure of Lysenkoism: Vladimir I. Zhadin’s forgotten theory of biological productivity (1940). Journal of the History of Biology, 53(1): 105–139, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09590-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
019-09590-5.pdf.

Roe:1989:JHBa

[RKMB89] Shirley A. Roe, Sharon E. Kingsland, Jane Maienschein, and Barbara G. Beddall. The J. H. B. bookshelf. Journal of the History of Biology, 22 (1):177–184, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209608.

Rheinberger:1984:DEN

[RM84] Hans-Jörg Rheinberger and Peter McLaughlin. Darwin’s experimental natural history. Journal of the History of Biology, 17(3):345–368, ???? 1984. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126368.

Roe:1988:JHBc

[RMR^{+}88] Shirley A. Roe, Jane Maienschein, Ronald Rainger, Elizabeth B. Keeney, and Donald Worster. The J. H. B. bookshelf. Journal of the History of Biology, 21(3): 521–526, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144094.

Rao:2011:JBH

[RN11] Veena Rao and Vidyanand Nanjundiah. J. B. S. Haldane, Ernst Mayr and the beanbag genetics dispute. Journal of the History of Biology, 44 (2):233–281, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9229-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9229-5.pdf.

Robin:2011:HRP

[Rob11] Nicolas Robin. Heritage of the romantic philosophy in post– Linnaean botany Reichenbach’s reception of Goethe’s metamorphosis of plants as a methodological and philosophical framework. Journal of the History of Biology, 44 (2):283–304, May 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9231-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9231-y.pdf.

Roe:1979:REC

[Roe79] Shirley A. Roe. Rationalism and embryology: Caspar Friedrich Wolff’s theory of epigenesis. Journal of the History of Biology, 12 (1):1–43, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00128134.

Roe:1982:JHB

[Roe82] Shirley A. Roe. The J. H. B. bookshelf. Journal of the History of Biology, 15(2): 317–323, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00233020.

Roe:1983:JHBb

[Roe83] Shirley A. Roe. The J. H.

B. bookshelf. Journal of the History of Biology, 16 (3):441–448, Fall 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00582411.

Roe:1984:JHBa

[Roe84] Shirley A. Roe. The J. H.

B. bookshelf. Journal of the History of Biology, 17(1): 147–152, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00397506.

Roe:1986:JHBa

[Roe86] Shirley A. Roe. The J. H.

B. bookshelf. Journal of the History of Biology, 19(1): 155–166, Spring 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00346620.

Roe:1989:JHBb

[Roe89] Shirley A. Roe. The J. H.

B. bookshelf. Journal of the History of Biology, 22 (3):507–511, Fall 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00401581.

Roe:1990:JHBa

[Roe90] Shirley A. Roe. The J. H. B. bookshelf. Journal of the History of Biology, 23(2): 335–339, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141476.

Roosth:2017:BR

[Roo17] Sophia Roosth. Book review. Journal of the History of Biology, 50(4):989– 991, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9496-5;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9496-5.pdf.

Ross:2018:BPD

[Ros18] Christian Ross. Brad Peyton, dir. Rampage. 2018. Film. Journal of the History of Biology, 51(4):897– 899, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9532-0;

 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9532-0.pdf.

Rothschild:1989:PPP

[Rot89] Lynn J. Rothschild. Protozoa, protista, protoctista: What’s in a name? Journal of the History of Biology, 22 (2):277–305, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139515.

Roe:1987:JHBc

[RPCH87] Shirley A. Roe, Philip J.

Pauly, John F. Cornell, and Joy Harvey. The J. H. B. bookshelf. Journal of the History of Biology, 20(3): 423–430, ???? 1987. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00139463.

Renner:2023:MRS

[RPM23] Susanne S. Renner, Ulrich Päßler, and Pierre Moret. “My reputation is at stake.” Humboldt’s Mountain Plant Geography in the making (1803–1825). Journal of the History of Biology, 56(1): 97–124, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09705-z.

Rader:2018:IE

[RR18] Karen Rader and Marsha Richmond. Inaugural editorial. Journal of the History of Biology, 51(1): 1–3, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9503-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9503-x.pdf.

Rader:2020:EMP

[RR20a] Karen Rader and Marsha Richmond. 2020 Everett Mendelsohn Prize. Journal of the History of Biology, 53 (1):1–3, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09593-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        020-09593-7.pdf.

Richmond:2020:AJT

[RR20b] Marsha Richmond and Karen Rader. Announcement of JHB topical collections. Journal of the History of Biology, 53 (3):319–320, September 2020. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09614-5.

Rader:2021:EMP

[RR21a] Karen Rader and Marsha Richmond. 2021 Everett Mendelsohn Prize. Journal of the History of Biology, 54 (2):147–149, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09626-9.

Rader:2021:JND

[RR21b] Karen Rader and Marsha Richmond. JHB’s “New developments in Darwin studies?” Redux. Journal of the History of Biology, 54(3):343– 344, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09656-3.

Rader:2022:EMP

[RR22a] Karen Rader and Marsha Richmond. 2022 Everett Mendelsohn Prize. Journal of the History of Biology, 55 (1):1–2, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09668-7.

Rader:2022:RWH

[RR22b] Karen Rader and Marsha Richmond. Reading and writing the history of biology at JHB. Journal of the History of Biology, 55(4):613– 614, December 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09700-w.

Richmond:2022:EJD

[RR22c] Marsha Richmond and Karen Rader. JHB ’s Darwin collection. Journal of the History of Biology, 55(2):379– 380, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09687-4.

Richmond:2023:EMP

[RR23] Marsha Richmond and Karen Rader. 2023 Everett Mendelsohn Prize. Journal of the History of Biology, 56(2): 211–213, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09701-9.

Roe:1985:JHB

[RRC^{+}85] Shirley A. Roe, Ronald Rainger, John F. Cornell, James J. Bono, Pietro Corsi, and William J. Haas. The J. H. B. bookshelf. Journal of the History of Biology, 18 (3):439–446, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138933.

Roe:1988:JHBa

[RRCB88] Shirley A. Roe, Ronald Rainger, Ralph Colp, and Keith R. Benson. The J. H. B. bookshelf. Journal of the History of Biology, 21(1): 165–171, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125798.

Roe:1987:JHBb

[RRE^{+}87] Shirley A. Roe, S. A. R., John T. Edsall, Muriel L. Blaisdell, and John F. Cornell. The J. H. B. bookshelf. Journal of the History of Biology, 20(2):281–287, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138438.

Russell:2005:BR

[RRM^{+}05] Edmund Russell, Mark C.

Russell, Jenny Marie, Jane Maienschein, Kim Fortun, Robert Olby, Joel B. Hagen, Richard W. Burkhardt, Charles H. Smith, Joe Cain, Keith R. Benson, Ulrike Hillemann, Peter L. Schmitthenner, Mark V. Barrow, Jr., and Ryan C. MacPherson. Book reviews. Journal of the History of Biology, 38(1): 153–184, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-004-6859-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-6859-5.pdf.

Roe:1983:JHBa

[RS83] Shirley A. Roe and John Scarborough. The J. H. B. bookshelf. Journal of the History of Biology, 16(1): 177–183, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00186680.

Rasmussen:2023:IE

[RS23] Nicolas Rasmussen and Vassiliki Betty Smocovitis. Inaugural editorial. Journal of the History of Biology, 56 (1):1–3, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09713-z.

Rasmussen:2024:E

[RS24] Nicolas Rasmussen and Vassiliki Betty Smocovitis. Editorial. Journal of the History of Biology, 57(1): 3–4, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09768-6.

Roe:1988:JHBb

[RSBM88] Shirley A. Roe, James A. Secord, Keith R. Benson, and Jane Maienschein. The J. H. B. bookshelf. Journal of the History of Biology, 21(2): 351–356, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146993.

Rudolph:1990:ISD

[Rud90] Emanuel D. Rudolph. Isaac Sprague, “delineator and naturalist”. Journal of the History of Biology, 23(1): 91–126, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00158156.

Rudge:2000:ERR

[Rud00] David Wÿss Rudge. Essay review: Recent introductory philosophy of biology texts. Journal of the History of Biology, 33(1): 181–187, March 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004765228726;
http://link.springer.com/
content/pdf/10.1023/A:1004765228726.
pdf.

Rudwick:2009:BR

[Rud09]  Martin Rudwick.  Book

review. Journal of the History of Biology, 42(2): 387–389, May 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9190-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9190-3.pdf.

Rudwick:2018:FMPb

[Rud18a] Martin J. S. Rudwick. The fate of the method of ‘paradigms’ in paleobiology. Journal of the History of Biology, 51(3):479– 533, September 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9501-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9501-z.pdf.

Rudwick:2018:FMPa

[Rud18b] Martin J. S. Rudwick. Functional morphology in paleobiology: Origins of the method of ‘paradigms’. Journal of the History of Biology, 51(1): 135–178, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9478-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9478-7.pdf.

Ruestow:1983:IIL

[Rue83] Edward G. Ruestow. Images and ideas: Leeuwenhoek’s perception of the spermatozoa. Journal of the History of Biology, 16(2): 185–224, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00124698.

Ruestow:1984:LCA

[Rue84] Edward G. Ruestow. Leeuwenhoek and the campaign against spontaneous generation. Journal of the History of Biology, 17(2):225– 248, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00143733.

Rumore:2010:BR

[Rum10] Gina Rumore. Book review.

Journal of the History of Biology, 43(2):407–409, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9233-9;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9233-9.pdf.

Rumore:2012:PSE

[Rum12] Gina Rumore. Preservation for science: The Ecological Society of America and the campaign for Glacier Bay National Monument. Journal of the History of Biology, 45(4): 613–650, November 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9301-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9301-9.pdf.

Rumore:2013:BR

[Rum13]  Gina Rumore.  Book re-

view. Journal of the History of Biology, 46(1):163– 165, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9350-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9350-3.pdf.

Ruse:2004:RCR

[Rus04] Michael Ruse. The romantic conception of Robert J. Richards. Journal of the History of Biology, 37(1): 3–23, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020389.
71636.cc;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020389.
71636.cc.pdf.

Ruse:2005:DRS

[Rus05] Michael Ruse. The Darwinian revolution, as seen in 1979 and as seen twenty–five years later in 2004. Journal of the History of Biology, 38 (1):3–17, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-6506-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
004-6506-1.pdf.

Ruse:2007:ERR

[Rus07] Michael Ruse. Essay review:

Restroom reading. Journal of the History of Biology, 40(1): 179–184, March 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9120-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9120-6.pdf.

Ruse:2012:BR

[Rus12]  Michael Ruse.  Book re-

view. Journal of the History of Biology, 45(1):175– 177, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9316-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9316-2.pdf.

Russell:2015:TPE

[Rus15] Doug Russell. Toward a pragmatist epistemology: Arthur O. Lovejoy’s and H. S. Jennings’s biophilosophical responses to neovitalism, 1909– 1914. Journal of the History of Biology, 48(1):37– 66, February 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9388-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9388-x.pdf.

Rieppel:2013:ANF

[RWE13] Olivier Rieppel, David M. Williams, and Malte C. Ebach. Adolf Naef (1883– 1949): On foundational concepts and principles of systematic morphology. Journal of the History of Biology, 46(3): 445–510, August 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9338-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9338-4.pdf.

Sackman:2009:BRS

[Sac09] Douglas Cazaux Sackman. Book review: Susanne Freidberg, Fresh: A Perishable History (Cambridge: Belknap Press of Harvard University Press, 2009). Journal of the History of Biology, 42(4):825– 827, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9213-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
009-9213-0.pdf.

Sandler:1973:RES

[San73]  Iris Sandler.  The re-

examination of Spallanzani’s interpretation of the role of the spermatic animalcules in fertilization. Journal of the History of Biology, 6(2):193– 223, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127608.

Sandler:1983:PLM

[San83] Iris Sandler. Pierre Louis Moreau de Maupertuis — a precursor of Mendel? Journal of the History of Biology, 16 (1):101–136, ???? 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00186677.

Santesmases:2006:PPB

[San06] María Jesús Santesmases.

Peace propaganda and biomedical experimentation: Influential uses of radioisotopes in endocrinology and molecular genetics in Spain (1947– 1971). Journal of the History of Biology, 39(4):765– 794, November 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9112-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-9112-6.pdf.

Sankaran:2010:MBF

[San10] Neeraja Sankaran. Mutant bacteriophages, Frank Macfarlane Burnet, and the changing nature of “genespeak” in the 1930s. Journal of the History of Biology, 43(3): 571–599, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9201-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9201-4.pdf.

Sankaran:2013:BR

[San13] Neeraja Sankaran. Book review. Journal of the History of Biology, 46(3):547– 549, August 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9361-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9361-0.pdf.

Santesmases:2016:BCW

[San16] María Jesús Santesmases. The bacterial cell wall in the antibiotic era: An ontology in transit between morphology and metabolism, 1940s– 1960s. Journal of the History of Biology, 49(1):3– 36, February 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9417-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9417-4.pdf.

Sapp:1983:SAF

[Sap83] Jan Sapp. The struggle for authority in the field of heredity, 1900–1932: New perspectives on the rise of genetics. Journal of the History of Biology, 16(3):311–342, Fall 1983. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00582405.

Sapp:1992:ERP

[Sap92] Jan Sapp. Essay review:

Portraying molecular biology. Journal of the History of Biology, 25(1):149– 155, Spring 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01947508.

Satishchandran:2020:BRO

[Sat20] Siddharth Satishchandran. Book review: Oren Harman and Michael Dietrich, Dreamers, Visionaries, and the Revolutionaries in the Life Sciences, The University of Chicago Press, Chicago, 2018, 324 pp., 18 illus. $40.00, £30.00 Paper. ISBN: 978-0-226-56990-1. Journal of the History of Biology, 53(4):659– 661, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-019-09578-1.

Stoltzfus:2014:MMF

[SC14] Arlin Stoltzfus and Kele Cable. Mendelian-mutationism: The forgotten evolutionary synthesis. Journal of the History of Biology, 47(4):501– 546, November 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9383-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9383-2.pdf.

Scarborough:1978:THH

[Sca78] John Scarborough. Theophrastus on herbals and herbal remedies. Journal of the History of Biology, 11(2): 353–385, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00389304.

Schaffner:1969:TEB

[Sch69] Kenneth F. Schaffner. Theories and explanations in biology. Journal of the History of Biology, 2(1):19– 33, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137265.

Schaffner:1974:PRD

[Sch74a] Kenneth F. Schaffner. The peripherality of reductionism in the development of molecular biology. Journal of the History of Biology, 7(1): 111–139, Spring 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00179296.

Schwartz:1974:CDD

[Sch74b] Joel S. Schwartz. Charles Darwin’s debt to Malthus and Edward Blyth. Journal of the History of Biology, 7(2): 301–318, Fall 1974. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351207.

Schweber:1977:OOR

[Sch77] Silvan S. Schweber. The origin of the Origin revisited. Journal of the History of Biology, 10(2):229–316, Fall 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00572644.

Schweber:1979:YD

[Sch79] Silvan S. Schweber. The young Darwin. Journal of the History of Biology, 12(1): 175–192, ???? 1979. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00128139.

Schweber:1980:ERE

[Sch80a] S. S. Schweber. Essay review early Victorian science: Science in culture. Journal of the History of Biology, 13(1): 121–140, ???? 1980. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125356.

Schweber:1980:DPE

[Sch80b] Silvan S. Schweber. Darwin and the political economists: Divergence of character. Journal of the History of Biology, 13(2):195–289, ???? 1980. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125744.

Schwartz:1984:DWD

[Sch84] Joel S. Schwartz. Darwin, Wallace, and the descent of man. Journal of the History of Biology, 17(2):271– 289, ???? 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00143735.

Schweber:1988:ERC

[Sch88] Silvan S. Schweber. Essay review: The correspondence of the young Darwin. Journal of the History of Biology, 21 (3):501–519, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00144093.

Schweber:1989:JHC

[Sch89] S. S. Schweber. John Herschel and Charles Darwin: a study in parallel lives. Journal of the History of Biology, 22(1):1–71, ???? 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209603.

Schwartz:1990:DWH

[Sch90] Joel S. Schwartz. Darwin, Wallace, and Huxley, and vestiges of the natural history of creation. Journal of the History of Biology, 23(1):

127–153, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00158157.

Schwartz:1995:GJR

[Sch95] Joel S. Schwartz. George John Romanes’s defense of Darwinism: The correspondence of Charles Darwin and his chief disciple. Journal of the History of Biology, 28(2): 281–316, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059191.

Schloegel:1999:AUT

[Sch99a] Judy Johns Schloegel. From anomaly to unification: Tracy sonneborn and the species problem in protozoa, 1954– 1957. Journal of the History of Biology, 32(1):93– 132, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004464509024;
http://link.springer.com/
content/pdf/10.1023/A:1004464509024.
pdf.

Schwartz:1999:RCT

[Sch99b] Joel S. Schwartz. Robert Chambers and Thomas Henry Huxley, science correspondents: The popularization and dissemination of Nineteenth Century natural science. Journal of the History of Biology, 32(2):343– 383, September 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004616403134;
http://link.springer.com/
content/pdf/10.1023/A:1004616403134.
pdf.

Schmidgen:2004:PPL

[Sch04] Henning Schmidgen. Pictures, preparations, and living processes: The production of immediate visual perception (Anschauung) in late-19th-Century physiology. Journal of the History of Biology, 37(3):477–513, October 2004. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2286-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2286-x.pdf.

Scharf:2009:IKN

[Sch09] Sara T. Scharf. Identification keys, the “natural method,” and the development of plant identification manuals. Journal of the History of Biology, 42(1):73–117, February 2009. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-008-9161-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9161-0.pdf.

Schmitt:2010:LSC

[Sch10] Stephane Schmitt. Lacepède’s syncretic contribution to the debates on natural history in France around 1800. Journal of the History of Biology, 43(3):429–457, August 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9227-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9227-7.pdf.

Schneider:2012:MBP

[Sch12] Laurence Schneider. Michurinist biology in the People’s Republic of China, 1948– 1956. Journal of the History of Biology, 45(3):525– 556, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9290-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9290-8.pdf.

Scholthof:2014:MVV

[Sch14] Karen-Beth G. Scholthof. Making a virus visible: Francis O. Holmes and a biological assay for tobacco mosaic virus. Journal of the History of Biology, 47(1):107– 145, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9353-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9353-0.pdf.

Scheffler:2015:PEE

[Sch15] Robin Wolfe Scheffler. The power of exercise and the exercise of power: The Harvard Fatigue Laboratory, distance running, and the disappearance of work, 1919– 1947. Journal of the History of Biology, 48(3):391– 423, August 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9392-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9392-1.pdf.

Schleper:2017:CCM

[Sch17] Simone Schleper. Conservation compromises: The MAB and the legacy of the International Biological Program, 1964–1974. Journal of the History of Biology, 50(1):133– 167, February 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9433-4;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9433-4.pdf.

Scheffler:2018:BRS

[Sch18] Robin Wolfe Scheffler. Book review: Stephen Hilgartner, Reordering Life: Knowledge and Control in the Genomics Revolution (Cambridge, MA: The MIT Press, 2017), 368 pp., $35.00 Hardcover ISBN: 978-0-262-03586-6. Journal of the History of Biology, 51(4): 879–881, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9529-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9529-8.pdf.

Schwartz:2023:EMO

[Sch23]  Joel Schwartz.  Everett

Mendelsohn, one colleague’s remembrances. Journal of the History of Biology, 56 (4):??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09752-6.

Scheffler:2025:AGU

[Sch25] Robin Wolfe Scheffler. Asilomar goes underground: The long legacy of recombinant DNA hazard debates for the Greater Boston Area biotechnology industry. Journal of the History of Biology, 58 (1):49–66, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09806-x.

Scriven:1969:EBS

[Scr69] Michael Scriven. Explanation in the biological sciences. Journal of the History of Biology, 2(1):187–198, ???? 1969. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137273.

Suarez-Diaz:2014:LWR

[SD14] Edna Suárez-Díaz. The long and winding road of molecular data in phylogenetic analysis. Journal of the History of Biology, 47(3):443– 478, August 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9373-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9373-9.pdf.

Suarez-Diaz:2019:BRB

[SD19a] Edna Suárez-Díaz. Book review: Bruno J. Strasser, Collecting Experiments. Making Big Data Biology (University of Chicago Press, 2019), 386 40pp., $37.62 Paper, ISBN: 978-0-226-63504-0. Journal of the History of Biology, 52(4): 733–735, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09588-z;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09588-z.pdf.

Suarez-Diaz:2019:EMB

[SD19b] Edna Suárez-Díaz. Erratum to: The Molecular Basis of Evolution and Disease: a Cold War Alliance. Journal of the History of Biology, 52(2):347, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9479-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9479-6.pdf. See [SD19c].

Suarez-Diaz:2019:MBE

[SD19c] Edna Suárez-Díaz. The molecular basis of evolution and disease: a Cold War alliance. Journal of the History of Biology, 52(2):325–346, June 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-017-9476-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9476-9.pdf. See erratum

[SD19b].

Stannard:1971:JHB

[SE71] Jerry Stannard and J. T. Edsall. The J. H. B. bookshelf. Journal of the History of Biology, 4(1):219– 220, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356984.

Selya:2003:ERD

[Sel03]  Rena Selya.  Essay re-

view: Defined by DNA: The intertwined lives of James Watson and Rosalind Franklin. Journal of the History of Biology, 36(3):591– 597, September 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004575.
16967.f3;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004575.
16967.f3.pdf.

Selya:2005:BRE

[Sel05] Rena Selya. Book reviews: Errol C. Friedberg, The Writing Life of James Watson (Cold Spring Harbor, N.Y.: Cold Spring Harbor Laboratory Press, 2005), xvii + 193 pp., illus., $25. Journal of the History of Biology, 38(2): 385–386, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-005-4226-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-4226-9.pdf.

Selya:2012:DSF

[Sel12] Rena Selya. Defending scientific freedom and democracy: The genetics society of America’s response to Lysenko. Journal of the History of Biology, 45(3):415– 442, August 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9288-2;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9288-2.pdf.

Selya:2023:EMS

[Sel23] Rena Selya. Everett Mendelsohn: A splendid mentor, primary source, and champion. Journal of the History of Biology, 56(4):??, December 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09749-1.

Sepkoski:2005:SJG

[Sep05] David Sepkoski. Stephen Jay Gould, Jack Sepkoski, and the ‘quantitative revolution’ in American paleobiology. Journal of the History of Biology, 38(2):209–237, Summer 2005. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2084-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
004-2084-5.pdf.

Sepkoski:2006:WCR

[Sep06] David Sepkoski. Worldviews in collision: Recent literature on the creation–evolution divide. Journal of the History of Biology, 39(3):607– 635, September 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-9000-0;
http://link.springer.com/
content/pdf/10.1007/s10739-
006-9000-0.pdf.

Sepkoski:2011:BR

[Sep11] David Sepkoski. Book review. Journal of the History of Biology, 44(4):797– 799, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9312-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9312-6.pdf.

Sepkoski:2013:TNH

[Sep13] David Sepkoski. Towards “a natural history of data”: Evolving practices and epistemologies of data in paleontology, 1800–2000. Journal of the History of Biology, 46(3): 401–444, August 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9336-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9336-6.pdf.

Sepkoski:2015:BR

[Sep15] David Sepkoski. Book review. Journal of the History of Biology, 48(3):487– 490, August 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9414-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9414-7.pdf.

Sepkoski:2019:USP

[Sep19] David Sepkoski. The unfinished synthesis?: Paleontology and evolutionary biology in the 20th Century. Journal of the History of Biology, 52 (4):687–703, December 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9537-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9537-8.pdf.

Serpente:2016:MTM

[Ser16] Norberto Serpente. More than a mentor: Leonard Darwin’s contribution to the assimilation of Mendelism into eugenics and Darwinism. Journal of the History of Biology, 49(3): 461–494, August 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9423-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9423-6.pdf.

Serpico:2021:CRI

[Ser21] Davide Serpico. The cyclical return of the IQ controversy: Revisiting the lessons of the resolution on genetics, race and intelligence. Journal of the History of Biology, 54 (2):199–228, June 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09637-6.

Shavit:2018:SSG

[SG18] Ayelet Shavit and James R.

Griesemer. Science and sentiment: Grinnell’s fact-based philosophy of biodiversity conservation. Journal of the History of Biology, 51(2): 283–318, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-017-9489-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        017-9489-4.pdf.

Shapere:1969:BUS

[Sha69] Dudley Shapere. Biology and the unity of science. Journal of the History of Biology, 2(1):3–18, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137264.

Shaw:1972:MCS

[Sha72] James Rochester Shaw. Models for cardiac structure and function in Aristotle. Journal of the History of Biology, 5(2):355–388, Fall 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00346664.

Shank:1985:GUH

[Sha85] Michael H. Shank. From Galen’s ureters to Harvey’s veins. Journal of the History of Biology, 18(3):331– 355, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138929.

Shapiro:2008:CBO

[Sha08] Adam R. Shapiro. Civic biology and the origin of the school antievolution movement. Journal of the History of Biology, 41(3):409– 433, September 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9148-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9148-2.pdf.

Shapiro:2011:BR

[Sha11] Adam Shapiro. Book review.

Journal of the History of Biology, 44(2):369–372, May 2011. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9297-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9297-1.pdf.

Summers:1992:JHB

[SHB^{+}92] William C. Summers, Joel B.

Hagen, Mark V. Barrow, Jr., Lynn Nyhart, and M. Susan Lindee. The J. H. B. bookshelf. Journal of the History of Biology, 25(2): 335–342, ???? 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00162845.

Shin:2018:BRS

[Shi18] Youjung Shin. Book review:

Sabina Leonelli, Data-Centric Biology: a Philosophical Study (Chicago: The University of Chicago Press, 2016), 288 pp., $35.00, Paperback, ISBN: 978-0-226-41647-2. Journal of the History of Biology, 51(4): 887–889, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9544-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9544-9.pdf.

Shmuely:2020:AWB

[Shm20] Shira Shmuely. Alfred Wallace’s baby orangutan: Game, pet, specimen. Journal of the History of Biology, 53(3):321– 343, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09611-8.

Shmuely:2023:BRR

[Shm23a] Shira Shmuely. Book review: Rob Boddice, Humane Professions: The Defence of Experimental Medicine, 1876– 1914, Cambridge UK: Cambridge University Press, 2021, ISBN: 978-1-108-49009-2, 204 pp. Journal of the History of Biology, 56(2):399– 400, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.

1007/s10739-023-09728-6.

See correction [Shm23b].

Shmuely:2023:CBR

[Shm23b] Shira Shmuely. Correction:

Book review: Rob Boddice, Humane Professions: The Defence of Experimental Medicine, 1876–1914, Cambridge UK: Cambridge University Press, 2021, ISBN: 978-1-108-49009-2, 204 pp. Journal of the History of Biology, 56(2):401, August 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09732-w.

See [Shm23a].

Shotwell:2013:RVS

[Sho13] R. Allen Shotwell. The revival of vivisection in the sixteenth century. Journal of the History of Biology, 46(2): 171–197, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9326-8;
http://link.springer.com/
content/pdf/10.1007/s10739-
012-9326-8.pdf.

Shtilmark:1992:ECA

[Sht92] F. R. Shtil’mark. The evolution of concepts about the preservation of nature in Soviet literature. Journal of the History of Biology, 25 (3):429–447, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00352001.

Silverstein:1994:HVE

[Sil94] Arthur M. Silverstein. The heuristic value of experimental systems: The case of immune hemolysis. Journal of the History of Biology, 27 (3):437–447, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058993.

Sinding:1989:HRR

[Sin89]  Christiane Sinding.  The

history of resistant rickets: a model for understanding the growth of biomedical knowledge. Journal of the History of Biology, 22(3): 461–495, Fall 1989. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00401578.

Singleton:1997:HCF

[Sin97] Rivers Singleton, Jr. Heterotrophic CO_{2}-fixation, mentors, and students: The Wood–Werkman reactions. Journal of the History of Biology, 30(1):91–120, March 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http://

link.springer.com/article/

10.1023/A:1004270612446;
http://link.springer.com/
content/pdf/10.1023/A:1004270612446.
pdf.

Singleton:2000:BBA

[Sin00] Rivers Singleton, Jr. From bacteriology to biochemistry: Albert Jan Kluyver and Chester Werkman at Iowa State. Journal of the History of Biology, 33(1):141– 180, March 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004775817881;
http://link.springer.com/
content/pdf/10.1023/A:1004775817881.
pdf.

Smocovitis:1993:JHB

[SJB^{+}93] Vassiliki Betty Smocovitis, John R. Jungck, Giulio Barsanti, Pamela M. Henson, Mark V. Barrow, Jr., Christoph H. Lüthy, and Charlotte M. Porter. The J. H. B. bookshelf. Journal of the History of Biology, 26 (3):571–587, Fall 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062062.

Sharpley:2022:FCL

[SK22] Christopher F. Sharpley and Clemens Koehn. Frequency and content of the last fifty years of papers on Aristotle’s writings on biological phenomena. Journal of the History of Biology, 55(3):585– 607, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09683-8.

Sterner:2018:MPS

[SL18] Beckett Sterner and Scott Lidgard. Moving past the systematics wars. Journal of the History of Biology, 51 (1):31–67, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9471-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9471-1.pdf.

Slack:2003:RSN

[Sla03] Nancy G. Slack. Are research schools necessary? Contrasting models of 20th Century research at Yale led by Ross Granville Harrison, Grace E. Pickford and G. Evelyn Hutchinson. Journal of the History of Biology, 36(3):501– 529, September 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000004573.
47187.76;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000004573.
47187.76.pdf.

Slavet:2008:FLP

[Sla08] Eliza Slavet. Freud’s ‘Lamarckism’ and the politics of racial science. Journal of the History of Biology, 41 (1):37–80, March 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9138-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
007-9138-4.pdf.

Sloan:1972:JLJ

[Slo72] Phillip R. Sloan. John Locke, John Ray, and the problem of the natural system. Journal of the History of Biology, 5(1):1–53, Spring 1972. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02113485.

Sloan:1985:ERE

[Slo85] Phillip R. Sloan. Essay review: Ernst Mayr on the history of biology. Journal of the History of Biology, 18(1): 145–153, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127960.

Sloan:1986:DVM

[Slo86] Phillip R. Sloan. Darwin, vital matter, and the transformism of species. Journal of the History of Biology, 19 (3):369–445, ???? 1986. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138286.

Sloan:2007:TNV

[Slo07] Phillip R. Sloan. Two new volumes of Darwin’s work: Essay review. Journal of the History of Biology, 40(2): 363–367, June 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9127-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
007-9127-7.pdf.

Stamhuis:2007:KBT

[SM07] Ida H. Stamhuis and Arve Monsen. Kristine Bonnevie, Tine Tammes and Elisabeth Schiemann in early genetics: Emerging chances for a university career for women. Journal of the History of Biology, 40(3):427– 466, September 2007. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-007-9132-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
007-9132-x.pdf.

Smith:1978:CDO

[Smi78] C. U. M. Smith. Charles Darwin, the origin of consciousness, and panpsychism. Journal of the History of Biology, 11(2):245–267, Fall 1978. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00389301.

Smith:1982:EPM

[Smi82] C. U. M. Smith. Evolution and the problem of mind: Part II. John Hughlings Jackson. Journal of the History of Biology, 15(2):241–262, ???? 1982. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00233016.

Smith:1990:CDI

[Smi90] Fabienne Smith. Charles Darwin’s ill health. Journal of the History of Biology, 23(3): 443–459, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00136378.

Smith:1992:CDH

[Smi92] Fabienne Smith. Charles Darwin’s health problems: The allergy hypothesis. Journal of the History of Biology, 25 (2):285–306, ???? 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00162843.

Smith:1999:CTL

[Smi99] C. U. M. Smith. Coleridge’s “theory of life”. Journal of the History of Biology, 32 (1):31–50, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004456223137;
http://link.springer.com/
content/pdf/10.1023/A:1004456223137.
pdf.

Smith:2015:ARW

[Smi15] Charles H. Smith. Alfred Russel Wallace and the road to natural selection, 1844–1858. Journal of the History of Biology, 48(2):279–300, May 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9397-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9397-9.pdf.

Smith:2025:WR

[Smi25]  Charles H. Smith.  Wal-

lace reconsidered. Journal of the History of Biology, 58 (2):155–156, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09821-y.

Smocovitis:1991:ERP

[Smo91] Vassiliki Betty Smocovitis.

Essay review: The politics of writing biology. Journal of the History of Biology, 24(3): 521–527, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00156323.

Smocovitis:1992:UBE

[Smo92] V. B. Smocovitis. Unifying biology: The evolutionary synthesis and evolutionary biology. Journal of the History of Biology, 25(1): 1–65, Spring 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01947504.

Smocovitis:1994:OEF

[Smo94] Vassiliki Betty Smocovitis.

Organizing evolution: Founding the society for the study of evolution (1939–1950). Journal of the History of Biology, 27(2):241–309, Summer 1994. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01062564.

Smocovitis:1999:ERT

[Smo99a] Vassiliki Betty Smocovitis.

Essay review: The tormenting desire for unity. Journal of the History of Biology, 32 (2):385–394, September 1999. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004620515855;
http://link.springer.com/
content/pdf/10.1023/A:1004620515855.
pdf.

Smocovitis:1999:LYB

[Smo99b] Vassiliki Betty Smocovitis. Living with your biographical subject: Special problems of distance, privacy and trust in the biography of G. Ledyard Stebbins, Jr. Journal of the History of Biology, 32 (3):421–438, December 1999. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004731419724;
http://link.springer.com/
content/pdf/10.1023/A:1004731419724.
pdf.

Smocovitis:2005:ERW

[Smo05a] Vassiliki Betty Smocovitis. Essay review: “what made Ernst unique?”. Journal of the History of Biology, 38(3): 609–614, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4046-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4046-y.pdf;   https:

//en.wikipedia.org/wiki/
Ernst_Mayr.

Smocovitis:2005:IAT

[Smo05b] Vassiliki Betty Smocovitis.

“It ain’t over ‘til it’s over”: Rethinking the Darwinian revolution. Journal of the History of Biology, 38(1): 33–49, ???? 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-6508-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-6508-z.pdf.

Smocovitis:2024:DDM

[Smo24] Vassiliki Betty Smocovitis.

Darwin’s “Dark matter” and the history of biology: An editorial introduction. Journal of the History of Biology, 57 (4):499–500, December 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09799-z.

Smocovitis:2025:EMP

[Smo25a] Vassiliki Betty Smocovitis.

The 2025 Everett Mendelsohn Prize. Journal of the History of Biology, 58 (1):??, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09810-1.

Smocovits:2025:IWW

[Smo25b] Vassiliki Betty Smocovits. The industry of Wallace and the Wallace Industry. Journal of the History of Biology, 58(2):??, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-025-09820-z.

Soderqvist:1994:DON

[Söd94] Thomas Söderqvist. Darwinian overtones: Niels K. Jerne and the origin of the selection theory of antibody formation. Journal of the History of Biology, 27(3): 481–529, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058995.

Soderqvist:2011:SSS

[Söd11]  Thomas Söderqvist.  The

seven sisters: Subgenres of bioi of contemporary life scientists. Journal of the History of Biology, 44(4):633– 650, November 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9272-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9272-x.pdf.

Sommer:2005:AHT

[Som05] Marianne Sommer. Ancient hunters and their modern representatives: William Sollas’s (1849–1936) anthropology from disappointed bridge to trunkless tree and the instrumentalisation of racial conflict. Journal of the History of Biology, 38(2):327– 365, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-5428-2;
http://link.springer.com/
content/pdf/10.1007/s10739-
004-5428-2.pdf.

Sommer:2008:HGN

[Som08] Marianne Sommer. History in the gene: Negotiations between molecular and organismal anthropology. Journal of the History of Biology, 41 (3):473–528, September 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9150-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
008-9150-3.pdf.

Sommer:2021:BRE

[Som21] Marianne Sommer. Book review: Erika Lorraine Milam, Creatures of Cain: The Hunt for Human Nature in Cold War America, (Princeton: Princeton University Press, 2019), 408 pp., 33 b/w illus., $29.95 Cloth, ISBN: 978-0-691-18188-2. Journal of the History of Biology, 54(1): 143–145, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09635-8.

Sommer:2024:BRE

[Som24] Marianne Sommer. [Book review:] Emma Kowal, Haunting Biology: Science and Indigeneity in Australia, Durham, NC: Duke University Press, 2023, ISBN: 978-1-4780-2753-9, 264 pp. Journal of the History of Biology, 57(2): 339–341, June 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09778-4.

Soon:2021:BRJ

[Soo21] Wayne Soon. Book review: Jenny Bangham. Blood Relations: Transfusion and the Making of Human Genetics. Journal of the History of Biology, 54(3):541– 543, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09645-6.

Sheets-Pyenson:1981:DDH

[SP81] Susan Sheets-Pyenson. Darwin’s data: His reading of natural history journals, 1837– 1842. Journal of the History of Biology, 14(2):231– 248, ???? 1981. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00141093.

Stuessy:2025:ABB

[SP25] Tod F. Stuessy and Ronald Pilatowski. Atomic botany: The botanical career of Janice Carson Beatley, and the flora, vegetation, and ecology of the Nevada Test Site in Cold War United States (1959–1973). Journal of the History of Biology, 58(2):247–273, June 2025. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09819-6.

Smocovitis:2023:EMF

[SR23a] Vassiliki Betty Smocovitis and Nicolas Rasmussen. Everett Mendelsohn (1931– 2023): Founding Editor of the Journal of the History of Biology. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09744-6.

Smocovitis:2023:GAH

[SR23b] Vassiliki Betty Smocovitis and Nicolas Rasmussen. Garland E. Allen (1936–2023), historian of life science. Journal of the History of Biology, 56(2): 215–217, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09723-x.

Smocovitis:2024:EMP

[SR24] Betty Smocovitis and Nicolas Rasmussen. The 2024 Everett Mendelsohn Prize. Journal of the History of Biology, 57(1):1–2, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-024-09767-7.

Soderqvist:1999:ERH

[SS99] Thomas Söderqvist and Craig Stillwell. Essay review: The historiography of immunology is still in its infancy. Journal of the History of Biology, 32 (1):205–215, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004654415985;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004654415985.
                                        pdf.

Singleton:2017:ROF

[SS17] Rivers Singleton, Jr. and David R. Singleton. Remembering our forebears: Albert Jan Kluyver and the unity of life. Journal of the History of Biology, 50(1):169– 218, February 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9438-7;
http://link.springer.com/
content/pdf/10.1007/s10739-
016-9438-7.pdf.

Stamhuis:1995:FCE

[Sta95] Ida H. Stamhuis. A female contribution to early genetics: Tine Tammes and Mendel’s laws for continuous characters. Journal of the History of Biology, 28 (3):495–531, Fall 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059390.

Stamos:1996:WDR

[Sta96] David N. Stamos. Was Darwin really a species nominalist? Journal of the History of Biology, 29(1):127–144, ???? 1996. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129698.

Stamhuis:2003:RHV

[Sta03] Ida H. Stamhuis. The reactions on Hugo de Vries’s intracellular pangenesis; the discussion with August Weismann. Journal of the History of Biology, 36(1):119– 152, March 2003. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022561530690;
http://link.springer.com/
content/pdf/10.1023/A:1022561530690.
pdf.

Stanley:2023:FHW

[Sta23] Matthew Stanley. A few hours a week: Everett Mendelsohn as teacher, mentor, and exemplar. Journal of the History of Biology, 56(4): ??, December 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09751-7.

Stack:2024:ARW

[Sta24] David Stack. Alfred Russel Wallace’s Darwinian opposition to eugenics. Journal of the History of Biology, 57(4): 557–579, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09792-6.

Stark:2025:IMN

[Sta25] James F. Stark. Imperial microbiology: The national collection of type cultures and the management of microorganisms, 1916–1922. Journal of the History of Biology, 58 (2):163–213, June 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09809-8.

Stewart:1976:FBO

[Ste76] Larry Stewart. Freud before Oedipus: Race and heredity in the origins of psychoanalysis. Journal of the History of Biology, 9(2):215–228, ???? 1976. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00209882.

Stevens:1984:HPC

[Ste84] P. F. Stevens. Haüy and A.-P. Candolle: Crystallography, botanical systematics, and comparative morphology, 1780–1840. Journal of the History of Biology, 17(1): 49–82, Spring 1984. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00397502.

Stenhouse:1990:DCF

[Ste90] John Stenhouse. Darwin’s captain: F. W. Hutton and the nineteenth-century Darwinian debates. Journal of the History of Biology, 23(3): 411–442, ???? 1990. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF00136377.

Stein:1997:AMA

[Ste97] Barbara R. Stein. Annie M. Alexander: Extraordinary patron. Journal of the History of Biology, 30(2): 243–266, June 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004235719745;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004235719745.
                                        pdf.

Steigerwald:2002:GMU

[Ste02] Joan Steigerwald. Goethe’s morphology: Urphänomene and aesthetic appraisal. Journal of the History of Biology, 35(2):291–328, June 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016028812658;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1016028812658.
                                        pdf.

Stevens:2004:BRA

[Ste04] P. F. Stevens. Book review: Augustin-Pyramus de Candolle, Memoires et Souvenirs (1878–1841), Jean-Daniel Candaux and Jean-Marc Drouin, eds., with the aid of Patrick Bungener and René Sigrist, Biliothèque d”Histoire des Sciences 5 (Genève: Georg, 2003), xv + 591 pp., illus., EUR33.00 paper. Journal of the History of Biology, 37(3):603– 604, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2095-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2095-2.pdf.

Sterling:2005:BRR

[Ste05] Keir B. Sterling. Book reviews: Russell M. Lawson, The Land Between the Rivers: Thomas Nuttall’s Ascent of the Arkansas 1819 (Ann Arbor: University of Michigan Press, 2004), 152 pp., maps, $27.00. Journal of the History of Biology, 38(2):393– 395, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4232-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4232-y.pdf.

Steffes:2007:POS

[Ste07] David M. Steffes. Panpsychic organicism: Sewall Wright’s philosophy for understanding complex genetic systems. Journal of the History of Biology, 40(2):327–361, June 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-006-9105-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        006-9105-5.pdf.

Stevens:2014:BR

[Ste14] Hallam Stevens. Book review. Journal of the History of Biology, 47(1):177– 179, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9374-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9374-3.pdf.

Stevens:2015:BR

[Ste15] Hallam Stevens. Book review. Journal of the History of Biology, 48(2):353–355, May 2015. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9404-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9404-9.pdf.

Stevens:2018:GGO

[Ste18] Hallam Stevens. Globalizing genomics: The origins of the International Nucleotide Sequence Database Collaboration. Journal of the History of Biology, 51(4):657– 691, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9490-y;
http://link.springer.com/
content/pdf/10.1007/s10739-
017-9490-y.pdf.

Steinert:2019:BRJ

[Ste19] Beatrice Steinert. Book review: Janina Wellmann, The Form of Becoming: Embryology and the Epistemology of Rhythm, 1760–1830. Journal of the History of Biology, 52 (3):493–495, September 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09576-3;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-09576-3.pdf.

Stevens:2024:BRM

[Ste24] Hallam Stevens. Book review: Miguel García-Sancho and James Lowe, A History of Genomics Across Species, Communities, and Projects, London: Palgrave Macmillan, 2023, ISBN: 978-3-031-06129-5, 380 pp. Journal of the History of Biology, 57(1): 161–163, March 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09764-w.

Stillwell:1994:TES

[Sti94] Craig R. Stillwell. Thymectomy as an experimental system in immunology. Journal of the History of Biology, 27 (3):379–401, Fall 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058991.

Stokland:2015:FSA

[Sto15] Håkon B. Stokland. Field studies in absentia: Counting and monitoring from a distance as technologies of government in Norwegian wolf management (1960s– 2010s). Journal of the History of Biology, 48(1):1– 36, February 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9393-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9393-0.pdf.

Struchkov:1992:NPM

[Str92] Anton Yu. Struchkov. Nature protection as moral duty: The ethical trend in the Russian conservation movement. Journal of the History of Biology, 25(3):413–428, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00352000.

Strick:1999:DOL

[Str99] James Strick. Darwinism and the origin of life: The role of H. C. Bastian in the British spontaneous generation debates, 1868–1873. Journal of the History of Biology, 32 (1):51–92, March 1999. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004460408116;
http://link.springer.com/
content/pdf/10.1023/A:1004460408116.
pdf.

Strick:2000:ERC

[Str00]  James Strick.  Essay re-

view: The Cambrian explosion (of books on the origin of life). Journal of the History of Biology, 33(2):371– 384, September 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004825832754;
http://link.springer.com/
content/pdf/10.1023/A:1004825832754.
pdf.

Strick:2004:CCD

[Str04] James E. Strick. Creating a cosmic discipline: The crystallization and consolidation of exobiology, 1957–1973. Journal of the History of Biology, 37(1):131–180, March 2004. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1023/B:HIST.0000020279.
                                        73895.f2;   http://link.
                                        springer.com/content/pdf/
                                        10.1023/B:HIST.0000020279.
                                        73895.f2.pdf.

Strasser:2010:CCC

[Str10] Bruno J. Strasser. Collecting, comparing, and computing sequences: The making of Margaret O. Dayhoff’s atlas of protein sequence and structure, 1954–1965. Journal of the History of Biology, 43(4): 623–660, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-009-9221-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9221-0.pdf.

Strick:2019:BRS

[Str19] James Strick. Book review: Scott Lidgard and Lynn Nyhart, eds., Biological Individuality: Integrating Scientific, Philosophical and Historical Perspectives. Journal of the History of Biology, 52(4):741– 742, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09587-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09587-0.pdf.

Strzadala:2020:NBE

[Str20] Agata Strzadała. The new biology as an example of Newspeak: The case of Polish zoology, 1948–1956. Journal of the History of Biology, 53 (1):141–157, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-020-09594-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
020-09594-6.pdf.

Strick:2023:BRD

[Str23]  James Strick.  Book re-

view: David P.D. Munns and Kärin Nickelsen. Far Beyond the Moon: A History of Life–Support Systems in the Space Age, Pittsburgh: University of Pittsburgh Press, 2021, ISBN 978-0-8229-4654-0, 216 pp. Journal of the History of Biology, 56(1): 205–206, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09711-1.

Straetmanns:2024:LPT

[Str24] Vera Maximilia Straetmanns.

The lady and the plants: Two notions of teleology in Agnes Arber’s philosophy of plants. Journal of the History of Biology, 57(4):533– 555, December 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09793-5.

Stuhrmann:2023:SSC

[Stu23] Cora Stuhrmann. Sociobiology on screen. The controversy through the lens of Sociobiology: Doing What Comes Naturally. Journal of the History of Biology, 56(2): 365–397, August 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09719-7.

Stuhrmann:2024:BRN

[Stu24]  Cora Stuhrmann.  [Book

review:] Nadine Weidman, Killer Instinct: The Popular Science of Human Nature in Twentieth-Century America, Cambridge: Harvard University Press, 2021 , ISBN 978-0-674-98347-2, 368 pp. Journal of the History of Biology, 57 (2):345–348, June 2024. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09780-w.

Sugden:2023:BRG

[Sug23]  Bill Sugden.  Book re-

view: Gregory Morgan. Virus Hunters: a History of Tumor Virology, Baltimore: Johns Hopkins University Press, 2021, ISBN 1421444011, xiv + 373 pp. Journal of the History of Biology, 56(1): 201–203, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09712-0.

Sulloway:1982:DHF

[Sul82a] Frank J. Sulloway. Darwin and his finches: The evolution of a legend. Journal of the History of Biology, 15 (1):1–53, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00132004.

Sulloway:1982:DCB

[Sul82b] Frank J. Sulloway. Darwin’s conversion: The Beagle voyage and its aftermath. Journal of the History of Biology, 15 (3):325–396, ???? 1982. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00133143.

Sulloway:1983:FRD

[Sul83] Frank J. Sulloway. Further remarks on Darwin’s spelling habits and the dating of Beagle voyage manuscripts. Journal of the History of Biology, 16(3):361–390, Fall 1983. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00582407.

Sulloway:2009:TTD

[Sul09] Frank J. Sulloway. Tantalizing tortoises and the Darwin– Galápagos legend. Journal of the History of Biology, 42(1): 3–31, February 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9173-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9173-9.pdf.

Sullivan:2011:INS

[Sul11] Gregory Sullivan. The instinctual nation-state: Non– Darwinian theories, state science and ultra-nationalism in Oka Asajiro’s evolution and human life. Journal of the History of Biology, 44(3): 547–586, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9258-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9258-0.pdf.

Summers:1991:COO

[Sum91] William C. Summers. From culture as organism to organism as cell: Historical origins of bacterial genetics. Journal of the History of Biology, 24 (2):171–190, ???? 1991. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF00209428.

Summers:1993:HBC

[Sum93] William C. Summers. How bacteriophage came to be used by the phage group. Journal of the History of Biology, 26(2):255–267, Summer 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061969.

Summers:2004:BRS

[Sum04] William C. Summers. Book review: Santiago Ramón y Cajal, advice for a young investigator. Journal of the History of Biology, 37(3):595– 596, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-2091-6;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-2091-6.pdf.

Sunderland:2010:RTH

[Sun10] Mary Evelyn Sunderland. Regeneration: Thomas Hunt Morgan’s window into development. Journal of the History of Biology, 43(2): 325–361, May 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9203-2;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9203-2.pdf.

Sunderland:2013:MNH

[Sun13] Mary E. Sunderland. Modernizing natural history: Berkeley’s museum of vertebrate zoology in transition. Journal of the History of Biology, 46(3): 369–400, August 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-012-9339-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        012-9339-3.pdf.

Suri:2022:BSC

[Sur22] Tara Suri. Between simians and cell lines: Rhesus monkeys, polio research, and the geopolitics of tissue culture (1934–1954). Journal of the History of Biology, 55(1): 115–146, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09666-9.

Steere-Williams:2012:BR

[SW12] Jacob Steere-Williams. Book review. Journal of the History of Biology, 45(1):163– 165, February 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1007/s10739-011-9313-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
011-9313-5.pdf.

Swazey:1968:JHB

[Swa68a] Judith P. Swazey. The J.

H. B. bookshelf. Journal of the History of Biology, 1 (2):333–340, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351925.

Swazey:1968:SCI

[Swa68b]  Judith P. Swazey.  Sher-

rington’s concept of integrative action. Journal of the History of Biology, 1(1): 57–89, ???? 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00149776.

Swazey:1970:APA

[Swa70a] Judith P. Swazey. Action propre and action commune: The localization of cerebral function. Journal of the History of Biology, 3(2):213– 234, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137352.

Swazey:1970:JHBa

[Swa70b] Judith P. Swazey. The J.

H. B. bookshelf. Journal of the History of Biology, 3(1): 183–187, Spring 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00569313.

Swazey:1970:JHBb

[Swa70c] Judith P. Swazey. The J. H. B. bookshelf. Journal of the History of Biology, 3(2): 361–365, ???? 1970. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00137358.

Swazey:1971:JHB

[Swa71] Judith P. Swazey. The J. H. B. bookshelf. Journal of the History of Biology, 4(2): 373–375, ???? 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138318.

Swetlitz:1995:JHE

[Swe95] Marc Swetlitz. Julian Huxley and the end of evolution. Journal of the History of Biology, 28(2):181– 217, Summer 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01059188.

Swiatczak:2014:IBD

[Swi14] Bartlomiej Swiatczak. Immune balance: The development of the idea and its applications. Journal of the History of Biology, 47(3): 411–442, August 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9370-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9370-z.pdf.

Tabery:2004:ESG

[Tab04] James G. Tabery. The “evolutionary synthesis” of George Udny Yule. Journal of the History of Biology, 37(1): 73–101, March 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:HIST.0000020390.
75208.ac;   http://link.
springer.com/content/pdf/
10.1023/B:HIST.0000020390.
75208.ac.pdf.

Tabery:2008:RFL

[Tab08]  James Tabery.  R. A.

Fisher, Lancelot Hogben, and the origin(s) of genotype– environment interaction. Journal of the History of Biology, 41(4):717–761, December 2008. CODEN JH-BIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10739-008-9155-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9155-y.pdf.

Tammone:1995:CDL

[Tam95] William Tammone. Competition, the division of labor, and Darwin’s principle of divergence. Journal of the History of Biology, 28(1): 109–131, Spring 1995. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01061248.

Tamborini:2015:PDT

[Tam15] Marco Tamborini. Paleontology and Darwin’s theory of evolution: The subversive role of statistics at the end of the 19th Century. Journal of the History of Biology, 48(4):575– 612, November 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9402-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9402-y.pdf.

Tamborini:2017:BR

[Tam17]  Marco Tamborini.  Book

review. Journal of the History of Biology, 50(2): 465–468, May 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9475-x;
http://link.springer.com/
content/pdf/10.1007/s10739-
017-9475-x.pdf.

Tamborini:2020:CAP

[Tam20a]  Marco Tamborini.  Chal-

lenging the adaptationist paradigm: Morphogenesis, constraints, and constructions. Journal of the History of Biology, 53(2):269– 294, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09604-7.

Tamborini:2020:TCD

[Tam20b]  Marco Tamborini.  The

Twentieth-Century desire for morphology. Journal of the History of Biology, 53(2): 211–216, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09605-6.

Tanghe:2017:HTO

[Tan17] Koen B. Tanghe. A historical taxonomy of Origin of Species problems and its relevance to the historiography of evolutionary thought. Journal of the History of Biology, 50(4): 927–987, November 2017. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-016-9453-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
016-9453-8.pdf.

Tanghe:2020:BRB

[Tan20a] Koen B. Tanghe. Book review: Bill Jenkins, Evolution Before Darwin. Theories of the Transmutation of Species in Edinburgh, 1804–1834. Journal of the History of Biology, 53(1):203–207, March 2020. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-020-09597-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
020-09597-3.pdf.

Tanghe:2020:BRM

[Tan20b] Koen B. Tanghe. Book review: Michel Morange, The Black Box of Biology. A History of the Molecular Revolution. Trans. by M. Cobb (Cambridge, MA: Harvard University Press, 2020), 528 pp., $45.00, £36.95, EUR40.50 Hardback, ISBN: 978-0-674-28136-3. Journal of the History of Biology, 53(3):489– 492, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09615-4.

Tauber:1997:HPP

[Tau97] Alfred I. Tauber. Historical and philosophical perspectives concerning immune cognition. Journal of the History of Biology, 30(3):419– 440, September 1997. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004247922979;
http://link.springer.com/
content/pdf/10.1023/A:1004247922979.
pdf.

Taylor:1988:TOH

[Tay88] Peter J. Taylor. Technocratic optimism, H. T. Odum, and the partial transformation of ecological metaphor after World War II. Journal of the History of Biology, 21 (2):213–244, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00146987.

Taylor:1998:MSP

[Tay98] Joseph E. Taylor III. Making salmon: The political economy of fishery science and the road not taken. Journal of the History of Biology, 31 (1):33–59, March 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004253405353;
http://link.springer.com/

                                        content/pdf/10.1023/A:1004253405353.
                                        pdf.

Taylor:2021:VPM

[Tay21] N. A. J. Taylor. The visual politics of Maralinga: Experiences, (re)presentations, and vulnerabilities. Journal of the History of Biology, 54 (1):95–106, April 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09632-x.

Teich:2007:HLR

[Tei07] Mikulás Teich. Haldane and Lysenko revisited. Journal of the History of Biology, 40 (3):557–563, September 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-007-9125-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        007-9125-9.pdf.

Testa:2023:STS

[Tes23] Caden Testa. Species transformation and social reform: The role of the will in Jean-Baptiste Lamarck’s transformist theory. Journal of the History of Biology, 56(1): 125–151, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09707-x.

Theunissen:1996:BDT

[The96] Bert Theunissen. The beginnings of the “Delft Tradition” revisited: Martinus Beijerinck and the genetics of microorganisms. Journal of the History of Biology, 29(2): 197–228, Summer 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00571082.

Theunissen:2008:BMT

[The08] Bert Theunissen. Breeding without Mendelism: Theory and practice of dairy cattle breeding in The Netherlands 1900–1950. Journal of the History of Biology, 41(4):637– 676, December 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9153-0;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9153-0.pdf.

Theunissen:2012:DHP

[The12] Bert Theunissen. Darwin and his pigeons. the analogy between artificial and natural selection revisited. Journal of the History of Biology, 45 (2):179–212, May 2012. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9310-8;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9310-8.pdf.

Thieffry:2001:REE

[Thi01] Denis Thieffry. Rationalizing early embryogenesis in the 1930s: Albert Dalcq on gradients and fields. Journal of the History of Biology, 34(1): 149–181, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1010389907221;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1010389907221.
                                        pdf.

Thomas:2005:AJN

[Tho05] Marion Thomas. Are animals just noisy machines?: Louis Boutan and the co-invention of animal and child psychology in the French Third Republic. Journal of the History of Biology, 38(3):425– 460, November 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-0555-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-0555-y.pdf.

Thurtle:2002:HHG

[Thu02] Phillip Thurtle. Harnessing heredity in Gilded Age America: Middle class mores and industrial breeding in a cultural context. Journal of the History of Biology, 35 (1):43–78, March 2002. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014519404949;
http://link.springer.com/
content/pdf/10.1023/A:1014519404949.
pdf.

Tobbell:2009:BRC

[Tob09] Dominique A. Tobbell. Book review: Catherine Brady, Elizabeth Blackburn and the Story of Telomeres: Deciphering the Ends of DNA (Cambridge, MA: The MIT Press, 2009), 408 pp., illus., $15.95 (paper). Journal of the History of Biology, 42(4):811– 814, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9217-9;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9217-9.pdf.

Tompkins:2022:DDM

[Tom22] Joshua D. Tompkins. Discovering DNA methylation, the history and future of the writing on DNA. Journal of the History of Biology, 55(4):865– 887, December 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10739-022-09691-8.

Tonn:2018:HTH

[Ton18] Jenna Tonn. The handmaid’s tale. Hulu. Season 1 (April– June 2017). Television. Journal of the History of Biology, 51(2):415–417, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-018-9516-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        018-9516-0.pdf.

Tauber:1994:FMB

[TP94] Alfred I. Tauber and Scott H. Podolsky. Frank Macfarlane Burnet and the immune self. Journal of the History of Biology, 27(3):531–573, Fall 1994. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01058996.

Trumpler:1997:CIT

[Tru97] Maria Trumpler. Converging images: Techniques of intervention and forms of representation of sodium-channel proteins in nerve cell membranes. Journal of the History of Biology, 30(1):55–89, March 1997. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004218611538;

http://link.springer.com/
content/pdf/10.1023/A:1004218611538.
pdf.

Tjossem:1996:JHB

[TS^{+}96] Sara F. Tjossem, Vassiliki Betty Smocovitis, et al. The J. H. B. bookshelf. Journal of the History of Biology, 29(1):145–154, ???? 1996. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00129699.

Turbil:2018:MHM

[Tur18] Cristiano Turbil. Making heredity matter: Samuel Butler’s idea of unconscious memory. Journal of the History of Biology, 51(1): 7–29, March 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9469-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9469-8.pdf.

Tyrrell:2023:BRR

[Tyr23] Ian Tyrrell. Book review:

Raf De Bont, Nature’s Diplomats: Science, Internationalism & Preservation, 1920– 1960. Pittsburgh: University of Pittsburgh Press, 2021, ISBN: 978-0-8229-4661-8, x + 373 pp. Journal of the History of Biology, 56(1): 197–200, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09715-x.

Uchida:1993:BSJ

[Uch93] Hisao Uchida. Building a science in Japan: The formative decades of molecular biology. Journal of the History of Biology, 26(3):499–517, Fall 1993. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF01062059.

Ulett:2013:BR

[Ule13] Mark Ulett. Book review. Journal of the History of Biology, 46(2):315–317, May 2013. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-013-9355-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        013-9355-y.pdf.

Ulett:2014:BR

[Ule14]  Mark Ulett.  Book re-

view. Journal of the History of Biology, 47(1):185– 188, February 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-014-9376-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        014-9376-1.pdf.

VanDenWijngaard:1994:FBC

[Van94a] Marianne Van Den Wijngaard. Feminism and the biological construction of female and male behavior. Journal of the History of Biology, 27 (1):61–90, Spring 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058627.

VanHelvoort:1994:CBB

[Van94b]  Ton Van Helvoort.  The

construction of bacteriophage as bacterial virus: Linking endogenous and exogenous thought styles. Journal of the History of Biology, 27(1): 91–139, Spring 1994. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01058628.

Vanpaemel:2010:BR

[Van10] Geert Vanpaemel. Book review. Journal of the History of Biology, 43(3):613– 615, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9252-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9252-6.pdf.

Varma:2009:TGT

[Var09] Charissa S. Varma. Threads that guide or ties that bind: William Kirby and the essentialism story. Journal of the History of Biology, 42(1): 119–149, February 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-008-9156-x;
 http://link.springer.com/
content/pdf/10.1007/s10739-
008-9156-x.pdf.

vandenBerg:2020:ANP

[vdBD20] Hein van den Berg and Boris Demarest. Axiomatic natural philosophy and the emergence of biology as a science. Journal of the History of Biology, 53(3):379– 422, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09609-2.

vanderValk:2014:FET

[vdV14] Arnold G. van der Valk. From formation to ecosystem: Tansley’s response to Clements’ climax. Journal of the History of Biology, 47 (2):293–321, May 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9363-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9363-y.pdf.

Vetter:2006:WOL

[Vet06]  Jeremy Vetter.  Wallace’s

other line: Human biogeography and field practice in the Eastern colonial tropics. Journal of the History of Biology, 39(1):89–123, Spring 2006. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-6543-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-6543-4.pdf.

Vetter:2012:LFR

[Vet12] Jeremy Vetter. Labs in the field? Rocky Mountain biological stations in the early Twentieth Century. Journal of the History of Biology, 45 (4):587–611, November 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9302-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9302-8.pdf.

Veuille:2019:CVS

[Veu19] Michel Veuille. Chance, variation and shared ancestry: Population genetics after the synthesis. Journal of the History of Biology, 52(4):537– 567, December 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-019-09584-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09584-3.pdf.

vonHardenberg:2019:CFI

[vH19] Wilko Graf von Hardenberg. Climate, fascism, and Ibex: Experiments in using population dynamics modeling as a historiographical tool. Journal of the History of Biology, 52 (3):463–483, September 2019. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09579-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09579-0.pdf.

Veigl:2020:FMD

[VHL20] Sophie Juliane Veigl, Oren Harman, and Ehud Lamm. Friedrich Miescher’s discovery in the historiography of genetics: From contamination to confusion, from nuclein to DNA. Journal of the History of Biology, 53(3):451– 484, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09608-3.

vanHelvoort:2019:HSB

[vHS19] Ton van Helvoort and Neeraja Sankaran. How seeing became knowing: The role of the electron microscope in shaping the modern definition of viruses. Journal of the History of Biology, 52(1): 125–160, March 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9530-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9530-2.pdf.

Vicedo:2018:DAN

[Vic18] Marga Vicedo. The ‘disadapted’ animal: Niko Tinbergen on human nature and the human predicament. Journal of the History of Biology, 51(2):191–221, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9485-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9485-8.pdf.

Vicedo:2023:BID

[Vic23]  Marga Vicedo.  Beyond

the instinct debate: Daniel Lehrman’s contributions to animal behavior studies. Journal of the History of Biology, 56(2):251–284, August 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09706-y.

Vinarski:2022:PPE

[Vin22] Maxim V. Vinarski. Pattern without process: Eugen Smirnov and the earliest project of numerical taxonomy (1923–1938). Journal of the History of Biology, 55(3): 559–583, October 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-022-09688-3.

vonMaltzahn:1971:NFO

[vM71] Kraft E. von Maltzahn. New formation of organs in plants — the foundation of plant morphogenesis. Journal of the History of Biology, 4(2): 307–317, ???? 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138315.

Vorzimmer:1969:DQA

[Vor69] Peter J. Vorzimmer. Darwin’s questions about the breeding of animals (1839). Journal of the History of Biology, 2(1): 269–281, ???? 1969. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00137278.

Vorzimmer:1977:DRN

[Vor77] Peter J. Vorzimmer. The Darwin reading notebooks (1838– 1860). Journal of the History of Biology, 10(1):107– 153, ???? 1977. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00126097.

Vorontsov:1992:NPG

[Vor92] Nikolai Nikolaevich Vorontsov.

Nature protection and government in the USSR. Journal of the History of Biology, 25 (3):369–383, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351998.

vanPutten:2020:TCS

[vP20] Kees van Putten. Trees, coral, and seaweed: An interpretation of sketches found in Darwin’s papers. Journal of the History of Biology, 53 (1):5–44, March 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-09591-4;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-09591-4.pdf.

vanPutten:2021:TDP

[vP21] Kees van Putten. Threedimensional phylogeny in two dimensions: How Darwin and other Nineteenth-Century naturalists created three-dimensional figures of the natural system by combining trees of life and maps of affinity. Journal of the History of Biology, 54(4):639– 687, December 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09662-5.

vanWyhe:2014:DAW

[vW14] John van Wyhe. A delicate adjustment: Wallace and Bates on the Amazon and “the problem of the origin of species”. Journal of the History of Biology, 47(4):627– 659, November 2014. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9378-z;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9378-z.pdf.

Ville:2020:CMC

[VWP20a] Simon Ville, Claire Wright, and Jude Philp. Correction to: Macleay’s Choice: Transacting the Natural History Trade in the Nineteenth Century. Journal of the History of Biology, 53(3):377– 378, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09613-6.

See [VWP20b].

Ville:2020:MCT

[VWP20b] Simon Ville, Claire Wright, and Jude Philp. Macleay’s choice: Transacting the natural history trade in the Nineteenth Century. Journal of the History of Biology, 53(3):345– 375, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09610-9.

See correction [VWP20a].

Waddell:2011:BR

[Wad11] Mark A. Waddell. Book review. Journal of the History of Biology, 44(3): ??, August 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9308-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9308-2.pdf.

Waisbren:1988:IME

[Wai88] Steven James Waisbren. The importance of morphology in the evolutionary synthesis as demonstrated by the contributions of the Oxford group: Goodrich, Huxley, and de Beer. Journal of the History of Biology, 21(2): 291–330, ???? 1988. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF00146990.

Waller:2001:GMS

[Wal01] John C. Waller. Gentlemanly men of science: Sir Francis Galton and the professionalization of the British life–sciences. Journal of the History of Biology, 34(1): 83–114, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1010357526856;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1010357526856.
                                        pdf.

Walsh:2015:OMU

[Wal15] Sarah Walsh. “One of the most uniform races of the entire world”: Creole eugenics and the myth of Chilean racial homogeneity. Journal of the History of Biology, 48(4):613– 639, November 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-015-9403-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        015-9403-x.pdf.

Walsh:2020:BRJ

[Wal20] Patrick M. Walsh. Book review: James F. Stark, The Cult of Youth: Anti-ageing in Modern Britain (Cambridge: Cambridge University Press, 2020), 262 pp., $99 hardback, ISBN: 978-1-108-48415-2. Journal of the History of Biology, 53(3):485– 487, September 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09612-7.

Walsh:2021:MFC

[Wal21] Patrick M. Walsh. Making a French connection: Darwin, Brown-Séquard and the epilepsy studies. Journal of the History of Biology, 54 (3):375–401, September 2021. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09647-4.

Ward:2023:MMR

[War23] Zina B. Ward. Muscles or movements? Representation in the nascent brain sciences. Journal of the History of Biology, 56(1):5–34, March 2023. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-023-09703-1.

Waizbort:2021:CFB

[WdLdS21a] Ricardo Francisco Waizbort, Maurício Roberto Motta Pinto da Luz, and Helio Ricardo da Silva. Correction to: The first Brazilian thesis of evolution: Haeckel’s recapitulation theory and its relations with the idea of progress. Journal of the History of Biology, 54 (3):483, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09657-2.

See [WdLdS21b].

Waizbort:2021:FBT

[WdLdS21b] Ricardo Francisco Waizbort, Maurício Roberto Motta Pinto da Luz, and Helio Ricardo da Silva. The first Brazilian thesis of evolution: Haeckel’s recapitulation theory and its relations with the idea of progress. Journal of the History of Biology, 54(3):447– 481, September 2021. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-021-09651-8.

See correction [WdLdS21a].

Weber:2020:BRM

[Web20] Thomas P. Weber. Book review: Michael Worboys, Julie-Marie Strange and Neil Pemberton, The Invention of the Modern Dog: Breed and Blood in Victorian Britain (Baltimore: Johns Hopkins University Press, 2018), 282 pp., £25 (hardcover), ISBN: 978-1-4214-2658-7. Journal of the History of Biology, 53(2): 311–313, June 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-020-09606-5.

Weir:1968:AMH

[Wei68] J. A. Weir. Agassiz, Mendel, and heredity. Journal of the History of Biology, 1(2): 179–203, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351919.

Weinstein:1977:HUW

[Wei77] Alexander Weinstein. How unknown was Mendel’s paper? Journal of the History of Biology, 10(2):341–364, Fall 1977. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00572646.

Weingart:1987:RSB

[Wei87] Peter Weingart. The rationalization of sexual behavior: The institutionalization of eugenic thought in Germany. Journal of the History of Biology, 20(2):159–193, ???? 1987. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00138436.

Weiner:1992:DE

[Wei92] Douglas R. Weiner. Demythologizing environmentalism. Journal of the History of Biology, 25(3): 385–411, Fall 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00351999.

Weidman:1996:PPA

[Wei96] Nadine Weidman. Psychobiology, progressivism, and the anti-progressive tradition. Journal of the History of Biology, 29(2):267– 308, Summer 1996. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00571085.

Weiss:2005:ERR

[Wei05] Sheila Weiss. “Essay review: Racial science and genetics at the Kaiser Wilhelm Society”. Journal of the History of Biology, 38(2):367– 379, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-3560-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-3560-2.pdf.

Weiss:2006:HGP

[Wei06] Sheila Faith Weiss. Human genetics and politics as mutually beneficial resources: The case of the Kaiser Wilhelm Institute for Anthropology, Human Heredity and Eugenics during the Third Reich. Journal of the History of Biology, 39 (1):41–88, Spring 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-6532-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-6532-7.pdf.

Weidman:2009:BRW

[Wei09] Nadine Weidman. Book review: William H. Tucker, The Cattell Controversy: Race, Science, and Ideology (Urbana and Chicago: University of Illinois Press, 2009) xi + 254 pp. Journal of the History of Biology, 42(4):815– 817, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9218-8;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9218-8.pdf.

Weissman:2010:OSD

[Wei10] Charlotte Weissman. The origins of species: The debate between August Weismann and Moritz Wagner. Journal of the History of Biology, 43(4): 727–766, December 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10739-009-9207-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        009-9207-y.pdf.

Weikart:2011:BRS

[Wei11] Richard Weikart. Book review: Sheila Faith Weiss, The Nazi Symbiosis: Human Genetics and Politics in the Third Reich (Chicago: University of Chicago Press, 2010), 392 pp., $45.00. Journal of the History of Biology, 44(1):159–161, February 2011. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9279-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9279-3.pdf.

Wells:1971:SWL

[Wel71]  Kentwood D. Wells.  Sir

William Lawrence (1783– 1867) a study of pre-Darwinian ideas on heredity and variation. Journal of the History of Biology, 4(2):319– 361, ???? 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00138316.

Wells:1973:HCN

[Wel73] Kentwood D. Wells. The historical context of natural selection: The case of Patrick Matthew. Journal of the History of Biology, 6(2): 225–258, ???? 1973. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00127609.

Werskey:1971:HHF

[Wer71] Paul Gary Werskey. Haldane and Huxley: The first appraisals. Journal of the History of Biology, 4(1):171– 183, Spring 1971. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00356981.

Werner:2002:JLP

[WH02] Petra Werner and Frederic L.

Holmes. Justus Liebig and the plant physiologists. Journal of the History of Biology, 35 (3):421–441, September 2002. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021182316792;
http://link.springer.com/
content/pdf/10.1023/A:1021182316792.
pdf.

Wheeler:1981:TPB

[Whe81] A. G. Wheeler, Jr. The tarnished plant bug: Cause of potato rot? — an episode in mid-nineteenth-century entomology and plant pathology. Journal of the History of Biology, 14(2):317–338, ???? 1981.

CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF00141097.

Whitney:2019:BRJ

[Whi19] Kristoffer Whitney. Book review: Joeri Bruyninckx, Listening in the Field: Recording and the Science of Birdsong. Journal of the History of Biology, 52(3):491– 492, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09571-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09571-8.pdf.

Whitney:2020:VSB

[Whi20] Kristoffer Whitney. Valuing shorebirds: Bureaucracy, natural history, and expertise in North American conservation. Journal of the History of Biology, 53(4):631– 652, December 2020. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-020-09616-3.

White:2022:MLD

[Whi22] Paul White. The many lives of Darwin’s letters. Journal of the History of Biology, 55(2): 395–401, August 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09685-6.

Wieland:2006:STA

[Wie06] Thomas Wieland. Scientific theory and agricultural practice: Plant breeding in Germany from the late 19th to the early 20th Century. Journal of the History of Biology, 39 (2):309–343, July 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-006-0006-4;
 http://link.springer.com/
content/pdf/10.1007/s10739-
006-0006-4.pdf.

Wightman:1969:MMS

[Wig69] William P. D. Wightman.

Myth and method in seventeenthcentury biological thought. Journal of the History of Biology, 2(2):321–336, ???? 1969. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF00125022.

Willis:1985:HBC

[Wil85] R. J. Willis. The historical bases of the concept of allelopathy. Journal of the History of Biology, 18(1): 71–102, ???? 1985. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF00127958.

Wilson:2004:BRN

[Wil04] Robert M. Wilson. Book review: Nancy Langston, Where Land and Water Meet: a Western Landscape Transformed. Journal of the History of Biology, 37(3):606– 608, October 2004. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-004-2097-0;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        004-2097-0.pdf.

Wilson:2012:MMM

[Wil12] Emily K. Wilson. Modeling man: The monkey colony at the Carnegie institution of Washington’s department of embryology, 1925–1971. Journal of the History of Biology, 45(2):213–251, May 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9282-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9282-8.pdf.

Williams:2016:IME

[Wil16] Nicola Williams. Irene Manton, Erwin Schrödinger and the puzzle of chromosome structure. Journal of the History of Biology, 49(3):

425–459, August 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9424-5;
http://link.springer.com/
content/pdf/10.1007/s10739-
015-9424-5.pdf.

Wilbanks:2019:SRS

[Wil19] Rebecca Wilbanks. Sophia Roosth, synthetic: How life got made. Journal of the History of Biology, 52(2): 349–352, June 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-019-9563-1;
http://link.springer.com/
content/pdf/10.1007/s10739-
019-9563-1.pdf.

Winther:2000:DVH

[Win00] Rasmus G. Winther. Darwin on variation and heredity. Journal of the History of Biology, 33(3):425– 455, December 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004834008068;
http://link.springer.com/
content/pdf/10.1023/A:1004834008068.
pdf.

Winsor:2001:PSE

[Win01a] Mary P. Winsor. The practitioner of science: Everyone her own historian. Journal of the History of Biology, 34 (2):229–245, June 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1010398900038;
http://link.springer.com/
content/pdf/10.1023/A:1010398900038.
pdf.

Winther:2001:AWG

[Win01b] Rasmus G. Winther. August Weismann on germ– plasm variation. Journal of the History of Biology, 34(3): 517–555, December 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012950826540;
http://link.springer.com/
content/pdf/10.1023/A:1012950826540.
pdf.

Winsor:2024:RUE

[Win24] Mary P. Winsor. Reconsidering utter extinction. Journal of the History of Biology, 57 (4):501–506, December 2024. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-024-09800-9.

Witteveen:2016:SSH

[Wit16] Joeri Witteveen. Suppressing synonymy with a homonym: The emergence of the nomenclatural type concept in Nineteenth Century natural history. Journal of the History of Biology, 49(1):135– 189, February 2016. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-015-9410-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
015-9410-y.pdf.

Welk-Joerger:2022:BBR

[WJ22] Nicole Welk-Joerger. The business with “bugs”: Ruminology and the commercial feed industry in the United States. Journal of the History of Biology, 55(1): 89–113, March 2022. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-022-09674-9.

Wood:2005:SBC

[WO05] Roger J. Wood and Vítezslav Orel. Scientific breeding in Central Europe during the Early Nineteenth Century: Background to Mendel’s later work. Journal of the History of Biology, 38(2):239– 272, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-004-5427-3;
 http://link.springer.com/
content/pdf/10.1007/s10739-
004-5427-3.pdf.

Woiak:2010:BR

[Woi10] Joanne Woiak. Book review. Journal of the History of Biology, 43(2):423–428, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9237-5;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9237-5.pdf.

Wolfe:2001:ERE

[Wol01] Audra J. Wolfe. Essay review: ELSI’s revenge. Journal of the History of Biology, 34(1): 183–193, March 2001. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1010332811017;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1010332811017.
                                        pdf.

Wolfe:2011:OMA

[Wol11] Audra J. Wolfe. The organization man and the archive: a look at the Bentley glass papers. Journal of the History of Biology, 44(1):147– 151, February 2011. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-011-9276-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        011-9276-6.pdf.

Wolfe:2012:CWC

[Wol12] Audra J. Wolfe. The Cold War context of the Golden Jubilee, or, why we think of Mendel as the father of genetics. Journal of the History of Biology, 45(3):389–414, August 2012. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-011-9291-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
011-9291-7.pdf.

Wolfe:2013:VRE

[Wol13] Charles T. Wolfe. Vitalism and the resistance to experimentation on life in the eighteenth century. Journal of the History of Biology, 46(2): 255–282, May 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-012-9349-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
012-9349-1.pdf.

Weber:2006:IOP

[WP06] Bruce H. Weber and John N.

Prebble. An issue of originality and priority: The correspondence and theories of oxidative phosphorylation of Peter Mitchell and Robert J. P. Williams, 1961–1980. Journal of the History of Biology, 39 (1):125–163, Spring 2006. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-005-3052-4;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        005-3052-4.pdf.

Wu:2019:BRP

[Wu19] Harry Yi-Jui Wu. Book review: Patrick Manning and Mat Savelli, eds., Global Transformations in the Life Sciences, 1945–1980 (Pittsburgh: University of Pittsburgh Press, 2018), 366 pp., $45.00 Cloth, ISBN: 978-0-8229-4527-7. Journal of the History of Biology, 52(3):497– 499, September 2019. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-019-09580-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        019-09580-7.pdf.

Winthrop-Young:2023:SPK

[WY23] Geoffrey Winthrop-Young. The social politics of Karl Escherich’s 1933 Inaugural Presidential Lecture. Journal of the History of Biology, 56 (1):65–95, March 2023. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-023-09702-2.

Yang:2013:BR

[Yan13] Andrew S. Yang. Book review. Journal of the History of Biology, 46(1):159– 162, February 2013. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-013-9351-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
013-9351-2.pdf.

Yang:2018:BRS

[Yan18] Andrew S. Yang. Book review: Susan Merrill Squier, Epigenetic Landscapes: Drawings as Metaphor (Durham: Duke University Press, 2017), 280 pp., 21 b&w illus., $25.95 Paperback, ISBN: 978-0-8223-6872-4. Journal of the History of Biology, 51(4):875– 877, December 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-018-9526-y;
 http://link.springer.com/
content/pdf/10.1007/s10739-
018-9526-y.pdf.

Yudell:2000:ERS

[YD00] Michael Yudell and Rob De-Salle. Essay review: Sociobiology: Twenty-five years later. Journal of the History of Biology, 33(3):577– 584, December 2000. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1004845822189;
                                        http://link.springer.com/
                                        content/pdf/10.1023/A:1004845822189.
                                        pdf.

Yi:2008:CVM

[Yi08] Doogab Yi. Cancer, viruses, and mass migration: Paul Berg’s venture into eukaryotic biology and the advent of recombinant DNA research and technology, 1967– 1980. Journal of the History of Biology, 41(4):589– 636, December 2008. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-008-9149-9;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        008-9149-9.pdf.

Yi:2021:DMS

[Yi21] Jongsik Christian Yi. Dialectical materialism serves voluntarist productivism: The epistemic foundation of Lysenkoism in Socialist China and North Vietnam. Journal of the History of Biology, 54 (3):513–539, September 2021. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10739-021-09652-7.

Yi:2025:AGC

[Yi25] Doogab Yi. Asilomar, gene cloning’s origins, and its commercial fate. Journal of the History of Biology, 58 (1):9–20, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10739-025-09803-0.

Young:1998:DRD

[You98] Christian C. Young. Defining the range: The development of carrying capacity in management practice. Journal of the History of Biology, 31(1):61–83, March 1998. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1004205522191;
http://link.springer.com/
content/pdf/10.1023/A:1004205522191.
pdf.

Young:2009:BRR

[You09] Chris Young. Book review:

Ronald L. Numbers, ed., Galileo Goes to Jail and Other Myths about Science and Religion (Cambridge, MA: Harvard University, 2009), x + 302 pp. $27.95. Journal of the History of Biology, 42(4): 823–824, November 2009. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-009-9212-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
009-9212-1.pdf.

Young:2010:BRa

[You10a] Chris Young. Book review. Journal of the History of Biology, 43(2):415–417, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9235-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9235-7.pdf.

Young:2010:BRb

[You10b]  Chris Young.  Book re-

view. Journal of the History of Biology, 43(3):605– 607, August 2010. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10739-010-9250-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10739-
                                        010-9250-8.pdf.

Young:2025:BRR

[You25] Christian C. Young. Book review: Roberta L. Millstein, The Land Is Our Community: Aldo Leopold’s Environmental Ethic for the New Millennium, Chicago: University of Chicago Press, 2024, ISBN: 978-0-226-83448-1, 183 pp. Journal of the History of Biology, 58(1):147– 149, March 2025. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10739-025-09814-x.

Yudell:2010:BR

[Yud10] Michael Yudell. Book review.

Journal of the History of Biology, 43(2):419–422, May 2010. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-010-9236-6;
http://link.springer.com/
content/pdf/10.1007/s10739-
010-9236-6.pdf.

Zallen:1992:RFS

[Zal92]  Doris T. Zallen.  The

Rockefeller Foundation and spectroscopy research: The programs at Chicago and Utrecht. Journal of the History of Biology, 25(1): 67–89, Spring 1992. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01947505.

Zallen:1993:LOJ

[Zal93a] Doris T. Zallen. The “light”

organism for the job: Green algae and photosynthesis research. Journal of the History of Biology, 26(2):269– 279, Summer 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01061970.

Zallen:1993:RBM

[Zal93b] Doris T. Zallen. Redrawing the boundaries of molecular biology: The case of photosynthesis. Journal of the History of Biology, 26(1): 65–87, Spring 1993. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF01060680.

Zallen:2005:BRL

[Zal05] Doris T. Zallen. Book review: Lenny Moss, What Genes Can’t Do, Series on Basic Bioethics, no. 6 (Cambridge, Mass.: The MIT Press, 2003), xx + 228 pp., illus., $34.95, $20.00 paper. Journal of the History of Biology, 38(2): 383–384, Summer 2005. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-005-4212-2;
 http://link.springer.com/
content/pdf/10.1007/s10739-
005-4212-2.pdf.

Zirkle:1968:RLH

[Zir68] Conway Zirkle. The role of Liberty Hyde Bailey and Hugo de Vries in the rediscovery of Mendelism. Journal of the History of Biology, 1(2):205–218, Fall 1968. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF00351920.

Zottoli:2015:MBL

[ZS15] Steven J. Zottoli and Ernst-August Seyfarth. The Marine Biological Laboratory (Woods Hole) and the scientific advancement of women in the early 20th Century: The example of Mary Jane Hogue (1883–1962). Journal of the History of Biology, 48(1):137– 167, February 2015. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-014-9384-1;
 http://link.springer.com/
content/pdf/10.1007/s10739-
014-9384-1.pdf.

Zwier:2018:MAT

[Zwi18] Karen R. Zwier. Methodology in Aristotle’s theory of spontaneous generation. Journal of the History of Biology, 51 (2):355–386, June 2018. CO-DEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10739-017-9494-7;
 http://link.springer.com/
content/pdf/10.1007/s10739-
017-9494-7.pdf.